A Complete Bibliography of The Journal of the American Statistical Association: 1960–1969

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

09 November 2023 Version 1.06

Title word cross-reference

2^{p} [PZ64, PZ68]. 2^{p−q} [Dan62a, Dan62b]. 2k − 1 [Mai66]. 2 × 2 [Koc69a]. χ^{2} [Mey67]. χ_{r}^{2} [Ger69]. D_{mn}, m ≤ n [Kim69]. F [BC68a, Cac65, Don68, Gir65, PP68, Pra68, The65b, Tik67, Tik68a]. J [LaV68c]. K [Con68b, Goo62b, LaV68c, Fis66a, Fis67a, Mae66, Mai66, SD63]. M [HC63, Sen65]. Pr{Y < X} [OCH64]. R [FA65]. r + c = m [FR65]. r ≠ c [FR65]. r × r [IKK69]. s^{c} [FR65]. s^{r} [FR65]. T [Sen69, Boy69, Cac65, DM65, Efr69, Kri67a, Kri67b, Kri68, Kri69, Owe65a, Owe65b, Pre69, SS63a, SS63b, Ste69a, Ste69b, The65b]. T^{2} [HD67, HC63]. v < 1000 [FR65]. v = s^{m} [FR65]. X [GJW68, OCH64]. Y [OCH64].

-Charts [GJW68]. -Class [Fis66a, Fis67a, Mae66]. -Dependent [Sen65]. -Distribution [Kri67b, Kri68, SS63b, Kri67a, Kri69, SS63a]. -Distributions [DM65, The65b, Cac65]. -Method [Sen69]. -Ratio [Pre69, Don68]. -Sample [Con68b]. -Statistics [SD63]. -Test [Efr69, Ger69, Owe65b, Owe65a, BC68a, Don68, Gir65, Tik67]. -Tests [Ste69b, Ste69a].

106 [Whi66b]. 119th [Ano60n]. 120th [Ano61o]. 121st [Ano62l]. 122nd [Ano63l]. 123rd [Ano64l]. 124th [Ano65n]. 125th [Ano66r]. 126th [Ano67o]. 127th [Ano68s]. 128th [Ano69s]. 1949 [M.67e]. 1953 [Bor60b]. 1959 [Ano60n]. 1960 [Ano61o]. 1961 [Ano62l]. 1962 [Ano63l, Sch66a]. 1963 [Ano63k, Ano64l]. 1964 [Ano65n]. 1965 [Ano66r]. 1966 [Ano67o]. 1967 [Ano68s]. 1968 [Ano69s, ABH^{+}69].

60 [Ano66f, Sch63].

Ability [FA64]. Absolute [Her65, TGD67]. Accelerated [Wal66a]. Acceptance [Ell64, FM64, GG61b, Sha65]. Accident [BH65, Joh67b, MSGM65]. Accidents [EG61]. Account [Bro64, Ury66]. Accounting [Dop67]. Accounts [AdJ66, Sch66a]. Accuracy [BS69c, Fle66, NM69, Sla68, SB68, Tho68a]. Acheson [Lav67]. Action [Kau68, Sch69]. Adam [Sch63]. Adaptive [CHG69, Wau66]. Addition [Kab63a]. Additivity [Man61]. Address [Bow64, Han61, Hau63, Lik60]. Adjacent [GW65]. Adjustment [Dav61a, Jor64, Jor65, Jor67b, Lov63, Lov66a, Mar65a, Ner65b, Ner66, Ros68a, Wau66, You68, Lov67]. Advances [Mur63]. Advisory [NS61]. Affected [Bec60b]. Affecting [BC68a]. Against [Bas65a, GP65, Ste69c]. Age [Kar62, KNS67, Maz63, Zel69b, CS62a]. Age-Order [Maz63]. Aged [Ros68b]. Aggregate [Raj64b]. Aggregates [Pow69]. Aggregation [Zel62]. Aging [BS69b]. Agrees [Key66, Key68]. Agriculture [Sch63]. Aircraft [Gal68]. Aitken [Pow69]. ALGOL [Gol63]. Algorithm [GJW68, Tra67]. Allan [Vaj66]. Allied [Rio67, Shu60]. Allocating [Pow69]. Allocation [Eri68, FA65]. Allowing [MB67b]. Almost [Hal61a]. also [Zys69]. Alternative [KG68, Lov66a, Lov67]. Alternatives [GP65, Hog65, Ste69c]. American [Ano60n, Ano61o, Ano62l, Ano63l, Ano64l, Ano65n, Ano66r, Ano67o, Ano68s, Ano69s, Ano61a, Bow64, Zel69b]. Among [Dun61b, Gla64, Han61, Mae66, MSGM65, Sch66b]. Amount [Rao68b]. Amplitude [Koo66]. Analysis [Abe62, AE69a, And64, AK65, Cra64b, Dav61a, Duv66, Dwy67b, FB68, FVCR61, Gab63, Gei63c, Gei64, Gho67, Gir65, Gla64, Gol61a, Gol64, Goo68, GG64, Has68, Hil63b, Hoa69, Jac60, JP61, Jon68a, Jor61b, Jor67b, JT66, Kab63a, KW64, Koc69b, Koe67, Kon62, Kon63, Lad64, Lor60, Lov63, Mal67a, Man61, MSD62, MS63, NG65, PB61, Qua67, Rao65, Rog65, RH65, She69, The63, The65a, Tho61a, Tra67, Win67a, ZT64, Zys63, Zys69, Dwy67a, FRCR61, GG56, GG67, Wol69, D.66, M.67b, S.66]. Analytical [GM67, Sed65, Sed67]. Analyzing [Bec60b, Den63, Fei60, Goo64, The68]. Anderson [Tin61]. Andrew [Gle66, S.66]. Angeles [Ano67o]. Annual [Ano60n, Ano61o, Ano62l, Ano63l, Ano64l, Ano65n, Ano66r, Ano67o, Ano68s, Ano69s]. Anscombe [Arm63]. Answer [War65]. Ante [Bro61c].

Anticipations [EG66, JS69]. AOQL [End69]. Applicable [EG61]. Application [Abe62, Bas63b, Bas64, Har64, JLA69, LB63, PRAL69, PB61, Rao65, Tak62, Tin67]. Applications [AZ69, Bai67, BL67b, Cor66, Dwy67a, Dwy67b, Ell64, Kin65a, Kin65b, Koc69c, MSGM65, Nai69, Sch68, TC65, TG65, Efr67, M.67a]. Applied [PZ64, PZ68]. Applying [WN66]. Appraisal [Lon67]. Approach [Bac69, Bha67, Bha69, BI64, Gei65, Joh61b, Mal67a, NG65, Roy68]. Appropriations [Gre65a]. Approximate [Fis67a, Goo69, GK63, Kim69, Leh61, TV68, TV69, Wal64b]. Approximately [FW64, Ser68]. Approximating [AB67, AB68, AB69a, AB69b]. Approximation [BKvE69, Cur67, FS64, PP68, Pra68, San60, Sla68]. Approximations [Geb69, Jos69, Rab69, SK69, You68]. April [ABH^{+}69]. Area [Fel65, GP68, Wha60, Woo66]. Areas [Mil67]. Argument [Dem64]. Arising [Rah64]. Arrangements [Sha67]. Articles [Ano66h]. Artificial [Shu60]. Asia [EB69]. Asian [NM69]. Aspects [Ehr63c, Kem66, Kem67, Koc69b, Oli66, Ril61]. Assay [Tsu67]. Assessing [Fle66]. Assessment [Win67a]. Assessors [Win69]. Assets [Fis63, Gut60]. Assistance [Dav62]. Associate [Cle69, Gir65]. Associated [PO61]. Association [Ano60n, Ano61o, Ano62l, Ano63l, Ano64l, Ano65n, Ano66r, Ano67o, Ano68s, Ano69s, Ber61a, GK63, Koc69a, Mos68, Ros66b, Ano61a, Bow64]. Assumption [Whi66b]. Assumptions [Cra66, Wag62b, War63b, War67]. Asymmetrical [Sin66]. Asymptotic [BH66, Cac67, Cap61a, Cap65, GP65, Gri62, HM68b, Hol67, Klo67a, Sha61a, SB69, War67, Gri64]. Asymptotically [NS65, SR67]. Attachment [Kor66a, Kor66b]. Attitude [BS69c, Mue63]. Attitudes [FA64]. Attitudinal [FA64, Kis62]. Attribute [Ehr62, Ehr63b]. Attrition [Kum67, Kum68]. Augment [Fie62]. Augmentations [Zys69]. August [Ano61o, Ano67o, Ano69s]. Authors [Ano60f, Ano60g, Ano60h, Ano60i, Ano61g, Ano61h, Ano61i, Ano61j, Ano62d, Ano62e, Ano62f, Ano62g, Ano63d, Ano63e, Ano63f, Ano63g, Ano64d, Ano64e, Ano64f, Ano64g, Ano65f, Ano65g, Ano65h, Ano65i, Ano66k, Ano66l, Ano66m, Ano66n, Ano67e, Ano67g, Ano67h, Ano67i, Ano67j, Ano68k, Ano68l, Ano68m, Ano68n, Ano69k, Ano69l, Ano69m, Ano69n, Ano65d, Ano68e, Ano69e]. Authorship [Bri63, MW63, Tia67]. Auto [RG69]. Auto-Correlated [RG69]. Autocorrelated [ZT64]. Automatic [FH67]. Automobiles [Hua63]. Autoregressive [Bai65, Tho67b]. Auxiliary [Raj65a, Wil62a]. Average [Cag64, Hay60, Leo62, SC62, TF63, Tay64]. Axiomatizations [Lov66a, Lov67]. Aylmer [You62].

B [Alb67, Sta67]. Baccarat [TW66]. Back [Ano60a, Ano60b, Ano61b, Ano61c, Ano61d, Ano61e, Ano62a, Ano62b, Ano63a, Ano63b, Ano63c, Ano64a, Ano64b, Ano65a, Ano65b, Ano65c, Ano66a, Ano66b, Ano66c, Ano66d, Ano67a, Ano67b, Ano67c, Ano67d, Ano68a, Ano68b, Ano68c, Ano68d, Ano69a, Ano69b, Ano69c, Ano69d]. Baggaley [Gle66]. Bakery [Wal65]. Balanced [Low69, Nai67]. Ballot [Tak62]. Banach [Cac67]. Bands [BG66, Dun68, Gaf64, GB67, HRHY67, HG68]. Barton [Rio67]. Baseball [Lin61]. Based [Ber53, GJW68, Gol64, GG61b, Her65, Kab67, Kin62a, Kum67, Kum68, MW67b, Sam66, Sen67, Sen68, Suz68, Swa62, Zac66]. Bases [Put67]. Basic [AZ69, Dal69, LaV68a]. Baye [Hil68a]. Bayes [BI64, Gei65, WD69, Zac69]. Bayesian [ZZK68, ZC69, AK65, Ant65, Bha67, Bha69, BT68b, Cor66, Hil63b, Hoa69, LJZ68, NG65, NH65, Win67a, ZT64, ZC65]. Be [Sam68c, GG56, GG67, Hud66, McE67]. Behavior [JS69, Ram69]. Behavioral [Reg68]. Behrens [Eat69, Pot65, Pre66]. Berge [M.67c]. Best [Gol62b, Gov66a, Har64, Mai66, McE67, Ogu69, TD62, ZE66a, Zys69]. Bet [TW66]. Beta [PP68, Pra68, Tan69]. Better [Ses63, Sha67]. Between [AK69, Agr64, Ber61a, Cac65, CR66, Don68, GG61a, GW65, Har67c, Har68b, Som62, The65b]. Bias [KNP62, KN68, Mor60b, Pea61, Tho66, Tom61, War65, Wau66, Zel62]. Bibliography [Gov64, Gov65, Jac60, Jac63, Shu60, M.67e]. Binary [Ber61a]. Binomial [AB67, AB68, AB69a, AB69b, Bli64, Bry65, FF68, Geb69, Gov62, Gov63a, Gov63b, Lau61, Pac60, PP68, Pra68, Rid62, Rut67, Sam66, San60, Sha61a, Sha66]. Bio [Tsu67]. Bio-Assay [Tsu67]. Bioassay [Ber53, Hsi69]. Births [Sin63, Sin68b]. Bivariate [Fre61b, Gil62, Gum60, Gum61, GG64, GM67, Hal67, HC63, Jen69, KJ63, MH67, Mus69b, Pla65, PRAL69, Ste69a, Ste69b]. Block [BC68a, FW61, Low69]. Blocks [Agr64, Noe63b]. BLS [Mar65a, Ner66, Ner65b, Ros68a, Stu66, You68]. BLUS [AK69, KA68, Neu69a, The68]. Board [Ano61a]. Bode [Goo69]. Book [Abe66, Abr69, Ada63, Ada69, Add69, Ade61a, Ade61b, Ade64, Air61, Ait68, Alb67, All60, All62, All66, And61a, And61b, And62a, And62b, AS64, And67, And60, And63b, Ano66j, Ano66i, Ant66, Ant69, Arn60, Arr61, Arr67, Ast65, Baa62, Bae61, Bai61, Bak69, Ban60, Ban61, Ban62, Ban65, Bar60, Bar65a, Bar66a, Bar67a, Bar65b, BN64, Bar67c, Bas65b, Bas60b, Bat68, Bat60, Bax61, Bea60, Bea64, Bee61, Bee62, Bel68, Ber62a, Ber62b, Ber63, Ber69a, BR61, Bic69, Bil60, Bir63, Bjo64, Ble63, Blo60, Blo65, Blu60, Blu61, Bly60, Bol61, Bon63, Bor60a, Bos64, Bou60, Bow63, Boy62, Bra63a, Bra63b, Bra65, Bra63c, Bre60, Bre61, Bre66, Bri69, Bri60, Bri64, Bro61a, Bro61b, Bro63a, Bro63b, Bro68, Bro60b, Bro61d, Bro61e, Bro63c]. Book [Bro67, Bru62, Bru66, Bru61, Bur62a, Bur61, Bur62b, Bur68a, Bus62, Byr65, C.61a, C.61b, C.61c, C.62a, C.62b, C.62c, C.62d, C.62e, C.62f, C.62g, C.63, C.69, Cad63, Cai68, Cal69, Cam61, Cap61b, Car61, Car62, Cha62, Che60, Chi62, Chi65a, Cho67, Cho64, Cho61, Chu66, Cla60, Cla61, Cli65, Coc61, Coc62, Coc64, Coh60a, Coh60b, Coh62a, Coh62b, Coh68, Col60a, Col62, Con68a, Con61, Coo62, Cor60, Cor67a, Cor67b, Cow60, Cow63, Cox61c, Cox61a, Cro69, Cro60, Cro68a, Cro68b, D.66, Dal67, Dar60, Dar63, Dav64, Dav65, Dav66, Dav61b, DeG66, Dea64, Dem63, Dem65, Dem66, Der61a, Der61b, Der62, Dhr64, Dia67, Die67, Doo63, Dop67, Dor60, Dor61a, Dor61b, DR69, Dru62, Dun60, Dun62a, Dun62b, Dun66, E.68, Eas61, Eas64, Eck61, Edw64, Edw67, Efr67, Ehr63a, Eic68]. Book [Eis62, Ela68, Ell60, Eme62, Eps63, Eps66, Est64, Est66, Eva61, Eva64, F.60a, F.60b, F.60c, F.60d, F.61a, F.61b, F.61c, F.61d, F.61e, F.61f, F.61g, F.61h, F.61i, F.62a, F.62b, F.62c, F.62d, F.62e, F.62f, F.63a, F.63b, Fab61, Fan64, Far63, Fei69, Fei61, Fei62, Fel64a, Fer62, Fin60, Fir61a, Fir61b, Fis61a, Fis65b, Fis65a, Fis68, Fis67b, Fis66b, Fob61, Fol61, Fol69, For60, For62, Fox61, Fox66, Fra66, Fra60, Fra63a, Fre69, Fre61a, Fre64, Fre65, Fri66, Fri65, Fri68, Fri62b, Fro63, Gar68, Gas69, Gei63a, Gei63b, Geo66, Ger61, Gim62, Git69, Gla60, Gla66, Gle66, Gle68, God63, God64, Gol62a, Gol69, Gol60, Gol61b, Gom61, Gor63, Gor61, Got68a, Got68b, Gov66b, Gra66, Gra60, Gre63a, Gre65b, Gre63b, Gre62, Gri61, Gri66, Gro60, Gue67]. Book [Gue65, Gue68, Gus60, H.60, H.61a, H.61b, H.61c, H.62a, H.62b, H.62c, H.63a, H.63b, H.64a, H.64b, H.64c, H.64d, H.64e, H.64f, Had66, Had63, Hae62, Ham68a, Ham65, Han63c, Har63, Har67a, Har69a, Har60b, Har60c, Har60d, Has60, Hau62, Hav67, Hea60, Hee60, Hen61a, Hen61b, Hen63, Hen61c, Hen64c, Her60, Her61, Hes64, Hil65a, Hil65b, Hil64a, Hil66, Hil68b, Hin65, Ho62, Hoc62, Hoc63a, Hoc60, Hoc63b, Hod67, Hoe62, Hof63, Hog67a, Hol61, Hol62, Hol63, Hol68a, Hoo65, Hoo67, Hop61, Hor69, Hou61, Hou62, How63, Hua64, Hun64, Hym61, Iji66, Int68, Ire61, J.63a, J.63b, J.64a, J.64b, J.64c, J.64d, J.64e, Jac66, Jac62, Jaf63, Jaf69, JLA69, Jes67, Jes64, Joh60, Joh61a, Joh62, Joh68, Joh69b, Jon68b, Jon69, Jor61a, Jor62, Jor67a, Jos61, Jos65, Jud62]. Book [Jud64, Kai65, Kal62, Kal65, Kan63, Kao60, Kap62, Kar63a, Kar63b, Kar68, Kat67, Kat61, Kat63, Kat69, Kau65, Kem62, Kem64, Key69, Kie60, Kim62, Kin66, Kin64b, Kin63, Kis61a, Kis61b, Kit60, Kla69, Kle60, Koe66, Koo62a, Koo64, Koo62b, Kos61a, Kos61b, Koz66, Kra60a, Kra65, Kra60b, Kre60, Kre64, Kru60, Kue65, KF63, Kuh60, Kul65, Kur60, Kur62, Lad63, Lam64, Lan67, Lat63, Lat61, Lav66, Lav67, Lea69, Lee64a, Lee68, Lei64, Ler62a, Ler62b, Lev62, Lie60, Lie61, Lie63, Lie64, Lin62a, Lin63a, Lin63b, Lin63c, Lin62b, Lit61, Lit63, Lit65, Loc60, Lon60, Lov66b, Lub61, Lyd61, M.60a, M.60b, M.61a, M.61b, M.61c, M.61d, M.61e, M.61f, M.61g, M.62a, M.62b, M.66a, M.66b, M.67e, M.67c, M.67a, M.67b, M.67d, M.69, Mad69a, Mad68, Mad69b, Mad60a, Mad60b]. Book [Mai61a, Mai61b, Mai63, Mal67b, Man60, Man64a, Man69a, Man63a, Man64b, Mar60a, Mar60b, Mar60c, Mar65b, Mar63a, Mar63b, Mar60d, Mar62b, Mas66, Mas64, Mas62, Mas63, Max63, May60, May61, May62, May63, McC69a, McC68, McC69b, McC69c, McD69, McF67, McK60, McK64, McK61, McK62a, McK62b, McK62c, Mee67, Mei60, Mes67, Mey61, Mey62, Mil69, Mil61a, Mil63a, Mil63b, Mil65, Mil61b, Mil62a, Mil62b, Min68, Min69b, Moh61, Mon67, Mon66, Moo64, Mor61a, Mor66a, Mor61b, Mor66b, Mor68, Mor61c, Mos62, Mos61, Mos63a, MH63, Mos63b, Mue60, Mue64, Mur66a, Mur66b, Mut60c, Mut60a, Mut62b, Mye66, N.61a, N.61b, N.61c, N.61d, N.61e, N.61f, N.61g, N.62, Nei60, Nei64, Nel61, Ner64, Ner65a, Net62, Net66, Net68, Neu65, Neu64, New61, New68, Nic66, Nic61, Nic62]. Book [Nie67, Nis61a, Nis61b, Noe63a, Noe64, Noe65, Nor65, Nou65, O.61a, O.61b, O.61c, O.62a, O.62b, O.62c, O.62d, O.63a, O.63b, O.63c, O.64a, O.64b, O.64c, O.64d, O.64e, O.64f, O.64g, O.64h, O.64i, O.66a, O.66b, O.66c, O.67a, O.67b, O.67c, O.68, Ode69, Oeh60, Oi62, Ols68, Ore61a, Ore61b, Ore62, Orr63b, Orr63a, Ost69, Owe69, P.62a, P.62b, P.62c, P.62d, P.62e, P.62f, P.62g, P.63a, P.63b, P.63c, P.63d, P.63e, P.64, P.65a, P.65b, P.66, P.67, Pag64, Pal67, Pal69, Pap60, Par61a, Par61b, Par62, Par64a, Par64b, Par66, Par69, Pat61, Pat69, Pav63, Pea69, Per61a, Per61b, Pes64, Pet62, Pet67, Phe62, Phi62, Pie68, Pin69, Pol60, Pol61, Pow64, Pow65, Pra61a, Pre65, Pro68, Pro64, Qua64, Qui65, Rao63a, Rao68a, Rap61, Rap63, Rap67, Rat62, Ray68, Red60, Rei67, Rei63, Rei66, Rem65, Res61]. Book [Res68, Rho62, Rho63, Ric64, Ric67, Rie67a, Rio67, Rob62b, Rob62a, Rob63a, Rob63c, Rob63b, Rob64a, Rob65b, Rol63, Roo61, Ros61a, Rug60, S.61, S.65a, S.65b, S.65c, S.66, Sam68a, San67, Sas64, Sav61, Sav62, Sav64, Sca61, Sca63, Sch61, Sch66a, Sch66c, Sch63, Sea69, Sed69, Sen61, Sev63, Sev68, Sha60, She67, Shr60, Shr64, Shu64, Shu65, Shu61, Sid68, Sid69, Sil66, Sim62a, Sim66, Sim62b, Sin69, Sit60a, Sit60b, Sla62, Smi61, Smi62a, Smi62b, Smi64c, Smi64b, Smi64a, Smi66, Smi67, Smi68, Sne61a, Sne61b, Sne62, Som65, Som69, Spa61, Spi64, Spi66, Spr64, Spu64, Sta61, Sta62, Sta64, Ste64a, Ste68, Ste64b, Ste62, Ste64d, Ste61b, Ste63a, Ste64g, Ste69d, Ste63b, Sti63, Sto66a, Sto66b, Sto60, Sto69, Str60, Str61a, Str61b, Sub67]. Book [Sup61, Syl69, Tae67, Tat66b, Tay62, Tay65, Tei61, Tel67b, Tem66, Tep68a, Ter63, Thi63, Tho61b, Tho62, Tho67a, Tho63b, Tho64, Tia67, Tie62, Til65, Tob69, Tri67, Tro63, Tur65, Vaj64, Vaj66, Van69, Van63, Vei64b, Vei64a, Vei68, Ven61, W.60, Wag60, Wag61, Wag62a, Wah69a, Wah69b, Wal61, Wal63, Wat60a, Wat60b, Wat63, Wat65, Web67, Wei61a, Wei61b, Wei60, Wei63a, Wei63b, Wei66, Wei67, Wel63, Wen65, Wes67, Wes60, Whi66a, vNW68, Whi66c, Wil61b, Wil65a, Wil65b, Wil67c, Win63, Win64, Win61, Wit62a, Wit62b, Wol60, Wol65, Wol61, Wor63, Wri60, Yan60, Yan62, Yen68, Yer61, Zab66, Zar60, Zar67, Zeh61, Zel65, Zie62, Zin61, Zor60]. Borrowing [Tho68a]. Both [Bai65, Don68, Hal64a, Mac64, Par67, Hal61b]. Bound [Ber61b, Col60b, Sam66]. Bounded [GC63, Hoe63]. Bounds [Agr64, Ajg68, BM65, BTM69, Har61, Jos69, Man69b, OCH64, Suz68]. Box [Cac67]. Bradley [RK68, RK67]. Brayfield [Edw67]. Breakdowns [Woo66]. British [Fit60]. Bulb [MP66a, MP66b]. Bureau [Woo63]. Busacker [M.67a]. Business [Bor60a]. Buying [Jus66].

C [Efr67, Jac66, Rei67]. Calculating [Hah69, Kni66, SC62]. California [Ano61o, Ano67o]. Can [Sam68c]. Canadian [FGP67]. Cancer [BE60a, Cor67a, Gre69, Hay65, BE60b]. Cannot [GG56, GG67]. Capacity [Hic64, Lee62]. Capital [BM67b]. Carlo [Gri68, Tho67b]. Cartes [Ber69b]. Case [Bas63a, CS62a, Gei63c, Kab64, NM69, Noe67, Smu60, Gei64]. Cases [Bas61, EB61, Kab63b, Wal64b]. Cash [LaV68a, LaV68b, LaV68c]. Categorical [Bha66, Gab66, Hoa69]. Categorized [Blu68]. Cauchy [Bar67b, Bar68, Bar66b, Blo66, Rid60a, RFT64]. Causal [Bas65a]. Censored [Hal60a, Hal60b, HM66a, HM66b, HM67a, HM67b, HM68a, HM68b, Tik68b]. Censoring [Gov66a, Man69b, Sho68]. Census [Ros68a, Tar62, WP69, Woo63, You68, Zac62, Zel64]. Censuses [EB61, FIM64, Kum67, Kum68]. Central [BA68, Gue64, Mey67, SS63a, SS63b]. Centrality [Mey67]. Century [Fit60]. Certain [Hog60a, Hol68b, Joh67b, Zys69]. Chain [Bas65a]. Chance [CC62, Sin68b]. Change [AG61, GP68, Lee64b, Moo69, Reg68]. Changes [Cox61b, Ein62, Mai60, Mus69a, WP69]. Changing [Han64, Han66, Ste67b]. Characteristics [Ber61a, BC68d, FFGM69]. Characterizing [Kot66]. Chart [Hil67b]. Charts [GJW68, Joh61b, Hil68c]. Chebyshev [Rab69]. Chi [Boy65, Gue64, Ham63, Jen69, Kot66, Man63b, MC66, Sla65, Sla66, Sla68, Boy66]. Chi-Square [Boy65, Gue64, Ham63, Jen69, Man63b, Sla65, Sla66, Sla68, Boy66]. Chicago [Ano65n]. Childless [EB61]. Children [MSGM65]. Choice [BI64, Edw67, Fis67a, Ham68b, Kim60, Wil64, Koz66]. Choices [Bow68]. Chronic [Cor67a]. Chung [McF67]. Cigarette [Gre69]. Circle [Gil62]. Circular [BH66, Har60a, Kam62, Mor59, Mor60a, Mor60b, Ste69c]. City [Ano62l]. Clarke [Vaj66]. Class [Ajg65, Bas65a, Cro64, Dav61a, EG61, Fis66a, Fis67a, Gir65, Hen64a, HM69, Jen69, Mae66, MSD62, Nai64, Pla65, Ram65, Sen67]. Classes [Ham63, Ham68b, Har68a]. Classical [Bas60a, Bas61, Bas63a, Cor66, Har68a, Tak69]. Classification [Bry65, Hai66, Har67c, Har68b, HG65, KN68]. Classifications [GK63, Hai66]. Classified [Goo68]. Claude [M.67c]. Cleveland [Ano64l]. Clinical [Cor66, CHG69, NG65, Zel69a]. Cloud [Bro60a]. Cluster [Nau65]. Clustering [Kin67, Rog65]. Clusters [EH69, Nau66]. Cochran [Kab63a]. Coding [FIM64]. Coefficient [Dav67, Gla62b, Kha68, Kra63, Ric68, Sea64, Sen68, Sha61b, TF63, Tay64]. Coefficients [AR61, DC69, Hai66, HH68b, LL61, Scl68, HH69, LL62]. Collecting [Hoc67]. Columns [FR65]. Combination [Hal61a, Kin62a, MB67b]. Combinations [MG69, Sio64]. Combinatorial [MS65]. Combined [Sha67]. Combining [Gei65]. Comment [And66]. Comments [Arm63, Fra69, NS61]. Committee [NS61]. Common [Agr64, FFGM69, PP68, Pra68, Zac66]. Community [PB67, PB69]. Commuting [Sch60]. Comovements [GG61a]. Company [Nel60]. Comparative [BS69a, SWC68]. Compare [Cag64]. Comparing [Ber61a]. Comparison [AK69, Hoc67, Laz61, Mar65a, McF67, Mor59, Ner65b, Ner66, Pre66, RK67, RK68, Rao69, Sla65, Sla66, SK69, Tin65, Mor60a]. Comparisons [Bru60, Cra64a, Dun61b, Gab66, Geb69, Mai66, Sch66b, Sen69, WD69]. Compartmental [BC68b]. Competing [BE60a, BE60b]. Competitors [Hol67]. Complete [HM66a, HM66b, HM67a, HM68a, Nat67, Tat61]. Completely [BC66, Har68b, Koc69b]. Completely-Random [Har68b]. Complex [Kis68]. Compliance [Adl63]. Components [Dun61a, Gho65, Hil65c, KMZ69, Low69, Mai60, Mas65]. Componentwise [RV61, RV63]. Composite [CR66, NS65, Ney69]. Composition [Aig65]. Compound [Hoa69, Pre68b]. Compounded [Dun69]. Compounding [Pre68b]. Compression [EP65, EP67]. Comprised [Rut66]. Compromise [Hur62]. Computation [End69, Joh67a, SD63, Sha61b, Joh69a]. Computational [BK69]. Computations [Gus61]. Computer [Klo66, Kni66, Lon67, M.67b, NBS67, NWS67, Stu66, Tei65a, Tei65b, Klo67b]. Computerized [O’R69]. Computers [FMR62, Sch68]. Computing [Cro64, LL61, LL62]. Concentration [Ein62, Gla62b, HT67, Kam68, Nel60]. Concept [Fel65, Lee62]. Concepts [CM69, Kul67, Jac66]. Concerned [Wil61a]. Concerning [BT68a, Bas63b, Bas64, Zei65]. Concomitant [CR66]. Concordance [Hay60]. Condition [McE67]. Conditional [Hog60b, Mar64, Tat66a]. Conditional-Normal [Tat66a]. Conditions [Efr69, Whi66b]. Conference [Blo69, Sch66a]. Confidence [AB67, AB68, AB69a, AB69b, Bir61, BG66, Bro69, Che66, DZ69, Dun68, FA67, Gaf64, GB67, HRHY67, HG68, Har61, Her65, JJ69, Kra63, LRT61, Mac64, Mad62, Man69b, MB67a, McC65a, Noe67, OCH64, Pac60, Pra61b, Pre66, PS68, Sid67, Wal67]. Confounding [Sha67]. Connection [Ban63]. Consider [MG69]. Consistency [JP61, vE64]. Constants [Hur62]. Constrained [MFLR68]. Constraints [CC62, HG68]. Constructing [Add63, Kra63, Sam68b, Ses63]. Construction [Agr66, BMN63, Kum67, Kum68, Mus69a]. Consumer [BS69c, Fer65, Fer66, FFGM69, Fis63, FA64, Hua63, Hur62, Jus66, Mue63, Ril61, Sin66, WN66, Wil67a, Gut60]. Consumption [BM67a]. Contains [Rao68b]. Contemporaneously [Par67]. Context [RG69]. Contingency [Goo64, Goo68, Ham68b, IKK69, Koc69a, MC66, Mos68, Nag69]. Continuous [Hil64b, Whi66b]. Contracts [FM64]. Contrasts [DM65, WGF63]. Contributions [Edw67, TF63]. Control [FIM64, Gas68, Joh61b, Lav67, Lor60, NS61, Dop67]. Controlled [Zel69a, GG56, GG67]. Controversy [Gre69, Spi60]. Convergence [Coa68]. Cooperation [Han61]. Correcting [Tay64]. Correction [AB69b, Bar68, Ber61b, Min69a]. Correlated [Gei65, Hil67a, Mad63, Par67, RG69]. Correlation [AR61, Cho66, Cra64b, Don68, DC69, Fie62, Gra63, GW65, Kra63, LL61, LL62, Rob66, TF63, Tay64]. Correlational [Gle66]. Correlations [Gus61]. Corrigenda [AB69a, AZ69, Ano60c, Ano60d, Ano65d, Ano66e, Ano66f, Ano67e, Ano68e, Ano68f, Ano69e, Bac69, BW66a, Bar67b, Bas64, Bas65c, BE60b, Ber69b, BK66a, Bha69, Bir66, Blo69, Bor60b, Boy66, BK69, Chi65b, CKMP69, Cho69, DH63, Dan62a, DN66, Dwy67a, Ede63, Ehr63b, EP67, Elk69, Fis62, FRCR61, Gei64, Gen66, GQ67, GG56, GG67, Goo62a, GH63, GH69, Gov63a, Gov65, Gri64, HS69, Hal60a, HB69, Han63a, Han66, HM66a, HM67a, HM68a, Hen64b, HH69, Hil68c, Joh69a, Jor65, Kem67, Kin64a, KH63, Kis69, Klo67b, Kon63, Kor66a, Kri67a, Kri69, Kum68, Lan62, Lan68a, Lau63a, LH69, LL62, Lil69a, Lov67, MP66a, Mar65a, Mat62a, Mic60, Mor60a, Mut62a, Nea67, Ner66, OF66b, Owe65a, PB69, Ram65, RK68, RV63, Ros66a, SS63a, Sch65]. Corrigenda [Shi64a, ST61, Sla66, Ste69a, SN64, Sve69, Tei65a, The65b, TG65, Tik68a, TV69, Wol69, Yas69, ZZK68, ZC69, dP64, God64]. Cost [And63a, Gal68, Hen64b]. Cost/Quantity [Gal68]. Costs [And66, O’R69]. Counting [Bue61]. Countries [EB69, NM69]. Couple [Sin63, Sin68b]. Course [AZ69]. Covariable [Bro64, Mal67a, Ury66]. Covariance [BG69, Con69a, Gei63c, Gei64, Gla67, Kab67, Lor60, Mal67a, Mar64, OF66b, OF66a, Qua67]. Covariances [Mar64]. Cramer [Tho66]. Credit [Ben66, Fis63, MF63]. Criminalistics [Kin65a, Kin65b, Mod63]. Criteria [Bha66, BS69b, FR67, Har64]. Criterion [BI64, HC63, Kim69, SC62, TVW68]. Critical [BL69, BI64, GW66, Hoc67, Mil64b, Ste69b, Ste69a]. Cross [GK63, Goo68, Hai66, Lee64c]. Cross-Classified [Goo68]. Cross-Section [Lee64c]. Crossroad [Bow68]. Cruon [Arr67]. Cubical [Elk68, Elk69]. Cumulant [TV68, TV69]. Cumulative [Jae64, Joh61b]. Cumulatives [KT68]. Current [Kar62]. Curriculum [Sch68]. Curtius [Bri63]. Curve [BR69, BC68d, Cra64b, Cra64a, Fis61b, Fis62, Hay65, Lec64]. Curve-Shape [BC68d]. Curves [Bir61, Hud66, Pik69]. Cycles [Bor60b, Bor60c, Bor60a].

D [Gol63, Rio67, GC64]. D.C [Ano60n, Ano68s]. Daniel [Gol63]. Danish [Gol64]. Data [Bha66, Blu68, Bro61c, BS69c, DAS69, EP65, EP67, Fei60, Fie62, FH67, FR67, Gab66, Goo68, Gre65a, Has68, Hoa69, Hoc67, Jae64, Jon68a, Jor67b, Key66, Key68, Kni66, Lad64, Lan68b, Lee64c, May68, Min69a, MS63, NM69, NW64, NG65, PO61, Pre68a, Rao65, Sch60, Ste67a, SB68, Stu66, Wha60, Lan68a]. David [Rio67, Tia67]. Death [Tom61]. Deaths [Ros66c]. Debt [Ben66]. Decapitated [Gov62, Gov63a, Rid62]. December [Ano60n, Ano62l, Ano65n, Ano68s]. Decision [All63, Ant65, Chu65, Gre65a, PRS64, Sto63, ZC65, ZZK68, ZC69, Edw67]. Decisions [LaV68a, LaV68b, LaV68c]. Defective [End69]. Degree [Man63b]. Delayed [And64]. Delivered [Ano60n, Ano61o, Ano62l, Ano63l, Ano64l, Ano65n, Ano66r, Ano67o, Ano68s, Ano69s]. Demand [Fer66, Lee64b, Lee64c, Wal65, Zar68]. Demands [Ney60]. Demographic [Maz63]. Denominator [Don68]. Densities [KT68]. Density [Elk68, GC64, Elk69]. Dependence [Har67c, Har68b, MM68]. Dependent [BW66a, BW66b, MFLR68, Sen65]. Deposits [Fer65, Fer66]. Depreciation [Wal66a]. Derivation [Gue64, Sär64]. Derivatives [Dwy67a, Dwy67b, TD69]. Description [Lau63a, Lau63b]. Design [Ant65, BC66, BC68a, BW66b, BM64, BHJ60, Ehr62, Ehr63b, Eri67, FGP67, FW61, GS65, GJW68, HG65, Jus66, Low69, NBS67, Sin68a, Tsu67, BW66a]. Designing [GP68, Sed67]. Designs [Agr64, Agr66, BK61, Cle69, FR65, FR66, Gir65, KW64, Nai64, RG64, Ros61b, Zac69, Ram65]. Destructive [Hil64b]. Detecting [Rog61]. Detection [GW65]. Determination [GJW68]. Determine [Mar62a]. Development [BT69, Bow68, FMR62, MF63, Nea66, Nea67]. Deviant [Sam68c]. Deviates [CC62]. Deviation [Col60b, Har61, Her65, LRT61]. Deviations [Fis61b, Fis62, Rid62]. Diagrammes [Ber69b]. Dichotomous [MS65]. Difference [CL69, Gla62b, O.64a]. Differences [Cur67, Dav61a, Dun61a, Edg61, KH59, KH63, Rog61]. Different [Ber61a, CA63, Fis66a, Sea63]. Differential [FM64, MSD62]. Differentiation [Neu69b]. Difficulties [Dem64]. Digest [EB69]. Digital [FMR62]. Directions [Ste69c]. Directors [Ano61a]. Dirichlet [TG65, CM69, TC65]. Disarray [Sha61b]. Discrepancy [AdJ66]. Discrete [BH65, Con69b, DN65, DN66, DM68a, Kin62a, Lan61, Lan62, McG69, Noe67, Tel67a]. Discrimination [CR66, Eas69, Gil68]. Discussion [ABH^{+}69, SBC^{+}62]. Disease [GBB67]. Diseases [Cor67a]. Disparities [BM67a]. Dispersion [Tra67, vE64]. Displaced [Sta67]. Disputed [Tia67]. Distributed [DN65, DN66, Fei60, Tin67]. Distribution [Arn68, BW65, BT68a, BP67, Bar66b, Bas63b, Bas64, Bas65c, Bas65d, BH65, Blo66, BKvE69, BA68, Cac67, Che66, Coh60c, Coh60d, CS62b, CM69, Con69b, Cox61b, Cur67, DN65, DN66, Dun69, Ell64, FF68, FS64, Fis61c, Fre61b, Geb69, Gho69, Gil62, Gla62a, GS68, Gue64, GG61b, Hah69, Hal67, Har64, HM67b, HM68b, Hil63b, Hil68a, Hoa69, HS68, Kha68, KJ63, Kim69, Kot66, Kri67b, Kri68, Lau63a, Lau63b, Lil69b, Mac64, Mah67, MW67a, MO67, Mat62a, Mat62b, McC65a, McG69, McN67, MH67, Mus69b, Nau65, Nau69, Pac60, PB67, PB69, Pre69, Ric68, Rid60a, Rid62, Rid64, Ros66b, Rut66, Saw69, SS63b, Sha66, Sho68, Shu68, SB69, Sta67, SC62, Tar66, TF63, Tei65a, TGD67, Tho69, TG65, TC65, Tik68b, Var69, Wal62, Wal64b, Wei62b, vE64]. distribution [Tei65b, Bar67b, Bar68, Kri67a, Kri69, SS63a]. Distribution-Free [Gla62a, Mac64, McC65a, McN67, Wal62, vE64]. Distributions [AR61, Arn67, BC64, BM65, Bar66c, Bas60a, Bli64, BFHH66, BC68d, Bur68b, Cac65, DM65, Eag68, EG61, FR68, Gum60, Gum61, GG64, GM67, HH68a, Has69, Hil63a, Hog62, HT63, HC63, Jen69, Kab63b, Kab64, KNS67, Kin62a, Lan61, Leh61, Lik67, Mus69b, Pla65, Rah64, Sam66, Sär64, Sha61a, Sid67, SB69, Tan69, TV68, TV69, Wil61a, Win67a, Zac66, Zei65, Lan62, The65b]. Distributive [Bre62]. Disturbance [Koe67]. Disturbances [Bai65, Par67, TN61, The65a, The68]. Diversification [And63a, And66, Hen64b, Nel60]. Dividend [Cox61b, FB68]. Divisible [Agr64, KW64]. Dixon [Nic66]. Do [Tuk60]. Documents [Boi69]. Doig [M.67e]. Domain [CW68]. Domestic [Ein62]. Dosage [Cra64a]. Double [Gov66a, HBA69, Sed65, SS65]. Doubly [BA68, HM68b, Kri67a, Kri67b, Kri68, Kri69, Sha66, Swa62]. Down [Dix65, Hsi69]. Drawn [Con66]. Dual [Lik60]. Dubins [DeG66]. Due [McN67, Wau66]. Dummy [War63b, War67]. Duncan [GJW68, Lav67].

Duplication [Rao68b]. Durable [Fis63, Lee64b]. Durbin [AK69]. During [Dav62, Lin61]. Dynamic [Con69a]. Dynamics [Syk69]. Dynamique [Arr67].

Early [All63, Chu65, Mil60]. Earnings [Bor66]. Ecology [PB67, PB69]. Econometric [Bas63b, Bas64, Con69a, GS61, Jor67b, LNF69, Ram69]. Economic [AR61, GJW68, Han63a, Han63b, Jor64, Jor65, Lov63, Moo69, NBS67, Ste67a, SB68, Wal66b, Ano60c]. Economica [Sto66a]. Economics [Sch69]. Edith [Die67]. Editing [FH67, Stu66]. Editor [AB69b]. Education [Bow65]. Educational [Dun61a]. Effect [BT68b, Gib64, Koc69a, Mor60b, NMR65, Sea66]. Effective [Ste64c]. Effects [Ban63, Har67c, Nai67]. Efficacy [EG66, Joi69]. Efficiencies [Wal64a, ZE66a]. Efficiency [Bec60a, BL67a, Cap65, CKMP68, GP65, Hol67, Kel69, Klo67a, KMZ69, McN67, Noe63b, PRAL69, Swa62, CKMP69]. Efficient [Cra64b, Par67, Zel62]. Efforts [Cla66]. Elasticities [Bre62]. Electronic [Lon67]. Elementary [PRS64]. Elements [Rut66, D.66, Sch66a]. Eliminating [War65]. Elimination [Agr66]. Ellipsoids [Sha65]. Emphasis [Fis63]. Empirical [BC66, BC68a, FB68, GG64, HC63, RB69]. Employed [Mil67]. Employing [Ful66]. Employment [Bor60b, Bor60c]. Engel [Cra64b]. Engineering [Rei67]. English [Die67]. Entries [EB61, Goo68]. Entry [Kor66a, Kor66b]. Enumeration [Zel64]. Enumerators [EB61]. Epidemic [Hil63b]. Epidemiological [Cor67a]. Equal [Add67, Nai69, Raj64b, Sla68, Zac66]. Equality [Bar66c, Hog62, HT63, Kha60, Koc69c, Wei62b]. Equation [Bas63a, Kab64, McG69]. Equations [CL69, Cra66, Kak67, Par67, Tel64, Zel63]. Equilibrium [Con69a]. Equivalence [Bha66, Jae64]. Equivalents [LaV68a, LaV68b, LaV68c]. Era [Tei65a, Tei65b]. Ergodic [DO67]. Error [Ajg68, BH66, Bor66, Ede61, Ede63, Fis66a, Gol61a, GC63, Hal61b, Hal64a, Har60a, Kam62, Min69a, Mor59, Mor60b, NMR65, O’R69, Put67, Sch62, TVW68, WTV69, WGF63, Wil67a, Zys69, Mor60a, Sch65, O.64a]. Error-Variance [Ajg68]. ’Errors [Lan66, Bai65, BW69, Bry65, Don68, DJ66, EB61, Hil67a, Kis68, Koc69a, May68, NW64, RG69, Rog61, Siv69, ZT64, Zel64]. Essentials [Yas69]. Establishment [Ein62]. Establishments [Rog65]. Estate [BMN63]. Estimate [Ede61, Ede63, GC63, GC64, Heg67, Mey67, Raj64a, She69]. Estimated [Hai66, Hud66]. Estimates [Cun64, Gei65, Gov66a, Hal61a, Kam62, Koe67, MM68, Mor59, Mor60a, Mor60b, MB67b, Rao69, Rob64b, Ros61b, Sen68, Sha61a, Sin68a, SK69, Swa62, Tom61, Woo66]. Estimating [Ans61a, Ber53, Bli64, Coh60c, Gri68, Hil63a, Kha68, Maz63, Pre68a, Sha66, Sho69, Tar62, Tho68b, Tho69, Tik68b, Wei63c, Zac62, Zel62]. Estimation [AE67, Aig65, Arn68, BC68b, Bha67, Bir64, Bir61, BTM69, Blo66, BC68c, BT68b, BFHH66, BHJ60, Coh60d, CS67, DJ68, DM65, Elk68, Elk69, FF68, Fel64b, FKCR62, Fis61c, Ful66, HM63, Hal64a, Hal64b, Hal65, Han63a, Han63b, Han64, Han66, Har68a, HM66a, HM66b, HM67a, HM67b, HM68a, HM68b, HRK69, Has69, Hen64a, Hic63, Hic64, HS68, Hog67b, KJ63, Kin62b, Koo66, KT68, Lau63a, Lau63b, LJZ68, Les63, MFLR68, Mos68, Neu69a, Oli66, Par67, Raj65b, RV61, RFT64, Rut66, Rut67, RB68, Sär68, ST69, SS69, Sea64, Sio64, Siv69, SC65, Tel64, Tho68a, Var69, Wal66a, WGF63, Wil62a, Zac66, Bha69, Kin64a, RV63]. Estimator [Ajg68, Ati69, Fis67a, Ogu69, PRAL69, RM67, Ric68, Sea66, SM69, Wil62b, Wil64]. Estimators [Ajg65, Bar66b, Bar67b, Bar68, Bas61, Bas63a, Bas63b, Bas64, BW69, BL67b, BH66, BTM69, CW68, Cra66, Fis66a, GH58, Har64, HH68b, HH69, Kab63b, Kab64, Kak67, KMZ69, KG68, Lan68a, Lan68b, Low69, Mae66, McE67, McG69, MG69, Mic59, Pow69, RB69, Sam66, Saw69, Scl68, Ses63, Sha67, SR67, Suk62, Tin65, Wil67b, ZE66a, ZE66b, Zel63, dP61, dP64, GH63, GH69, Mic60]. Eugene [O.66a]. Evaluation [Bro60a, Cab60, Fel65, Goo69, GS68, LaV68a, LaV68b, LaV68c, MF63, Ros68a, Tar62, Whi66b, Win69]. Evasive [War65]. Evidence [Lee64c]. Evolving [Net65]. Ex [Bro61c]. Exact [Bas61, Bas63a, Bas63b, Bas64, BG69, Con66, DN65, FS64, Goo60, Goo62a, Har61, Kab63b, Kab64, Kim69, Leh61, Man69b, MW67a, Ric68, Saw69, Tar66, WTV69, Zel63, DN66]. Examples [SK69]. Exceedance [DAS69]. Exchange [NO66]. Exchanging [LaV68c]. Exclusion [Tak67]. Expectancy [Wil63]. Expectations [Hog60b, Nau66, Wau66]. Expected [DS65, FW64, Rid62, Sla68, Wei62a]. Expenditures [Fis63, NW64]. Experience [BM67b, Stu66]. Experiment [Jus66, MB67b]. Experimental [Ehr62, Ehr63b, Kem66, Kem67, Li62, NBS67]. Experiments [Ban63, BK66b, GG56, GG67, Koc69b, MB67b, NBS67, NS65, PZ64, PZ68, RK67, RK68, BK66a]. ’Explicit [Bas65a]. Exploration [Ram69]. Exploratory [Mas65]. Exponential [Ans61a, Arn68, BP67, Bas65c, Bas65d, BE60b, BE60a, Bir54, Bir66, Fre61b, Gla67, Gov66a, Gum60, Has69, Hil63a, HM69, HT63, Lik67, Lil69b, MO67, Rut66, SC65, Var69]. Exponentially [Mut60b, Mut62a]. Exposition [PRS64]. Exposure [THS61]. Extended [McC65b, Mil64b]. Extension [Fre61b, Ger69, Hal60b, Kab63a, Hal60a]. Extensions [Man63b]. Extent [Gol64]. Extremal [GG64, GM67]. Extreme [Chu65, HM68b, Has68, PRAL69, Ros68b]. Extreme-Value [Has68]. Extremes [Wal64b].

F [Rio67]. F. [Arm63]. Factor [BS69a, Goo64, Nag69, Sed67, She69]. Factorial [Ban63, BK61, PZ64, PZ68, Sha67]. Factors [BC68a, Bec60b, Hah69, PZ64, PZ68]. Failure [BP67, EB61]. Failure-Time [MH67]. Failures [Mil60, Rut67]. Fall [Hen64a]. Fallible [Lor60]. Familial [GBB67]. Families [Dav62, GW65]. Family [BC64, Has69]. Favorable [TW66]. Federal [Bow64, Bow68, BT62]. Federalist [Tia67]. Fellegi [Bre67]. Female [EB69, Smu60]. Ferguson [S.66]. Fertility [Maz63]. Few [MW67b]. Fiducial [Dem64]. Field [CJ69]. Fifty [Bar65c]. Filters [Ner65b, Mar65a, Ner66]. Final [NS61]. Financial [Bor60a, Fer65, Fer66, FFGM69, May68]. Finite [Ajg68, Bas60a, Bas61, Bas63a, Bas63b, Bas64, BL69, Con69b, Dun68, Gri68, Has69, Hoa69, Kin62b, Kin64a, Mat62a, Mat62b, Mic59, Mic60, Nat68, Nat69, RM67, Roy68, San60, Tho67b, Zac69, Zel63, M.67a]. Finite-Time [Gri68]. Firm [Ste64e]. First [CKMP68, CKMP69, Cle69, Edg61, Gas68, HM68b, Kel69]. First- [Kel69]. First-Median [Gas68]. First-Order [CKMP68, CKMP69]. Fisher [Bar65c, Boy69, Dem64, Eat69, Pot65, Pre66, You62]. Fisher-Type [Pot65]. Fit [GP65, Rie67b, Sla65, Sla66, Sla68, Suz68]. Fitting [BR69, Cra64a, Fis61b, Fis62, Hal61b, Hud66, Jae64]. Five [Stu66]. Fixed [CS62b, DO67, Hal60a, Hal60b, Zac69]. Fletcher [Vaj66]. Folklore [Osb67]. Follow [CKMP68, CKMP69, Fei60, TK61]. Follow-Up [Fei60, TK61]. Food [Lad61, PB61]. Force [FGP67, Kor66a, Kor66b, Smu60]. Forecasting [BT62, BS69c, Moo69, Mue63, Ste61a, Ste67a, Ano60c]. Forecasts [Hic63, Mut60b, Mut62a]. Form [Coa68, Zar68]. Formation [Mai60]. Forming [Raj64b]. Forms [Koc69c]. Formula [Con66]. Formulae [Kab63a]. Formulas [Gla62b]. Formulation [Cab60]. Foundations [Bir62, PRS64, SBC^{+}62]. Four [Rao69]. Fourier [KT68, Net65]. Fractional [Add63, Ban63, Dan62a, Dan62b, PZ68, WA67]. Frame [Rao68b]. Framework [GAESH69]. Frederick [Tia67]. Free [Gla62a, Mac64, McC65a, McN67, Wal62, vE64]. Freedom [Man63b]. Frequencies [Gab63, Sla68]. Frequency [Add67, Bas61, Bas63a]. Friedman [Ger69]. Front [Ano60e, Ano61f, Ano62c, Ano64c, Ano65e, Ano66g, Ano67f, Ano68g, Ano68h, Ano68i, Ano68j, Ano69f, Ano69g, Ano69h, Ano69i, Ano60d]. Fruits [Wal65]. Fry [Rei67]. Function [Ber53, Cro64, Gib64, Lik60, Lik67, Oli66, Rid64, Rio67, Shu68, Tra67, Wal66a, War63a]. Functional [Zar68]. Functions [Aiy63, Bas61, Bas63a, Bas63b, Bas64, GS68, HM63, Hal64b, Hal65, McF67, Mus69b, ZE66a, ZE66b, Sve69]. Further [And66, Koe67, NS61]. Future [Bow68].

G [Alb67, M.67e, M.67a, Rio67, Sto66a]. Gains [BM67b]. Gamble [DeG66]. Game [Lin61]. Games [M.67c]. Gaming [Shu60]. Gamma [BW65, Dav67, GG61b, LB63, Ros66b]. Gauss [LO66]. Gaussian [Gri68, Shu68]. GCL [Bas64, Bas63b, Kab63b, Kab64]. General [BL67b, BC68d, Bur68b, FR65, Sch66b, Sha65, Suz68, Wal62]. Generalisations [Rah64]. Generalization [CM69, LO66, RK67, RK68, Sen69, Tak62]. Generalized [BW65, Bas60a, Bas61, Bas63a, Cac67, FR66, Gol62b, HH68a, RM67, Tan69]. Generate [OF66b, OF66a]. Generating [Cro64, Osb67]. Generation [Sho68]. Geoffrey [Vaj66]. Geometric [MW67a]. George [McF67, S.66]. German [Die67]. Gestion [Arr67]. Ghana [Kpe69]. Ghouila [M.67c]. Ghouila-Houri [M.67c]. Go [Tuk60]. Goals [Dop67]. Goodman [Dav67, Som62]. Goodness [GP65, Rie67b, Sla65, Sla66, Sla68]. Goodness-of-Fit [GP65, Sla65, Sla66]. Goods [BT62, Fis63, Lee64b]. Graphic [Sha61b]. Graphical [Bac69, Bac69]. Graphique [Ber69b].

Graphs [GT65, M.67a]. Greek [Sch63]. Greenberg [Rao65]. Griffin [Sha61b]. Gross [LNF69]. Group [Agr64, KW64, O.66a]. Grouped [DM68a, Fis61c]. Grouping [Cra64b, FR67, Lan68b, McN67, Vin69, War63a, Lan68a]. Groups [Laz61]. Growth [Kam68, Nel60, Oli66]. Guenther [Jac66]. Guide [Gol63].

H [Edw67, McF67, Kau68, Whi66b]. Haenszel [Cor67a, Man63b]. Handbook [BK69]. Hannan [Mar65a, Ner66, Ner65b]. Harmonic [Abe62]. Hazard [BM65]. Health [Bro65, Gre69]. Height [Kar62]. Henrici [D.66, O.64a]. Here [Tuk60]. Heterogeneity [Agr66]. Heterogeneous [MW67b]. Heteroscedastic [BW69, Lan68a, Lan68b, RB68]. Heteroskedasticity [Gle69]. Hierarchical [War63a]. High [PRAL69]. Higher [Bow65, EB69]. Hildebrand [McF67]. Hilsenrath [M.67b]. Hirsch [Sch66a]. Historical [Smu60]. Histories [THS61]. History [Tei65a, Tei65b]. Hoeffding [KS69]. Hog [Abe62]. Hold [Gut60]. Holdings [Fer65, Fer66]. Homogeneity [Con69b, IKK69, Mad63]. Homoscedasticity [GQ65, HC63, GQ67]. Horvitz [Heg67]. Hotelling [HD67]. Houri [M.67c]. Household [BT69, BM67a, Cag64, JP61, Mai60, NW64]. Households [Hoc67]. Human [BM67b, Coa68, GW65]. Hurwitz [Dem69, Fra69, Han69]. Hyper [BC64]. Hyper-Poisson [BC64]. Hyperbolic [HH68a]. Hypotheses [BT68a, Bas65c, Bas65d, Bha66, BK66a, BK66b, Bir61, Cor66, CR66, Gri62, GW66, Hog61, Hog65, Mus67, Nag69, NS65, Pag63]. Hypothesis [Gri64, Ney69, Qua60, Sch66b].

I. [ABH^{+}69]. Ideal [Sto63]. Identically [DN65, DN66]. Identifiability [Bas60a]. Identification [Ste64c]. Identified [Bas61, Kab63b]. If [DeG66]. II [AE67, BC68d, Hal64b, Key68, Kin65b, LaV68b, Pra68]. III [AE69a, Bur68b, GK63, Hal65, LaV68c]. Ijiri [Dop67]. Illinois [Ano65n]. Illustration [Tho67b]. Illustrative [SN63, SN64]. Improve [Cla66]. Improved [Scl68]. Improvements [How69]. Incentive [FM64]. Inclusion [Tak67]. Income [Adl63, Cox61b, Dav62, Dun61a, Lee64b, Lee64c]. Incomplete [Fei60, Low69, PO61, The63]. Incremental [LaV68b]. Independence [CM69, Gla62a, Goo68, MS65, TN61]. Independent [AK65, BW69, Ben62, DN65, DN66, Gla64, HG68]. Indeterminism [Ney60]. Index [AG61, Ano66h, Ano66j, Ano66i, BMN63, Ban63, Dac64, DS65, Hur62, Ril61, WN66, Wil67a, Ano66e, Ano66f]. Indexes [Kis68]. India [BM67a]. Indian [Ram69]. Indians [CS62a]. Indicated [Tom61]. Indifference [NH65]. Individual [Adl63]. Individuals [DC69]. Industrial [Lav67, Ste61a]. Industrialization [Spi60]. Industries [PB61]. Industry [Gal68, Kam68, NWS67, PB61]. Inequalities [Ben62, DeG66, Hoe63]. Inequality [Hal67, Hoc65, Jen69, JT66, KS69, MG69]. Inference [Bas63b, Bir62, Efr67, Gre69, GT65, Har68a, Hil65c, Kem66, MFLR68, MW63, Nai69, SBC^{+}62, Ste67b, Bas64, Kem67, Jac66, Tia67]. Inferences [Ell64, LB63, Rob67]. Information [Ano60o, Ano61p, Ano62m, Ano63-30, Ano64m, Ano65q, Ano66s, Ano66t, Ano67q, Ano67r, Ano68v, Ano68w, Ano69t, Bor66, Cla66, CR66, Eri65, Eri68, Hil63a, Kul67, LaV68a, LaV68b, LaV68c, Nat67, Raj65a, Ste67b, The63, Ano68f]. Informative [Rob67]. Inherent [Dem64]. Initial [Hua63]. Innovation [DJ68]. Insensitivity [Ant65]. Inspection [Ans61b, Min69a, O’R69]. Insufficient [Chi64, Chi65b]. Integer [Vin69]. Integral [Gil62]. Intelligence [Shu60]. Intensities [Ber61a]. Intentions [Fis63, Jus66]. Inter [MB67b]. Inter-Experiment [MB67b]. Interaction [Goo64, Nag69]. Interactions [Goo68, Sen69]. Intercensal [Tar62]. ’Interdependent [Bas65a]. Interest [Adl63, FA65]. Interindustry [McF67, Shi64a, Shi64b]. Intermediate [Gle66]. Internal [LL60, Siv69]. International [NM69, Shi64a, Shi64b, Bor60a]. Interpolation [Fri62a]. Interpretation [KNS67, Smu60, Som62]. Interrelated [BM64]. Intersection [Rob64b]. Interstate [McF67]. Interval [CS62b, GW66, HM63, Hal64a, Hal64b, Hal65, Har64, Pre66, ST69, Sio64, Tho68a]. Intervals [DZ69, Eas69, Hah69, Her65, JJ69, Kab67, LRT61, Mac64, Mad62, McC65a, Nau66, Noe67, Pra61b, Wal62]. Interviewer [Kis62]. Interviewers [Ste64c]. Interviews [NW64]. Intrablock [Gir65]. Introduction [Ano69j, BR66, M.67a, Nic66, Alb67]. Invariant [FR67, Nat67]. Inventory [Bro61c]. Inverse [Gov63b, MW67a, Shu68]. Inversion [Rao65, TV68, TV69]. Inverted [TG65, TC65]. Investigating [Put67]. Investigation [AEL68, HSB66, Spi60]. Investment [Bro61c, EG66, Gre65a, Hor63, Hua63, JS69]. Irregular [Ban63]. Item [FMR62]. Iterated [Hog62, HT63, Leo62, Mus67]. Iterative [Tel64]. IV [AE69b]. Ivan [Koz66].

J [ABH^{+}69, Arm63, DeG66, Lav67]. John [Nic66]. Join [Hud66]. Joint [FKCR62, SB69, Swa62]. Joseph [M.67b]. Judgment [Win67b].

Kaufmann [Arr67]. Kendall [M.67e, Rio67, Kni66, Sen68, Som62]. Kind [Ban63]. Kinds [Don68]. Known [AK65, Col60b, Kau68, Kha68, ST69, Sea64]. Kolmogorov [Lil69a, Sla66, Cap65, Kim69, Klo67a, Lil67, Lil69b, Sla65, Suz68]. Kronecker [Neu69b]. Kruskal [Dav67, Som62].

L [M.67a, Tia67]. Labor [Kor66b, Sch63, Smu60, Kor66a]. Labour [FGP67]. Lags [Tin67]. Lamp [BK66b, BK66a]. Lamst [Nag69]. Land [Cla66]. Large [AE69a, CR66, Has68, Lor60, Raj64c, Sea66, SK69]. Large-Sample [Lor60]. Larger [BC68c]. Largest [Nau66, ST69]. Lasorsa [Sto66a, Sto66a]. Last [Bar65c]. Latent [MSD62]. Latter [Som62]. Lattice [BK61, FR65, FR66]. Law [Ans61a, Goo69, Kot66]. Layouts [Koc69b, PS67]. Leading [Bas61, Bas63a, Fit60, Kab63b, Kab64]. Learning [Pik69, Wal66a]. Least [Ati69, Chi64, Chi65b, Cho66, Cun64, Duv66, GJ61, Gol61a, Kha67, Koc69c, Lon67, McE67, RH65, Saw69, Zys69]. Least-Squares [Duv66, McE67]. Lehmann [AEL68]. Lemma [Koc69c]. Length [Mad62, Pra61b]. Leonard [DeG66]. Leptokurtosis [HC63]. Lester [DeG66]. Letters [Bri63]. Level [DS65, Joi69]. Levels [GG56, GG67]. Life [BP67, Bha67, Bha69, Ehr62, Ehr63b, Gov64, Gov65, GG61b, Key66, Key68, Kpe69, Kum67, Kum68, Man69b, MW67b, Mil60, SN63, Var69, SN64]. Life-Testing [MW67b]. Light [MP66b, MP66a]. Likelihood [Bir64, BFHH66, Fra63b, HM66b, HM67b, HM68b, LJZ68, Mey67, Oli66, ZE66a, ZE66b, HM66a, HM67a, HM68a]. Likelihoods [SK69]. Limit [AB67, AB68, AB69a, AB69b]. Limiting [Bec60a]. Limits [BW65, Hil67b, How69, Kra63, MB67a, Mos64, OCH64, Pac60, Sin66]. Lindgren [Alb67]. Line [Dun68, FA67, Gaf64, Jae64, Nau65]. Linear [AE67, AE69a, AE69b, Aiy63, Ajg65, Bas60a, Bas61, Bas63a, BK66a, BK66b, Bir64, BL67b, BG66, Cho66, CL69, Dav61a, Den63, Eag68, Fis61b, Fis62, FA67, GS65, Gla64, Gol62b, Gov66a, GB67, Gri62, Gri64, HM63, Hal64a, Hal64b, Hal65, HRHY67, HG68, Hei68, HH68b, HH69, Jor64, Jor65, Kab63c, Kha67, Koo66, Les63, McE67, MFLR68, Pag63, Qua60, Sch66b, Sch69, Scl68, SB69, Sio64, Tel64, TVW68, Wag62b, WTV69, WN66, You68, Efr67]. Linearly [Hal61a]. Linearly-Optimum [Hal61a]. Lines [BR69, Hal61b]. Linkage [FS69, Tep68b]. Linking [Boi69]. Liquid [Gut60]. Lists [Raj61]. Literature [AE66, M.67e]. Liu [McF67]. Live [Gai62]. Living [And63a, And66, Hen64b]. Local [HM66b, HM66a, HM67a, HM68a]. Local-Maximum-Likelihood [HM66b, HM66a, HM67a, HM68a]. Location [Bar66b, Bar67b, Bar68, BL67b, BTM69, Blo66, CS67, Cun64, Fis61c, Hog60b, Nea66, Noe67, Pra64, PS68, SR67, Wei63c, Nea67]. Log [Fei60, Hei68, Nau69, Tik68b]. Log-Linear [Hei68]. Log-Normal [Nau69, Tik68b]. Log-Normally [Fei60]. Logarithm [Nau69]. Logarithmic [PB67, PB69]. Logistic [Ber53, Gum61, HM67b, Oli66, Tar66]. Logit [Gri62, Gri64]. Loglinear [BK66a, BK66b]. Lognormal [HM66a, HM66b, HM67a, HM68a, Hil63b, Kau68, Lau63a, Lau63b, MW67b, Tho69]. Loss [McN67]. Losses [BM67b]. Lots [Ans61b]. Lower [AB68, AB69a, AB69b, Ber61b]. Lung [BE60b, BE60a, Gre69].

M [M.67e, Rio67, Rut67]. MacIntyre [Die67]. Major [Fis63, Laz61]. Make [EB61]. Making [NO66]. Male [EB69]. Males [Kpe69]. Management [Hor63, Dop67, Vaj66]. Mann [BKvE69, Hal60a, Hal60b, Cap61a, Mil64b]. Manpower [Stu66]. Mantel [Man63b]. Manufacturer [Wal65]. Manufacturing [Bor60b, Bor60c, Ein62, JS69, McF67]. March [Ano61n, Ano63j]. Marginal [IKK69]. Marginals [Arn67]. Mark [Bri63]. Market [Kam68, Nel60, NO66]. Marketing [Lad61]. Markoff [CKMP69, CKMP68]. Markov [Cho69, LO66, Reg68]. Match [Cac67]. Matched [McC65b]. Matching [Nat67, Raj61]. Material [End69]. Mathematical [Bru60, Die67, LH69, Mae66, MSGM65, Sve69]. Mathematics [Vaj66, Koz66]. Matrices [DO67, Har67b, Koc69c, Rao65]. Matrix [Con69a, Dwy67a, Dwy67b, Mar64, Neu69b, OF66a, Syk69, TD69, Tra67, WA67, OF66b]. Matter [Ano60a, Ano60b, Ano60e, Ano61b, Ano61c, Ano61d, Ano61e, Ano61f, Ano62a, Ano62b, Ano62c, Ano63a, Ano63b, Ano63c, Ano64a, Ano64b, Ano64c, Ano65a, Ano65b, Ano65c, Ano65e, Ano66a, Ano66b, Ano66c, Ano66d, Ano66g, Ano67a, Ano67b, Ano67c, Ano67d, Ano67f, Ano68a, Ano68b, Ano68c, Ano68d, Ano68g, Ano68h, Ano68i, Ano68j, Ano69a, Ano69b, Ano69c, Ano69d, Ano69f, Ano69g, Ano69h, Ano69i, Ano60d]. Maxima [Gri68, TD69]. Maximum [Bir64, BFHH66, FW64, GM66, HM66a, HM66b, HM67a, HM67b, HM68a, HM68b, LJZ68, Mey67, Nau65, Oli66, Rid64, Wei62a, ZE66a, ZE66b]. Maximum-Likelihood [HM67b, HM68b]. McCracken [Gol63]. McElrath [Alb67]. Mean [AK65, CKMP68, CKMP69, Gla62b, Her65, Kab67, Kha68, KMZ69, Lil67, Lil69a, Lil69b, MG69, RM67, ST69, Sea66, Tho68a, Tho68b, Tho69, TVW68, WTV69, Zac66]. Means [BC68c, BT68b, Dun61b, HB68, HB69, Har68b, Kin64a, Mar64, Sid67, Sio64, Kin62b]. Measure [Hay60, Ros66b]. Measured [DC69]. Measurement [Gal68, Mus69a, NMR65, Rog61, Tho63a]. Measures [GK63, HT67, Joi69, Koc69a, Lad61]. Measuring [Leo62, Ros68b, Tat61]. Mechanism [Sin68b]. Median [BC68d, Gas68, Joi69, Mac64, McC65a, Rid60b, Rid60a]. Medians [HL67]. Medical [Ans63, Arm63, Col63]. Meeting [Ano60n, Ano61o, Ano62l, Ano63l, Ano64l, Ano65n, Ano66r, Ano67o, Ano68s, Ano69s]. Memorial [Han69, You62]. Mendenhall [Gov65, Gov64]. Messick [Edw67]. Method [BMN63, Ber53, BL67b, BI64, Bre67, Bue61, Dix65, Fei60, Hic64, Hsi69, Kel69, Kni66, Kum67, Lau63a, Lau63b, Li62, Mar62a, Raj65a, Rao65, Sen69, Sha61a, Tak67, Zac62, Zel62, Kum68]. Methodological [Win67b]. Methodology [Bar65c]. Methods [Agr66, Bir54, Bir66, Cra64a, Goo61, Goo64, GP68, Jae64, Jes69, KT68, Net65, RG69, TF63, WP69, Wil62a, You68, Gle66, O.64a]. Metodologia [Sto66a]. Metropolitan [Fel65, Mil67]. Migration [BM67b, Gol64, LL60, Mil67, Siv69, Tar62, Tho61a, Wil63, Zac62]. Military [Kar62]. Mills [Rab69]. Minima [Gri68, TD69]. Minimal [HG65, Wag62b]. Minimize [FW64, Wei62a]. Minimizing [May68]. Minimum [DH59, Fis61b, Fis62, Jor64, ZE66b, DH63, Jor65]. Minneapolis [Ano63l]. Minnesota [Ano63l]. Misclassified [Pre68a]. Mises [Tho66]. Mismatching [NMR65]. Missing [AE66, AE67, AE69a, AE69b, Coh60d, FW61, Gla64, Goo68, HS68, Kel69]. Misspecification [Wau66]. Mixed [Ans61a, Gho67, HG65, SM69]. Mixed-Exponential [Ans61a]. Mixtures [Bli64, Has69, Hil63a]. Model [AEGH67, Bec60b, Ben66, BT68b, BT62, Col63, GS61, GJW68, Gol62b, Goo61, GAESH69, Hay65, HH68b, HH69, Hil65c, Hil67a, HG65, JLA69, LNF69, Lan66, Man69b, Maz63, MSGM65, NWS67, PB67, PB69, RK67, Rao69, Reg68, Rob66, RB68, SC65, Syk69, Tep68b, Tho61a, ZT64, RK68]. Models [Bas65a, BC68b, BG66, Bro69, Bru60, Con69a, Cra66, FA67, Gho67, GB67, Hog65, Joh67b, MFLR68, Osb67, Sin63, Tat66a, Wau66, ZC65, ZZK68, ZC69]. Modifications [Gas65]. Modified [Coh60c, Duv66, Mar65a, Ner65b, Ner66, Pre68b]. Moment [Zei65]. Moments [Cle69, Gho69, Gov63b, Jos69, Kri67a, Kri67b, MW67a, Nat68, Nat69, Sch62, Sha61a, Tar66, Tik64, Sch65]. Money [Lee64c, Zar68]. Monotone [BM65, BP67]. Monte [Gri68, Tho67b]. Monthly [Woo63, Woo66]. Morgenstern [Bor60a]. Mortality [CA63, EB69, MSD62, Zel69b]. Mosteller [Tia67]. Movement [TK61]. Mover [Goo61]. Moving [Leo62, Tel67a]. Multi [AEGH67, KH59, KH63, Raj65a, SS69, Sed67, Sin68a, TD62]. Multi-Auxiliary [Raj65a]. Multi-Factor [Sed67]. Multi-Proportions [AEGH67]. Multi-Stage [KH59, KH63, SS69, Sin68a, TD62]. Multinomial [BK66a, BK66b, Blu68, Hoa69, KN68, Wei62b]. Multinormal [AK65]. Multiple [Dav61a, DJ66, Dun61b, DM65, Dun69, Gab66, Hai66, Hsi69, Jor61b, Kab63a, Kra63, Lov63, Pag63, SD63, Sen69, WD69]. Multiplicative [Bec60b]. Multipliers [MS61]. Multisample [Sen67]. Multistage [Eri68]. Multivariate [AE66, AE67, AE69a, AE69b, Arn67, Arn68, Boi69, Cha68, Che66, Cho66, Dwy67a, Dwy67b, Elk68, Elk69, Fle66, Gei63c, Ger69, HS68, Kab63c, Kab67, Kel69, Kha67, Mah67, MO67, PB67, PS68, RM67, Sha65, Sid67, SB69, Tan69, TD69, Tra67, War63b, Gei64, PB69]. Must [DeG66].

N [Han69, Rio67, FF68]. National [AdJ66, Woo66]. Native [Zel64]. Nature [Wal62]. Necessary [McE67, Ogu69]. Negative [Gov62, Gov63a, Lau61, MM68, Rid62, Sha61a, Tik64]. Negroes [Zel69b]. Neither [AK65]. Nested [MM68]. Net [Siv69, Tar62, Zac62]. Networks [M.67a, M.67c]. Nevada [TW66]. Nineteenth [Fit60]. Niven [Koz66]. No [Nag69]. Noise [Mil64a]. Nomogram [Boy65, Boy66, Boy69]. Nomograph [LL61, LL62]. Non [Ant65, BW66a, BW66b, BTM69, BA68, CL69, Fel63, FA64, Gib64, Gra63, Gue64, HM63, Hal64b, Hal65, Jes69, Koc69a, Man61, Mey67, Rao65, Ros65, Ros66a, SS63a, SS63b, Sen65, Sho68, Sto63, Wag62b]. Non-Additivity [Man61]. Non-Attitudinal [FA64]. Non-Central [BA68, Gue64, Mey67, SS63a, SS63b]. Non-Centrality [Mey67]. Non-Ideal [Sto63]. Non-Linear [HM63, Hal64b, Hal65, Wag62b]. Non-Normality [Gib64]. Non-Optimal [Ant65]. Non-Orthogonal [BW66a, BW66b, Rao65]. Non-Parametric [Ros65, Ros66a, Sen65, Sho68]. Non-Regular [BTM69]. Non-Replacement [Jes69]. Non-Rotating [Fel63]. Non-Sampling [Koc69a]. Non-Stationary [CL69, Gra63]. Noncentral [Kri67a, Kri67b, Kri68, Kri69]. Noncontrolled [NS65]. Nonhomogeneous [Dac64]. Nonparametric [GG61a, Hol67, Hol68b, OCH64, PS67, PS68, Sär64, ST60, ST61, S.66]. Nonresponse [Eri67, Rao68b]. Nonsample [Wal64b]. Nonwhite [Tom61]. Nor [AK65]. Normal [BC68c, CC62, Che66, Col60b, Cur67, Ell64, FKCR62, Gil62, Hah69, Har64, Her65, Hil63a, HS68, How69, Kha68, KJ63, Klo64, Mar64, Nau69, OCH64, PP68, Pra68, Pre68b, ST69, Sch69, Sid67, Swa62, Tat66a, TGD67, Tik68b, Zac66, Mar65c]. Normality [Cap61a, Gib64, Lil67, Lil69a, Put67, SWC68, War63b, War67]. Normally [Fei60]. Note [AE69b, Arn67, Bas61, Bas63a, Bas65a, Bec60a, Ber61b, Bha66, Bir64, Blo66, Bre67, BA68, Cap61a, Dun68, Eag68, Fis61b, Fis62, FW61, Gal68, GJ61, Hay60, Hei68, HBA69, Jae64, Joh67b, Kab63b, Kha68, Koo66, KS69, Lan66, Mah67, Neu69a, Ogu69, Pik69, Raj64a, Rog61, RFT64, Sha61b, SC62, Tan69, TK61, Tik64, Ury66, Zac62, Zys63, vE64]. Notes [Ano60f, Ano60g, Ano60h, Ano60i, Ano61g, Ano61h, Ano61i, Ano61j, Ano62d, Ano62e, Ano62f, Ano62g, Ano63d, Ano63e, Ano63f, Ano63g, Ano64d, Ano64e, Ano64f, Ano64g, Ano65f, Ano65g, Ano65h, Ano65i, Ano66k, Ano66l, Ano66m, Ano66n, Ano67g, Ano67h, Ano67i, Ano67j, Ano68k, Ano68l, Ano68m, Ano68n, Ano69k, Ano69l, Ano69m, Ano69n, Koe67, Ano65d, Ano67e, Ano68e, Ano69e]. Number [Agr64, Edg61, Ham63, Har68a, Sin63, Sin68b]. Numbers [Ban63, Har67c, Har68b, Pea61]. Numerator [Don68]. Numerical [D.66, Ehr63c, Geb69, Goo69, M.67b, MF63, OF66b, OF66a].

Obeys [Qua60]. ’Objective [Goo69, War63a]. Observation [Bai65, CS62b, Hen64a, Chi65b]. Observations [AE66, AE67, AE69a, AE69b, And64, Chi64, DZ69, Fle66, Gla62a, Gla64, Har67c, Har68b, HS68, Hog67b, Kab67, Kau68, Kel69, MW67b, Sea66, Tat61]. Obtaining [Tra67]. OC [Lec64]. Occasions [RG64, SS65]. Occupancy [Har68a]. Occupational [Dun61a]. Occur [Con66]. Off [MS61]. Offset [Gil62]. Ohio [Ano64l]. Old [Roy68]. Omission [Kab63a]. Omitted [Ano67e]. Omnibus [Bir61]. Omnitab [M.67b]. One [Col63, Hai66, HRK69, Har67c, Hil65c, Man63b, Mos64, SC65, Tho66]. One-Parameter [SC65]. One-Sample [Tho66]. One-Way [Har67c, Hil65c]. One-Way-Classification [Hai66]. Ones [Coh60d]. Operations [Bro60a]. Operators [TV68, TV69]. Optimal [Ant65, BM64, BL67b, Eri67, Eri68, GS65, Joh67a, Joh69a, Kau68, Mut60b, NS65, Ser68, Mut62a]. Optimality [Lec64]. Optimize [War63a]. Optimum [Eri65, FA65, Hal61a, Ham68b, Heg67, PS67, Ros61b, Tep68b]. Order [Aiy63, Bar66b, CKMP68, CKMP69, DM68a, Jos69, Kel69, Man69b, MW67a, Maz63, Mus69b, Tar66, Tat61, Bar67b, Bar68]. Ordered [BFHH66, Edg61, Pag63, Tat61, WGF63]. Ordinary [McE67, Saw69]. Ornstein [Gri68]. Orthogonal [BW66a, BW66b, Rao65]. Orthonormal [Put67]. Oskar [Bor60a, Tin61]. Other [Ber61b, Cor67a, Han61, Joi69, PP68, Pra68]. Outcome [Hic63, Nat67]. Outliers [Bas65c, Bas65d]. Output [Bre62, Ein62]. Over-Identified [Bas61].

Pair [McC65b]. Paired [Bru60, Gla62a, RK67, RK68]. Paired-Comparison [RK67, RK68]. Pairs [Nai67]. Pan [Sch63]. Paper [ABH^{+}69, Arm63, Kim60, Sha61b]. Papers [Ano60n, Ano61o, Ano62l, Ano63l, Ano64l, Ano65n, Ano66r, Ano67o, Ano68s, Ano69s, Sch66a].

Parabolic [Cun64]. Parallel [ZE66b]. Parameter [Arn68, BT68a, BC64, Blo66, FF68, GC64, HM66a, HM66b, HM67a, HM68a, Hay65, Hil63b, HM69, Mey67, Sha66, SC65, Var69]. Parameters [Bli64, BTM69, Coh60c, FKCR62, Fis61c, Gov66a, HM66a, HM66b, HM67a, HM67b, HM68a, HM68b, Hoc65, HS68, Kha60, KJ63, LB63, Noe67, Sam66, Sha61a, She69, Sho69, Swa62, Tik68b, Wei63c]. Parametric [Ede61, Ede63, HM63, Hal64b, Hal65, Ros65, Ros66a, Sen65, Sho68, Zys69]. Part [LaV68a, LaV68b, LaV68c]. Partial [BC68b, Dav67, Gus61, Kab67, LL61, LL62, Som62, Wau66]. Partially [Blu68, Tat61]. Particular [BE60b, BE60a]. Partition [LaV68c]. Partition- [LaV68c]. Partitioned [Rao65]. Partitioning [CA63]. Pascal [Mat62a, Mat62b]. Patient [CA63]. Patients [Hay65]. Pattern [Han64, Han66]. Patterns [Bac69, Net65, Zel69b]. PBIB [Cle69, Gir65]. Pearson [Sla66, Sla65]. Pennsylvania [Ano66r]. People [Kor66b, Kor66a]. Pepelasis [Sch63]. Percentage [EH69, SS63a, SS63b]. Percentile [Gas65]. Percentiles [Eas69, Hil68a]. Performance [Gen65, Gen66]. Permutation [BC66, BC68a, Sen67]. Persons [Bue61, Mil67]. Peter [D.66, O.64a]. Phase [Suk62]. Philadelphia [Ano66r]. Pittsburgh [Ano69s]. Planning [BS69a]. Plans [Add63, Ban63, BM64, FMR62, FW64, Hil64b, O’R69, Whi66b]. Play [Zel69a]. Plot [FW61, Koc69b]. Plotting [Kim60]. Point [AE67, Cun64, DO67, Hal60b, Hil63b, Lon67, Rut66, SK69, ZC65, ZZK68, ZC69, Hal60a]. Point-Source [Hil63b]. Points [EH69, Hud66, Nau65, Rob64b, SS63a, SS63b]. Poisson [Bir66, BC64, Bir54, Coh60c, Coh60d, Jor61b, Mah67, Pre68b, Sam66, Sta67, Tik64]. Policies [Joh67a, Joh69a]. Policy [FB68, Wal66b]. Polynomial [Rab69, Rob64b]. Pooling [HB68, HB69]. Population [Ajg68, Bry65, Coa68, Col60b, CA63, EB61, Gol64, Gov66a, Har61, HM68a, Hil68a, Kum67, Kum68, LRT61, Mat62b, Mic59, Mic60, PB67, PB69, Rid63, Rid64, Rob67, Ros68b, Roy68, Sea66, Syk69, THS61, Tom61, Mat62a]. Populations [Ber61a, Bir54, Con69b, DM68a, FKCR62, HM66a, HM66b, HM67a, Hoa69, How69, Kha60, Kin62b, Kin64a, Lik67, Rah64, RM67, Rid60b, San60, ST69, Swa62, TD62, Wil64, Zac69, Bir66]. Positions [Kim60]. Positive [Gov63b]. Possibilities [May68, Sch60]. Post [Bro61c, Ful66, Wil62b]. Post-Stratified [Wil62b]. Posterior [Hil68a]. Potential [Cho69]. Potentials [WN66]. Power [AK69, FR65, Gib64, Gri62, Gri64, KA68, Lik67, MC66, Owe65a, Owe65b, Rid64, Sla68, TGD67, Tik67, Tik68a, TV68, TV69]. Power-Function [Lik67, Rid64]. Power-Series [TV68, TV69]. Powers [WA67]. Practical [HSB66]. Precise [Ber53]. Precision [AK65, GH58, GH63, GH69, Kau68, Tho63a]. Predetermined [Pow69]. Prediction [Bai65, Che66, GS65, Gol62b, Hah69, Hal61b, Kab67, Rob65a, ZC65, Wol69, ZZK68, ZC69]. Predictive [FA64]. Predictor [NS65]. Preliminary [Hic63]. Prescribed [Sam66]. Presence [TK61]. Present [Bas65d, Cur67, Bas65c]. Presentation [Bac69]. Presented [Sch66a].

Presidential [Han61, Hau63, Lik60, Bow64]. Price [BMN63, Bre62, Dac64, FM64, Hur62, Mus69a, Osb67, Ril61, WN66, Wil67a]. Prices [FA64]. Prime [FR65]. Prime-Power [FR65]. Principal [Mas65]. Principles [Dal69, Fra63b]. Prior [DZ69, Eri65, Eri68, Tei65a, The63, Win67a, Tei65b]. Probabilistic [Rob65a]. Probabilities [Ajg65, Ajg68, BI64, Con66, Cro64, Fel63, GP68, Har60a, LJZ68, McG69, Nat67, Nau66, PP68, Pra68, Wei62b]. Probability [Alb67, Bec60a, Ben62, Bre67, CW68, Con66, DO67, Edg61, Elk68, Elk69, Gen65, Gen66, Har66, Hen64a, Hil67b, Hil68c, Hoe63, Jes69, Jus66, Kim60, Kin65a, Kin65b, KT68, Mod63, Nai69, PO61, PB67, PB69, Raj64c, Raj65b, Rao63b, RB69, Rut67, Sea63, Sin63, Sud66, Sve69, Win69, Yas69, Nic66, Rei67]. Probable [Kam62, Mor59, Mor60a, Mor60b]. Probit [Gri62, Gri64]. Problem [Boi69, Cac67, Eat69, Har68a, Hic63, JLA69, LB63, MW63, NBS67, Pra64, Tak62, Wil61a]. Problems [Aig65, Ehr62, Ehr63b, Gre69, Kau68, Lau63b, MS63, Nai69, Ney69, Pot65, Pre66, PS68, Sch69, Sch60, Shi64b, Tak69, WD69, ZC65, ZZK68, ZC69, Blo69, Lau63a, Shi64a, Sve69]. Procedure [Aig65, Col60b, CHG69, FW61, FH67, HT63, KA68, Man63b, NH65, OF66b, OF66a, PO61, Raj66, Sea64, ST60, ST61, TD62, The68]. Procedures [Bai68, BP67, BR66, Gab66, Gas66, Kin67, Pra64, Pre66, PS67, Rao63b, Ros68a, Sea63, Sha65, Sto63, Tho63a, Zys69]. Proceedings [Blo69]. Process [AK65, CKMP68, Gri68, Jor61b, Nat67, Pre68b, CKMP69]. Processes [Bir54, Bir66, Cho69, Koo66, Tel67a, DeG66]. Processing [Min69a, Wha60]. Produce [Woo66]. Producer [Sin66]. Product [And63a, And66, Dac64, Goo62b, LNF69, Nat68, Nat69, Nel60, Rah64, Rid64, Tar66, Hen64b]. Production [Abe62, McF67, Ste61a]. Productivity [Dac64, McF67]. Products [BG69, Goo60, Neu69b, Wal65, Goo62a]. Profitability [Ste64e]. Program [M.67b]. Programmation [Arr67]. Programmed [Nic66]. Programmes [Cab60]. Programming [Fis61b, Vin69, WN66, Fis62, Gol63, M.67c]. Programs [Lon67]. Progress [Lin61]. Progressive [Man69b]. Proneness [BH65, Joh67b]. Propagation [O.64a]. Propensities [Gut60]. Properties [AEL68, BH66, CL69, FR68, GM67, KG68, Mat62a, Mat62b, Mut60b, Mut62a, PB67, RG69, PB69]. Property [Heg67]. Proportional [Add67, BG66, Bre67, GP68, Sär68]. Proportionate [Raj64c, Raj65b]. Proportions [AEGH67, CM69, Gab63, Hil63a]. Proposal [MS63]. Proposed [Pre66]. Proving [Koc69c]. Pseudo [Pea61]. Pseudo-Random [Pea61]. Publications [Ano60j, Ano60k, Ano60l, Ano60m, Ano61n, Ano61k, Ano61l, Ano61m, Ano62h, Ano62i, Ano62j, Ano62k, Ano63j, Ano63k, Ano63h, Ano63i, Ano64h, Ano64i, Ano64j, Ano64k, Ano65j, Ano65k, Ano65l, Ano65m, Ano66o, Ano66p, Ano66q, Ano67k, Ano67l, Ano67m, Ano67n, Ano68o, Ano68p, Ano68q, Ano68r, Ano69o, Ano69p, Ano69q, Ano69r]. Purchase [Jus66]. Purchases [BT62].

Quadratic [Hoc65, Koc69c]. Qualitative [Gil68]. Quality [AG61, Bow65, FIM64, Hor63, Ste67b, Lav67]. Quantal [Ber53]. Quantification [Win67b]. Quantiles [BH66, Has68]. Quantitative [Osb67]. Quantity [Bow65, Gal68]. Quarterly [GS61]. Quasi [Goo68, HL67, Koo66, LRT61, Les63]. Quasi-Independence [Goo68]. Quasi-Linear [Koo66, Les63]. Quasi-Ranges [LRT61]. Question [GBB67, GAESH69]. Queues [Tak62]. Quick [Gra63, Nea66, Nea67]. Quintus [Bri63]. Quotas [Sud66]. Quotient [Rid64].

R [Arr67, Gle66, Nic66]. R. [Bar65c]. Radhakrishna [Efr67]. Radial [Sch62, Sch65]. Random [Bai67, Ben62, BG69, BT68b, Bro69, BFHH66, CJ69, DN65, DN66, Gho67, Goo62b, Har67c, Har68b, HH68b, HH69, Hil67a, Pea61, Sam68b, Sve69, Tsu67, Hoe63]. Randomization [Cle69]. Randomized [AEGH67, BC66, BC68a, FW61, GAESH69, Koc69b, MS61, NS65, Noe63b, Raj65b, War65]. Randomness [Ste64f, Ste69c]. Range [Aig65, Bur68b, Con69b, Dun68, Gho65, Hil67b, Hil68c]. Ranges [Har61, LRT61]. Rank [BL67a, Chu65, Cur67, Gas65, HS69, Klo64, McC65b, Mos64, Qua67, TF63, Tay64]. Ranking [BR66, Bru60, SC62]. Rankings [Tay64]. Ranks [DM68b, Pag63, ST60, ST61]. Rao [Efr67]. Rapidly [Ste67b]. Rate [BM65, BP67, Ede61, Mai60, Siv69, Ede63]. Rates [Maz63, Tar62, Tom61]. Ratio [Bec60a, Don68, Gen65, Gen66, GH58, GH63, GH69, GC63, KNP62, Mic59, Mic60, Pre69, Rab69, Raj64a, Rao69, RV61, RV63, Suk62, Tin65, Zac62, dP61, dP64]. Ratio-Type [GH58, GH63, GH69, Rao69, Suk62]. Rational [Bac69]. Rationality [Goo69]. Ratios [BC66, Bro69, KH59, KH63, Mar65c, Sho69]. Re [GP68]. Re-Designing [GP68]. Readings [LH69]. Real [BMN63, Rog61]. Received [Ano60j, Ano60k, Ano60l, Ano60m, Ano61n, Ano61k, Ano61l, Ano61m, Ano62h, Ano62i, Ano62j, Ano62k, Ano63j, Ano63k, Ano63h, Ano63i, Ano64h, Ano64i, Ano64j, Ano64k, Ano65j, Ano65k, Ano65l, Ano65m, Ano66o, Ano66p, Ano66q, Ano67k, Ano67l, Ano67m, Ano67n, Ano68o, Ano68p, Ano68q, Ano68r, Ano69o, Ano69p, Ano69q, Ano69r, Dav62]. Reciprocal [Gov62, Gov63a, Rid62]. Recognizing [GM66]. Record [FS69, Mue63, Nat67]. Recording [EB61]. Records [Tep68b]. Rectangles [FR65]. Rectangular [BK61, Har61, Kha60, Rah64, Sid67]. Rectifying [Ans61b]. Recurrence [Gov63b, McG69, Mus69b]. Recursive [Sho68]. Reduces [Sea66]. Reducing [Bai67]. Redundant [MW67a]. Reference [BE60b, BE60a, BH65, Gre69, Neu69b]. Regimes [Qua60]. Region [Sta67]. Regional [BM67a, Bor60c, Hic63, Sch66a, Bor60b]. Regions [Bro69, Che66, FA67, GW66, PS68, Sid67]. Register [Gol64]. Regression [AE67, AE69a, AE69b, BMN63, BW69, CJ69, Cra64b, Cun64, Dav61a, DJ66, Dun68, Eag68, Eas69, Gaf64, GS65, Gla64, Gol62b, Gus61, Hai66, Hal64a, HRHY67, HG68, Hei68, Hoc65, Jor61b, JT66, Kab63a, Kab63c, Kak67, Kel69, Kha67, Koe67, Kon62, Lad64, Lov63, Mah67, Mal67a, Mas65, Mic59, Mic60, Mil64a, Par67, Pot65, Qua60, RG69, RB68, Scl68, Sen68, SM69, Tat66a, Tel64, TN61, The63, The65a, The68, TVW68, Wag62b, WTV69, War63b, War67, Wil67b, Woo66, Zel63, ZT64, ZC65, ZZK68, ZC69, Kon63]. Regressions [Cho66, KG68, Rob64b, Wal64a, Zel62]. Regular [Agr64, BTM69, DO67]. Reinterview [Bai68]. Related [Bar66c, Fri62a, Gov64, Gov65, PP68, Pra68]. Relating [Bas65c, Bas65d]. Relation [Cac65, Joh67b, MW67a, Mus69b, The65b]. Relations [Gov63b]. Relationship [Bir64, Mae66]. Relationships [Gal68, Kar62]. Relative [Bec60a, EG66, Fis66a, GP65, GC63, O’R69, ST60, ST61]. Relatively [Ber53]. Relevance [Tei65a, Tei65b]. Reliability [Bha67, Bha69, Fer65, Fer66, LB63, MW67a, Mar62a, MB67a, Rut66, Var69, ZE66a, ZE66b]. Reliable [Man69b]. Religious [Laz61]. Remarks [Bas63b, Eat69, Raj66, Bas64]. Removed [End69]. Renyi [BL69]. Repeated [Gol64, RG64]. Repeatedness [MSGM65]. Replaced [End69]. Replacement [Ajg65, Bre67, Con66, Fel63, Har66, Jes69, Raj66, Rao63b]. Replicate [Add63]. Replicates [Dan62a, Dan62b]. Replications [PZ68]. Report [Ano61a, NS61]. Reported [Dav62]. Reporting [Adl63, SB68]. Reports [Bor66, Mar62a]. Representation [Har67b]. Representations [BA68, Kri68, Kri69]. Reproducible [Bue61]. Required [GC63, GC64]. Research [Bai68, Mas65]. Réseaux [Ber69b]. Reset [Joh67a, Joh69a]. Residence [THS61]. Residences [THS61]. Residual [FVCR61, Gol61a, Nai67, Zys63, FRCR61]. Residuals [CKMP68, CKMP69, Koo66, Put67]. Resolution [Hog61]. Respect [CA63, Sla65, Sla66]. Response [AEGH67, Ans61a, Ber53, Bor66, Cra64a, Fel64b, GAESH69, JP61, May68, NW64, NMR65, War65]. Responses [FA65, MS65]. Restraints [Hoc65]. Restricted [Ati69, Cun64, Hog65]. Restriction [Den63]. Restrictions [JT66, TVW68, WTV69]. Results [BC66, MW67b, Nat69, Zel63]. Retail [Rog65, Woo66]. Reversal [NO66, PO61]. Review [Abe66, Abr69, Ada63, Ada69, Add69, Ade61a, Ade61b, Ade64, AE66, Air61, Ait68, Alb67, All60, All62, All66, And61a, And61b, And62a, And62b, AS64, And67, And60, And63b, Ant66, Ant69, Arn60, Arr61, Arr67, Ast65, Baa62, Bae61, Bai61, Bak69, Ban60, Ban61, Ban62, Ban65, Bar60, Bar65a, Bar66a, Bar67a, Bar65b, BN64, Bar67c, Bas65b, Bas60b, Bat68, Bat60, Bax61, Bea60, Bea64, Bee61, Bee62, Bel68, Ber62a, Ber62b, Ber63, Ber69a, BR61, Bic69, Bil60, Bir63, Bjo64, Ble63, Blo60, Blo65, Blu60, Blu61, Bly60, Bol61, Bon63, Bor60a, Bos64, Bou60, Bow63, Boy62, Bra63a, Bra63b, Bra65, Bra63c, Bre60, Bre61, Bre66, Bri69, Bri60, Bri64, Bro61a, Bro61b, Bro63a, Bro63b, Bro68, Bro60b, Bro61d, Bro61e, Bro63c, Bro65]. Review [Bro67, Bru62, Bru66, Bru61, Bur62a, Bur61, Bur62b, Bur68a, Bus62, Byr65, C.61a, C.61b, C.61c, C.62a, C.62b, C.62c, C.62d, C.62e, C.62f, C.62g, C.63, C.69, Cad63, Cai68, Cal69, Cam61, Cap61b, Car61, Car62, Cha62, Che60, Chi62, Chi65a, Cho67, Cho64, Cho61, Chu66, Cla60, Cla61, Cli65, Coc61, Coc62, Coc64, Coh60a, Coh60b, Coh62a, Coh62b, Coh68, Col60a, Col62, Con68a, Con61, Coo62, Cor60, Cor67a, Cor67b, Cow60, Cow63, Cox61c, Cox61a, Cro69, Cro60, Cro68a, Cro68b, D.66, Dal67, Dar60, Dar63, Dav64, Dav65, Dav66, Dav61b, DeG66, Dea64, Dem63, Dem65, Dem66, Der61a, Der61b, Der62, Dhr64, Dia67, Die67, Doo63, Dop67, Dor60, Dor61a, Dor61b, DR69, Dru62, Dun60, Dun62a, Dun62b, Dun66, E.68, Eas61, Eas64, Eck61, Edw64, Edw67, Efr67, Ehr63a, Eic68]. Review [Eis62, Ela68, Ell60, Eme62, Eps63, Eps66, Est64, Est66, Eva61, Eva64, F.60a, F.60b, F.60c, F.60d, F.61a, F.61b, F.61c, F.61d, F.61e, F.61f, F.61g, F.61h, F.61i, F.62a, F.62b, F.62c, F.62d, F.62e, F.62f, F.63a, F.63b, Fab61, Fan64, Far63, Fei69, Fei61, Fei62, Fel64a, Fer62, Fin60, Fir61a, Fir61b, Fis61a, Fis65b, Fis65a, Fis68, Fis67b, Fis66b, Fob61, Fol61, Fol69, For60, For62, Fox61, Fox66, Fra66, Fra60, Fra63a, Fre69, Fre61a, Fre64, Fre65, Fri66, Fri65, Fri68, Fri62b, Fro63, Gar68, Gas69, Gei63a, Gei63b, Geo66, Ger61, Gim62, Git69, Gla60, Gla66, Gle66, Gle68, God63, God64, Gol62a, Gol69, Gol60, Gol61b, Gol63, Gom61, Gor63, Gor61, Got68a, Got68b, Gov66b, Gra66, Gra60, Gre63a, Gre65b, Gre63b, Gre62, Gri61, Gri66, Gro60]. Review [Gue67, Gue65, Gue68, Gus60, H.60, H.61a, H.61b, H.61c, H.62a, H.62b, H.62c, H.63a, H.63b, H.64a, H.64b, H.64c, H.64d, H.64e, H.64f, Had66, Had63, Hae62, Ham68a, Ham65, Han63c, Har63, Har67a, Har69a, Har60b, Har60c, Har60d, Has60, Hau62, Hav67, Hea60, Hee60, Hen61a, Hen61b, Hen63, Hen61c, Hen64c, Her60, Her61, Hes64, Hil65a, Hil65b, Hil64a, Hil66, Hil68b, Hin65, Ho62, Hoc62, Hoc63a, Hoc60, Hoc63b, Hod67, Hoe62, Hof63, Hog67a, Hol61, Hol62, Hol63, Hol68a, Hoo65, Hoo67, Hop61, Hor69, Hou61, Hou62, How63, Hua64, Hun64, Hym61, Iji66, Int68, Ire61, J.63a, J.63b, J.64a, J.64b, J.64c, J.64d, J.64e, Jac66, Jac62, Jaf63, Jaf69, Jes67, Jes64, Joh60, Joh61a, Joh62, Joh68, Joh69b, Jon68b, Jon69, Jor61a, Jor62, Jor67a, Jos61, Jos65, Jud62, Jud64]. Review [Kai65, Kal62, Kal65, Kan63, Kao60, Kap62, Kar63a, Kar63b, Kar68, Kat67, Kat61, Kat63, Kat69, Kau65, Kem62, Kem64, Key69, Kie60, Kim62, Kin66, Kin64b, Kin63, Kis61a, Kis61b, Kit60, Kla69, Kle60, Koe66, Koo62a, Koo64, Koo62b, Kos61a, Kos61b, Koz66, Kra60a, Kra65, Kra60b, Kre60, Kre64, Kru60, Kue65, KF63, Kuh60, Kul65, Kur60, Kur62, Lad63, Lam64, Lan67, Lat63, Lat61, Lav66, Lav67, Lea69, Lee64a, Lee68, Lei64, Ler62a, Ler62b, Lev62, Lie60, Lie61, Lie63, Lie64, Lin62a, Lin63a, Lin63b, Lin63c, Lin62b, Lit61, Lit63, Lit65, Loc60, Lon60, Lov66b, Lub61, Lyd61, M.60a, M.60b, M.61a, M.61b, M.61c, M.61d, M.61e, M.61f, M.61g, M.62a, M.62b, M.66a, M.66b, M.67e, M.67c, M.67a, M.67b, M.67d, M.69, Mad69a, Mad68, Mad69b, Mad60a, Mad60b, Mai61a]. Review [Mai61b, Mai63, Mal67b, Man60, Man64a, Man69a, Man63a, Man64b, Mar60a, Mar60b, Mar60c, Mar65b, Mar63a, Mar63b, Mar60d, Mar62b, Mas66, Mas64, Mas62, Mas63, Max63, May60, May61, May62, May63, McC69a, McC68, McC69b, McC69c, McD69, McF67, McK60, McK64, McK61, McK62a, McK62b, McK62c, Mee67, Mei60, Mes67, Mey61, Mey62, Mil69, Mil61a, Mil63a, Mil63b, Mil65, Mil61b, Mil62a, Mil62b, Min68, Min69b, Moh61, Mon67, Mon66, Moo64, Mor61a, Mor66a, Mor61b, Mor66b, Mor68, Mor61c, Mos62, Mos61, Mos63a, MH63, Mos63b, Mue60, Mue64, Mur66a, Mur66b, Mut60c, Mut60a, Mut62b, Mye66, N.61a, N.61b, N.61c, N.61d, N.61e, N.61f, N.61g, N.62, Nei60, Nei64, Nel61, Ner64, Ner65a, Net62, Net66, Net68, Neu65, Neu64, New61, New68, Nic66, Nic61, Nic62, Nie67]. Review [Nis61a, Nis61b, Noe63a, Noe64, Noe65, Nor65, Nou65, O.61a, O.61b, O.61c, O.62a, O.62b, O.62c, O.62d, O.63a, O.63b, O.63c, O.64a, O.64b, O.64c, O.64d, O.64e, O.64f, O.64g, O.64h, O.64i, O.66a, O.66b, O.66c, O.67a, O.67b, O.67c, O.68, Ode69, Oeh60, Oi62, Ols68, Ore61a, Ore61b, Ore62, Orr63b, Orr63a, Ost69, Owe69, P.62a, P.62b, P.62c, P.62d, P.62e, P.62f, P.62g, P.63a, P.63b, P.63c, P.63d, P.63e, P.64, P.65a, P.65b, P.66, P.67, Pag64, Pal67, Pal69, Pap60, Par61a, Par61b, Par62, Par64a, Par64b, Par66, Par69, Pat61, Pat69, Pav63, Pea69, Per61a, Per61b, Pes64, Pet62, Pet67, Phe62, Phi62, Pie68, Pin69, Pol60, Pol61, Pow64, Pow65, Pra61a, Pre65, Pro68, Pro64, Qua64, Qui65, Rao63a, Rao68a, Rap61, Rap63, Rap67, Rat62, Ray68, Red60, Rei67, Rei63, Rei66, Rem65, Res61, Res68]. Review [Rho62, Rho63, Ric64, Ric67, Rie67a, Rio67, Rob62b, Rob62a, Rob63a, Rob63c, Rob63b, Rob64a, Rob65b, Rol63, Roo61, Ros61a, Rug60, S.61, S.65a, S.65b, S.65c, S.66, Sam68a, San67, Sas64, Sav61, Sav62, Sav64, Sca61, Sca63, Sch61, Sch66a, Sch66c, Sch63, Sea69, Sed69, Sen61, Sev63, Sev68, Sha60, She67, Shr60, Shr64, Shu64, Shu65, Shu61, Sid68, Sid69, Sil66, Sim62a, Sim66, Sim62b, Sin69, Sit60a, Sit60b, Sla62, Smi61, Smi62a, Smi62b, Smi64c, Smi64b, Smi64a, Smi66, Smi67, Smi68, Sne61a, Sne61b, Sne62, Som65, Som69, Spa61, Spi64, Spi66, Spr64, Spu64, Sta61, Sta62, Sta64, Ste64a, Ste68, Ste64b, Ste62, Ste64d, Ste61b, Ste63a, Ste64g, Ste69d, Ste63b, Sti63, Sto66a, Sto66b, Sto60, Sto69, Str60, Str61a, Str61b, Sub67, Sup61]. Review [Syl69, Tae67, Tat66b, Tay62, Tay65, Tei61, Tel67b, Tem66, Tep68a, Ter63, Thi63, Tho61b, Tho62, Tho67a, Tho63b, Tho64, Tia67, Tie62, Til65, Tob69, Tri67, Tro63, Tur65, Vaj64, Vaj66, Van69, Van63, Vei64b, Vei64a, Vei68, Ven61, W.60, Wag60, Wag61, Wag62a, Wah69a, Wah69b, Wal61, Wal63, Wat60a, Wat60b, Wat63, Wat65, Web67, Wei61a, Wei61b, Wei60, Wei63a, Wei63b, Wei66, Wei67, Wel63, Wen65, Wes67, Wes60, Whi66a, vNW68, Whi66c, Wil61b, Wil65a, Wil65b, Wil67c, Win63, Win64, Win61, Wit62a, Wit62b, Wol60, Wol65, Wol61, Wor63, Wri60, Yan60, Yan62, Yen68, Yer61, Zab66, Zar60, Zar67, Zeh61, Zel65, Zie62, Zin61, Zor60]. Reviewers [Ano66i]. Reviews [Ano66j]. Revised [AdJ66, Ros68a]. Revisions [Ste67a]. Risks [BE60b, BE60a, CA63, Sin66]. Robert [M.67a]. Robust [BL67a, CS67, Gas66, Hog67b, SR67]. Robustness [BL67b, Don68, DM68b, HD67, Pra64, Sto63]. Role [Fis63]. Ronald [You62]. Rotating [Fel63, Woo63]. Rotation [RG64]. Rounded [MS61]. Rounded-Off [MS61]. Rows [FR65]. Ruin [Tak69]. Rule [WD69, Zel69a]. Rules [Rob67, Win69]. Runs [Dun69, Edg61].

S [Edw67, McF67, STM^{+}65]. Saaty [M.67a]. Sales [Ben66]. Same [DC69, Hal60a, Hal60b, Rut66]. Sample [AEL68, AE69a, All63, Bai67, Bas60a, Bas61, Bas63a, Bas63b, Bas64, BT69, BL67a, BL69, BC68d, Bur68b, CKMP68, CKMP69, Col60b, CW68, Con66, Con68b, CR66, Dav62, Eri67, FW64, Gas65, Gho69, Gla62a, GC63, GC64, Har61, Has68, Her65, HBA69, Hil67b, Hil68c, Hol67, Hsi69, Jac63, Joi69, Kab67, Kau68, Kim69, Klo64, Klo67a, KG68, Kon62, Kon63, LRT61, Lor60, Mac64, Mil64b, Mos64, Noe63b, OF66b, OF66a, Pra64, Raj65a, RG69, Ros65, Ros66a, Ros66b, Ros61b, RFT64, SK69, Tho66, Tho67b, Wau66, Wei62a, Wil64, Zel63]. Samples [Cra64a, DM68a, Dix65, Eri68, Fel63, GP68, Hah69, Hal60a, Hal60b, HM66b, HM67b, HM68b, Has68, KH59, Kis68, Lik67, Mad63, Raj61, Rid60b, Rid60a, Rid64, ST60, Swa62, Tel67a, Tik68b, Woo63, ZE66a, Zac66, Zac69, HM66a, HM67a, HM68a, KH63, ST61]. Sampling [Ajg65, Ajg68, BM64, BW69, Blu68, Bre67, Con66, Con69b, Cro64, Cur67, Ell64, Eri65, FMR62, Fel63, FA65, FW64, GK63, GG61b, Har66, Hil68a, Hil64b, Jes69, Kim69, Koc69a, MW67a, McC65a, MM68, MS61, Mur63, Raj64c, Raj65b, Raj66, Rao63b, RG64, Rao68b, RB69, Rid63, Roy68, Sam66, San60, Sär68, Saw69, Sch69, Sed65, Sha65, SS65, Sin68a, SC62, Sud66, Suk62, Tei65a, Wal67, Whi66b, Wil67a, dP61, Tei65b, Kis69, dP64]. Samuel [STM^{+}65]. Sarhan [Rao65]. Savage [DeG66]. Saving [Cag64, Ram69]. Scalar [GP65]. Scale [Fis61c, Raj64c, Sho69, Wei63c]. Scheffe [Eat69, Pot65]. Scheffé-Type [Pot65]. Scheme [Sed65]. Schenkman [O.66a]. School [SN63, SN64]. Schooling [BM67b]. Science [Ney60, Ney69]. Sciences [LH69]. Scientifique [Arr67]. Score [Lin61]. Scores [Klo64, TGD67]. Scoring [Win69]. Screening [Li62]. Seasonal [Abe62, Han63a, Han63b, Han64, Han66, Jor64, Jor65, Jor67b, Lad64, Leo62, Les63, Lov63, Lov66a, Lov67, Mar65a, Ner65b, Ner66, Net65, Ros68a, Ros66c, You68]. Seasonality [Ril61]. Secant [HH68a]. Sech [Fis61c]. Second [Cle69]. Section [Lee64c]. Seeding [Bro60a]. Seemingly [Kak67, KG68, Zel62, Zel63]. Segment [GB67]. Segmented [BR69, Hud66]. Selected [SB68]. Selecting [Col63]. Selection [BR66, FMR62, TD62, Wil64]. Semi [Agr64]. Semi-Regular [Agr64]. Seminologie [Ber69b]. Sensitivity [Cra66, DS65, Fis66a, GG56, GG67, Sch66b]. Separate [Qua60]. September [Ano63l, Ano64l, Ano66r, Ano63k]. Sequence [BM64, GM66, Sam68b]. Sequences [Dan62b, Nai67, Dan62a]. Sequential [And64, Ans63, Arm63, Bec60a, Cor66, CHG69, CR66, FMR62, Gen65, Gen66, Gho65, Gho67, HM69, Jac60, Li62, Wei62b, Wei62a, Zac69]. Sequentially [PZ64, PZ68]. Sequentielle [Arr67]. Serial [Gra63]. Serially [Par67]. Series [AR61, Dan62b, DJ68, Den63, Duv66, Edg61, Fri62a, GG61a, Gra63, Han63b, Jor64, Kri68, KT68, Lov63, MB67a, PB67, PB69, Sen65, TV68, TV69, ZE66b, Dan62a, Han63a, Jor65, Kri69]. Service [CS62b]. Services [BT62, Lad61, Wal65]. Set [Bai67, Den63, Tel64]. Setting [Har61, LRT61]. Several [Bec60b, Geb69, GP68, Hog62, HT63, RG69, Rid60b, Sho68, TD62]. Sexes [GW65]. Shape [BC68d]. Shared [Sch68]. Shares [Bre62]. Sheila [Die67]. Short [Moo69]. Short-Term [Moo69]. Shorter [DZ69, HRHY67]. Shrinkage [Tho68a, Tho68b]. Sibs [GW65]. Side [Ber61b, TW66]. Sided [Gas68, Hah69, How69, Wal62]. Sidney [Edw67]. Siegel [Edw67, Edw67]. Sign [Gib64, Mac64]. Signed [Cur67, McC65b]. Signed-Rank [Cur67]. Significance [DS65, Joi69, Lan61, Pag63, Lan62]. Significantly [GP68]. Signs [Edg61]. Similar [MB67b]. Similarity [Har67b, Som62]. Simple [AE67, AE69a, AE69b, Ber53, BI64, GS65, Goo64, GB67, Joh61b, Mae66, Raj66, Zys69]. Simplification [The68]. Simulation [Gri68, NBS67, NWS67, Shu60, Tei65b, Tei65a]. Simultaneous [BC68b, Cra66, Gab66, JJ69, Nai69, Tho63a]. Simultaneous-Equations [Cra66]. Single [Kum67, Kum68]. Singly [Swa62]. Singular [Mar64]. Sir [You62]. Size [Bre67, Cox61b, FW64, Gho69, GP68, GC63, GC64, Kau68, Kin62b, Nau65, Nau66, Raj64b, Raj64c, Raj65b, Rob67, Ste64e, Wei62a, Zac66, Zac69, Kin64a]. Sizes [BL69, GP68, Mac64]. Slippage [Con68b]. SMA [Fel65]. Small [AEL68, Cra64a, Dix65, Hil67b, Joi69, KG68, RG69, Rid60b, Sla68, Wau66, ZE66a, Zac66, Hil68c]. Small-Sample [RG69]. Smaller [Bai67]. Smallest [Nau66, WGF63]. Smirnov [Lil69a, Cap65, Kim69, Klo67a, Lil67, Lil69b, Suz68]. Smoking [BE60b, BE60a, Bro65, Gre69]. Snodgrass [Bri63]. Social [Cab60, LH69]. Societies [Han61]. Society [Hau63]. Sociological [AZ69]. Solution [Boi69, Eat69]. Some [Agr66, Arm63, BT68a, BC66, Bas65c, Bas65d, BL69, BH66, BS69a, BS69b, Coh60d, Cor66, Cra64a, Dal69, Den63, Dwy67b, Eat69, Ehr62, Ehr63c, FR68, FR67, Geb69, GQ65, GG61a, GT65, GM67, HH68b, Hog67b, HC63, How69, Jes69, Kam62, Kem66, Koc69b, Koe67, Lad61, Lee64c, LB63, Lik67, Mat62b, Mic59, Mos64, Mur63, Nai69, Nau66, Neu69b, Osb67, PB67, PB69, Pot65, Pra64, PS67, PS68, Rah64, Raj66, Rao68b, Ril61, Ros65, Ros66a, Sär64, SS63b, Sen65, Sen67, Suk62, Suz68, Syk69, TF63, Tho68b, Tin65, Wal62, Win67b, Zel63, vE64, EB69, Sed67, Dwy67a, Ehr63b, GQ67, HH69, Kem67, Mat62a, Mic60, SS63a]. Source [Hil63b]. Sources [Sch60]. South [EB69]. Southeast [NM69]. Space [EP65, EP67]. Spaces [Put67]. Spanish [Bue61]. Spatial [Rog65]. Special [Ban63, BH65, Gei63c, Gei64, Gre69, Neu69b, Rid60b]. Specific [Maz63]. Specification [Fis66a, Fis67a, Gol61a, Sha65, Sin66]. Specified [Arn67, Con66, Hen64a]. Spectra [Koo66]. Spectral [CL69, Ros68a]. Speed [SB68]. Spherical [Elk68, Elk69]. Split [Koc69b]. Spread [ST60, ST61]. SPRT [Gho69, Lec64]. Square [Boy65, FR66, Fis61c, Gue64, Ham63, Jen69, KMZ69, Kot66, Man63b, MC66, Sla65, Sla66, Sla68, TVW68, WTV69, Boy66]. Squared [DM68b]. Squares [Add67, Ati69, Chi64, Chi65b, Cho66, Cun64, Duv66, GJ61, Gol61a, Kha67, Koc69c, Lon67, McE67, RH65, Saw69, Zys69]. Stabilities [RB69]. Stabilization [TV68, TV69]. Stabilizing [Lau61]. Stable [Coa68, FR68]. Stage [KH59, KH63, RG69, Sam66, Saw69, Sch69, SS69, Sin68a, TD62]. Standard [Col60b, Ede61, Ede63, Har61, Kis68, LRT61, Rid62]. Standardization [Shi64a, Shi64b]. Stanford [Ano61o]. State [Adl63, LNF69, Reg68, Tar62]. States [Bor60b, Bor60c, GS61, Laz61, LL60, Mil67, Ros66c, THS61, Wil63]. Stationary [CKMP68, CKMP69, CL69, Cun64, DJ68, DJ66, Gra63, Gri68, Tel67a]. Statistic [Cap61a, Jac63, Leh61, Man69b, McC65b, Mil64b, Noe63b]. Statistica [Sto66a, Sto66a]. Statistical [AdJ66, Ano60n, Ano61a, Ano61o, Ano62l, Ano63l, Ano64l, Ano65n, Ano66r, Ano67o, Ano68s, Ano69s, Bar65c, Bas63b, Bas64, Bas65a, Bir54, Bir61, Bir62, Bow64, Bow68, Bri63, Bro60a, Dal69, Efr67, Goo61, Gre69, GT65, Han61, Har68a, Har67c, Har68b, Hog61, Jac66, JLA69, Koc69b, M.67e, M.67b, Mas65, MM68, Ney69, PB67, PB69, PB61, SBC^{+}62, Spi60, Ste67b, Tin61, Bir66]. Statistically [Ber53]. Statisticians [Fit60, Ney60]. Statistics [AE66, AE67, AE69a, AE69b, Aiy63, AZ69, Bar66b, Bar67b, Bar68, Bas60a, BL69, Bow64, Cab60, Cro64, DM68a, DS65, Ehr63c, EP65, EP67, Gai62, Hau63, Hog60b, Hog60a, Hol68b, HG65, Jos69, LL60, Lik60, Lik67, MP66b, MW67a, Mus69b, Rah64, SD63, Sch68, Sen67, Sho68, SB69, Smu60, Sve69, Tar66, Wal66b, Bac69, BK69, MP66a, Alb67, Lav67]. Stayer [Goo61]. Step [Kin67]. Step-Wise [Kin67]. Stepwise [GJ61, Gol61a, Kab63c, Wal64a]. Stimulas [GG56]. Stimulus [GG67]. Stochastic [Ben66, CL69, Cra66, DeG66, Koo66, Rog65, Syk69, Tel67a]. Stochastically [BFHH66]. Stock [FFGM69, FA64, Hua63, NO66, Osb67]. Stopping [Rob67]. Storage [JLA69]. Straight [BR69, FA67, Gaf64, Hal61b, Jae64]. Strata [FA65, Ful66, Raj64b]. Strategies [Hoc67]. Stratification [BHJ60, DH59, HSB66, Sär68, Ser68, SS65, DH63]. Stratified [Cha68, Eri65, Eri68, Jon68a, McC65a, Sär68, Wil62b, Wil64, dP61, dP64]. Stratum [HRK69, RB69]. Structural [Bir64, Mal67a, Ric68]. Structure [Joh67a, Joh69a, PB61]. Structures [Zys69]. Student [Owe65a, Ste69a, Boy69, Efr69, Hal67, Kot66, Owe65b, Ste69b]. Studies [Kis62, Sed67, Siv69, Tho67b]. Study [BC68a, BE60b, BE60a, Cor67a, MSGM65, NW64, RB69, Rob66, SWC68, Smu60]. Subclass [Har68b]. Subclasses [Har68b]. Subject [Bai65, Den63, Hal61b, Hal64a]. Subjective [Goo69]. Subsample [Har69b]. Subsamples [Jon68a]. Subsets [MFLR68]. Substitute [Har64]. Successive [Kum67, Kum68, SS65, Sin68a]. Sufficiency [Fra63b]. Sufficient [HG65, McE67]. Suggestions [Win67b]. Suitable [Gra63]. Sum [Ben62, Chu65, DN65, DN66, DM68b, Joh61b, Nau69, ST60, ST61]. Summaries [Ano60n, Ano61o, Ano62l, Ano63l, Ano64l, Ano65n, Ano66r, Ano67o, Ano68s, Ano69s]. Sums [Tel67a, Hoe63, Mar65c]. Supplement [Gov64, Gov65]. Supplier [DAS69]. Support [Tat61]. Surname [Bue61]. Surplus [Sch63]. Survey [Bor66, Jus66, MS63, Raj64c, Stu66, War65, FGP67, Kis69]. Surveys [BT69, BS69a, Cha68, CW68, Dal69, Fer65, Fer66, JP61, KNP62, Kon62, Mue63, Raj65a, SS69, Sed65, Stu66, Woo63, Kon63]. Survival [Ede61, Ede63, Fei60, Gla67, Hay65, Tar62, TK61, Zac62]. Symmetric [FR68, Gov66a, Ney69, WD69, Rio67]. Symmetry [Efr69, FS64, IKK69].

System [BC68d, Bur68b, Par67, Qua60, Rut66, Bow68]. Systematic [EP65, Har66, Mil64a, Raj64c, Raj65b, SB69, EP67]. Systems [CL69, MW67a, Mar62a, MB67a, MF63, NBS67, ZE66b].

t [BA68, Hal67]. Ta-Chung [McF67]. Table [Edg61, IKK69, Key66, Key68, Mac64, Mil64b, Tho69]. Tables [BM65, BL69, Goo64, Goo68, Hai66, Ham68b, Jac63, Koc69a, Kpe69, Kra63, Kum67, Kum68, McC65b, MC66, Mos68, Nag69, Pac60, Shi64b, SN63, SN64, Tik67, WTV69, Shi64a, Tik68a, Rio67]. Tail [PP68, Pra68]. Taking [Bro64, Ury66]. Tau [Hay60, Kni66, Sen68, Sha61b, Som62, SC62]. Tax [Bre62, Adl63]. Taxpayer [Adl63]. Technique [Bir61, War65, Woo66]. Techniques [Add63, Den63, Duv66, FMR62, Tho68b]. Teen [CS62a]. Teen-age [CS62a]. Telemetry [EP65, Mar62a, EP67]. Ten [Mue63]. Term [Moo69]. Terry [RK68, RK67]. Test [AK69, AEL68, All63, BC68a, BP67, Bas60a, Bec60a, Bha66, Boy69, Bri63, Cap65, Chu65, Cle69, Col60b, Con69b, Cor66, CR66, DAS69, DS65, Don68, DM68b, Efr69, EH69, FS64, Gab66, Gas68, Gen65, Ger69, Gib64, Gir65, Gla62a, Gle69, Goo69, Gra63, Hal60a, Hal60b, Ham63, HBA69, Hol67, Hol68b, Joi69, Klo67a, Lil67, Lil69a, Lil69b, Mac64, MS65, Nai69, Nea66, Nea67, Ney69, Owe65b, Pag63, PO61, Pre66, Sho68, Sin66, Sla68, Tho66, TGD67, Tik67, TVW68, WTV69, Wei62b, Gen66, Owe65a, Tik68a]. Testability [Bas65a]. Testing [Bar66c, Bha67, Bha69, Bir61, Blo69, Ehr62, Gov64, Gov65, Hil64b, HT63, Kha60, MW67b, Mil60, Sho69, Ste64f, TN61, Var69, Ehr63b]. Tests [BT68a, BW66a, BW66b, Bas65c, Bas65d, BK66a, BK66b, BL67a, Con68b, FM64, Gas65, GP65, Gho65, GQ65, GQ67, GG61a, Gri62, Gri64, GW66, GG61b, HM69, Hog62, Kin62a, Klo64, Lan61, Lan62, Mad63, Man63b, McN67, MC66, Mos64, Mus67, NS65, Osb67, Pot65, Qua60, Ros65, Sch66b, Sen65, Sen67, SWC68, Sla65, Sla66, Ste69b, Ste69c, Suz68, Wei62a, Zel62, vE64, HS69, Ros66a, Ste69a]. Textile [NWS67]. Theil [SM69]. Their [Ban63, KH59, KH63, Mue63, Put67, MW67a]. Theorem [Cho66, Kha67, LO66, Zei65, Hil68a]. Theorems [Ell64, Neu69b]. Theoretical [GAESH69, Joh61b]. Theory [Ant65, BH65, Cho69, FR65, FS69, GK63, HS69, Kin65a, Kin65b, LaV68a, MS61, Mur63, PRAL69, Rao68b, Roy68, Sve69, Tak62, Vin69, O.66a]. There [FA65]. Thomas [M.67a]. Thomson [Heg67]. Thornton [Rei67]. Three [BT68a, Bas63a, Goo64, HM66a, HM66b, HM67a, HM68a, Hil63b, Hoc67, Kab64, Man69b, Nag69, Rao63b, Sch60]. Three- [HM68a]. Three-Equation [Bas63a, Kab64]. Three-Factor [Goo64]. Three-Order-Statistic [Man69b]. Three-Parameter [BT68a, HM66a, HM66b, HM67a, Hil63b]. Ties [Klo66, Klo67b, Leh61, RK67, Tay64, RK68]. Time [AR61, CS62b, DJ68, Den63, Duv66, Fer65, Fri62a, GG61a, Gra63, Gri68, Han63a, Han63b, Jor64, Jor65, Lov63, MH67, Sch68, Sen65]. Time-Shared [Sch68]. Time/Failure [MH67]. Tolerance [BW65, Che66, How69, Wal62].

Tool [Pow69]. Topics [Gov64, Shu60, Gov65]. Total [CS62b]. Totals [BC68b, Kin62b, Kin64a]. Tournament [Sea63]. Trade [NM69, Woo66]. Trading [Osb67]. Transactions [Bor60a]. Transformations [Gri62, TV68, Gri64, TV69]. Transforms [Ehr63c]. Transition [DO67, LJZ68]. Translation [Lau63a, Lau63b]. Transportation [M.67c]. Treated [Hay65]. Treatments [Agr64, BK61, Col63, FR65, Pag63]. Trees [Har67b]. Trend [Les63, S.66]. Trends [Tom61]. Trial [Zel69a]. Trials [Ans63, Arm63, Cor66, CHG69, NG65, Rut67]. Triangular [Rid63]. Trivariate [Eag68]. True [GC64, Sea66]. Truncated [Coh60d, Fis61c, Gen65, Gen66, KJ63, Sha61a, Sha66, Swa62, Tar66, Tik64]. Truncation [DAS69]. Trust [Hor63]. Tukey [Nea67, Nea66]. Twain [Bri63]. Twins [Rob66]. Two [AEL68, Agr66, All63, BC64, Bas61, Bha66, BL67a, BC68c, BS69a, BFHH66, BHJ60, Cle69, Col63, Con66, Con68b, CR66, Cro64, Ell64, FKCR62, Gas65, Gas68, Gir65, Hah69, Har68b, Hay65, Hol67, How69, HG65, Jac63, Jae64, Kab63b, Kau68, Kim69, Klo64, Klo67a, Koc69c, Kul67, Kum67, Kum68, Man61, Mil64b, Mos64, Nau69, Noe63b, Pra64, Pre66, PS67, Qua60, RB69, RG69, Reg68, Rob64b, Ros65, Ros66a, Sam66, Saw69, Sch69, Ste69c, Suk62, Var69, Wal62, Wil62a, Wil64, Zac66, Ber61a]. Two-Action [Kau68, Sch69]. Two-Associate [Cle69]. Two-Factor [BS69a]. Two-Parameter [BC64, Hay65]. Two-Phase [Suk62]. Two-Sample [All63, BL67a, Jac63, Klo64, Mil64b, Mos64, Noe63b, Pra64, Ros65, Ros66a]. Two-Sided [Gas68, Hah69, How69, Wal62]. Two-Stage [RG69, Sam66, Saw69, Sch69]. Two-State [Reg68]. Two-Variable [Cro64]. Two-Way [Agr66, BHJ60, Har68b, HG65, Man61, PS67, Wil64]. Type [BL69, GH58, GH63, GH69, Koc69c, Pot65, Rao69, Suk62]. Typical [Har69b].

U [McF67, Sen67]. U-Statistics [Sen67]. U.S. [AdJ66, JS69, Tom61]. Uhlenbeck [Gri68]. Unbalanced [Har67c]. Unbiased [Gol62b, GH58, GH63, GH69, Hai66, Hal61a, Har68a, Jor64, Jor65, McE67, MG69, Mic59, Mic60, RV61, Sam68b, Wil62a, ZE66a, ZE66b, Zac66, dP61, RV63, dP64]. Unbiasedness [Kak67]. Uncertainty [LaV68a, LaV68b, LaV68c, PRS64]. Uncontrolled [Mal67a]. Uncorrelated [Aiy63, Hog60a, Hol68b]. Underdeveloped [Wha60]. Unequal [CW68, Con66, Gab63, Har66, HC63, Rao63b, RB69]. Ungrouped [Kni66]. Unified [RH65, Sär64]. Uniform [Bar66c, BG66, Con69b, GC64, Mar65c]. Uniformly [Ses63, Sha67]. Uniqueness [Joh67b, Zei65]. Unit [HRK69, Ste64f]. United [Bor60b, Bor60c, GS61, Laz61, LL60, Mil67, Ros66c, THS61, Wil63]. Units [Con66, FA65, GC64, RB69]. Universe [Nat68, Nat69]. Unknown [Hai66, HB68, Kin62b, Kin64a, Lil67, Lil69b, MG69, Rao68b, HB69, Lil69a]. Unpaired [ST60, ST61]. Unrelated [GAESH69, Kak67, KG68, Zel62, Zel63]. Unrestricted [End69]. untitled [Abe66, Abr69, Ada63, Ada69, Add69, Ade61a, Ade61b, Ade64, Air61, Ait68, All60, All62, All66, And61a, And61b, And62a, And62b, AS64, And67, And60, And63b, Ano63m, Ano63n, Ano63o, Ano63p, Ano63q, Ano63r, Ano63s, Ano63t, Ano63u, Ano63v, Ano63w, Ano63x, Ano63y, Ano63z, Ano63-27, Ano63-28, Ano63-29, Ano65o, Ano65p, Ano67p, Ano68t, Ano68u, Ant66, Ant69, Arn60, Arr61, Ast65, Baa62, Bae61, Bai61, Bak69, Ban60, Ban61, Ban62, Ban65, Bar60, Bar65a, Bar66a, Bar67a, Bar65b, BN64, Bar67c, Bas65b, Bas60b, Bat68, Bat60, Bax61, Bea60, Bea64, Bee61, Bee62, Bel68, Ber62a, Ber62b, Ber63, Ber69a, BR61, Bic69, Bil60, Bir63, Bjo64, Ble63, Blo60, Blo65, Blu60, Blu61, Bly60, Bol61, Bon63, Bos64, Bou60, Bow63, Boy62, Bra63a]. untitled [Bra63b, Bra65, Bra63c, Bre60, Bre61, Bre66, Bri69, Bri60, Bri64, Bro61a, Bro61b, Bro63a, Bro63b, Bro68, Bro60b, Bro61d, Bro61e, Bro63c, Bro67, Bru62, Bru66, Bru61, Bur62a, Bur61, Bur62b, Bur68a, Bus62, Byr65, C.61a, C.61b, C.61c, C.62a, C.62b, C.62c, C.62d, C.62e, C.62f, C.62g, C.63, C.69, Cad63, Cai68, Cal69, Cam61, Cap61b, Car61, Car62, Cha62, Che60, Chi62, Chi65a, Cho67, Cho64, Cho61, Chu66, Cla60, Cla61, Cli65, Coc61, Coc62, Coc64, Coh60a, Coh60b, Coh62a, Coh62b, Coh68, Col60a, Col62, Con68a, Con61, Coo62, Cor60, Cor67b, Cow60, Cow63, Cox61c, Cox61a, Cro69, Cro60, Cro68a, Cro68b, Dal67, Dar60, Dar63, Dav64, Dav65, Dav66, Dav61b, Dea64, Dem63, Dem65, Dem66, Der61a, Der61b, Der62, Dhr64, Dia67, Doo63, Dor60, Dor61a, Dor61b]. untitled [DR69, Dru62, Dun60, Dun62a, Dun62b, Dun66, E.68, Eas61, Eas64, Eck61, Edw64, Ehr63a, Eic68, Eis62, Ela68, Ell60, Eme62, Eps63, Eps66, Est64, Est66, Eva61, Eva64, F.60a, F.60b, F.60c, F.60d, F.61a, F.61b, F.61c, F.61d, F.61e, F.61f, F.61g, F.61h, F.61i, F.62a, F.62b, F.62c, F.62d, F.62e, F.62f, F.63a, F.63b, Fab61, Fan64, Far63, Fei69, Fei61, Fei62, Fel64a, Fer62, Fin60, Fir61a, Fir61b, Fis61a, Fis65b, Fis65a, Fis68, Fis67b, Fis66b, Fob61, Fol61, Fol69, For60, For62, Fox61, Fox66, Fra66, Fra60, Fra63a, Fre69, Fre61a, Fre64, Fre65, Fri66, Fri65, Fri68, Fri62b, Fro63, Gar68, Gas69, Gei63a, Gei63b, Geo66, Ger61, Gim62, Git69, Gla60, Gla66, Gle68, God63, Gol62a, Gol69, Gol60, Gol61b, Gom61, Gor63, Gor61, Got68a, Got68b]. untitled [Gov66b, Gra66, Gra60, Gre63a, Gre65b, Gre63b, Gre62, Gri61, Gri66, Gro60, Gue67, Gue65, Gue68, Gus60, H.60, H.61a, H.61b, H.61c, H.62a, H.62b, H.62c, H.63a, H.63b, H.64a, H.64b, H.64c, H.64d, H.64e, H.64f, Had66, Had63, Hae62, Ham68a, Ham65, Han63c, Har63, Har67a, Har69a, Har60b, Har60c, Har60d, Has60, Hau62, Hav67, Hea60, Hee60, Hen61a, Hen61b, Hen63, Hen61c, Hen64c, Her60, Her61, Hes64, Hil65a, Hil65b, Hil64a, Hil66, Hil68b, Hin65, Ho62, Hoc62, Hoc63a, Hoc60, Hoc63b, Hod67, Hoe62, Hof63, Hog67a, Hol61, Hol62, Hol63, Hol68a, Hoo65, Hoo67, Hop61, Hor69, Hou61, Hou62, How63, Hua64, Hun64, Hym61, Iji66, Int68, Ire61, J.63a, J.63b, J.64a, J.64b, J.64c, J.64d, J.64e, Jac62, Jaf63, Jaf69, Jes67, Jes64, Joh60, Joh61a, Joh62, Joh68]. untitled [Joh69b, Jon68b, Jon69, Jor61a, Jor62, Jor67a, Jos61, Jos65, Jud62, Jud64, Kai65, Kal62, Kal65, Kan63, Kao60, Kap62, Kar63a, Kar63b, Kar68, Kat67, Kat61, Kat63, Kat69, Kau65, Kem62, Kem64, Key69, Kie60, Kim62, Kin66, Kin64b, Kin63, Kis61a, Kis61b, Kit60, Kla69, Kle60, Koe66, Koo62a, Koo64, Koo62b, Kos61a, Kos61b, Kra60a, Kra65, Kra60b, Kre60, Kre64, Kru60, Kue65, KF63, Kuh60, Kul65, Kur60, Kur62, Lad63, Lam64, Lan67, Lat63, Lat61, Lav66, Lea69, Lee64a, Lee68, Lei64, Ler62a, Ler62b, Lev62, Lie60, Lie61, Lie63, Lie64, Lin62a, Lin63a, Lin63b, Lin63c, Lin62b, Lit61, Lit63, Lit65, Loc60, Lon60, Lov66b, Lub61, Lyd61, M.60a, M.60b, M.61a, M.61b, M.61c, M.61d, M.61e, M.61f, M.61g, M.62a, M.62b, M.66a, M.66b, M.67d, M.69, Mad69a]. untitled [Mad68, Mad69b, Mad60a, Mad60b, Mai61a, Mai61b, Mai63, Mal67b, Man60, Man64a, Man69a, Man63a, Man64b, Mar60a, Mar60b, Mar60c, Mar65b, Mar63a, Mar63b, Mar60d, Mar62b, Mas66, Mas64, Mas62, Mas63, Max63, May60, May61, May62, May63, McC69a, McC68, McC69b, McC69c, McD69, McK60, McK64, McK61, McK62a, McK62b, McK62c, Mee67, Mei60, Mes67, Mey61, Mey62, Mil69, Mil61a, Mil63a, Mil63b, Mil65, Mil61b, Mil62a, Mil62b, Min68, Min69b, Moh61, Mon67, Mon66, Moo64, Mor61a, Mor66a, Mor61b, Mor66b, Mor68, Mor61c, Mos62, Mos61, Mos63a, MH63, Mos63b, Mue60, Mue64, Mur66a, Mur66b, Mut60c, Mut60a, Mut62b, Mye66, N.61a, N.61b, N.61c, N.61d, N.61e, N.61f, N.61g, N.62, Nei60, Nei64, Nel61, Ner64, Ner65a, Net62, Net66, Net68, Neu65, Neu64, New61, New68]. untitled [Nic61, Nic62, Nie67, Nis61a, Nis61b, Noe63a, Noe64, Noe65, Nor65, Nou65, O.61a, O.61b, O.61c, O.62a, O.62b, O.62c, O.62d, O.63a, O.63b, O.63c, O.64b, O.64c, O.64d, O.64e, O.64f, O.64g, O.64h, O.64i, O.66b, O.66c, O.67a, O.67b, O.67c, O.68, Ode69, Oeh60, Oi62, Ols68, Ore61a, Ore61b, Ore62, Orr63b, Orr63a, Ost69, Owe69, P.62a, P.62b, P.62c, P.62d, P.62e, P.62f, P.62g, P.63a, P.63b, P.63c, P.63d, P.63e, P.64, P.65a, P.65b, P.66, P.67, Pag64, Pal67, Pal69, Pap60, Par61a, Par61b, Par62, Par64a, Par64b, Par66, Par69, Pat61, Pat69, Pav63, Pea69, Per61a, Per61b, Pes64, Pet62, Pet67, Phe62, Phi62, Pie68, Pin69, Pol60, Pol61, Pow64, Pow65, Pra61a, Pre65, Pro68, Pro64, Qua64, Qui65, Rao63a, Rao68a, Rap61, Rap63, Rap67, Rat62, Ray68, Red60, Rei63, Rei66, Rem65, Res61]. untitled [Res68, Rho62, Rho63, Ric64, Ric67, Rie67a, Rob62b, Rob62a, Rob63a, Rob63c, Rob63b, Rob64a, Rob65b, Rol63, Roo61, Ros61a, Rug60, S.61, S.65a, S.65b, S.65c, Sam68a, San67, Sas64, Sav61, Sav62, Sav64, Sca61, Sca63, Sch61, Sch66c, Sea69, Sed69, Sen61, Sev63, Sev68, Sha60, She67, Shr60, Shr64, Shu64, Shu65, Shu61, Sid68, Sid69, Sil66, Sim62a, Sim66, Sim62b, Sin69, Sit60a, Sit60b, Sla62, Smi61, Smi62a, Smi62b, Smi64c, Smi64b, Smi64a, Smi66, Smi67, Smi68, Sne61a, Sne61b, Sne62, Som65, Som69, Spa61, Spi64, Spi66, Spr64, Spu64, Sta61, Sta62, Sta64, Ste64a, Ste68, Ste64b, Ste62, Ste64d, Ste61b, Ste63a, Ste64g, Ste69d, Ste63b, Sti63, Sto66b, Sto60, Sto69, Str60, Str61a, Str61b, Sub67, Sup61, Syl69, Tae67, Tat66b]. untitled [Tay62, Tay65, Tei61, Tel67b, Tem66, Tep68a, Ter63, Thi63, Tho61b, Tho62, Tho67a, Tho63b, Tho64, Tie62, Til65, Tob69, Tri67, Tro63, Tur65, Vaj64, Van69, Van63, Vei64b, Vei64a, Vei68, Ven61, W.60, Wag60, Wag61, Wag62a, Wah69a, Wah69b, Wal61, Wal63, Wat60a, Wat60b, Wat63, Wat65, Web67, Wei61a, Wei61b, Wei60, Wei63a, Wei63b, Wei66, Wei67, Wel63, Wen65, Wes67, Wes60, Whi66a, vNW68, Whi66c, Wil61b, Wil65a, Wil65b, Wil67c, Win63, Win64, Win61, Wit62a, Wit62b, Wol60, Wol65, Wol61, Wor63, Wri60, Yan60, Yan62, Yen68, Yer61, Zab66, Zar60, Zar67, Zeh61, Zel65, Zie62, Zin61, Zor60]. Up-and-Down [Dix65, Hsi69]. Upper [AB67, Col60b, OCH64]. Use [Cab60, Cag64, Cla66, CW68, Ehr63c, Gra63, Har61, JLA69, LRT61, Put67, Raj64c, San60, Sär68, Tat61, The63, Woo63, Woo66, Fie62, Leo62]. Used [Cro64, EP65, EP67]. Useful [GT65, Koc69c, Rob66]. User [Lon67]. Uses [Rei67]. Using [CR66, DZ69, DM65, Eri65, Eri68, FMR62, Gil68, Gri62, Gri64, Har69b, Raj65a, She69, Sin68a]. Utilization [Sea64]. Utilizing [GP68].

Validation [FFGM69]. Validity [Dav62, Sla65, Sla66]. Value [GC64, Has68, PRAL69]. Values [BL69, HM68b, Har69b, Mil64b, Rid62, Rid64, Ste69a, Ste69b]. Variability [Ste64e]. Variable [Bai65, Cro64, Gov62, Gov63b, Lan66, Lor60, MFLR68, Rid62, Sär68, Tin67, Gov63a]. Variables [Ben62, BG69, BFHH66, DN65, FA64, Gil68, Gla64, Goo62b, Hal61b, Hal64a, HG68, Kis62, Li62, Mar64, NS65, Hoe63, Mar65c, DN66]. Variance [Ajg68, BC66, BW69, Bro69, DH59, DH63, DJ68, Fel64b, Gab63, Gei63c, Gei64, Gho65, Gho67, Gla62b, Goo60, Goo62a, Goo62b, GW66, HB68, HB69, HRK69, Hil65c, HC63, Jon68a, Jor64, Jor65, Kis62, KMZ69, Lil67, Lil69a, Low69, Man61, MM68, Raj64a, Raj65b, RB69, Rid60b, Rid60a, Ros61b, Sam66, TV68, TV69, WGF63, Wil62b, Wil67b, ZE66b, Zys69, BTM69]. Variances [Bec60b, GC63, JJ69, KH59, KH63, Lau61, MG69, Nau66, Put67, ST69, Sha61a, War67]. Variate [Kab63a, Mey67, Tik64, War67]. Variates [Nau69, War63b, Wil62a]. Variation [Abe62, Han63a, Han63b, Kha68, Les63, MB67b, Ros66c, Sea64, Sin63, Sin68b]. Variations [Leo62]. Various [SWC68]. Varying [Ajg65, Ajg68, Fel63, Hor63]. Vector [Cho66, DO67]. Vectors [Ste64f]. Vegetables [Wal65]. Version [Gas68, Ros66b]. Versions [Syk69]. Very [Has68]. View [Lon67, ZC65, ZZK68, ZC69]. Vocabulary [Die67]. Vol [Ano67e]. Volume [Ano60o, Ano61p, Ano62m, Ano63-30, Ano64m, Ano65q, Ano66s, Ano66t, Ano67q, Ano67r, Ano68v, Ano68w, Ano69t, Ano68f]. Volumes [Ano66f]. vs [SB68].

W [Alb67]. Wage [Bre62]. Wald [Bec60a, Lec64]. Walk [Tsu67]. Wallace [Tia67]. Warning [MH67]. Warning-Time [MH67]. Warning-Time/Failure-Time [MH67]. Washington [Ano60n, Ano68s]. Watson [AK69]. Way [Agr66, BHJ60, Hai66, Har67c, Har68b, Hil65c, HG65, Man61, PS67, Wil64]. Wealth [Cag64, Lee64c]. Weather [NS61]. Weibull [BT68a, Man69b]. Weight [Kar62, Tin67]. Weight-Height [Kar62]. Weighted [BW69, Mut60b, Mut62a, Wil67b]. Weighting [Wil61a, Wil62b]. Welfare [Dav62]. Werner [Sch66a]. Where [Tuk60]. Whether [GBB67]. Which [FW64, MFLR68, Wei62a, Bec60b, MG69, Sea66, Zys69]. Whites [Zel64].

Whitney [Cap61a, Hal60a, BKvE69, Hal60b, Mil64b]. who [Dav62]. Whose [Hud66]. Widely [Cro64]. Widows [CS62a]. Width [BG66, Ham63]. Wilcoxon [Hal60a, Klo67b, All63, BKvE69, Cap61a, FS64, Hal60b, Hol67, Jac63, Klo66, Leh61, McC65b, Mil64b, Noe63b, Noe67]. Wilks [STM^{+}65]. Will [Con66, Hen64a]. William [Cor67a, Fra69, Jac66, Dem69, Han69]. Winner [Zel69a]. Winning [Sea63]. Wisconsin [Adl63]. Wise [Kin67]. within [GC64]. Without [Ajg65, Con66, Fel63, Goo68, Har66, Raj66, Rao63b, Bre67]. Witte [Die67]. Work [Tin61]. Working [Kpe69]. World [Ste67b]. Worst [Whi66b].

Years [Bar65c, Mue63, Stu66]. Yiji [Dop67]. York [Ano62l]. Yotopoulos [Sch63]. Young [Kor66a, Kor66b]. Youths [Kar62].

Z [Sch66a]. Zellner [Kak67]. Zero [Cur67, EB61, Rut67, Tra67]. Zeros [Coh60d].

References

Anderson:1967:AUB

[AB67] T. W. Anderson and Herman Burstein. Approximating the upper binomial confidence limit. Journal of the American Statistical Association, 62(319):857–861, September 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

           jstor.org/stable/2283675.

Anderson:1968:ALB

[AB68] T. W. Anderson and Herman Burstein. Approximating the lower binomial confidence limit. Journal of the American Statistical Association, 63(324):1413–1415, December 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www. jstor.org/stable/2285894. See corrigenda [AB69a] and Editor’s correction [AB69b].

Anderson:1969:CAL

[AB69a] T. W. Anderson and Herman Burstein. Corrigenda: “Approximating the Lower Binomial Confidence Limit”. Journal of the American Statistical Association, 64(328):1699, December 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

           http://www.jstor.org/stable/2286121. See [AB68].

Anderson:1969:ECA

[AB69b] T. W. Anderson and Herman Burstein. Editor’s correction to Approximating the Lower Binomial Confidence Limit. Journal of the American Statistical Association, 64(326):669, June 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283652. See [AB68].

Abel:1962:HAS

[Abe62] Martin E. Abel. Harmonic analysis of seasonal variation with an application to hog production. Journal of the American Statistical Association, 57(299):655–667, September 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282403.

Abelson:1966:BRB

[Abe66] Robert P. Abelson. Book review: [untitled]. Journal of the American Statistical Association, 61(315):875, September 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282802.

Anscombe:1969:DPJ

[ABH^{+}69] F. J. Anscombe, Irwin D. J. Bross, H. O. Hartley, Rolf E. Bargmann, H. A. David, Marvin Zelen, R. L. Anderson, Miles Davis, and I. J. Good. Discussion of paper by I. J. Good, April 9, 1968. Journal of the American Statistical Association, 64(325): 50–66, March 1969. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/

2283719. See [Goo69].

Abrahamson:1969:BRB

[Abr69] Innis G. Abrahamson. Book review: [untitled]. Journal of the American Statistical Association, 64(325):401, March 1969. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283758.

Adams:1963:BRB

[Ada63] Earl W. Adams. Book review: [untitled]. Journal of the American Statistical Association, 58(303):866–867, September 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282754.

Adams:1969:BRB

[Ada69] F. Gerard Adams. Book review: [untitled]. Journal of the American Statistical Association, 64(328):1673–1674, December 1969. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2286105.

Addelman:1963:TCF

[Add63] Sidney Addelman. Techniques for constructing fractional replicate plans. Journal of the American Statistical Association, 58 (301):45–71, March 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2282954.

Addelman:1967:EPF

[Add67] Sidney Addelman. Equal and proportional frequency squares. Journal of the American Statistical Association, 62(317):226– 240, March 1967. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2282925.

Addelman:1969:BRB

[Add69] Sidney Addelman. Book review: [untitled]. Journal of the American Statistical Association, 64(327):1096–1097, September 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283497.

Adelman:1961:BRBa

[Ade61a] Irma Adelman. Book review: [untitled]. Journal of the American Statistical Association, 56(295):757–758, September 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282107.

Adelman:1961:BRBb

[Ade61b] Irma Adelman. Book review: [untitled]. Journal of the American Statistical Association, 56(295):762–763, September 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282113.

Adelman:1964:BRB

[Ade64] M. A. Adelman. Book review: [untitled]. Journal of the American Statistical Association, 59(306):602–604, June 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283013.

Adams:1966:SDR

[AdJ66] F. Gerard Adams and Peter E. de Janosi. On the statistical discrepancy in the revised U.S. national accounts. Journal of the American Statistical Association, 61(316):1219–1229, December 1966. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283211.

Adler:1963:TCR

[Adl63] Norman Adler. Taxpayer compliance in reporting interest income under the Wisconsin State Individual Income Tax. Journal of the American Statistical Association, 58(302):487–496, June 1963. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283283.

Afifi:1966:MOM

[AE66] A. A. Afifi and R. M. Elashoff. Missing observations in multivariate statistics: I. review of the literature. Journal of the American Statistical Association, 61(315):595–604, September 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282773.

Afifi:1967:MOM

[AE67] A. A. Afifi and R. M. Elashoff. Missing observations in multivariate statistics: II. point estimation in simple linear regression. Journal of the American Statistical Association, 62(317):10–29, March 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2282906.

Afifi:1969:MOMa

[AE69a] A. A. Afifi and R. M. Elashoff. Missing observations in multivariate statistics. III: Large sample analysis of simple linear regression. Journal of the American Statistical Association, 64(325): 337–358, March 1969. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2283744.

Afifi:1969:MOMb

[AE69b] A. A. Afifi and R. M. Elashoff. Missing observations in multivariate statistics. IV: a note on simple linear regression. Journal of the American Statistical Association, 64(325):359–365, March 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2283745.

Abul-Ela:1967:MPR

[AEGH67] Abdel-Latif A. Abul-Ela, Bernard G. Greenberg, and Daniel G. Horvitz. A multi-proportions randomized response model. Journal of the American Statistical Association, 62(319):990–1008, September 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2283687.

Afifi:1968:ISS

[AEL68] A. A. Afifi, R. M. Elashoff, and P. G. Langley. An investigation into the small sample properties of a two sample test of Lehmann’s. Journal of the American Statistical Association, 63 (321):345–352, March 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2283854.

Adelman:1961:IQC

[AG61] Irma Adelman and Zvi Griliches. On an index of quality change. Journal of the American Statistical Association, 56(295):535–548, September 1961. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2282078.

Agrawal:1964:BNC

[Agr64] Hiralal Agrawal. On the bounds of the number of common treatments between blocks of semi-regular group divisible designs. Journal of the American Statistical Association, 59(307):867–871, September 1964. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2283104.

Agrawal:1966:SMC

[Agr66] Hiralal Agrawal. Some methods of construction of designs for twoway elimination of heterogeneity, 1. Journal of the American Statistical Association, 61(316):1153–1171, December 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283205.

Aigner:1965:EPR

[Aig65] Dennis J. Aigner. An estimation procedure for range composition problems. Journal of the American Statistical Association, 60(309):308–319, March 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283155.

Aird:1961:BRB

[Air61] John S. Aird. Book review: [untitled]. Journal of the American Statistical Association, 56(294):443, June 1961. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282290.

Aitkin:1968:BRB

[Ait68] M. A. Aitkin. Book review: [untitled]. Journal of the American Statistical Association, 63(321):379–380, March 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283865.

Aiyar:1963:ULF

[Aiy63] K. R. Aiyar. On uncorrelated linear functions of order statistics. Journal of the American Statistical Association, 58(301):245– 246, March 1963. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2282968.

Ajgaonkar:1965:CLE

[Ajg65] S. G. Prabhu Ajgaonkar. On a class of linear estimators in sampling with varying probabilities without replacement. Journal of the American Statistical Association, 60(310):637–642, June 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282697.

Ajgaonkar:1968:BEV

[Ajg68] S. G. Prabhu Ajgaonkar. Bounds for the error-variance of an estimator in sampling with varying probabilities from a finite population. Journal of the American Statistical Association, 63(321): 91–98, March 1968. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2283830.

Ando:1965:BAI

[AK65] Albert Ando and G. M. Kaufman. Bayesian analysis of the independent multinormal process. neither mean nor precision known. Journal of the American Statistical Association, 60(309):347– 358, March 1965. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2283159.

Abrahamse:1969:CBP

[AK69] A. P. J. Abrahamse and J. Koerts. A comparison between the power of the Durbin–Watson test and the power of the BLUS test. Journal of the American Statistical Association, 64(327):938–948, September 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2283474.

Albert:1967:BRB

[Alb67] George E. Albert. Book review: Introduction to Probability and Statistics, by B. W. Lindgren; G. W. McElrath. Journal of the American Statistical Association, 62(317):305–306, March 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282942.

Allen:1960:BRB

[All60] Terrence M. Allen. Book review: [untitled]. Journal of the American Statistical Association, 55(292):774–776, December 1960. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281620.

Allen:1962:BRB

[All62] Robert Loring Allen. Book review: [untitled]. Journal of the American Statistical Association, 57(297):245, March 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282477.

Alling:1963:EDW

[All63] David W. Alling. Early decision in the Wilcoxon two-sample test. Journal of the American Statistical Association, 58(303):713–720, September 1963. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2282719.

Allen:1966:BRB

[All66] W. R. Allen. Book review: [untitled]. Journal of the American Statistical Association, 61(313):290–292, March 1966. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2283080.

Ando:1960:BRB

[And60] Albert Ando. Book review: [untitled]. Journal of the American Statistical Association, 55(292):764–765, December 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281611.

Anderson:1961:BRBa

[And61a] R. L. Anderson. Book review: [untitled]. Journal of the American Statistical Association, 56(294):414–415, June 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282271.

Anderson:1961:BRBb

[And61b] T. W. Anderson. Book review: [untitled]. Journal of the American Statistical Association, 56(294):419–421, June 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282275.

Anderson:1962:BRBa

[And62a] Henry Anderson. Book review: [untitled]. Journal of the American Statistical Association, 57(298):521–522, June 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281672.

Anderson:1962:BRBb

[And62b] W. H. Locke Anderson. Book review: [untitled]. Journal of the American Statistical Association, 57(298):523–524, June 1962. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281674.

Anderson:1963:PDC

[And63a] Henry Anderson. Product diversification and the cost of living. Journal of the American Statistical Association, 58(303):807–819, September 1963. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/

2282728. See corrigenda [Hen64b] and comment [And66].

Andrews:1963:BRB

[And63b] William H. Andrews. Book review: [untitled]. Journal of the American Statistical Association, 58(303):834–835, September 1963. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282734.

Anderson:1964:SAD

[And64] T. W. Anderson. Sequential analysis with delayed observations. Journal of the American Statistical Association, 59(308):1006– 1015, December 1964. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2282620.

Anderson:1966:PDL

[And66] Henry Anderson. Product diversification and living costs: a further comment. Journal of the American Statistical Association, 61 (315):788–799, September 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2282787. See [And63a].

Anderson:1967:BRB

[And67] Theodore R. Anderson. Book review: [untitled]. Journal of the American Statistical Association, 62(320):1502–1503, December 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2283798.

Anonymous:1960:BMa

[Ano60a] Anonymous. Back matter. Journal of the American Statistical Association, 55(289):244, March 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2282205.

Anonymous:1960:BMb

[Ano60b] Anonymous. Back matter. Journal of the American Statistical Association, 55(290):402–414, June 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281774.

Anonymous:1960:CEF

[Ano60c] Anonymous. Corrigenda: Economic Forecasting. Journal of the American Statistical Association, 55(292):756, December 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281604.

Anonymous:1960:CFM

[Ano60d] Anonymous. Corrigenda: Front Matter. Journal of the American Statistical Association, 55(292):756, December 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281605.

Anonymous:1960:FM

[Ano60e] Anonymous. Front matter. Journal of the American Statistical Association, 55(292):iii–v, December 1960. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281585.

Anonymous:1960:NAAa

[Ano60f] Anonymous. Notes about authors. Journal of the American Statistical Association, 55(289):212–216, March 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282188.

Anonymous:1960:NAAb

[Ano60g] Anonymous. Notes about authors. Journal of the American Statistical Association, 55(290):349, June 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281748.

Anonymous:1960:NAAc

[Ano60h] Anonymous. Notes about authors. Journal of the American Statistical Association, 55(291):581–582, September 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281916.

Anonymous:1960:NAAd

[Ano60i] Anonymous. Notes about authors. Journal of the American Statistical Association, 55(292):752–753, December 1960. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2281598.

Anonymous:1960:PRa

[Ano60j] Anonymous. Publications received. Journal of the American Statistical Association, 55(289):243, March 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282204.

Anonymous:1960:PRb

[Ano60k] Anonymous. Publications received. Journal of the American Statistical Association, 55(290):400–401, June 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281773.

Anonymous:1960:PRc

[Ano60l] Anonymous. Publications received. Journal of the American Statistical Association, 55(291):621–624, September 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281948.

Anonymous:1960:PRd

[Ano60m] Anonymous. Publications received. Journal of the American Statistical Association, 55(292):790–793, December 1960. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2281639.

Anonymous:1960:SPD

[Ano60n] Anonymous. Summaries of papers delivered at the 119th Annual Meeting of the American Statistical Association, Washington, D.C., December 27–30, 1959. Journal of the American Statistical Association, 55(290):350–374, June 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281749.

Anonymous:1960:VI

[Ano60o] Anonymous. Volume information. Journal of the American Statistical Association, 55(292):i–800, December 1960. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2281584.

Anonymous:1961:ASA

[Ano61a] Anonymous. American Statistical Association: Report of the Board of Directors, 1960. Journal of the American Statistical Association, 56(294):474–491, June 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282323.

Anonymous:1961:BMa

[Ano61b] Anonymous. Back matter. Journal of the American Statistical Association, 56(293):222, March 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2282391.

Anonymous:1961:BMb

[Ano61c] Anonymous. Back matter. Journal of the American Statistical Association, 56(294):492, June 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2282324.

Anonymous:1961:BMc

[Ano61d] Anonymous. Back matter. Journal of the American Statistical Association, 56(295):782, September 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282133.

Anonymous:1961:BMd

[Ano61e] Anonymous. Back matter. Journal of the American Statistical Association, 56(296):1044, December 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282042.

Anonymous:1961:FM

[Ano61f] Anonymous. Front matter. Journal of the American Statistical Association, 56(296):v–vii, December 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281993.

Anonymous:1961:NAAa

[Ano61g] Anonymous. Notes about authors. Journal of the American Statistical Association, 56(293):160–162, March 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282343.

Anonymous:1961:NAAb

[Ano61h] Anonymous. Notes about authors. Journal of the American Statistical Association, 56(294):384–387, June 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282265.

Anonymous:1961:NAAc

[Ano61i] Anonymous. Notes about authors. Journal of the American Statistical Association, 56(295):737–739, September 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282094.

Anonymous:1961:NAAd

[Ano61j] Anonymous. Notes about authors. Journal of the American Statistical Association, 56(296):1001–1004, December 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282012.

Anonymous:1961:PRa

[Ano61k] Anonymous. Publications received. Journal of the American Statistical Association, 56(294):471–473, June 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282322.

Anonymous:1961:PRb

[Ano61l] Anonymous. Publications received. Journal of the American Statistical Association, 56(295):778–781, September 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282132.

Anonymous:1961:PRc

[Ano61m] Anonymous. Publications received. Journal of the American Statistical Association, 56(296):1031–1033, December 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282041.

Anonymous:1961:PRM

[Ano61n] Anonymous. Publications received march 1961. Journal of the American Statistical Association, 56(293):219–221, March 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282390.

Anonymous:1961:SPD

[Ano61o] Anonymous. Summaries of papers delivered at the 120th Annual Meeting of the American Statistical Association, Stanford, California, August 23–6, 1960. Journal of the American Statistical Association, 56(294):388–410, June 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2282266.

Anonymous:1961:VI

[Ano61p] Anonymous. Volume information. Journal of the American Statistical Association, 56(296):i–1043, December 1961. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2281992.

Anonymous:1962:BMa

[Ano62a] Anonymous. Back matter. Journal of the American Statistical Association, 57(298):533–550, June 1962. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281689.

Anonymous:1962:BMb

[Ano62b] Anonymous. Back matter. Journal of the American Statistical Association, 57(300):951, December 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281856.

Anonymous:1962:FM

[Ano62c] Anonymous. Front matter. Journal of the American Statistical Association, 57(300):v–vii, December 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281803.

Anonymous:1962:NAAa

[Ano62d] Anonymous. Notes about authors. Journal of the American Statistical Association, 57(297):196–198, March 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282451.

Anonymous:1962:NAAb

[Ano62e] Anonymous. Notes about authors. Journal of the American Statistical Association, 57(298):478–480, June 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281653.

Anonymous:1962:NAAc

[Ano62f] Anonymous. Notes about authors. Journal of the American Statistical Association, 57(299):690–691, September 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282407.

Anonymous:1962:NAAd

[Ano62g] Anonymous. Notes about authors. Journal of the American Statistical Association, 57(300):914–916, December 1962. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2281819.

Anonymous:1962:PRa

[Ano62h] Anonymous. Publications received. Journal of the American Statistical Association, 57(297):265–268, March 1962. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282492.

Anonymous:1962:PRb

[Ano62i] Anonymous. Publications received. Journal of the American Statistical Association, 57(298):528–532, June 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281688.

Anonymous:1962:PRc

[Ano62j] Anonymous. Publications received. Journal of the American Statistical Association, 57(299):724–726, September 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282434.

Anonymous:1962:PRd

[Ano62k] Anonymous. Publications received. Journal of the American Statistical Association, 57(300):949–950, December 1962. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2281855.

Anonymous:1962:SPD

[Ano62l] Anonymous. Summaries of papers delivered at the 121st Annual Meeting of the American Statistical Association, New York City, December 27–30, 1961. Journal of the American Statistical Association, 57(298):481–501, June 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281654.

Anonymous:1962:VI

[Ano62m] Anonymous. Volume information. Journal of the American Statistical Association, 57(300):i–959, December 1962. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2281802.

Anonymous:1963:BMa

[Ano63a] Anonymous. Back matter. Journal of the American Statistical Association, 58(302):598–610, June 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283318.

Anonymous:1963:BMb

[Ano63b] Anonymous. Back matter. Journal of the American Statistical Association, 58(303):877–878, September 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282771.

Anonymous:1963:BMc

[Ano63c] Anonymous. Back matter. Journal of the American Statistical Association, 58(304):1209, December 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283394.

Anonymous:1963:NAAa

[Ano63d] Anonymous. Notes about authors. Journal of the American Statistical Association, 58(301):247–249, March 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282969.

Anonymous:1963:NAAb

[Ano63e] Anonymous. Notes about authors. Journal of the American Statistical Association, 58(302):538–540, June 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283290.

Anonymous:1963:NAAc

[Ano63f] Anonymous. Notes about authors. Journal of the American Statistical Association, 58(303):820–823, September 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282729.

Anonymous:1963:NAAd

[Ano63g] Anonymous. Notes about authors. Journal of the American Statistical Association, 58(304):1158–1160, December 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283340.

Anonymous:1963:PRa

[Ano63h] Anonymous. Publications received. Journal of the American Statistical Association, 58(302):595–597, June 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283317.

Anonymous:1963:PRb

[Ano63i] Anonymous. Publications received. Journal of the American Statistical Association, 58(304):1205–1208, December 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283393.

Anonymous:1963:PRM

[Ano63j] Anonymous. Publications received. march, 1963. Journal of the American Statistical Association, 58(301):271–274, March 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282991.

Anonymous:1963:PRS

[Ano63k] Anonymous. Publications received, September, 1963. Journal of the American Statistical Association, 58(303):873–876, September 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2282770.

Anonymous:1963:SPD

[Ano63l] Anonymous. Summaries of papers delivered at the 122nd Annual Meeting of the American Statistical Association, Minneapolis, Minnesota, September 7–10, 1962. Journal of the American Statistical Association, 58(302):541–564, June 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283291.

Anonymous:1963:Ua

[Ano63m] Anonymous. [untitled]. Journal of the American Statistical Association, 58(302):593, June 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283313.

Anonymous:1963:Ub

[Ano63n] Anonymous. [untitled]. Journal of the American Statistical Association, 58(302):593, June 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283314.

Anonymous:1963:Uc

[Ano63o] Anonymous. [untitled]. Journal of the American Statistical Association, 58(302):593–594, June 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283315.

Anonymous:1963:Ud

[Ano63p] Anonymous. [untitled]. Journal of the American Statistical Association, 58(302):594, June 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283316.

Anonymous:1963:Ue

[Ano63q] Anonymous. [untitled]. Journal of the American Statistical Association, 58(303):869, September 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2282757.

Anonymous:1963:Uf

[Ano63r] Anonymous. [untitled]. Journal of the American Statistical Association, 58(303):869, September 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2282758.

Anonymous:1963:Ug

[Ano63s] Anonymous. [untitled]. Journal of the American Statistical Association, 58(303):869, September 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2282759.

Anonymous:1963:Uh

[Ano63t] Anonymous. [untitled]. Journal of the American Statistical Association, 58(303):869–870, September 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282760.

Anonymous:1963:Ui

[Ano63u] Anonymous. [untitled]. Journal of the American Statistical Association, 58(303):870, September 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2282761.

Anonymous:1963:Uj

[Ano63v] Anonymous. [untitled]. Journal of the American Statistical Association, 58(303):870, September 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2282762.

Anonymous:1963:Uk

[Ano63w] Anonymous. [untitled]. Journal of the American Statistical Association, 58(303):870, September 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2282763.

Anonymous:1963:Ul

[Ano63x] Anonymous. [untitled]. Journal of the American Statistical Association, 58(303):870, September 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2282764.

Anonymous:1963:Um

[Ano63y] Anonymous. [untitled]. Journal of the American Statistical Association, 58(303):870–871, September 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282765.

Anonymous:1963:Un

[Ano63z] Anonymous. [untitled]. Journal of the American Statistical Association, 58(303):871, September 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2282766.

Anonymous:1963:Uo

[Ano63-27] Anonymous. [untitled]. Journal of the American Statistical Association, 58(303):871, September 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2282767.

Anonymous:1963:Up

[Ano63-28] Anonymous. [untitled]. Journal of the American Statistical Association, 58(303):871, September 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2282768.

Anonymous:1963:Uq

[Ano63-29] Anonymous. [untitled]. Journal of the American Statistical Association, 58(303):872, September 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2282769.

Anonymous:1963:VI

[Ano63-30] Anonymous. Volume information. Journal of the American Statistical Association, 58(304):i–1219, December 1963. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2283319.

Anonymous:1964:BMa

[Ano64a] Anonymous. Back matter. Journal of the American Statistical Association, 59(306):633–644, June 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283040.

Anonymous:1964:BMb

[Ano64b] Anonymous. Back matter. Journal of the American Statistical Association, 59(308):1337, December 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282677.

Anonymous:1964:FM

[Ano64c] Anonymous. Front matter. Journal of the American Statistical Association, 59(308):985–986, December 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282618.

Anonymous:1964:NAAa

[Ano64d] Anonymous. Notes about authors. Journal of the American Statistical Association, 59(305):273–277, March 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282875.

Anonymous:1964:NAAb

[Ano64e] Anonymous. Notes about authors. Journal of the American Statistical Association, 59(306):572–574, June 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283010.

Anonymous:1964:NAAc

[Ano64f] Anonymous. Notes about authors. Journal of the American Statistical Association, 59(307):961–964, September 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283114.

Anonymous:1964:NAAd

[Ano64g] Anonymous. Notes about authors. Journal of the American Statistical Association, 59(308):1292–1295, December 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282642.

Anonymous:1964:PRa

[Ano64h] Anonymous. Publications received. Journal of the American Statistical Association, 59(305):316–318, March 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282902.

Anonymous:1964:PRb

[Ano64i] Anonymous. Publications received. Journal of the American Statistical Association, 59(306):629–632, June 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283039.

Anonymous:1964:PRc

[Ano64j] Anonymous. Publications received. Journal of the American Statistical Association, 59(307):982–984, September 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283132.

Anonymous:1964:PRd

[Ano64k] Anonymous. Publications received. Journal of the American Statistical Association, 59(308):1327–1329, December 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282676.

Anonymous:1964:SPD

[Ano64l] Anonymous. Summaries of papers delivered at the 123rd Annual Meeting of the American Statistical Association, Cleveland, Ohio, September 4–7, 1963. Journal of the American Statistical Association, 59(306):575–599, June 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283011.

Anonymous:1964:VI

[Ano64m] Anonymous. Volume information. Journal of the American Statistical Association, 59(308):i–1336, December 1964. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2282617.

Anonymous:1965:BMa

[Ano65a] Anonymous. Back matter. Journal of the American Statistical Association, 60(309):393–394, March 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283182.

Anonymous:1965:BMb

[Ano65b] Anonymous. Back matter. Journal of the American Statistical Association, 60(310):685–698, June 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282708.

Anonymous:1965:BMc

[Ano65c] Anonymous. Back matter. Journal of the American Statistical Association, 60(311):936, September 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283269.

Anonymous:1965:CNA

[Ano65d] Anonymous. Corrigenda: Notes About Authors. Journal of the American Statistical Association, 60(312):1249, December 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283448.

Anonymous:1965:FM

[Ano65e] Anonymous. Front matter. Journal of the American Statistical Association, 60(312):??, December 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283396.

Anonymous:1965:NAAa

[Ano65f] Anonymous. Notes about authors. Journal of the American Statistical Association, 60(309):359–362, March 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283160.

Anonymous:1965:NAAb

[Ano65g] Anonymous. Notes about authors. Journal of the American Statistical Association, 60(310):643–684, June 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282698.

Anonymous:1965:NAAc

[Ano65h] Anonymous. Notes about authors. Journal of the American Statistical Association, 60(311):920–922, September 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283258.

Anonymous:1965:NAAd

[Ano65i] Anonymous. Notes about authors. Journal of the American Statistical Association, 60(312):1246–1248, December 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283444.

Anonymous:1965:PRa

[Ano65j] Anonymous. Publications received. Journal of the American Statistical Association, 60(309):390–392, March 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283181.

Anonymous:1965:PRb

[Ano65k] Anonymous. Publications received. Journal of the American Statistical Association, 60(310):682–684, June 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282707.

Anonymous:1965:PRc

[Ano65l] Anonymous. Publications received. Journal of the American Statistical Association, 60(311):934–935, September 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283268.

Anonymous:1965:PRd

[Ano65m] Anonymous. Publications received. Journal of the American Statistical Association, 60(312):1244–1245, December 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283443.

Anonymous:1965:SPD

[Ano65n] Anonymous. Summaries of papers delivered at the 124th Annual Meeting of the American Statistical Association, Chicago, Illinois, December 27–30, 1964. Journal of the American Statistical Association, 60(310):646–674, June 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2282699.

Anonymous:1965:Ua

[Ano65o] Anonymous. [untitled]. Journal of the American Statistical Association, 60(311):933, September 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283266.

Anonymous:1965:Ub

[Ano65p] Anonymous. [untitled]. Journal of the American Statistical Association, 60(311):934, September 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283267.

Anonymous:1965:VI

[Ano65q] Anonymous. Volume information. Journal of the American Statistical Association, 60(312):i–1258, December 1965. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2283395.

Anonymous:1966:BMa

[Ano66a] Anonymous. Back matter. Journal of the American Statistical Association, 61(313):303–304, March 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283089.

Anonymous:1966:BMb

[Ano66b] Anonymous. Back matter. Journal of the American Statistical Association, 61(314):560–575, June 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282855.

Anonymous:1966:BMc

[Ano66c] Anonymous. Back matter. Journal of the American Statistical Association, 61(315):896, September 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282819.

Anonymous:1966:BMd

[Ano66d] Anonymous. Back matter. Journal of the American Statistical Association, 61(316):1253–1254, December 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283239.

Anonymous:1966:CI

[Ano66e] Anonymous. Corrigenda: Index. Journal of the American Statistical Association, 61(316):1247–1248, December 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283231.

Anonymous:1966:CIV

[Ano66f] Anonymous. Corrigenda: Index to Volumes 51–60. Journal of the American Statistical Association, 61(316):1248, December 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283232.

Anonymous:1966:FM

[Ano66g] Anonymous. Front matter. Journal of the American Statistical Association, 61(314):??, June 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283183.

Anonymous:1966:IA

[Ano66h] Anonymous. Index of articles. Journal of the American Statistical Association, 61(314):1–23, June 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283184.

Anonymous:1966:IBRb

[Ano66i] Anonymous. Index of book reviewers. Journal of the American Statistical Association, 61(314):59–65, June 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283186.

Anonymous:1966:IBRa

[Ano66j] Anonymous. Index of book reviews. Journal of the American Statistical Association, 61(314):24–58, June 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283185.

Anonymous:1966:NAAa

[Ano66k] Anonymous. Notes about authors. Journal of the American Statistical Association, 61(313):300–302, March 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283088.

Anonymous:1966:NAAb

[Ano66l] Anonymous. Notes about authors. Journal of the American Statistical Association, 61(314):538–540, June 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282853.

Anonymous:1966:NAAc

[Ano66m] Anonymous. Notes about authors. Journal of the American Statistical Association, 61(315):887–891, September 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282817.

Anonymous:1966:NAAd

[Ano66n] Anonymous. Notes about authors. Journal of the American Statistical Association, 61(316):1244–1245, December 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283222.

Anonymous:1966:PRa

[Ano66o] Anonymous. Publications received. Journal of the American Statistical Association, 61(313):297–299, March 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283087.

Anonymous:1966:PRb

[Ano66p] Anonymous. Publications received. Journal of the American Statistical Association, 61(314):536–537, June 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282852.

Anonymous:1966:PRc

[Ano66q] Anonymous. Publications received. Journal of the American Statistical Association, 61(315):892–895, September 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282818.

Anonymous:1966:SPD

[Ano66r] Anonymous. Summaries of papers delivered at the 125th Annual Meeting of the American Statistical Association, Philadelphia, Pennsylvania, September 8–11, 1965. Journal of the American Statistical Association, 61(314):541–559, June 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282854.

Anonymous:1966:VIa

[Ano66s] Anonymous. Volume information. Journal of the American Statistical Association, 61(313):v, March 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283041.

Anonymous:1966:VIb

[Ano66t] Anonymous. Volume information. Journal of the American Statistical Association, 61(316):1255–1261, December 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283187.

Anonymous:1967:BMa

[Ano67a] Anonymous. Back matter. Journal of the American Statistical Association, 62(317):319–320, March 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282950.

Anonymous:1967:BMb

[Ano67b] Anonymous. Back matter. Journal of the American Statistical Association, 62(318):747–762, June 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284015.

Anonymous:1967:BMc

[Ano67c] Anonymous. Back matter. Journal of the American Statistical Association, 62(319):1101–1104, September 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283715.

Anonymous:1967:BMd

[Ano67d] Anonymous. Back matter. Journal of the American Statistical Association, 62(320):1524–1525, December 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283822.

Anonymous:1967:CNA

[Ano67e] Anonymous. Corrigenda: Notes about Authors. Omitted from Vol. 61 (1966). Journal of the American Statistical Association, 62 (320):1522–1523, December 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283821.

Anonymous:1967:FM

[Ano67f] Anonymous. Front matter. Journal of the American Statistical Association, 62(317):??, March 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2282904.

Anonymous:1967:NAAa

[Ano67g] Anonymous. Notes about authors. Journal of the American Statistical Association, 62(317):315–318, March 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282949.

Anonymous:1967:NAAb

[Ano67h] Anonymous. Notes about authors. Journal of the American Statistical Association, 62(318):716–720, June 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284012.

Anonymous:1967:NAAc

[Ano67i] Anonymous. Notes about authors. Journal of the American Statistical Association, 62(319):1094–1098, September 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283713.

Anonymous:1967:NAAd

[Ano67j] Anonymous. Notes about authors. Journal of the American Statistical Association, 62(320):1511–1515, December 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283808.

Anonymous:1967:PRa

[Ano67k] Anonymous. Publications received. Journal of the American Statistical Association, 62(317):311–314, March 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282948.

Anonymous:1967:PRb

[Ano67l] Anonymous. Publications received. Journal of the American Statistical Association, 62(318):721–722, June 1967. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284013.

Anonymous:1967:PRc

[Ano67m] Anonymous. Publications received. Journal of the American Statistical Association, 62(319):1099–1100, September 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283714.

Anonymous:1967:PRd

[Ano67n] Anonymous. Publications received. Journal of the American Statistical Association, 62(320):1516–1517, December 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283809.

Anonymous:1967:SPD

[Ano67o] Anonymous. Summaries of papers delivered at the 126th Annual Meeting of the American Statistical Association, Los Angeles, California, August 15–19, 1966. Journal of the American Statistical Association, 62(318):723–746, June 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284014.

Anonymous:1967:U

[Ano67p] Anonymous. [untitled]. Journal of the American Statistical Association, 62(320):1510, December 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283807.

Anonymous:1967:VIa

[Ano67q] Anonymous. Volume information. Journal of the American Statistical Association, 62(317):??, March 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282903.

Anonymous:1967:VIb

[Ano67r] Anonymous. Volume information. Journal of the American Statistical Association, 62(320):1527–1543, December 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283763.

Anonymous:1968:BMa

[Ano68a] Anonymous. Back matter. Journal of the American Statistical Association, 63(321):392–393, March 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283872.

Anonymous:1968:BMb

[Ano68b] Anonymous. Back matter. Journal of the American Statistical Association, 63(322):768–791, June 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284060.

Anonymous:1968:BMc

[Ano68c] Anonymous. Back matter. Journal of the American Statistical Association, 63(323):1073–1075, September 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283914.

Anonymous:1968:BMd

[Ano68d] Anonymous. Back matter. Journal of the American Statistical Association, 63(324):1552–1554, December 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285924.

Anonymous:1968:CNA

[Ano68e] Anonymous. Corrigenda: Notes About Authors. Journal of the American Statistical Association, 63(324):1550, December 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2285920.

Anonymous:1968:CVI

[Ano68f] Anonymous. Corrigenda: Volume Information. Journal of the American Statistical Association, 63(324):1549, December 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2285915.

Anonymous:1968:FMa

[Ano68g] Anonymous. Front matter. Journal of the American Statistical Association, 63(321):??, March 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283824.

Anonymous:1968:FMb

[Ano68h] Anonymous. Front matter. Journal of the American Statistical Association, 63(322):??, June 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284016.

Anonymous:1968:FMc

[Ano68i] Anonymous. Front matter. Journal of the American Statistical Association, 63(323):??, September 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283873.

Anonymous:1968:FMd

[Ano68j] Anonymous. Front matter. Journal of the American Statistical Association, 63(324):??, December 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285872.

Anonymous:1968:NAAa

[Ano68k] Anonymous. Notes about authors. Journal of the American Statistical Association, 63(321):386–390, March 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283870.

Anonymous:1968:NAAb

[Ano68l] Anonymous. Notes about authors. Journal of the American Statistical Association, 63(322):739–742, June 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284057.

Anonymous:1968:NAAc

[Ano68m] Anonymous. Notes about authors. Journal of the American Statistical Association, 63(323):1066–1069, September 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283912.

Anonymous:1968:NAAd

[Ano68n] Anonymous. Notes about authors. Journal of the American Statistical Association, 63(324):1542–1546, December 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2285913.

Anonymous:1968:PRa

[Ano68o] Anonymous. Publications received. Journal of the American Statistical Association, 63(321):391, March 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283871.

Anonymous:1968:PRb

[Ano68p] Anonymous. Publications received. Journal of the American Statistical Association, 63(322):743–744, June 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284058.

Anonymous:1968:PRc

[Ano68q] Anonymous. Publications received. Journal of the American Statistical Association, 63(323):1070–1072, September 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283913.

Anonymous:1968:PRd

[Ano68r] Anonymous. Publications received. Journal of the American Statistical Association, 63(324):1547–1548, December 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2285914.

Anonymous:1968:SPD

[Ano68s] Anonymous. Summaries of papers delivered at the 127th Annual Meeting of the American Statistical Association, Washington, D.C., December 27–30, 1967. Journal of the American Statistical Association, 63(322):745–767, June 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284059.

Anonymous:1968:Ua

[Ano68t] Anonymous. [untitled]. Journal of the American Statistical Association, 63(321):382, March 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283868.

Anonymous:1968:Ub

[Ano68u] Anonymous. [untitled]. Journal of the American Statistical Association, 63(324):1541, December 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285912.

Anonymous:1968:VIa

[Ano68v] Anonymous. Volume information. Journal of the American Statistical Association, 63(321):??, March 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283823.

Anonymous:1968:VIb

[Ano68w] Anonymous. Volume information. Journal of the American Statistical Association, 63(324):1555–1574, December 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2285871.

Anonymous:1969:BMa

[Ano69a] Anonymous. Back matter. Journal of the American Statistical Association, 64(325):411–413, March 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283762.

Anonymous:1969:BMb

[Ano69b] Anonymous. Back matter. Journal of the American Statistical Association, 64(326):710–725, June 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283669.

Anonymous:1969:BMc

[Ano69c] Anonymous. Back matter. Journal of the American Statistical Association, 64(327):1107–1109, September 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283504.

Anonymous:1969:BMd

[Ano69d] Anonymous. Back matter. Journal of the American Statistical Association, 64(328):1704–1706, December 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286147.

Anonymous:1969:CNA

[Ano69e] Anonymous. Corrigenda: Notes About Authors. Journal of the American Statistical Association, 64(328):1702, December 1969.

CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2286142.

Anonymous:1969:FMa

[Ano69f] Anonymous. Front matter. Journal of the American Statistical Association, 64(325):??, March 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283716.

Anonymous:1969:FMb

[Ano69g] Anonymous. Front matter. Journal of the American Statistical Association, 64(326):??, June 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283629.

Anonymous:1969:FMc

[Ano69h] Anonymous. Front matter. Journal of the American Statistical Association, 64(327):??, September 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283457.

Anonymous:1969:FMd

[Ano69i] Anonymous. Front matter. Journal of the American Statistical Association, 64(328):??, December 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286051.

Anonymous:1969:I

[Ano69j] Anonymous. [introduction]. Journal of the American Statistical Association, 64(328):1121, December 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286052.

Anonymous:1969:NAAa

[Ano69k] Anonymous. Notes about authors. Journal of the American Statistical Association, 64(325):403–407, March 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283760.

Anonymous:1969:NAAb

[Ano69l] Anonymous. Notes about authors. Journal of the American Statistical Association, 64(326):684–687, June 1969. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283666.

Anonymous:1969:NAAc

[Ano69m] Anonymous. Notes about authors. Journal of the American Statistical Association, 64(327):1100–1104, September 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283502.

Anonymous:1969:NAAd

[Ano69n] Anonymous. Notes about authors. Journal of the American Statistical Association, 64(328):1687–1695, December 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286118.

Anonymous:1969:PRa

[Ano69o] Anonymous. Publications received. Journal of the American Statistical Association, 64(325):408–410, March 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283761.

Anonymous:1969:PRb

[Ano69p] Anonymous. Publications received. Journal of the American Statistical Association, 64(326):688, June 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283667.

Anonymous:1969:PRc

[Ano69q] Anonymous. Publications received. Journal of the American Statistical Association, 64(327):1105–1106, September 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283503.

Anonymous:1969:PRd

[Ano69r] Anonymous. Publications received. Journal of the American Statistical Association, 64(328):1696–1698, December 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286119.

Anonymous:1969:SPD

[Ano69s] Anonymous. Summaries of papers delivered at the 128th Annual Meeting of the American Statistical Association, Pittsburgh, August 20–23, 1968. Journal of the American Statistical Association, 64(326):689–709, June 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283668.

Anonymous:1969:VI

[Ano69t] Anonymous. Volume information. Journal of the American Statistical Association, 64(328):1707–1728, December 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286050.

Anscombe:1961:EME

[Ans61a] F. J. Anscombe. Estimating a mixed-exponential response law. Journal of the American Statistical Association, 56(295):493–502, September 1961. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2282075.

Anscombe:1961:RIL

[Ans61b] F. J. Anscombe. Rectifying inspection of lots. Journal of the American Statistical Association, 56(296):807–823, December 1961. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281996.

Anscombe:1963:SMT

[Ans63] F. J. Anscombe. Sequential medical trials. Journal of the American Statistical Association, 58(302):365–383, June 1963. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2283272. See comments [Arm63].

Antelman:1965:INO

[Ant65] Gordon R. Antelman. Insensitivity to non-optimal design in Bayesian decision theory. Journal of the American Statistical Association, 60(310):584–601, June 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282693.

Antelman:1966:BRB

[Ant66] Gordon Antelman. Book review: [untitled]. Journal of the American Statistical Association, 61(313):265–267, March 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283063.

Antelman:1969:BRB

[Ant69] Gordon Antelman. Book review: [untitled]. Journal of the American Statistical Association, 64(328):1682–1683, December 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2286115.

Ames:1961:DCC

[AR61] Edward Ames and Stanley Reiter. Distributions of correlation coefficients in economic time series. Journal of the American Statistical Association, 56(295):637–656, September 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282085.

Armitage:1963:SMT

[Arm63] P. Armitage. Sequential medical trials: Some comments on F. J. Anscombe’s paper. Journal of the American Statistical Association, 58(302):384–387, June 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283273. See [Ans63].

Arnold:1960:BRB

[Arn60] K. J. Arnold. Book review: [untitled]. Journal of the American Statistical Association, 55(292):760–761, December 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281608.

Arnold:1967:NMD

[Arn67] Barry C. Arnold. A note on multivariate distributions with specified marginals. Journal of the American Statistical Association, 62 (320):1460–1461, December 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283790.

Arnold:1968:PEM

[Arn68] Barry C. Arnold. Parameter estimation for a multivariate exponential distribution. Journal of the American Statistical Association, 63(323):848–852, September 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283877.

Arrow:1961:BRB

[Arr61] Kenneth J. Arrow. Book review: [untitled]. Journal of the American Statistical Association, 56(295):760, September 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282110.

Arrow:1967:BRB

[Arr67] Kenneth J. Arrow. Book review: La Programmation Dynamique: Gestion Scientifique Sequentielle, by A. Kaufmann; R. Cruon. Journal of the American Statistical Association, 62(317):294– 295, March 1967. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2282935.

Anderson:1964:BRB

[AS64] W. H. Locke Anderson and G. Stigler. Book review: [untitled]. Journal of the American Statistical Association, 59(305): 297, March 1964. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2282888.

Astrachan:1965:BRB

[Ast65] Max Astrachan. Book review: [untitled]. Journal of the American Statistical Association, 60(309):388–389, March 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283180.

Atiqullah:1969:RLS

[Ati69] M. Atiqullah. On a restricted least squares estimator. Journal of the American Statistical Association, 64(327):964–968, September 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2283477.

Anderson:1969:CBC

[AZ69] T. R. Anderson and M. Zelditch, Jr. Corrigenda: “A Basic Course in Statistics, with Sociological Applications”. Journal of the American Statistical Association, 64(328):1699, December 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286129.

Bulgren:1968:NRD

[BA68] W. G. Bulgren and D. E. Amos. A note on representations of the doubly non-central t distribution. Journal of the American Statistical Association, 63(323):1013–1019, September 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283892.

Baar:1962:BRB

[Baa62] Jacob Baar. Book review: [untitled]. Journal of the American Statistical Association, 57(300):942–943, December 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281841.

Bachi:1969:CGR

[Bac69] Roberto Bachi. Corrigenda: Graphical Rational Patterns. A New Approach to Graphical Presentation of Statistics. Journal of the American Statistical Association, 64(328):1699, December 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2286128.

Baerncopf:1961:BRB

[Bae61] David A. Baerncopf. Book review: [untitled]. Journal of the American Statistical Association, 56(294):451–453, June 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282298.

Bailey:1961:BRB

[Bai61] Norman T. J. Bailey. Book review: [untitled]. Journal of the American Statistical Association, 56(296):1015–1016, December 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282020.

Bailey:1965:PAV

[Bai65] Martin J. Bailey. Prediction of an autoregressive variable subject both to disturbances and to errors of observation. Journal of the American Statistical Association, 60(309):164–181, March 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2283143.

Bain:1967:RRS

[Bai67] Lee J. Bain. Reducing a random sample to a smaller set, with applications. Journal of the American Statistical Association, 62 (318):510–519, June 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2283978.

Bailar:1968:RRR

[Bai68] Barbara A. Bailar. Recent research in reinterview procedures. Journal of the American Statistical Association, 63(321):41–63, March 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2283827.

Baker:1969:BRB

[Bak69] Frank B. Baker. Book review: [untitled]. Journal of the American Statistical Association, 64(326):681–682, June 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283664.

Bancroft:1960:BRB

[Ban60] Gertrude Bancroft. Book review: [untitled]. Journal of the American Statistical Association, 55(292):767–768, December 1960. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281614.

Bancroft:1961:BRB

[Ban61] T. A. Bancroft. Book review: [untitled]. Journal of the American Statistical Association, 56(293):184–185, March 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282358.

Bancroft:1962:BRB

[Ban62] T. A. Bancroft. Book review: [untitled]. Journal of the American Statistical Association, 57(299):717–718, September 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282424.

Banerjee:1963:INF

[Ban63] K. S. Banerjee. Index numbers for factorial effects and their connection with a special kind of irregular fractional plans of factorial experiments. Journal of the American Statistical Association, 58 (302):497–508, June 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2283284.

Bancroft:1965:BRB

[Ban65] T. A. Bancroft. Book review: [untitled]. Journal of the American Statistical Association, 60(309):386–387, March 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283178.

Barclay:1960:BRB

[Bar60] George W. Barclay. Book review: [untitled]. Journal of the American Statistical Association, 55(290):397–398, June 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281771.

Bargmann:1965:BRB

[Bar65a] Rolf E. Bargmann. Book review: [untitled]. Journal of the American Statistical Association, 60(312):1220–1222, December 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283424.

Barna:1965:BRB

[Bar65b] Tibor Barna. Book review: [untitled]. Journal of the American Statistical Association, 60(312):1240–1241, December 1965. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283436.

Bartlett:1965:RFL

[Bar65c] M. S. Bartlett. R. A. Fisher and the last fifty years of statistical methodology. Journal of the American Statistical Association, 60(310):395–409, June 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2282678.

Bargmann:1966:BRB

[Bar66a] Rolf E. Bargmann. Book review: [untitled]. Journal of the American Statistical Association, 61(316):1239–1241, December 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283220.

Barnett:1966:OSE

[Bar66b] V. D. Barnett. Order statistics estimators of the location of the Cauchy distribution. Journal of the American Statistical Association, 61(316):1205–1218, December 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www. jstor.org/stable/2283210. See corrigenda [Bar67b, Bar68].

Barr:1966:TEU

[Bar66c] D. R. Barr. On testing the equality of uniform and related distributions. Journal of the American Statistical Association, 61 (315):856–864, September 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2282795.

Bargmann:1967:BRB

[Bar67a] Rolf E. Bargmann. Book review: [untitled]. Journal of the American Statistical Association, 62(319):1080–1081, September 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283702.

Barnett:1967:COS

[Bar67b] V. D. Barnett. Corrigenda: Order Statistics Estimators of the Location of the Cauchy Distribution. Journal of the American Statistical Association, 62(320):1518, December 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283810. See [Bar66b].

Bartko:1967:BRB

[Bar67c] John J. Bartko. Book review: [untitled]. Journal of the American Statistical Association, 62(320):1507–1508, December 1967. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283802.

Barnett:1968:COS

[Bar68] V. D. Barnett. Correction: Order Statistics Estimators of the Location of the Cauchy Distribution. Journal of the American Statistical Association, 63(321):383–385, March 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283869. See [Bar66b].

Basmann:1960:FSD

[Bas60a] R. L. Basmann. On finite sample distributions of generalized classical linear identifiability test statistics. Journal of the American Statistical Association, 55(292):650–659, December 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281588.

Bassie:1960:BRB

[Bas60b] V. Lewis Bassie. Book review: [untitled]. Journal of the American Statistical Association, 55(290):378–379, June 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281753.

Basmann:1961:NEF

[Bas61] R. L. Basmann. A note on the exact finite sample frequency functions of generalized classical linear estimators in two leading overidentified cases. Journal of the American Statistical Association, 56(295):619–636, September 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2282084.

Basmann:1963:NEF

[Bas63a] R. L. Basmann. A note on the exact finite sample frequency functions of generalized classical linear estimators in a leading threeequation case. Journal of the American Statistical Association, 58(301):161–171, March 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2282960.

Basmann:1963:RCA

[Bas63b] R. L. Basmann. Remarks concerning the application of exact finite sample distribution functions of GCL estimators in econometric statistical inference. Journal of the American Statistical Association, 58(304):943–976, December 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283325. See corrigenda [Bas64].

Basmann:1964:CRC

[Bas64] R. L. Basmann. Corrigenda: Remarks Concerning the Application of Exact Finite Sample Distribution Functions of GCL Estimators in Econometric Statistical Inference. Journal of the American Statistical Association, 59(308):1296, December 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282644. See [Bas63b].

Basmann:1965:NST

[Bas65a] R. L. Basmann. A note on the statistical testability of ’explicit causal chains’ against the class of ’interdependent’ models. Journal of the American Statistical Association, 60(312):1080–1093, December 1965. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2283407.

Basmann:1965:BRB

[Bas65b] Robert L. Basmann. Book review: [untitled]. Journal of the American Statistical Association, 60(311):924–927, September 1965. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283260.

Basu:1965:CST

[Bas65c] A. P. Basu. Corrigenda: On Some Tests of Hypotheses Relating to the Exponential Distribution When Some Outliers Are Present. Journal of the American Statistical Association, 60(312): 1249, December 1965. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/

2283445. See [Bas65d].

Basu:1965:STH

[Bas65d] A. P. Basu. On some tests of hypotheses relating to the exponential distribution when some outliers are present. Journal of the American Statistical Association, 60(310):548–559, June 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282690. See [Bas65c].

Baten:1960:BRB

[Bat60] W. D. Baten. Book review: [untitled]. Journal of the American Statistical Association, 55(292):787–788, December 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281634.

Bateman:1968:BRB

[Bat68] Merrill J. Bateman. Book review: [untitled]. Journal of the American Statistical Association, 63(324):1538–1540, December 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2285909.

Baxter:1961:BRB

[Bax61] Glen E. Baxter. Book review: [untitled]. Journal of the American Statistical Association, 56(293):182–183, March 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282356.

Bardwell:1964:TPF

[BC64] George E. Bardwell and Edwin L. Crow. A two-parameter family of hyper-Poisson distributions. Journal of the American Statistical Association, 59(305):133–141, March 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282864.

Baker:1966:SER

[BC66] Frank B. Baker and Raymond O. Collier, Jr. Some empirical results on variance ratios under permutation in the completely randomized design. Journal of the American Statistical Association, 61(315):813–820, September 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2282790.

Baker:1968:ESF

[BC68a] Frank B. Baker and R. O. Collier. An empirical study into factors affecting the F -test under permutation for the randomized block design. Journal of the American Statistical Association, 63 (323):902–911, September 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283882.

Beauchamp:1968:SEP

[BC68b] John J. Beauchamp and Richard G. Cornell. Simultaneous estimation by partial totals for compartmental models. Journal of the American Statistical Association, 63(322):573–583, June 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2284028.

Blumenthal:1968:ELT

[BC68c] Saul Blumenthal and Arthur Cohen. Estimation of the larger of two normal means. Journal of the American Statistical Association, 63 (323):861–876, September 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283879.

Burr:1968:GSDa

[BC68d] Irving W. Burr and Peter J. Cislak. On a general system of distributions: I. Its curve-shape characteristics; II. The sample median. Journal of the American Statistical Association, 63(322):627–635, June 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2284033.

Berkson:1960:CER

[BE60a] Joseph Berkson and Lila Elveback. Competing exponential risks, with particular reference to the study of smoking and lung cancer. Journal of the American Statistical Association, 55(291):415– 428, September 1960. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/

2281905. See corrigenda [BE60b].

Berkson:1960:CCE

[BE60b] Joseph Berkson and Lila Elveback. Corrigenda: Competing Exponential Risks, with Particular Reference to the Study of Smoking and Lung Cancer. Journal of the American Statistical Association, 55(292):754, December 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281599. See [BE60a].

Beach:1960:BRB

[Bea60] E. F. Beach. Book review: [untitled]. Journal of the American Statistical Association, 55(290):386–387, June 1960. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2281759.

Bear:1964:BRB

[Bea64] Donald V. T. Bear. Book review: [untitled]. Journal of the American Statistical Association, 59(305):296–297, March 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282887.

Bechhofer:1960:NLR

[Bec60a] Robert Bechhofer. A note on the limiting relative efficiency of the Wald sequential probability ratio test. Journal of the American Statistical Association, 55(292):660–663, December 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281589.

Bechhofer:1960:MMA

[Bec60b] Robert E. Bechhofer. A multiplicative model for analyzing variances which are affected by several factors. Journal of the American Statistical Association, 55(290):245–264, June 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281738.

Beegle:1961:BRB

[Bee61] J. Allan Beegle. Book review: [untitled]. Journal of the American Statistical Association, 56(294):442, June 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282289.

Beem:1962:BRB

[Bee62] Eugene R. Beem. Book review: [untitled]. Journal of the American Statistical Association, 57(298):519–520, June 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281671.

Belsley:1968:BRB

[Bel68] David A. Belsley. Book review: [untitled]. Journal of the American Statistical Association, 63(323):1054–1057, September 1968. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283902.

Bennett:1962:PIS

[Ben62] George Bennett. Probability inequalities for the sum of independent random variables. Journal of the American Statistical Association, 57(297):33–45, March 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2282438.

Benishay:1966:SMC

[Ben66] Haskel Benishay. A stochastic model of credit sales debt. Journal of the American Statistical Association, 61(316):1010–1028, December 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2283196.

Berkson:1953:SPR

[Ber53] Joseph Berkson. A statistically precise and relatively simple method of estimating the bioassay with quantal response, based on the logistic function. Journal of the American Statistical Association, 48(263):565–599, September 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281010. See comments [Ber61b].

Berger:1961:CIA

[Ber61a] Agnes Berger. On comparing intensities of association between two binary characteristics in two different populations. Journal of the American Statistical Association, 56(296):889–908, December 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282002.

Berkson:1961:OSL

[Ber61b] Joseph Berkson. The other side of the lower bound. A note with a correction. Journal of the American Statistical Association, 56 (295):670–674, September 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2282087. See [Ber53].

Berman:1962:BRBa

[Ber62a] Simeon Berman. Book review: [untitled]. Journal of the American Statistical Association, 57(298):502–503, June 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281655.

Berman:1962:BRBb

[Ber62b] Simeon M. Berman. Book review: [untitled]. Journal of the American Statistical Association, 57(299):692–693, September 1962. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282408.

Berman:1963:BRB

[Ber63] Simeon M. Berman. Book review: [untitled]. Journal of the American Statistical Association, 58(302):565–566, June 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283292.

Berman:1969:BRB

[Ber69a] Abraham J. Berman. Book review: [untitled]. Journal of the American Statistical Association, 64(326):671–672, June 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283654.

Bertin:1969:CSG

[Ber69b] Jacques Bertin. Corrigenda: Seminologie Graphique: Les Diagrammes — Les Réseaux — Les Cartes. Journal of the American Statistical Association, 64(328):1699, December 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286122.

Brunk:1966:MLE

[BFHH66] H. D. Brunk, W. E. Franck, D. L. Hanson, and R. V. Hogg. Maximum likelihood estimation of the distributions of two stochastically ordered random variables. Journal of the American Statistical Association, 61(316):1067–1080, December 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283201.

Bowden:1966:CBU

[BG66] David C. Bowden and Franklin A. Graybill. Confidence bands of uniform and proportional width for linear models. Journal of the American Statistical Association, 61(313):182–198, March 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283053.

Bohrnstedt:1969:ECP

[BG69] George W. Bohrnstedt and Arthur S. Goldberger. On the exact covariance of products of random variables. Journal of the American Statistical Association, 64(328):1439–1442, December 1969. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2286081.

Bhattacharya:1965:DDS

[BH65] S. K. Bhattacharya and M. S. Holla. On a discrete distribution with special reference to the theory of accident proneness. Journal of the American Statistical Association, 60(312):1060–1066, December 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2283405.

Blischke:1966:APS

[BH66] W. R. Blischke and A. H. Halpin. Asymptotic properties of some estimators of quantiles of circular error. Journal of the American Statistical Association, 61(315):618–632, September 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282775.

Bhapkar:1966:NET

[Bha66] V. P. Bhapkar. A note on the equivalence of two test criteria for hypotheses in categorical data. Journal of the American Statistical Association, 61(313):228–235, March 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283057.

Bhattacharya:1967:BAL

[Bha67] Samir Kumar Bhattacharya. Bayesian approach to life testing and reliability estimation. Journal of the American Statistical Association, 62(317):48–62, March 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor. org/stable/2282909. See corrigenda [Bha69].

Bhattacharya:1969:CBA

[Bha69] Samir Kumar Bhattacharya. Corrigenda: Bayesian Approach to Life Testing and Reliability Estimation. Journal of the American Statistical Association, 64(328):1699, December 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286120. See [Bha67].

Bryant:1960:DET

[BHJ60] Edward C. Bryant, H. O. Hartley, and R. J. Jessen. Design and estimation in two-way stratification. Journal of the American Statistical Association, 55(289):105–124, March 1960. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2282182.

Braga-Illa:1964:SAB

[BI64] Alvise Braga-Illa. A simple approach to the Bayes choice criterion: The method of critical probabilities. Journal of the American Statistical Association, 59(308):1227–1230, December 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282640.

Bickel:1969:BRB

[Bic69] P. J. Bickel. Book review: [untitled]. Journal of the American Statistical Association, 64(325):397–399, March 1969. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2283753.

Billingsley:1960:BRB

[Bil60] Patrick Billingsley. Book review: [untitled]. Journal of the American Statistical Association, 55(290):377–378, June 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281752.

Birnbaum:1954:SMP

[Bir54] Allan Birnbaum. Statistical methods for Poisson processes and exponential populations. Journal of the American Statistical Association, 49(266):254–266, June 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor. org/stable/2280934. See corrigenda [Bir66].

Birnbaum:1961:CCO

[Bir61] Allan Birnbaum. Confidence curves: An omnibus technique for estimation and testing statistical hypotheses. Journal of the American Statistical Association, 56(294):246–249, June 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282249.

Birnbaum:1962:FSI

[Bir62] Allan Birnbaum. On the foundations of statistical inference. Journal of the American Statistical Association, 57(298):269–306, June 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2281640. See dis-

cussion [SBC^{+}62].

Birnbaum:1963:BRB

[Bir63] Z. W. Birnbaum. Book review: [untitled]. Journal of the American Statistical Association, 58(304):1164, December 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283352.

Birch:1964:NML

[Bir64] M. W. Birch. A note on the maximum likelihood estimation of a linear structural relationship. Journal of the American Statistical Association, 59(308):1175–1178, December 1964. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2282631.

Birnbaum:1966:CSM

[Bir66] Allan Birnbaum. Corrigenda: Statistical Methods for Poisson Processes and Exponential Populations. Journal of the American Statistical Association, 61(316):1246, December 1966. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2283225. See [Bir54].

Bjork:1964:BRB

[Bjo64] Gordon C. Bjork. Book review: [untitled]. Journal of the American Statistical Association, 59(305):301–303, March 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282890.

Brenna:1961:FTR

[BK61] Leroy Stanley Brenna and Clyde Young Kramer. Factorial treatments in rectangular lattice designs. Journal of the American Statistical Association, 56(294):368–378, June 1961. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2282263.

Bhat:1966:CLT

[BK66a] B. R. Bhat and S. R. Kulkarni. Corrigenda: LAMP Tests of Linear and Loglinear Hypotheses in Multinomial Experiments. Journal of the American Statistical Association, 61(316):1246, December 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2283224. See

[BK66b].

Bhat:1966:LTL

[BK66b] B. R. Bhat and S. R. Kulkarni. Lamp tests of linear and loglinear hypotheses in multinomial experiments. Journal of the American Statistical Association, 61(313):236–245, March 1966. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283058. See corrigenda

[BK66a].

Bruning:1969:CCH

[BK69] James L. Bruning and B. L. Kintz. Corrigenda: Computational Handbook of Statistics. Journal of the American Statistical Association, 64(328):1699, December 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286125.

Buckle:1969:AWM

[BKvE69] N. Buckle, C. Kraft, and C. van Eeden. An approximation to the Wilcoxon–Mann–Whitney distribution. Journal of the American Statistical Association, 64(326):591–599, June 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283641.

Birnbaum:1967:ERT

[BL67a] Allan Birnbaum and Eugene Laska. Efficiency robust two-sample rank tests. Journal of the American Statistical Association, 62 (320):1241–1251, December 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283773.

Birnbaum:1967:ORG

[BL67b] Allan Birnbaum and Eugene Laska. Optimal robustness: a general method, with applications to linear estimators of location. Journal of the American Statistical Association, 62(320):1230–1240, December 1967. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2283772.

Birnbaum:1969:TCV

[BL69] Z. W. Birnbaum and B. P. Lientz. Tables of critical values of some Renyi type statistics for finite sample sizes. Journal of the American Statistical Association, 64(327):870–877, September 1969. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283468.

Bleile:1963:BRB

[Ble63] George W. Bleile. Book review: [untitled]. Journal of the American Statistical Association, 58(301):266–268, March 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282983.

Blischke:1964:EPM

[Bli64] W. R. Blischke. Estimating the parameters of mixtures of binomial distributions. Journal of the American Statistical Association, 59(306):510–528, June 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283005.

Bloomfield:1960:BRB

[Blo60] Arthur I. Bloomfield. Book review: [untitled]. Journal of the American Statistical Association, 55(291):619–620, September 1960. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281947.

Bloomfield:1965:BRB

[Blo65] Arthur I. Bloomfield. Book review: [untitled]. Journal of the American Statistical Association, 60(312):1241–1242, December 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283437.

Bloch:1966:NEL

[Blo66] Daniel Bloch. A note on the estimation of the location parameter of the Cauchy distribution. Journal of the American Statistical Association, 61(315):852–855, September 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282794.

Bloom:1969:CPC

[Blo69] Benjamin S. Bloom. Corrigenda: Proceedings of the 1967 Conference on Testing Problems. Journal of the American Statistical Association, 64(328):1699, December 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286126.

Blumberg:1960:BRB

[Blu60] Mark S. Blumberg. Book review: [untitled]. Journal of the American Statistical Association, 55(290):396–397, June 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281770.

Blumen:1961:BRB

[Blu61] Isadore Blumen. Book review: [untitled]. Journal of the American Statistical Association, 56(296):1024–1025, December 1961. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282028.

Blumenthal:1968:MSP

[Blu68] Saul Blumenthal. Multinomial sampling with partially categorized data. Journal of the American Statistical Association, 63 (322):542–551, June 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284025.

Blyth:1960:BRB

[Bly60] Colin R. Blyth. Book review: [untitled]. Journal of the American Statistical Association, 55(290):399, June 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281772.

Beightler:1964:DOS

[BM64] C. S. Beightler and L. G. Mitten. Design of an optimal sequence of interrelated sampling plans. Journal of the American Statistical Association, 59(305):96–104, March 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282861.

Barlow:1965:TBD

[BM65] Richard E. Barlow and Albert W. Marshall. Tables of bounds for distributions with monotone hazard rate. Journal of the American Statistical Association, 60(311):872–890, September 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283254.

Bhattacharya:1967:RDH

[BM67a] N. Bhattacharya and B. Mahalanobis. Regional disparities in household consumption in India. Journal of the American Statistical Association, 62(317):143–161, March 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282918.

Bowman:1967:SEG

[BM67b] Mary Jean Bowman and Robert G. Myers. Schooling, experience, and gains and losses in human capital through migration. Journal of the American Statistical Association, 62(319):875–898, September 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2283677.

Bailey:1963:RMR

[BMN63] Martin J. Bailey, Richard F. Muth, and Hugh O. Nourse. A regression method for real estate price index construction. Journal of the American Statistical Association, 58(304):933–942, December 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2283324.

Barndorff-Nielsen:1964:BRB

[BN64] O. Barndorff-Nielsen. Book review: [untitled]. Journal of the American Statistical Association, 59(307):970–972, September 1964. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283120.

Bois:1969:SPL

[Boi69] N. S. D’Andrea Du Bois, Jr. A solution to the problem of linking multivariate documents. Journal of the American Statistical Association, 64(325):163–174, March 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283726.

Boles:1961:BRB

[Bol61] James N. Boles. Book review: [untitled]. Journal of the American Statistical Association, 56(293):202–203, March 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282371.

Bonnen:1963:BRB

[Bon63] James T. Bonnen. Book review: [untitled]. Journal of the American Statistical Association, 58(301):260–262, March 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282979.

Borts:1960:BRB

[Bor60a] George H. Borts. Book review: International Financial Transactions and Business Cycles, by Oskar Morgenstern. Journal of the American Statistical Association, 55(289):223–228, March 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282193.

Borts:1960:CRC

[Bor60b] George H. Borts. Corrigenda: Regional Cycles of Manufacturing Employment in the United States 1914–1953. Journal of the American Statistical Association, 55(292):755, December 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281600. See [Bor60c].

Borts:1960:RCM

[Bor60c] George H. Borts. Regional cycles of manufacturing employment in the United States, 1914–1953. Journal of the American Statistical Association, 55(289):151–211, March 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282187. See corrigenda [Bor60b].

Borus:1966:RES

[Bor66] Michael E. Borus. Response error in survey reports of earnings information. Journal of the American Statistical Association, 61 (315):729–738, September 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2282783.

Bossons:1964:BRB

[Bos64] John Bossons. Book review: [untitled]. Journal of the American Statistical Association, 59(307):977–979, September 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283125.

Boulding:1960:BRB

[Bou60] Kenneth E. Boulding. Book review: [untitled]. Journal of the American Statistical Association, 55(291):616–617, September 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2281943.

Bowlby:1963:BRB

[Bow63] Roger L. Bowlby. Book review: [untitled]. Journal of the American Statistical Association, 58(304):1184–1185, December 1963. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283368.

Bowman:1964:ASA

[Bow64] Raymond T. Bowman. The American Statistical Association and Federal statistics: Presidential address. Journal of the American Statistical Association, 59(305):1–17, March 1964. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2282856.

Bowker:1965:QQH

[Bow65] Albert H. Bowker. Quality and quantity in higher education. Journal of the American Statistical Association, 60(309):1–15, March 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2283133.

Bowman:1968:CCF

[Bow68] Raymond T. Bowman. Crossroad choices for the future development of the Federal Statistical System. Journal of the American Statistical Association, 63(323):801–816, September 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283874.

Boyd:1962:BRB

[Boy62] Harper W. Boyd, Jr. Book review: [untitled]. Journal of the American Statistical Association, 57(298):510–511, June 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281661.

Boyd:1965:NCS

[Boy65] William C. Boyd. A nomogram for chi-square. Journal of the American Statistical Association, 60(309):344–346, March 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283158. See corrigenda [Boy66].

Boyd:1966:CNC

[Boy66] William C. Boyd. Corrigenda: A Nomogram for Chi-Square. Journal of the American Statistical Association, 61(316):1246, December 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2283226. See

[Boy65].

Boyd:1969:NSF

[Boy69] William C. Boyd. A nomogram for the “Student”–Fisher t test. Journal of the American Statistical Association, 64(328):1664– 1667, December 1969. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2286098.

Barlow:1967:ELT

[BP67] Richard E. Barlow and Frank Proschan. Exponential life test procedures when the distribution has monotone failure rate. Journal of the American Statistical Association, 62(318):548–560, June 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283982.

Bharucha-Reid:1961:BRB

[BR61] A. T. Bharucha-Reid. Book review: [untitled]. Journal of the American Statistical Association, 56(294):444, June 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282292.

Barr:1966:IRS

[BR66] D. R. Barr and M. Haseeb Rizvi. An introduction to ranking and selection procedures. Journal of the American Statistical Association, 61(315):640–646, September 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2282777.

Bellman:1969:CFS

[BR69] Richard Bellman and Robert Roth. Curve fitting by segmented straight lines. Journal of the American Statistical Association, 64 (327):1079–1084, September 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283487.

Brabb:1963:BRB

[Bra63a] George J. Brabb. Book review: [untitled]. Journal of the American Statistical Association, 58(304):1195–1196, December 1963. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283376.

Brady:1963:BRB

[Bra63b] Dorothy S. Brady. Book review: [untitled]. Journal of the American Statistical Association, 58(301):253–254, March 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282973.

Brazer:1963:BRB

[Bra63c] Harvey E. Brazer. Book review: [untitled]. Journal of the American Statistical Association, 58(303):832–833, September 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282733.

Brady:1965:BRB

[Bra65] Dorothy S. Brady. Book review: [untitled]. Journal of the American Statistical Association, 60(309):373–374, March 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283168.

Brems:1960:BRB

[Bre60] Hans Brems. Book review: [untitled]. Journal of the American Statistical Association, 55(291):595–597, September 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281926.

Brennan:1961:BRB

[Bre61] Michael J. Brennan. Book review: [untitled]. Journal of the American Statistical Association, 56(295):763–764, September 1961. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282114.

Brems:1962:WPT

[Bre62] Hans Brems. Wage, price, and tax elasticities of output and distributive shares. Journal of the American Statistical Association, 57(299):607–621, September 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2282398.

Brewer:1966:BRB

[Bre66] James K. Brewer. Book review: [untitled]. Journal of the American Statistical Association, 61(313):282–283, March 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283073.

Brewer:1967:NFM

[Bre67] K. R. W. Brewer. A note on Fellegi’s method of sampling without replacement with probability proportional to size. Journal of the American Statistical Association, 62(317):79–85, March 1967. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282911.

Brimmer:1960:BRB

[Bri60] Andrew F. Brimmer. Book review: [untitled]. Journal of the American Statistical Association, 55(292):781, December 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281625.

Brinegar:1963:MTQ

[Bri63] Claude S. Brinegar. Mark Twain and the Quintus Curtius Snodgrass Letters: a statistical test of authorship. Journal of the American Statistical Association, 58(301):85–96, March 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282956.

Brittain:1964:BRB

[Bri64] John A. Brittain. Book review: [untitled]. Journal of the American Statistical Association, 59(305):311–313, March 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282894.

Brillinger:1969:BRB

[Bri69] David R. Brillinger. Book review: [untitled]. Journal of the American Statistical Association, 64(325):390–391, March 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283749.

Brownlee:1960:SEC

[Bro60a] K. A. Brownlee. Statistical evaluation of cloud seeding operations. Journal of the American Statistical Association, 55(291): 446–453, September 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/ stable/2281907. See comments [NS61].

Brownlee:1960:BRB

[Bro60b] O. H. Brownlee. Book review: [untitled]. Journal of the American Statistical Association, 55(291):585–586, September 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281919.

Brock:1961:BRB

[Bro61a] Paul Brock. Book review: [untitled]. Journal of the American Statistical Association, 56(294):466, June 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282310.

Bronfenbrenner:1961:BRB

[Bro61b] M. Bronfenbrenner. Book review: [untitled]. Journal of the American Statistical Association, 56(296):1026–1027, December 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282031.

Brown:1961:EAE

[Bro61c] Murray Brown. Ex ante and ex post data in inventory investment. Journal of the American Statistical Association, 56(295): 518–534, September 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2282077.

Brownlee:1961:BRBa

[Bro61d] K. A. Brownlee. Book review: [untitled]. Journal of the American Statistical Association, 56(294):421–422, June 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282276.

Brownlee:1961:BRBb

[Bro61e] O. H. Brownlee. Book review: [untitled]. Journal of the American Statistical Association, 56(294):424–425, June 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282279.

Bronfenbrenner:1963:BRB

[Bro63a] Martin Bronfenbrenner. Book review: [untitled]. Journal of the American Statistical Association, 58(302):585–587, June 1963. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283308.

Bross:1963:BRB

[Bro63b] Irwin D. J. Bross. Book review: [untitled]. Journal of the American Statistical Association, 58(304):1171–1173, December 1963. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283356.

Brownlee:1963:BRB

[Bro63c] K. A. Brownlee. Book review: [untitled]. Journal of the American Statistical Association, 58(304):1181–1182, December 1963. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283365.

Bross:1964:TCA

[Bro64] Irwin D. J. Bross. Taking a covariable into account. Journal of the American Statistical Association, 59(307):725–736, September 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2283094.

Brownlee:1965:RSH

[Bro65] K. A. Brownlee. A review of “smoking and health”. Journal of the American Statistical Association, 60(311):722–739, September 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2283241.

Brownlee:1967:BRB

[Bro67] K. A. Brownlee. Book review: [untitled]. Journal of the American Statistical Association, 62(318):690–694, June 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283998.

Brown:1968:BRB

[Bro68] Murray Brown. Book review: [untitled]. Journal of the American Statistical Association, 63(321):367–370, March 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283856.

Broemeling:1969:CRV

[Bro69] L. D. Broemeling. Confidence regions for variance ratios of random models. Journal of the American Statistical Association, 64 (326):660–664, June 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2283650.

Brunk:1960:MMR

[Bru60] H. D. Brunk. Mathematical models for ranking from paired comparisons. Journal of the American Statistical Association, 55(291): 503–520, September 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2281911.

Brunner:1961:BRB

[Bru61] Karl Brunner. Book review: [untitled]. Journal of the American Statistical Association, 56(294):454–455, June 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282301.

Brunk:1962:BRB

[Bru62] H. D. Brunk. Book review: [untitled]. Journal of the American Statistical Association, 57(297):238–239, March 1962. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2282470.

Brunk:1966:BRB

[Bru66] H. D. Brunk. Book review: [untitled]. Journal of the American Statistical Association, 61(313):295, March 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283083.

Bryson:1965:ECB

[Bry65] Marion R. Bryson. Errors of classification in a binomial population. Journal of the American Statistical Association, 60(309): 217–224, March 1965. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2283147.

Booth:1969:PST

[BS69a] Gordon Booth and J. Sedransk. Planning some two-factor comparative surveys. Journal of the American Statistical Association, 64 (326):560–573, June 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2283638.

Bryson:1969:SCA

[BS69b] Maurice C. Bryson and M. M. Siddiqui. Some criteria for aging. Journal of the American Statistical Association, 64(328): 1472–1483, December 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286084.

Burch:1969:FAC

[BS69c] S. W. Burch and H. O. Stekler. The forecasting accuracy of consumer attitude data. Journal of the American Statistical Association, 64(328):1225–1233, December 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286063.

Brown:1962:FMF

[BT62] Murray Brown and Paul Taubman. A forecasting model of Federal purchases of goods and services. Journal of the American Statistical Association, 57(299):633–647, September 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282401.

Bain:1968:STH

[BT68a] Lee J. Bain and Darrel R. Thoman. Some tests of hypotheses concerning the three-parameter Weibull distribution. Journal of the American Statistical Association, 63(323):853–860, September 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2283878.

Box:1968:BEM

[BT68b] G. E. P. Box and G. C. Tiao. Bayesian estimation of means for the random effect model. Journal of the American Statistical Association, 63(321):174–181, March 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283838.

Bershad:1969:DHS

[BT69] Max A. Bershad and Benjamin J. Tepping. The development of household sample surveys. Journal of the American Statistical Association, 64(328):1134–1140, December 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286055.

Blischke:1969:NRE

[BTM69] W. R. Blischke, A. J. Truelove, and P. B. Mundle. On non-regular estimation. I. Variance bounds for estimators of location parameters. Journal of the American Statistical Association, 64(327): 1056–1072, September 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2283485.

Buechley:1961:RMC

[Bue61] Robert W. Buechley. A reproducible method of counting persons of Spanish surname. Journal of the American Statistical Association, 56(293):88–97, March 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2282333.

Burkholder:1961:BRB

[Bur61] D. L. Burkholder. Book review: [untitled]. Journal of the American Statistical Association, 56(295):759–760, September 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282109.

Burke:1962:BRB

[Bur62a] C. J. Burke. Book review: [untitled]. Journal of the American Statistical Association, 57(298):506–507, June 1962. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2281658.

Burkholder:1962:BRB

[Bur62b] D. L. Burkholder. Book review: [untitled]. Journal of the American Statistical Association, 57(297):205–207, March 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282455.

Burr:1968:BRB

[Bur68a] Irving W. Burr. Book review: [untitled]. Journal of the American Statistical Association, 63(322):732–733, June 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284051.

Burr:1968:GSDb

[Bur68b] Irving W. Burr. On a general system of distributions: III. the sample range. Journal of the American Statistical Association, 63 (322):636–643, June 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284034.

Bush:1962:BRB

[Bus62] Robert R. Bush. Book review: [untitled]. Journal of the American Statistical Association, 57(299):696–698, September 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282410.

Bain:1965:TLG

[BW65] Lee J. Bain and David L. Weeks. Tolerance limits for the generalized gamma distribution. Journal of the American Statistical Association, 60(312):1142–1152, December 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283412.

Ballas:1966:CDT

[BW66a] J. A. Ballas and J. T. Webster. Corrigenda: On Dependent Tests from a Non-Orthogonal Design. Journal of the American Statistical Association, 61(316):1246, December 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283223. See [BW66b].

Ballas:1966:DTN

[BW66b] J. A. Ballas and J. T. Webster. On dependent tests from a nonorthogonal design. Journal of the American Statistical Association, 61(315):803–812, September 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor. org/stable/2282789. See corrigenda [BW66a].

Bement:1969:VWR

[BW69] T. R. Bement and J. S. Williams. Variance of weighted regression estimators when sampling errors are independent and heteroscedastic. Journal of the American Statistical Association, 64 (328):1369–1382, December 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286075.

Byrnes:1965:BRB

[Byr65] James C. Byrnes. Book review: [untitled]. Journal of the American Statistical Association, 60(309):370–371, March 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283166.

C:1961:BRBa

[C.61a] J. C. Book review: [untitled]. Journal of the American Statistical Association, 56(294):469, June 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282318.

C:1961:BRBb

[C.61b] J. C. Book review: [untitled]. Journal of the American Statistical Association, 56(294):470, June 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282321.

C:1961:BRBc

[C.61c] J. C. Book review: [untitled]. Journal of the American Statistical Association, 56(295):775, September 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282124.

C:1962:BRBa

[C.62a] J. C. Book review: [untitled]. Journal of the American Statistical Association, 57(297):264, March 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282490.

C:1962:BRBb

[C.62b] J. C. Book review: [untitled]. Journal of the American Statistical Association, 57(300):944, December 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281844.

C:1962:BRBc

[C.62c] J. C. Book review: [untitled]. Journal of the American Statistical Association, 57(300):944, December 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281845.

C:1962:BRBd

[C.62d] J. C. Book review: [untitled]. Journal of the American Statistical Association, 57(300):945–946, December 1962. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281847.

C:1962:BRBe

[C.62e] J. C. Book review: [untitled]. Journal of the American Statistical Association, 57(300):947, December 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281852.

C:1962:BRBf

[C.62f] J. C. Book review: [untitled]. Journal of the American Statistical Association, 57(300):947, December 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281853.

C:1962:BRBg

[C.62g] J. C. Book review: [untitled]. Journal of the American Statistical Association, 57(300):948, December 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281854.

C:1963:BRB

[C.63] J. C. Book review: [untitled]. Journal of the American Statistical Association, 58(301):269, March 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282988.

C:1969:BRB

[C.69] E. L. C. Book review: [untitled]. Journal of the American Statistical Association, 64(325):402, March 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283759.

Cutler:1963:PPP

[CA63] Sidney J. Cutler and Lillian M. Axtell. Partitioning of a patient population with respect to different mortality risks. Journal of the American Statistical Association, 58(303):701–712, September 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2282718.

Cabello:1960:USF

[Cab60] Octavio Cabello. The use of statistics in the formulation and evaluation of social programmes. Journal of the American Statistical Association, 55(291):454–468, September 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281908.

Cacoullos:1965:RBD

[Cac65] Theophilos Cacoullos. A relation between t and F -distributions. Journal of the American Statistical Association, 60(310):528–531, June 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2282687. See

corrigenda [The65b].

Cacoullos:1967:ADG

[Cac67] T. Cacoullos. Asymptotic distribution for a generalized Banach match box problem. Journal of the American Statistical Association, 62(320):1252–1257, December 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283774.

Cady:1963:BRB

[Cad63] Foster B. Cady. Book review: [untitled]. Journal of the American Statistical Association, 58(303):865–866, September 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282753.

Cagan:1964:UWC

[Cag64] Philip Cagan. The use of wealth to compare households’ average saving. Journal of the American Statistical Association, 59 (307):737–745, September 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283095.

Cain:1968:BRB

[Cai68] George L. Cain, Jr. Book review: [untitled]. Journal of the American Statistical Association, 63(322):737, June 1968. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2284054.

Calvin:1969:BRB

[Cal69] Lyle D. Calvin. Book review: [untitled]. Journal of the American Statistical Association, 64(327):1085–1087, September 1969. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283488.

Campbell:1961:BRB

[Cam61] Howard E. Campbell. Book review: [untitled]. Journal of the American Statistical Association, 56(294):467, June 1961. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2282312.

Capon:1961:NAN

[Cap61a] Jack Capon. A note on the asymptotic normality of the Mann– Whitney–Wilcoxon statistic. Journal of the American Statistical Association, 56(295):687–691, September 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282089.

Capron:1961:BRB

[Cap61b] William M. Capron. Book review: [untitled]. Journal of the American Statistical Association, 56(295):768, September 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282118.

Capon:1965:AEK

[Cap65] Jack Capon. On the asymptotic efficiency of the Kolmogorov– Smirnov test. Journal of the American Statistical Association, 60 (311):843–853, September 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283250.

Carter:1961:BRB

[Car61] H. O. Carter. Book review: [untitled]. Journal of the American Statistical Association, 56(294):454, June 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282300.

Carter:1962:BRB

[Car62] Anne P. Carter. Book review: [untitled]. Journal of the American Statistical Association, 57(297):247–248, March 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282479.

Charnes:1962:CCN

[CC62] A. Charnes and W. W. Cooper. Chance constraints and normal deviates. Journal of the American Statistical Association, 57 (297):134–148, March 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2282444.

Chakravarty:1962:BRB

[Cha62] S. Chakravarty. Book review: [untitled]. Journal of the American Statistical Association, 57(298):513–515, June 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281664.

Chatterjee:1968:MSS

[Cha68] Samprit Chatterjee. Multivariate stratified surveys. Journal of the American Statistical Association, 63(322):530–534, June 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2284023.

Chenery:1960:BRB

[Che60] Hollis B. Chenery. Book review: [untitled]. Journal of the American Statistical Association, 55(290):387–388, June 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281760.

Chew:1966:CPT

[Che66] Victor Chew. Confidence, prediction, and tolerance regions for the multivariate normal distribution. Journal of the American Statistical Association, 61(315):605–617, September 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282774.

Cornfield:1969:APS

[CHG69] Jerome Cornfield, Max Halperin, and Samuel W. Greenhouse. An adaptive procedure for sequential clinical trials. Journal of the American Statistical Association, 64(327):759–770, September 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2283459.

Chipman:1962:BRB

[Chi62] John S. Chipman. Book review: [untitled]. Journal of the American Statistical Association, 57(297):241–242, March 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282473.

Chipman:1964:LSI

[Chi64] John S. Chipman. On least squares with insufficient observations. Journal of the American Statistical Association, 59(308): 1078–1111, December 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2282625. See [Chi65b].

Chipman:1965:BRB

[Chi65a] John S. Chipman. Book review: [untitled]. Journal of the American Statistical Association, 60(312):1215–1217, December 1965. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283421.

Chipman:1965:CLS

[Chi65b] John S. Chipman. Corrigenda: On Least Squares with Insufficient Observation. Journal of the American Statistical Association, 60(312):1249, December 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283447. See [Chi64].

Chow:1961:BRB

[Cho61] Gregory C. Chow. Book review: [untitled]. Journal of the American Statistical Association, 56(293):209–210, March 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282378.

Chovitz:1964:BRB

[Cho64] Bernard Chovitz. Book review: [untitled]. Journal of the American Statistical Association, 59(308):1318–1319, December 1964. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282665.

Chow:1966:TLS

[Cho66] Gregory C. Chow. A theorem on least squares and vector correlation in multivariate linear regressions. Journal of the American Statistical Association, 61(314):413–414, June 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282831.

Chover:1967:BRB

[Cho67] Joshua Chover. Book review: [untitled]. Journal of the American Statistical Association, 62(319):1070–1071, September 1967. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283695.

Chover:1969:CMP

[Cho69] Joshua Chover. Corrigenda: Markov processes and potential theory. Journal of the American Statistical Association, 64(328): 1699, December 1969. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2286123.

Chun:1965:ERS

[Chu65] D. Chun. On an extreme rank sum test with early decision. Journal of the American Statistical Association, 60(311):859–863, September 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2283252.

Church:1966:BRB

[Chu66] Donald E. Church. Book review: [untitled]. Journal of the American Statistical Association, 61(313):280, March 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283070.

Cohen:1969:RRF

[CJ69] Ayala Cohen and Richard H. Jones. Regression on a random field. Journal of the American Statistical Association, 64(328): 1172–1182, December 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286060.

Chipman:1968:ESM

[CKMP68] John S. Chipman, Koteswara Rao Kadiyala, Albert Madansky, and John W. Pratt. Efficiency of the sample mean when residuals follow a first-order stationary Markoff process. Journal of the American Statistical Association, 63(324):1237–1246, December 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2285880. See

corrigenda [CKMP69].

Chipman:1969:CES

[CKMP69] John S. Chipman, Koteswara Rao Kadiyala, Albert Madansky, and John W. Pratt. Corrigenda: Efficiency of the Sample Mean when Residuals Follow a First-Order Stationary Markoff Process. Journal of the American Statistical Association, 64(328):1700, December 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2286134. See

[CKMP68].

Chow:1969:SPN

[CL69] Gregory C. Chow and Richard E. Levitan. Spectral properties of non-stationary systems of linear stochastic difference equations. Journal of the American Statistical Association, 64(326):581–590, June 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2283640.

Clark:1960:BRB

[Cla60] Lincoln Clark. Book review: [untitled]. Journal of the American Statistical Association, 55(291):617–618, September 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281944.

Clark:1961:BRB

[Cla61] Paul G. Clark. Book review: [untitled]. Journal of the American Statistical Association, 56(295):750–751, September 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282102.

Clawson:1966:REI

[Cla66] Marion Clawson. Recent efforts to improve land use information. Journal of the American Statistical Association, 61(315): 647–657, September 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2282778.

Cleroux:1969:FSM

[Cle69] Robert Cleroux. First and second moments of the randomization test in two-associate PBIB designs. Journal of the American Statistical Association, 64(328):1424–1433, December 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286079.

Cliff:1965:BRB

[Cli65] Norman Cliff. Book review: [untitled]. Journal of the American Statistical Association, 60(312):1225–1227, December 1965. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283426.

Connor:1969:CIP

[CM69] Robert J. Connor and James E. Mosimann. Concepts of independence for proportions with a generalization of the Dirichlet distribution. Journal of the American Statistical Association, 64 (325):194–206, March 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2283728.

Coale:1968:CHP

[Coa68] Ansley J. Coale. Convergence of a human population to a stable form. Journal of the American Statistical Association, 63(322): 395–435, June 1968. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2284017.

Cochran:1961:BRB

[Coc61] William G. Cochran. Book review: [untitled]. Journal of the American Statistical Association, 56(294):418–419, June 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282274.

Cochran:1962:BRB

[Coc62] William G. Cochran. Book review: [untitled]. Journal of the American Statistical Association, 57(298):503–505, June 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281656.

Cochran:1964:BRB

[Coc64] W. G. Cochran. Book review: [untitled]. Journal of the American Statistical Association, 59(308):1315–1316, December 1964. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282663.

Cohen:1960:BRBa

[Coh60a] Jacob Cohen. Book review: [untitled]. Journal of the American Statistical Association, 55(289):229–230, March 1960. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2282195.

Cohen:1960:BRBb

[Coh60b] Morris Cohen. Book review: [untitled]. Journal of the American Statistical Association, 55(291):611–613, September 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281939.

Cohen:1960:EPM

[Coh60c] A. Clifford Cohen, Jr. Estimating the parameters of a modified Poisson distribution. Journal of the American Statistical Association, 55(289):139–143, March 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2282184.

Cohen:1960:ETP

[Coh60d] A. Clifford Cohen, Jr. Estimation in the truncated Poisson distribution when zeros and some ones are missing. Journal of the American Statistical Association, 55(290):342–348, June 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281747.

Cohen:1962:BRBa

[Coh62a] A. Clifford Cohen. Book review: [untitled]. Journal of the American Statistical Association, 57(299):700–701, September 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282413.

Cohen:1962:BRBb

[Coh62b] A. Clifford Cohen. Book review: [untitled]. Journal of the American Statistical Association, 57(299):701, September 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282414.

Cohen:1968:BRB

[Coh68] Arthur Cohen. Book review: [untitled]. Journal of the American Statistical Association, 63(321):373–374, March 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283860.

Colm:1960:BRB

[Col60a] Gerhard Colm. Book review: [untitled]. Journal of the American Statistical Association, 55(289):236–237, March 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282200.

Colton:1960:TPS

[Col60b] Theodore Colton. A test procedure with a sample from a normal population when an upper bound to the standard deviation is known. Journal of the American Statistical Association, 55 (289):94–104, March 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2282181.

Colm:1962:BRB

[Col62] Gerhard Colm. Book review: [untitled]. Journal of the American Statistical Association, 57(299):693–696, September 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282409.

Colton:1963:MSO

[Col63] Theodore Colton. A model for selecting one of two medical treatments. Journal of the American Statistical Association, 58(302): 388–400, June 1963. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2283274.

Conrad:1961:BRB

[Con61] Alfred H. Conrad. Book review: [untitled]. Journal of the American Statistical Association, 56(295):752–753, September 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282104.

Connor:1966:EFP

[Con66] W. S. Connor. An exact formula for the probability that two specified sampling units will occur in a sample drawn with unequal probabilities and without replacement. Journal of the American Statistical Association, 61(314):384–390, June 1966. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2282827.

Conolly:1968:BRB

[Con68a] B. W. Conolly. Book review: [untitled]. Journal of the American Statistical Association, 63(322):725, June 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284045.

Conover:1968:TSS

[Con68b] W. J. Conover. Two K-sample slippage tests. Journal of the American Statistical Association, 63(322):614–626, June 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284032.

Conlisk:1969:ECM

[Con69a] John Conlisk. The equilibrium covariance matrix of dynamic econometric models. Journal of the American Statistical Association, 64(325):277–279, March 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283735.

Connor:1969:SDR

[Con69b] Robert J. Connor. The sampling distribution of the range from discrete uniform finite populations and a range test for homogeneity. Journal of the American Statistical Association, 64(328): 1443–1458, December 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286082.

Cootner:1962:BRB

[Coo62] Paul H. Cootner. Book review: [untitled]. Journal of the American Statistical Association, 57(298):511–512, June 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281662.

Cornell:1960:BRB

[Cor60] Francis G. Cornell. Book review: [untitled]. Journal of the American Statistical Association, 55(290):381–382, June 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281755.

Cornfield:1966:BTS

[Cor66] Jerome Cornfield. A Bayesian test of some classical hypotheses, with applications to sequential clinical trials. Journal of the American Statistical Association, 61(315):577–594, September 1966. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282772.

Cornell:1967:BRBa

[Cor67a] Richard G. Cornell. Book review: Epidemiological Study of Cancer and Other Chronic Diseases, by William Haenszel. Journal of the American Statistical Association, 62(317):293–294, March 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2282934.

Cornell:1967:BRBb

[Cor67b] Richard G. Cornell. Book review: [untitled]. Journal of the American Statistical Association, 62(319):1085, September 1967. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283706.

Cowden:1960:BRB

[Cow60] Dudley J. Cowden. Book review: [untitled]. Journal of the American Statistical Association, 55(289):218–220, March 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282191.

Cowden:1963:BRB

[Cow63] Dudley J. Cowden. Book review: [untitled]. Journal of the American Statistical Association, 58(303):867–868, September 1963. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282755.

Cox:1961:BRBb

[Cox61a] Edwin B. Cox. Book review: [untitled]. Journal of the American Statistical Association, 56(295):774–775, September 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282123.

Cox:1961:CSD

[Cox61b] Edwin B. Cox. Changes in the size distribution of dividend income. Journal of the American Statistical Association, 56(294):250–259, June 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2282250.

Cox:1961:BRBa

[Cox61c] Gertrude M. Cox. Book review: [untitled]. Journal of the American Statistical Association, 56(294):416–417, June 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282272.

Cox:1966:LSS

[CR66] C. Philip Cox and Thomas D. Roseberry. A large sample sequential test, using concomitant information, for discrimination between two composite hypotheses. Journal of the American Statistical Association, 61(314):357–367, June 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282824.

Cramer:1964:SCM

[Cra64a] Elliot M. Cramer. Some comparisons of methods of fitting the dosage response curve for small samples. Journal of the American Statistical Association, 59(307):779–793, September 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283098.

Cramer:1964:EGR

[Cra64b] J. S. Cramer. Efficient grouping, regression and correlation in Engel curve analysis. Journal of the American Statistical Association, 59(305):233–250, March 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2282872.

Cragg:1966:SSE

[Cra66] J. G. Cragg. On the sensitivity of simultaneous-equations estimators to the stochastic assumptions of the models. Journal of the American Statistical Association, 61(313):136–151, March 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283050.

Cromarty:1960:BRB

[Cro60] William A. Cromarty. Book review: [untitled]. Journal of the American Statistical Association, 55(291):609–610, September 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2281936.

Cronholm:1964:TVG

[Cro64] James N. Cronholm. A two-variable generating function for computing the sampling probabilities of a class of widely used statistics. Journal of the American Statistical Association, 59(306):487–491, June 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2283003.

Cross:1968:BRB

[Cro68a] John G. Cross. Book review: [untitled]. Journal of the American Statistical Association, 63(324):1533–1536, December 1968. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2285907.

Croteau:1968:BRB

[Cro68b] John T. Croteau. Book review: [untitled]. Journal of the American Statistical Association, 63(324):1529–1531, December 1968. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2285904.

Crofts:1969:BRB

[Cro69] A. E. Crofts, Jr. Book review: [untitled]. Journal of the American Statistical Association, 64(325):400–401, March 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283756.

Coale:1962:CIT

[CS62a] Ansley J. Coale and Frederick F. Stephan. The case of the Indians and the teen-age widows. Journal of the American Statistical Association, 57(298):338–347, June 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281643.

Connor:1962:DTS

[CS62b] W. S. Connor and Norman C. Severo. Distribution of total service time for a fixed observation interval. Journal of the American Statistical Association, 57(298):376–386, June 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281646.

Crow:1967:REL

[CS67] Edwin L. Crow and M. M. Siddiqui. Robust estimation of location. Journal of the American Statistical Association, 62(318):353–389, June 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2283968.

Cunia:1964:LSE

[Cun64] T. Cunia. Least squares estimates and parabolic regression with restricted location for the stationary point. Journal of the American Statistical Association, 59(306):564–571, June 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283009.

Cureton:1967:NAS

[Cur67] Edward E. Cureton. The normal approximation to the signed-rank sampling distribution when zero differences are present. Journal of the American Statistical Association, 62(319):1068–1069, September 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2283694.

Conner:1968:UDE

[CW68] James R. Conner and Delane E. Welsch. Use of domain estimators with unequal probability in sample surveys. Journal of the American Statistical Association, 63(323):984–992, September 1968. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283888.

D:1966:BRB

[D.66] H. A. D. Book review: Elements of Numerical Analysis, by Peter Henrici. Journal of the American Statistical Association, 61 (313):295–296, March 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2283084.

Dacy:1964:PPI

[Dac64] Douglas C. Dacy. A price and productivity index for a nonhomogeneous product. Journal of the American Statistical Association, 59(306):469–480, June 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283001.

Dalenius:1967:BRB

[Dal67] Tore Dalenius. Book review: [untitled]. Journal of the American Statistical Association, 62(320):1497–1498, December 1967. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283795.

Daly:1969:SBP

[Dal69] Joseph F. Daly. Some basic principles of statistical surveys. Journal of the American Statistical Association, 64(328):1129–1133, December 1969. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2286054.

Daniel:1962:CSF

[Dan62a] Cuthbert Daniel. Corrigenda: Sequences of Fractional Replicates in the 2^{p−q} Series. Journal of the American Statistical Association, 57(300):919, December 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281823. See [Dan62b].

Daniel:1962:SFR

[Dan62b] Cuthbert Daniel. Sequences of fractional replicates in the 2^{p−q} series. Journal of the American Statistical Association, 57(298): 403–429, June 1962. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/

2281648. See corrigenda [Dan62a].

Darling:1960:BRB

[Dar60] Donald A. Darling. Book review: [untitled]. Journal of the American Statistical Association, 55(292):786–787, December 1960. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281632.

Darling:1963:BRB

[Dar63] D. A. Darling. Book review: [untitled]. Journal of the American Statistical Association, 58(303):839–840, September 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282739.

Deely:1969:ETT

[DAS69] J. J. Deely, D. E. Amos, and G. P. Steck. The exceedance test for truncation of a Supplier’s data. Journal of the American Statistical Association, 64(327):823–829, September 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283463.

Davies:1961:MLR

[Dav61a] M. Davies. Multiple linear regression analysis with adjustment for class differences. Journal of the American Statistical Association, 56(295):729–735, September 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2282092.

Davis:1961:BRB

[Dav61b] Lance E. Davis. Book review: [untitled]. Journal of the American Statistical Association, 56(295):747–749, September 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282099.

David:1962:VIR

[Dav62] Martin David. The validity of income reported by a sample of families who received welfare assistance during 1959. Journal of the American Statistical Association, 57(299):680–685, September 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2282405.

David:1964:BRB

[Dav64] Martin David. Book review: [untitled]. Journal of the American Statistical Association, 59(307):970, September 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283119.

David:1965:BRB

[Dav65] Martin David. Book review: [untitled]. Journal of the American Statistical Association, 60(311):923–924, September 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283259.

David:1966:BRB

[Dav66] H. T. David. Book review: [untitled]. Journal of the American Statistical Association, 61(315):882–883, September 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282811.

Davis:1967:PCG

[Dav67] James A. Davis. A partial coefficient for Goodman and Kruskal’s gamma. Journal of the American Statistical Association, 62(317): 189–193, March 1967. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2282922.

Dunn:1969:CCM

[DC69] Olive Jean Dunn and Virginia Clark. Correlation coefficients measured on the same individuals. Journal of the American Statistical Association, 64(325):366–377, March 1969. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283746.

Deane:1964:BRB

[Dea64] Phyllis Deane. Book review: [untitled]. Journal of the American Statistical Association, 59(305):278–279, March 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282876.

DeGroot:1966:BRB

[DeG66] Morris H. DeGroot. Book review: How to Gamble If You Must: Inequalities for Stochastic Processes, by Lester E. Dubins; Leonard J. Savage. Journal of the American Statistical Association, 61 (314):531–533, June 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2282848.

Deming:1963:BRB

[Dem63] W. Edwards Deming. Book review: [untitled]. Journal of the American Statistical Association, 58(304):1164–1166, December 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283353.

Dempster:1964:DIF

[Dem64] A. P. Dempster. On the difficulties inherent in Fisher’s fiducial argument. Journal of the American Statistical Association, 59 (305):56–66, March 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2282858.

Deming:1965:BRB

[Dem65] W. Edwards Deming. Book review: [untitled]. Journal of the American Statistical Association, 60(312):1228–1230, December 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283428.

Dempster:1966:BRB

[Dem66] A. P. Dempster. Book review: [untitled]. Journal of the American Statistical Association, 61(316):1233–1235, December 1966. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283215.

Deming:1969:WH

[Dem69] W. Edwards Deming. On William Hurwitz. Journal of the American Statistical Association, 64(328):1150–1151, December 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2286057. See comments

[Fra69].

Denton:1963:STA

[Den63] Frank T. Denton. Some techniques for analyzing a set of time series subject to a linear restriction. Journal of the American Statistical Association, 58(302):513–518, June 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283286.

Derber:1961:BRBa

[Der61a] Milton Derber. Book review: [untitled]. Journal of the American Statistical Association, 56(295):765–766, September 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282116.

Derber:1961:BRBb

[Der61b] Milton Derber. Book review: [untitled]. Journal of the American Statistical Association, 56(295):766–768, September 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282117.

Derrick:1962:BRB

[Der62] Lucile Derrick. Book review: [untitled]. Journal of the American Statistical Association, 57(297):244, March 1962. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2282476.

Dalenius:1959:MVS

[DH59] Tore Dalenius and Joseph L. Hodges, Jr. Minimum variance stratification. Journal of the American Statistical Association, 54(285): 88–101, March 1959. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/

2282141. See corrigenda [DH63].

Dalenius:1963:CMV

[DH63] Tore Dalenius and Joseph L. Hodges, Jr. Corrigenda: Minimum Variance Stratification. Journal of the American Statistical Association, 58(304):1161, December 1963. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283341. See [DH59].

Dhrymes:1964:BRB

[Dhr64] Phoebus J. Dhrymes. Book review: [untitled]. Journal of the American Statistical Association, 59(306):613–614, June 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283021.

Diamond:1967:BRB

[Dia67] Peter Diamond. Book review: [untitled]. Journal of the American Statistical Association, 62(319):1073–1075, September 1967. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283697.

Dieter:1967:BRB

[Die67] Ulrich Dieter. Book review: German–English Mathematical Vocabulary, by Sheila MacIntyre; Edith Witte. Journal of the American Statistical Association, 62(317):303, March 1967. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2282939.

Dixon:1965:MSS

[Dix65] W. J. Dixon. The up-and-down method for small samples. Journal of the American Statistical Association, 60(312):967–978, December 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2283398.

Duncan:1966:MRS

[DJ66] David B. Duncan and Richard H. Jones. Multiple regression with stationary errors. Journal of the American Statistical Association, 61(316):917–928, December 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283189.

Davis:1968:EIV

[DJ68] Herbert T. Davis and Richard H. Jones. Estimation of the innovation variance of a stationary time series. Journal of the American Statistical Association, 63(321):141–149, March 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283835.

Dunn:1965:EMC

[DM65] Olive Jean Dunn and Frank J. Massey, Jr. Estimation of multiple contrasts using t-distributions. Journal of the American Statistical Association, 60(310):573–583, June 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282692.

David:1968:OSD

[DM68a] H. A. David and R. S. Mishriky. Order statistics for discrete populations and for grouped samples. Journal of the American Statistical Association, 63(324):1390–1398, December 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2285892.

Duran:1968:RSS

[DM68b] Benjamin S. Duran and Paul W. Mielke, Jr. Robustness of sum of squared ranks test. Journal of the American Statistical Association, 63(321):338–344, March 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283853.

David:1965:EDS

[DN65] H. A. David and J. E. Norman, Jr. Exact distribution of the sum of independent identically distributed discrete random variables. Journal of the American Statistical Association, 60(311): 837–842, September 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/ stable/2283249. See corrigenda [DN66].

David:1966:CED

[DN66] H. A. David and J. E. Norman, Jr. Corrigenda: Exact Distribution of the Sum of Independent Identically Distributed Discrete Random Variables. Journal of the American Statistical Association, 61 (316):1246–1247, December 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283227. See [DN65].

Decell:1967:FPP

[DO67] Henry P. Decell, Jr. and P. L. Odell. On the fixed point probability vector of regular or ergodic transition matrices. Journal of the American Statistical Association, 62(318):600–602, June 1967.

CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283986.

Donaldson:1968:RTE

[Don68] Theodore S. Donaldson. Robustness of the F -test to errors of both kinds and the correlation between the numerator and denominator of the F -ratio. Journal of the American Statistical Association, 63(322):660–676, June 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284037.

Doob:1963:BRB

[Doo63] J. L. Doob. Book review: [untitled]. Journal of the American Statistical Association, 58(304):1201, December 1963. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2283382.

Dopuch:1967:BRB

[Dop67] Nicholas Dopuch. Book review: Management Goals and Accounting for Control, by Yiji Ijiri. Journal of the American Statistical Association, 62(317):300–301, March 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282937.

Dorn:1960:BRB

[Dor60] Harold F. Dorn. Book review: [untitled]. Journal of the American Statistical Association, 55(291):586–587, September 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281920.

Dorn:1961:BRBa

[Dor61a] Harold F. Dorn. Book review: [untitled]. Journal of the American Statistical Association, 56(294):429–430, June 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282282.

Dorn:1961:BRBb

[Dor61b] Harold F. Dorn. Book review: [untitled]. Journal of the American Statistical Association, 56(294):439–440, June 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282286.

dePascual:1961:URE

[dP61] Jose Nieto de Pascual. Unbiased ratio estimators in stratified sampling. Journal of the American Statistical Association, 56(293): 70–87, March 1961. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/

2282332. See corrigenda [dP64].

dePascual:1964:CUR

[dP64] Jose Nieto de Pascual. Corrigenda: Unbiased Ratio Estimators in Stratified Sampling. Journal of the American Statistical Association, 59(308):1298–1299, December 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282649. See [dP61].

Drane:1969:BRB

[DR69] J. Wanzer Drane and Joan S. Reisch. Book review: [untitled]. Journal of the American Statistical Association, 64(328):1679–1680, December 1969. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2286112.

Drummond:1962:BRB

[Dru62] Ian M. Drummond. Book review: [untitled]. Journal of the American Statistical Association, 57(297):210–211, March 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282457.

Dempster:1965:ESL

[DS65] A. P. Dempster and M. Schatzoff. Expected significance level as a sensitivity index for test statistics. Journal of the American Statistical Association, 60(310):420–436, June 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282680.

Duncan:1960:BRB

[Dun60] Otis Dudley Duncan. Book review: [untitled]. Journal of the American Statistical Association, 55(292):757–758, December 1960. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281606.

Duncan:1961:OCE

[Dun61a] Otis Dudley Duncan. Occupational components of educational differences in income. Journal of the American Statistical Association, 56(296):783–792, December 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281994.

Dunn:1961:MCA

[Dun61b] Olive Jean Dunn. Multiple comparisons among means. Journal of the American Statistical Association, 56(293):52–64, March 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282330.

Duncan:1962:BRBa

[Dun62a] Acheson J. Duncan. Book review: [untitled]. Journal of the American Statistical Association, 57(297):231–232, March 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282467.

Duncan:1962:BRBb

[Dun62b] Otis Dudley Duncan. Book review: [untitled]. Journal of the American Statistical Association, 57(300):941–942, December 1962. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281840.

Dunn:1966:BRB

[Dun66] Olive Jean Dunn. Book review: [untitled]. Journal of the American Statistical Association, 61(315):876–877, September 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282804.

Dunn:1968:NCB

[Dun68] Olive Jean Dunn. A note on confidence bands for a regression line over a finite range. Journal of the American Statistical Association, 63(323):1028–1033, September 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283894.

Dunn:1969:CMR

[Dun69] James E. Dunn. A compounded multiple runs distribution. Journal of the American Statistical Association, 64(328):1415–1423, December 1969. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2286078.

Duvall:1966:TSA

[Duv66] Richard M. Duvall. Time series analysis by modified least-squares techniques. Journal of the American Statistical Association, 61 (313):152–165, March 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2283051.

Dwyer:1967:CSA

[Dwy67a] Paul S. Dwyer. Corrigenda: Some Applications of Matrix Derivatives in Multivariate Analysis. Journal of the American Statistical Association, 62(320):1518, December 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283811. See [Dwy67b].

Dwyer:1967:SAM

[Dwy67b] Paul S. Dwyer. Some applications of matrix derivatives in multivariate analysis. Journal of the American Statistical Association, 62(318):607–625, June 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor. org/stable/2283988. See corrigenda [Dwy67a].

Deely:1969:SCI

[DZ69] J. J. Deely and W. J. Zimmer. Shorter confidence intervals using prior observations. Journal of the American Statistical Association, 64(325):378–386, March 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283747.

E:1968:BRB

[E.68] W. A. E. Book review: [untitled]. Journal of the American Statistical Association, 63(322):738, June 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284056.

Eagleson:1968:NLR

[Eag68] G. K. Eagleson. A note on linear regression in trivariate distributions. Journal of the American Statistical Association, 63(323): 1042–1044, September 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2283897.

Easterlin:1961:BRB

[Eas61] Richard A. Easterlin. Book review: [untitled]. Journal of the American Statistical Association, 56(293):178, March 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282352.

Easterlin:1964:BRB

[Eas64] Richard A. Easterlin. Book review: [untitled]. Journal of the American Statistical Association, 59(305):279–283, March 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282877.

Easterling:1969:DIP

[Eas69] Robert G. Easterling. Discrimination intervals for percentiles in regression. Journal of the American Statistical Association, 64 (327):1031–1041, September 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283483.

Eaton:1969:SRS

[Eat69] Morris L. Eaton. Some remarks on Scheffe’s solution to the Behrens–Fisher problem. Journal of the American Statistical Association, 64(328):1318–1322, December 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286070.

El-Badry:1961:FEM

[EB61] M. A. El-Badry. Failure of enumerators to make entries of zero: Errors in recording childless cases in population censuses. Journal of the American Statistical Association, 56(296):909–924, December 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2282003.

El-Badry:1969:HFT

[EB69] M. A. El-Badry. Higher female than male mortality in some countries of South Asia: a digest. Journal of the American Statistical Association, 64(328):1234–1244, December 1969. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2286064.

Eckstein:1961:BRB

[Eck61] Otto Eckstein. Book review: [untitled]. Journal of the American Statistical Association, 56(294):427–428, June 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282281.

Ederer:1961:PES

[Ede61] Fred Ederer. A parametric estimate of the standard error of the survival rate. Journal of the American Statistical Association, 56 (293):111–118, March 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/ stable/2282336. See corrigenda [Ede63].

Ederer:1963:CPE

[Ede63] Fred Ederer. Corrigenda: A Parametric Estimate of the Standard Error of the Survival Rate. Journal of the American Statistical Association, 58(304):1161, December 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283342. See [Ede61].

Edgington:1961:PTN

[Edg61] Eugene S. Edgington. Probability table for number of runs of signs of first differences in ordered series. Journal of the American Statistical Association, 56(293):156–159, March 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282342.

Edwards:1964:BRB

[Edw64] Alfred L. Edwards. Book review: [untitled]. Journal of the American Statistical Association, 59(307):967, September 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283116.

Edwards:1967:BRB

[Edw67] Ward Edwards. Book review: Decision and Choice: Contributions of Sidney Siegel, by Sidney Siegel; S. Messick; A. H. Brayfield. Journal of the American Statistical Association, 62(317):291– 293, March 1967. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2282933.

Efron:1967:BRB

[Efr67] Bradley Efron. Book review: Linear Statistical Inference and its Applications, by C. Radhakrishna Rao. Journal of the American Statistical Association, 62(317):290–291, March 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282932.

Efron:1969:STU

[Efr69] Bradley Efron. Student’s t-test under symmetry conditions. Journal of the American Statistical Association, 64(328):1278–1302, December 1969. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2286068.

Edwards:1961:CDA

[EG61] Carol B. Edwards and John Gurland. A class of distributions applicable to accidents. Journal of the American Statistical Association, 56(295):503–517, September 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2282076.

Evans:1966:REI

[EG66] Michael K. Evans and Edward W. Green. The relative efficacy of investment anticipations. Journal of the American Statistical Association, 61(313):104–116, March 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283047.

Engelman:1969:PPT

[EH69] L. Engelman and J. A. Hartigan. Percentage points of a test for clusters. Journal of the American Statistical Association, 64 (328):1647–1648, December 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286096.

Ehrenfeld:1962:SED

[Ehr62] Sylvain Ehrenfeld. Some experimental design problems in attribute life testing. Journal of the American Statistical Association, 57 (299):668–679, September 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor. org/stable/2282404. See corrigenda [Ehr63b].

Ehrenfeld:1963:BRB

[Ehr63a] S. Ehrenfeld. Book review: [untitled]. Journal of the American Statistical Association, 58(304):1177–1178, December 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283361.

Ehrenfeld:1963:CSE

[Ehr63b] Sylvain Ehrenfeld. Corrigenda: Some Experimental Design Problems in Attribute Life Testing. Journal of the American Statistical Association, 58(304):1161–1162, December 1963. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2283343. See [Ehr62].

Ehrenfeld:1963:SNA

[Ehr63c] Sylvain Ehrenfeld. Some numerical aspects of the use of transforms in statistics. Journal of the American Statistical Association, 58 (304):879–893, December 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283320.

Eicker:1968:BRB

[Eic68] F. Eicker. Book review: [untitled]. Journal of the American Statistical Association, 63(322):722–723, June 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284043.

Einhorn:1962:CCD

[Ein62] Henry Adler Einhorn. Changes in concentration of domestic manufacturing establishment output: 1939–1958. Journal of the American Statistical Association, 57(300):797–803, December 1962. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281810.

Eisenhart:1962:BRB

[Eis62] Churchill Eisenhart. Book review: [untitled]. Journal of the American Statistical Association, 57(297):233–237, March 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282468.

Elashoff:1968:BRB

[Ela68] Janet D. Elashoff. Book review: [untitled]. Journal of the American Statistical Association, 63(323):1059–1060, September 1968. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283905.

Elkins:1968:CSE

[Elk68] Thomas A. Elkins. Cubical and spherical estimation of multivariate probability density. Journal of the American Statistical Association, 63(324):1495–1513, December 1968. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285898. See corrigenda [Elk69].

Elkins:1969:CCS

[Elk69] Thomas A. Elkins. Corrigenda: Cubical and Spherical Estimation of Multivariate Probability Density. Journal of the American Statistical Association, 64(328):1700, December 1969. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2286135. See [Elk68].

Ellison:1960:BRB

[Ell60] B. E. Ellison. Book review: [untitled]. Journal of the American Statistical Association, 55(290):392–393, June 1960. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2281765.

Ellison:1964:TTI

[Ell64] Bob E. Ellison. Two theorems for inferences about the normal distribution with applications in acceptance sampling. Journal of the American Statistical Association, 59(305):89–95, March 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282860.

Emerson:1962:BRB

[Eme62] John Philip Emerson. Book review: [untitled]. Journal of the American Statistical Association, 57(299):702–704, September 1962. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282416.

Endres:1969:CUA

[End69] Allen C. Endres. The computation of the unrestricted AOQL when defective material is removed but not replaced. Journal of the American Statistical Association, 64(326):665–668, June 1969. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283651.

Eisenberger:1965:SSU

[EP65] Isidore Eisenberger and Edward C. Posner. Systematic statistics used for data compression in space telemetry. Journal of the American Statistical Association, 60(309):97–133, March 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283140. See corrigenda [EP67].

Eisenberger:1967:CSS

[EP67] Isidore Eisenberger and Edward C. Posner. Corrigenda: Systematic Statistics used for Data Compression in Space Telemetry. Journal of the American Statistical Association, 62(320):1518, December 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2283812. See

[EP65].

Epstein:1963:BRB

[Eps63] Benjamin Epstein. Book review: [untitled]. Journal of the American Statistical Association, 58(303):837–838, September 1963. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282737.

Epstein:1966:BRB

[Eps66] Benjamin Epstein. Book review: [untitled]. Journal of the American Statistical Association, 61(313):268, March 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283064.

Ericson:1965:OSS

[Eri65] W. A. Ericson. Optimum stratified sampling using prior information. Journal of the American Statistical Association, 60(311): 750–771, September 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2283243.

Ericson:1967:OSD

[Eri67] W. A. Ericson. Optimal sample design with nonresponse. Journal of the American Statistical Association, 62(317):63–78, March 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2282910.

Ericson:1968:OAS

[Eri68] W. A. Ericson. Optimal allocation in stratified and multistage samples using prior information. Journal of the American Statistical Association, 63(323):964–983, September 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283887.

Estes:1964:BRB

[Est64] W. K. Estes. Book review: [untitled]. Journal of the American Statistical Association, 59(306):600–601, June 1964. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2283012.

Estrup:1966:BRB

[Est66] Hector Estrup. Book review: [untitled]. Journal of the American Statistical Association, 61(316):1241–1243, December 1966. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283221.

Evans:1961:BRB

[Eva61] Richard V. Evans. Book review: [untitled]. Journal of the American Statistical Association, 56(293):213–214, March 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282381.

Evans:1964:BRB

[Eva64] Robert Evans, Jr. Book review: [untitled]. Journal of the American Statistical Association, 59(306):614–615, June 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283022.

F:1960:BRBa

[F.60a] R. F. Book review: [untitled]. Journal of the American Statistical Association, 55(292):788, December 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281635.

F:1960:BRBb

[F.60b] R. F. Book review: [untitled]. Journal of the American Statistical Association, 55(292):788, December 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281636.

F:1960:BRBc

[F.60c] R. F. Book review: [untitled]. Journal of the American Statistical Association, 55(292):788–789, December 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281637.

F:1960:BRBd

[F.60d] R. F. Book review: [untitled]. Journal of the American Statistical Association, 55(292):789, December 1960. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281638.

F:1961:BRBa

[F.61a] R. F. Book review: [untitled]. Journal of the American Statistical Association, 56(293):215, March 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282384.

F:1961:BRBb

[F.61b] R. F. Book review: [untitled]. Journal of the American Statistical Association, 56(293):216, March 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282385.

F:1961:BRBc

[F.61c] R. F. Book review: [untitled]. Journal of the American Statistical Association, 56(293):217, March 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282388.

F:1961:BRBd

[F.61d] R. F. Book review: [untitled]. Journal of the American Statistical Association, 56(294):468, June 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282315.

F:1961:BRBe

[F.61e] R. F. Book review: [untitled]. Journal of the American Statistical Association, 56(295):775, September 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282125.

F:1961:BRBf

[F.61f] R. F. Book review: [untitled]. Journal of the American Statistical Association, 56(296):1027, December 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282032.

F:1961:BRBg

[F.61g] R. F. Book review: [untitled]. Journal of the American Statistical Association, 56(296):1027, December 1961. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282033.

F:1961:BRBh

[F.61h] R. F. Book review: [untitled]. Journal of the American Statistical Association, 56(296):1027–1028, December 1961. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2282034.

F:1961:BRBi

[F.61i] R. F. Book review: [untitled]. Journal of the American Statistical Association, 56(296):1028, December 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282035.

F:1962:BRBa

[F.62a] R. F. Book review: [untitled]. Journal of the American Statistical Association, 57(298):527, June 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281684.

F:1962:BRBb

[F.62b] R. F. Book review: [untitled]. Journal of the American Statistical Association, 57(299):722–723, September 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282431.

F:1962:BRBc

[F.62c] R. F. Book review: [untitled]. Journal of the American Statistical Association, 57(299):723, September 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282432.

F:1962:BRBd

[F.62d] R. F. Book review: [untitled]. Journal of the American Statistical Association, 57(300):946, December 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281848.

F:1962:BRBe

[F.62e] R. F. Book review: [untitled]. Journal of the American Statistical Association, 57(300):946, December 1962. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281849.

F:1962:BRBf

[F.62f] R. F. Book review: [untitled]. Journal of the American Statistical Association, 57(300):946–947, December 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281851.

F:1963:BRBa

[F.63a] R. F. Book review: [untitled]. Journal of the American Statistical Association, 58(301):268, March 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282985.

F:1963:BRBb

[F.63b] R. F. Book review: [untitled]. Journal of the American Statistical Association, 58(301):270, March 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282990.

Friend:1964:PAC

[FA64] Irwin Friend and F. Gerard Adams. The predictive ability of consumer attitudes, stock prices, and non-attitudinal variables. Journal of the American Statistical Association, 59(308):987–1005, December 1964. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2282619.

Folks:1965:OAS

[FA65] John Leroy Folks and Charles E. Antle. Optimum allocation of sampling units to strata when there are R responses of interest. Journal of the American Statistical Association, 60(309):225– 233, March 1965. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2283148.

Folks:1967:SLC

[FA67] John Leroy Folks and Charles E. Antle. Straight line confidence regions for linear models. Journal of the American Statistical Association, 62(320):1365–1374, December 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283783.

Fabens:1961:BRB

[Fab61] Augustus J. Fabens. Book review: [untitled]. Journal of the American Statistical Association, 56(295):745–746, September 1961. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282097.

Fand:1964:BRB

[Fan64] David I. Fand. Book review: [untitled]. Journal of the American Statistical Association, 59(306):609–610, June 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283018.

Farrell:1963:BRB

[Far63] Roger H. Farrell. Book review: [untitled]. Journal of the American Statistical Association, 58(304):1179–1180, December 1963. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283363.

Fama:1968:DPE

[FB68] Eugene F. Fama and Harvey Babiak. Dividend policy: An empirical analysis. Journal of the American Statistical Association, 63 (324):1132–1161, December 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2285874.

Feinleib:1960:MAL

[Fei60] Manning Feinleib. A method of analyzing log-normally distributed survival data with incomplete follow-up. Journal of the American Statistical Association, 55(291):534–545, September 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281913.

Feinstein:1961:BRB

[Fei61] Amiel Feinstein. Book review: [untitled]. Journal of the American Statistical Association, 56(294):417–418, June 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282273.

Feinstein:1962:BRB

[Fei62] Amiel Feinstein. Book review: [untitled]. Journal of the American Statistical Association, 57(298):507–508, June 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281659.

Feinleib:1969:BRB

[Fei69] Manning Feinleib. Book review: [untitled]. Journal of the American Statistical Association, 64(326):682–683, June 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283665.

Fellegi:1963:SVP

[Fel63] Ivan P. Fellegi. Sampling with varying probabilities without replacement: Rotating and non-rotating samples. Journal of the American Statistical Association, 58(301):183–201, March 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282963.

Feldman:1964:BRB

[Fel64a] Dorian Feldman. Book review: [untitled]. Journal of the American Statistical Association, 59(306):621–622, June 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283028.

Fellegi:1964:RVE

[Fel64b] I. P. Fellegi. Response variance and its estimation. Journal of the American Statistical Association, 59(308):1016–1041, December 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2282621.

Feldt:1965:MAC

[Fel65] Allan G. Feldt. The metropolitan area concept: An evaluation of the 1950 SMA’s. Journal of the American Statistical Association, 60(310):617–636, June 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2282696.

Ferguson:1962:BRB

[Fer62] Allen R. Ferguson. Book review: [untitled]. Journal of the American Statistical Association, 57(297):220–221, March 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282459.

Ferber:1965:RCS

[Fer65] Robert Ferber. The reliability of consumer surveys of financial holdings: Time deposits. Journal of the American Statistical Association, 60(309):148–163, March 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283142.

Ferber:1966:RCS

[Fer66] Robert Ferber. The reliability of consumer surveys of financial holdings: Demand deposits. Journal of the American Statistical Association, 61(313):91–103, March 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283046.

Feldman:1968:EPB

[FF68] Dorian Feldman and Martin Fox. Estimation of the parameter n in the binomial distribution. Journal of the American Statistical Association, 63(321):150–158, March 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283836.

Ferber:1969:VCF

[FFGM69] Robert Ferber, John Forsythe, Harold W. Guthrie, and E. Scott Maynes. Validation of consumer financial characteristics: Common stock. Journal of the American Statistical Association, 64 (326):415–432, June 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2283630.

Fellegi:1967:NDC

[FGP67] I. P. Fellegi, G. B. Gray, and R. Platek. The new design of the Canadian Labour Force Survey. Journal of the American Statistical Association, 62(318):421–453, June 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283973.

Freund:1967:PAD

[FH67] R. J. Freund and H. O. Hartley. A procedure for automatic data editing. Journal of the American Statistical Association, 62(318): 341–352, June 1967. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2283967.

Fiering:1962:UCA

[Fie62] Myron B. Fiering. On the use of correlation to augment data. Journal of the American Statistical Association, 57(297):20–32, March 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2282437.

Fasteau:1964:CQC

[FIM64] Herman H. Fasteau, J. Jack Ingram, and George Minton. Control of quality of coding in the 1960 censuses. Journal of the American Statistical Association, 59(305):120–132, March 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282863.

Finkner:1960:BRB

[Fin60] A. L. Finkner. Book review: [untitled]. Journal of the American Statistical Association, 55(292):784–785, December 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281629.

Firestone:1961:BRBa

[Fir61a] John M. Firestone. Book review: [untitled]. Journal of the American Statistical Association, 56(293):189–190, March 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282362.

Firestone:1961:BRBb

[Fir61b] O. J. Firestone. Book review: [untitled]. Journal of the American Statistical Association, 56(295):755–757, September 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282106.

Fisher:1961:BRB

[Fis61a] Janet A. Fisher. Book review: [untitled]. Journal of the American Statistical Association, 56(294):438, June 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282285.

Fisher:1961:NCF

[Fis61b] Walter D. Fisher. A note on curve fitting with minimum deviations by linear programming. Journal of the American Statistical Association, 56(294):359–362, June 1961. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282261. See corrigenda [Fis62].

Fisk:1961:ELS

[Fis61c] P. R. Fisk. Estimation of location and scale parameters in a truncated grouped sech square distribution. Journal of the American Statistical Association, 56(295):692–702, September 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282090.

Fisher:1962:CNC

[Fis62] Walter D. Fisher. Corrigenda: A Note on Curve Fitting with Minimum Deviations by Linear Programming. Journal of the American Statistical Association, 57(300):917, December 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281820. See [Fis61b].

Fisher:1963:CDG

[Fis63] Janet A. Fisher. Consumer durable goods expenditures, with major emphasis on the role of assets, credit and intentions. Journal of the American Statistical Association, 58(303):648–657, September 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2282712.

Fisher:1965:BRBb

[Fis65a] Gordon Fisher. Book review: [untitled]. Journal of the American Statistical Association, 60(311):928–930, September 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283262.

Fisher:1965:BRBa

[Fis65b] Walter D. Fisher. Book review: [untitled]. Journal of the American Statistical Association, 60(309):384–385, March 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283176.

Fisher:1966:RSS

[Fis66a] Franklin M. Fisher. The relative sensitivity to specification error of different k-class estimators. Journal of the American Statistical Association, 61(314):345–356, June 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282823.

Fisk:1966:BRB

[Fis66b] George Fisk. Book review: [untitled]. Journal of the American Statistical Association, 61(313):280–281, March 1966. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2283071.

Fisher:1967:ASC

[Fis67a] Franklin M. Fisher. Approximate specification and the choice of a k-class estimator. Journal of the American Statistical Association, 62(320):1265–1276, December 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283776.

Fishlow:1967:BRB

[Fis67b] Albert Fishlow. Book review: [untitled]. Journal of the American Statistical Association, 62(318):711–713, June 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284009.

Fisher:1968:BRB

[Fis68] Walter D. Fisher. Book review: [untitled]. Journal of the American Statistical Association, 63(323):1062–1063, September 1968. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283908.

FitzPatrick:1960:LBS

[Fit60] Paul J. FitzPatrick. Leading British statisticians of the Nineteenth Century. Journal of the American Statistical Association, 55(289):38–70, March 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2282178.

Fields:1962:JEP

[FKCR62] R. I. Fields, C. Y. Kramer, and C. W. Clunies-Ross. Joint estimation of the parameters of two normal populations. Journal of the American Statistical Association, 57(298):446–454, June 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281651.

Fleiss:1966:AAM

[Fle66] Joseph L. Fleiss. Assessing the accuracy of multivariate observations. Journal of the American Statistical Association, 61(314):

403–412, June 1966. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2282830.

Flehinger:1964:ICP

[FM64] Betty J. Flehinger and James Miller. Incentive contracts and price differential acceptance tests. Journal of the American Statistical Association, 59(305):149–159, March 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282866.

Fan:1962:DSP

[FMR62] C. T. Fan, Mervin E. Muller, and Ivan Rezucha. Development of sampling plans by using sequential (item by item) selection techniques and digital computers. Journal of the American Statistical Association, 57(298):387–402, June 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281647.

Fobes:1961:BRB

[Fob61] Melcher P. Fobes. Book review: [untitled]. Journal of the American Statistical Association, 56(294):467–468, June 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282313.

Foley:1961:BRB

[Fol61] Donald L. Foley. Book review: [untitled]. Journal of the American Statistical Association, 56(294):463–465, June 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282308.

Folks:1969:BRB

[Fol69] J. Leroy Folks. Book review: [untitled]. Journal of the American Statistical Association, 64(327):1091–1092, September 1969. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283493.

Forsythe:1960:BRB

[For60] George E. Forsythe. Book review: [untitled]. Journal of the American Statistical Association, 55(290):391–392, June 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281764.

Forsythe:1962:BRB

[For62] George E. Forsythe. Book review: [untitled]. Journal of the American Statistical Association, 57(300):924–926, December 1962. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281829.

Fox:1961:BRB

[Fox61] Martin Fox. Book review: [untitled]. Journal of the American Statistical Association, 56(293):183–184, March 1961. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2282357.

Fox:1966:BRB

[Fox66] Martin Fox. Book review: [untitled]. Journal of the American Statistical Association, 61(315):881–882, September 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282810.

Federer:1965:GTP

[FR65] W. T. Federer and B. L. Raktoe. General theory of prime-power lattice designs: Lattice rectangles for v = s^{m} treatments in s^{r} rows and s^{c} columns for r + c = m, r ≠ c, and v < 1000. Journal of the American Statistical Association, 60(311):891–904, September 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2283255.

Federer:1966:GLS

[FR66] W. T. Federer and B. L. Raktoe. Generalized lattice square designs. Journal of the American Statistical Association, 61(315): 821–832, September 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2282791.

Friedman:1967:SIC

[FR67] H. P. Friedman and J. Rubin. On some invariant criteria for grouping data. Journal of the American Statistical Association, 62(320): 1159–1178, December 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2283767.

Fama:1968:SPS

[FR68] Eugene F. Fama and Richard Roll. Some properties of symmetric stable distributions. Journal of the American Statistical Association, 63(323):817–836, September 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283875.

Frankel:1960:BRB

[Fra60] Marvin Frankel. Book review: [untitled]. Journal of the American Statistical Association, 55(289):234–236, March 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282199.

Frankel:1963:BRB

[Fra63a] Marvin Frankel. Book review: [untitled]. Journal of the American Statistical Association, 58(301):258–259, March 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282977.

Fraser:1963:SLP

[Fra63b] D. A. S. Fraser. On the sufficiency and likelihood principles. Journal of the American Statistical Association, 58(303):641–647, September 1963. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2282711.

Frank:1966:BRB

[Fra66] Ronald Frank. Book review: [untitled]. Journal of the American Statistical Association, 61(313):285, March 1966. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2283075.

Frankel:1969:CWH

[Fra69] Lester R. Frankel. Comments: On William Hurwitz. Journal of the American Statistical Association, 64(328):1152–1153, December 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2286058. See

[Dem69].

Freund:1961:CRA

[FRCR61] Rudolf J. Freund, W. Richard, and C. W. Clunies-Ross. Corrigenda: Residual Analysis. Journal of the American Statistical Association, 56(296):1005, December 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282013. See [FVCR61].

Freimer:1961:BRB

[Fre61a] Marshall Freimer. Book review: [untitled]. Journal of the American Statistical Association, 56(294):411–412, June 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282268.

Freund:1961:BEE

[Fre61b] John E. Freund. A bivariate extension of the exponential distribution. Journal of the American Statistical Association, 56(296): 971–977, December 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2282007.

Freimer:1964:BRB

[Fre64] Marshall Freimer. Book review: [untitled]. Journal of the American Statistical Association, 59(305):293, March 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282884.

Freimer:1965:BRB

[Fre65] Marshall Freimer. Book review: [untitled]. Journal of the American Statistical Association, 60(309):383–384, March 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283175.

Freiberger:1969:BRB

[Fre69] Walter F. Freiberger. Book review: [untitled]. Journal of the American Statistical Association, 64(328):1668–1669, December 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2286100.

Friedman:1962:ITS

[Fri62a] Milton Friedman. The interpolation of time series by related series. Journal of the American Statistical Association, 57(300):729– 757, December 1962. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2281805.

Friend:1962:BRB

[Fri62b] Irwin Friend. Book review: [untitled]. Journal of the American Statistical Association, 57(300):933–935, December 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281835.

Friedman:1965:BRB

[Fri65] James W. Friedman. Book review: [untitled]. Journal of the American Statistical Association, 60(309):369–370, March 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283165.

Friedlaender:1966:BRB

[Fri66] Ann F. Friedlaender. Book review: [untitled]. Journal of the American Statistical Association, 61(316):1238–1239, December 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283219.

Friedman:1968:BRB

[Fri68] Herman P. Friedman. Book review: [untitled]. Journal of the American Statistical Association, 63(321):375, March 1968. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283862.

Fromm:1963:BRB

[Fro63] Gary Fromm. Book review: [untitled]. Journal of the American Statistical Association, 58(304):1168–1171, December 1963. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283355.

Fellingham:1964:AED

[FS64] S. A. Fellingham and D. J. Stoker. An approximation for the exact distribution of the Wilcoxon test for symmetry. Journal of the American Statistical Association, 59(307):899–905, September 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2283109.

Fellegi:1969:TRL

[FS69] Ivan P. Fellegi and Alan B. Sunter. A theory for record linkage. Journal of the American Statistical Association, 64(328): 1183–1210, December 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286061.

Fuller:1966:EEP

[Ful66] Wayne A. Fuller. Estimation employing post strata. Journal of the American Statistical Association, 61(316):1172–1183, December 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2283206.

Freund:1961:RA

[FVCR61] Rudolf J. Freund, Richard W. Vail, and C. W. Clunies-Ross. Residual analysis. Journal of the American Statistical Association, 56 (293):98–104, March 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/ stable/2282334. See corrigenda [FRCR61].

Folks:1961:NMP

[FW61] John Leroy Folks and Del Lon West. Note on the missing plot procedure in a randomized block design. Journal of the American Statistical Association, 56(296):933–941, December 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282005.

Freeman:1964:SPW

[FW64] David Freeman and Lionel Weiss. Sampling plans which approximately minimize the maximum expected sample size. Journal of the American Statistical Association, 59(305):67–88, March 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282859.

Gabriel:1963:AVP

[Gab63] K. R. Gabriel. Analysis of variance of proportions with unequal frequencies. Journal of the American Statistical Association, 58 (304):1133–1157, December 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283339.

Gabriel:1966:STP

[Gab66] K. R. Gabriel. Simultaneous test procedures for multiple comparisons on categorical data. Journal of the American Statistical Association, 61(316):1081–1096, December 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283202.

Greenberg:1969:UQR

[GAESH69] Bernard G. Greenberg, Abdel-Latif A. Abul-Ela, Walt R. Simmons, and Daniel G. Horvitz. The unrelated question randomized response model: Theoretical framework. Journal of the American Statistical Association, 64(326):520–539, June 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283636.

Gafarian:1964:CBS

[Gaf64] A. V. Gafarian. Confidence bands in straight line regression. Journal of the American Statistical Association, 59(305):182– 213, March 1964. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2282869.

Gainsbrugh:1962:SWL

[Gai62] Martin R. Gainsbrugh. Statistics we live by. Journal of the American Statistical Association, 57(297):1–9, March 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282435.

Gallant:1968:NMC

[Gal68] A. R. Gallant. A note on the measurement of cost/quantity relationships in the aircraft industry. Journal of the American Statistical Association, 63(324):1247–1252, December 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2285881.

Gart:1968:BRB

[Gar68] John J. Gart. Book review: [untitled]. Journal of the American Statistical Association, 63(324):1540, December 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2285910.

Gastwirth:1965:PMT

[Gas65] Joseph L. Gastwirth. Percentile modifications of two sample rank tests. Journal of the American Statistical Association, 60(312): 1127–1141, December 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2283411.

Gastwirth:1966:RP

[Gas66] Joseph L. Gastwirth. On robust procedures. Journal of the American Statistical Association, 61(316):929–948, December 1966. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283190.

Gastwirth:1968:FMT

[Gas68] J. L. Gastwirth. The first-median test: a two-sided version of the control median test. Journal of the American Statistical Association, 63(322):692–706, June 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284039.

Gastwirth:1969:BRB

[Gas69] Joseph Gastwirth. Book review: [untitled]. Journal of the American Statistical Association, 64(327):1090–1091, September 1969. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283492.

Graybill:1967:LSC

[GB67] Franklin A. Graybill and David C. Bowden. Linear segment confidence bands for simple linear models. Journal of the American Statistical Association, 62(318):403–408, June 1967. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2283971.

Gold:1967:QWD

[GBB67] Ruth Z. Gold, Simeon M. Berman, and Agnes Berger. On the question of whether a disease is familial. Journal of the American Statistical Association, 62(318):409–420, June 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283972.

Graybill:1963:SSR

[GC63] Franklin A. Graybill and Terrence L. Connell. Sample size required to estimate the ratio of variances with bounded relative error. Journal of the American Statistical Association, 58(304): 1044–1047, December 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2283330.

Graybill:1964:SSR

[GC64] Franklin A. Graybill and Terrence L. Connell. Sample size required to estimate the parameter in the uniform density within d units of the true value. Journal of the American Statistical Association, 59(306):550–556, June 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283007.

Gebhardt:1969:SNC

[Geb69] Friedrich Gebhardt. Some numerical comparisons of several approximations to the binomial distribution. Journal of the American Statistical Association, 64(328):1638–1646, December 1969. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2286095.

Geisler:1963:BRB

[Gei63a] M. A. Geisler. Book review: [untitled]. Journal of the American Statistical Association, 58(304):1197–1198, December 1963. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283379.

Geisser:1963:BRB

[Gei63b] Seymour Geisser. Book review: [untitled]. Journal of the American Statistical Association, 58(304):1175–1176, December 1963. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283359.

Geisser:1963:MAV

[Gei63c] Seymour Geisser. Multivariate analysis of variance for a special covariance case. Journal of the American Statistical Association, 58(303):660–669, September 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor. org/stable/2282714. See corrigenda [Gei64].

Geisser:1964:CMA

[Gei64] Seymour Geisser. Corrigenda: Multivariate Analysis of Variance for a Special Covariance Case. Journal of the American Statistical Association, 59(308):1296, December 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282645. See [Gei63c].

Geisser:1965:BAC

[Gei65] Seymour Geisser. A Bayes approach for combining correlated estimates. Journal of the American Statistical Association, 60(310): 602–607, June 1965. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2282694.

Genizi:1965:PTS

[Gen65] Abraham Genizi. On the performance of the truncated sequential probability ratio test. Journal of the American Statistical Association, 60(312):979–984, December 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283399. See [Gen66].

Genizi:1966:CPT

[Gen66] Abraham Genizi. Corrigenda: On the Performance of the Truncated Sequential Probability Ratio Test. Journal of the American Statistical Association, 61(316):1247, December 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283228. See [Gen65].

Georgiadis:1966:BRB

[Geo66] Hourmouzis George Georgiadis. Book review: [untitled]. Journal of the American Statistical Association, 61(313):276–280, March 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2283069.

Gerschenkron:1961:BRB

[Ger61] Alexander Gerschenkron. Book review: [untitled]. Journal of the American Statistical Association, 56(296):1006–1008, December 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2282015.

Gerig:1969:MEF

[Ger69] Thomas M. Gerig. A multivariate extension of Friedman’s χ_{r}^{2}- test. Journal of the American Statistical Association, 64(328): 1595–1608, December 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286091.

Golub:1956:CAS

[GG56] Abraham Golub and Frank E. Grubbs. Corrigenda: Analysis of Sensitivity Experiments when the Levels of Stimulas Cannot be Controlled. Journal of the American Statistical Association, 51 (276):650–651, December 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281482. See [GG67].

Goodman:1961:SNT

[GG61a] Leo A. Goodman and Yehuda Grunfeld. Some nonparametric tests for comovements between time series. Journal of the American Statistical Association, 56(293):11–26, March 1961. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2282326.

Gupta:1961:GDA

[GG61b] Shanti S. Gupta and Phyllis A. Groll. Gamma distribution in acceptance sampling based on life tests. Journal of the American Statistical Association, 56(296):942–970, December 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282006.

Gumbel:1964:AEB

[GG64] E. J. Gumbel and Neil Goldstein. Analysis of empirical bivariate extremal distributions. Journal of the American Statistical Association, 59(307):794–816, September 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283099.

Golub:1967:CAS

[GG67] Abraham Golub and Frank E. Grubbs. Corrigenda: Analysis of Sensitivity Experiments when the Levels of Stimulus Cannot be Controlled. Journal of the American Statistical Association, 62 (320):1519, December 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2283814. See [GG56].

Goodman:1958:PUR

[GH58] Leo A. Goodman and H. O. Hartley. The precision of unbiased ratio-type estimators. Journal of the American Statistical Association, 53(282):491–508, June 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor. org/stable/2281870. See corrigenda [GH63, GH69].

Goodman:1963:CPU

[GH63] Leo A. Goodman and H. O. Hartley. Corrigenda: The Precision of Unbiased Ratio-Type Estimators. Journal of the American Statistical Association, 58(304):1162, December 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283344. See [GH58].

Goodman:1969:CPU

[GH69] Leo Goodman and H. O. Hartley. Corrigenda: The Precision of Unbiased Ratio-Type Estimators. Journal of the American Statistical Association, 64(328):1700, December 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286136. See [GH58].

Ghosh:1965:SRT

[Gho65] B. K. Ghosh. Sequential range tests for components of variance. Journal of the American Statistical Association, 60(311):826–836, September 1965. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2283248.

Ghosh:1967:SAV

[Gho67] B. K. Ghosh. Sequential analysis of variance under random and mixed models. Journal of the American Statistical Association, 62 (320):1401–1417, December 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283786.

Ghosh:1969:MDS

[Gho69] B. K. Ghosh. Moments of the distribution of sample size in a SPRT. Journal of the American Statistical Association, 64(328): 1560–1575, December 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286089.

Gibbons:1964:ENN

[Gib64] Jean D. Gibbons. Effect on non-normality on the power function of the sign test. Journal of the American Statistical Association, 59(305):142–148, March 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2282865.

Gilliland:1962:IBN

[Gil62] Dennis C. Gilliland. Integral of the bivariate normal distribution over an offset circle. Journal of the American Statistical Association, 57(300):758–768, December 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281806.

Gilbert:1968:DUQ

[Gil68] Ethel S. Gilbert. On discrimination using qualitative variables. Journal of the American Statistical Association, 63(324):1399– 1412, December 1968. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2285893.

Gimbutas:1962:BRB

[Gim62] Marija Gimbutas. Book review: [untitled]. Journal of the American Statistical Association, 57(299):716–717, September 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282423.

Giri:1965:TIA

[Gir65] N. Giri. On the F -test in the intrablock analysis of a class of two associate PBIB designs. Journal of the American Statistical Association, 60(309):285–293, March 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283153.

Gittelsohn:1969:BRB

[Git69] Alan M. Gittelsohn. Book review: [untitled]. Journal of the American Statistical Association, 64(328):1678–1679, December 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2286111.

Goldberger:1961:NSL

[GJ61] Arthur S. Goldberger and D. B. Jochems. Note on stepwise least squares. Journal of the American Statistical Association, 56(293): 105–110, March 1961. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2282335.

Goel:1968:ADE

[GJW68] A. L. Goel, S. C. Jain, and S. M. Wu. An algorithm for the determination of the economic design of X-charts based on Duncan’s model. Journal of the American Statistical Association, 63 (321):304–320, March 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2283850.

Goodman:1963:MAC

[GK63] Leo A. Goodman and William H. Kruskal. Measures of association for cross classifications III: Approximate sampling theory. Journal of the American Statistical Association, 58(302):310–364, June 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2283271.

Glaser:1960:BRB

[Gla60] Daniel Glaser. Book review: [untitled]. Journal of the American Statistical Association, 55(291):589–590, September 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281922.

Glasser:1962:DFT

[Gla62a] Gerald J. Glasser. A distribution-free test of independence with a sample of paired observations. Journal of the American Statistical Association, 57(297):116–133, March 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282443.

Glasser:1962:VFM

[Gla62b] Gerald J. Glasser. Variance formulas for the mean difference and coefficient of concentration. Journal of the American Statistical Association, 57(299):648–654, September 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282402.

Glasser:1964:LRA

[Gla64] M. Glasser. Linear regression analysis with missing observations among the independent variables. Journal of the American Statistical Association, 59(307):834–844, September 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283101.

Glaser:1966:BRB

[Gla66] R. George Glaser. Book review: [untitled]. Journal of the American Statistical Association, 61(316):1235–1236, December 1966. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283216.

Glasser:1967:ESC

[Gla67] Marvin Glasser. Exponential survival with covariance. Journal of the American Statistical Association, 62(318):561–568, June 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283983.

Gleser:1966:BRB

[Gle66] Leon Jay Gleser. Book review: Intermediate Correlational Methods, by Andrew R. Baggaley. Journal of the American Statistical Association, 61(314):527–529, June 1966. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282844.

Gleser:1968:BRB

[Gle68] Leon Jay Gleser. Book review: [untitled]. Journal of the American Statistical Association, 63(324):1536–1538, December 1968. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2285908.

Glejser:1969:NTH

[Gle69] H. Glejser. A new test for heteroskedasticity. Journal of the American Statistical Association, 64(325):316–323, March 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283741.

Gilbert:1966:RMS

[GM66] John P. Gilbert and Frederick Mosteller. Recognizing the maximum of a sequence. Journal of the American Statistical Association, 61(313):35–73, March 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283044.

Gumbel:1967:SAP

[GM67] E. J. Gumbel and C. K. Mustafi. Some analytical properties of bivariate extremal distributions. Journal of the American Statistical Association, 62(318):569–588, June 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283984.

Godwin:1963:BRB

[God63] H. J. Godwin. Book review: [untitled]. Journal of the American Statistical Association, 58(303):838–839, September 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282738.

Godwin:1964:BRB

[God64] H. J. Godwin. Book review: Corrigenda: [Book Review]. Journal of the American Statistical Association, 59(308):1297, December 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2282646.

Golden:1960:BRB

[Gol60] Philip Golden. Book review: [untitled]. Journal of the American Statistical Association, 55(292):770, December 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281616.

Goldberger:1961:SLS

[Gol61a] Arthur S. Goldberger. Stepwise least squares: Residual analysis and specification error. Journal of the American Statistical Association, 56(296):998–1000, December 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282011.

Goldsmith:1961:BRB

[Gol61b] Harold F. Goldsmith. Book review: [untitled]. Journal of the American Statistical Association, 56(294):441–442, June 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282288.

Gold:1962:BRB

[Gol62a] Ruth Z. Gold. Book review: [untitled]. Journal of the American Statistical Association, 57(298):517, June 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281667.

Goldberger:1962:BLU

[Gol62b] Arthur S. Goldberger. Best linear unbiased prediction in the generalized linear regression model. Journal of the American Statistical Association, 57(298):369–375, June 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281645.

Golub:1963:RBG

[Gol63] Gene H. Golub. Review: A Guide to ALGOL Programming, by Daniel D. McCracken. Journal of the American Statistical Association, 58(304):1202, December 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283384.

Goldstein:1964:ERM

[Gol64] Sidney Goldstein. The extent of repeated migration: An analysis based on the Danish population register. Journal of the American Statistical Association, 59(308):1121–1132, December 1964. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282627.

Goldberg:1969:BRB

[Gol69] Samuel Goldberg. Book review: [untitled]. Journal of the American Statistical Association, 64(325):399–400, March 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283755.

Gomory:1961:BRB

[Gom61] Ralph E. Gomory. Book review: [untitled]. Journal of the American Statistical Association, 56(295):761–762, September 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282112.

Goodman:1960:EVP

[Goo60] Leo A. Goodman. On the exact variance of products. Journal of the American Statistical Association, 55(292):708–713, December 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2281592. See

corrigenda [Goo62a].

Goodman:1961:SMM

[Goo61] Leo A. Goodman. Statistical methods for the Mover–Stayer model. Journal of the American Statistical Association, 56(296):841–868, December 1961. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2281999.

Goodman:1962:CEV

[Goo62a] Leo A. Goodman. Corrigenda: On the Exact Variance of Products. Journal of the American Statistical Association, 57(300): 917, December 1962. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/

2281821. See [Goo60].

Goodman:1962:VPR

[Goo62b] Leo A. Goodman. The variance of the product of K random variables. Journal of the American Statistical Association, 57(297): 54–60, March 1962. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2282440.

Goodman:1964:SMA

[Goo64] Leo A. Goodman. Simple methods for analyzing three-factor interaction in contingency tables. Journal of the American Statistical Association, 59(306):319–352, June 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282992.

Goodman:1968:ACC

[Goo68] Leo A. Goodman. The analysis of cross-classified data: Independence, quasi-independence, and interactions in contingency tables with or without missing entries. Journal of the American Statistical Association, 63(324):1091–1131, December 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2285873.

Good:1969:SEB

[Goo69] I. J. Good. A subjective evaluation of Bode’s law and an ’objective’ test for approximate numerical rationality. Journal of the American Statistical Association, 64(325):23–49, March 1969. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2283718. See discussion [ABH^{+}69].

Gort:1961:BRB

[Gor61] Michael Gort. Book review: [untitled]. Journal of the American Statistical Association, 56(293):198–201, March 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282369.

Goreux:1963:BRB

[Gor63] L. M. Goreux. Book review: [untitled]. Journal of the American Statistical Association, 58(303):868, September 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282756.

Gottlieb:1968:BRBa

[Got68a] Manuel Gottlieb. Book review: [untitled]. Journal of the American Statistical Association, 63(321):381, March 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283867.

Gottlieb:1968:BRBb

[Got68b] Manuel Gottlieb. Book review: [untitled]. Journal of the American Statistical Association, 63(322):736, June 1968. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284053.

Govindarajulu:1962:RDN

[Gov62] Zakkula Govindarajulu. The reciprocal of the decapitated negative binomial variable. Journal of the American Statistical Association, 57(300):906–913, December 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor. org/stable/2281818. See corrigenda [Gov63a].

Govindarajulu:1963:CRD

[Gov63a] Zakkula Govindarajulu. Corrigenda: The Reciprocal of the Decapitated Negative Binomial Variable. Journal of the American Statistical Association, 58(304):1162, December 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283345. See [Gov62].

Govindarajulu:1963:RRI

[Gov63b] Zakkula Govindarajulu. Recurrence relations for the inverse moments of the positive binomial variable. Journal of the American Statistical Association, 58(302):468–473, June 1963. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2283280.

Govindarajulu:1964:SMB

[Gov64] Zakkula Govindarajulu. A supplement to Mendenhall’s bibliography on life testing and related topics. Journal of the American Statistical Association, 59(308):1231–1291, December 1964. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282641. See [Gov65].

Govindarajulu:1965:CSM

[Gov65] Zakkula Govindarajulu. Corrigenda: A Supplement to Mendenhall’s Bibliography on Life Testing and Related Topics. Journal of the American Statistical Association, 60(312):1249–1250, December 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2283449. See

[Gov64].

Govindarajulu:1966:BLE

[Gov66a] Zakkula Govindarajulu. Best linear estimates under symmetric censoring of the parameters of double exponential population. Journal of the American Statistical Association, 61(313):

248–258, March 1966. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2283060.

Govindarajulu:1966:BRB

[Gov66b] Zakkula Govindarajulu. Book review: [untitled]. Journal of the American Statistical Association, 61(316):1237–1238, December 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2283218.

Gelzer:1965:ARE

[GP65] Joseph Gelzer and Ronald Pyke. The asymptotic relative efficiency of goodness-of-fit tests against scalar alternatives. Journal of the American Statistical Association, 60(310):410–419, June 1965. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282679.

Gray:1968:SMR

[GP68] G. B. Gray and R. Platek. Several methods of re-designing area samples utilizing probabilities proportional to size when the sizes change significantly. Journal of the American Statistical Association, 63(324):1280–1297, December 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285884.

Goldfeld:1965:STH

[GQ65] Stephen M. Goldfeld and Richard E. Quandt. Some tests for homoscedasticity. Journal of the American Statistical Association, 60 (310):539–547, June 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/ stable/2282689. See corrigenda [GQ67].

Goldfeld:1967:CST

[GQ67] Stephen M. Goldfeld and Richard E. Quandt. Corrigenda: Some Tests for Homoscedasticity. Journal of the American Statistical Association, 62(320):1518–1519, December 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283813. See [GQ65].

Graybill:1960:BRB

[Gra60] Franklin A. Graybill. Book review: [untitled]. Journal of the American Statistical Association, 55(291):604–605, September 1960. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281932.

Granger:1963:QTS

[Gra63] C. W. J. Granger. A quick test for serial correlation suitable for use with non-stationary time series. Journal of the American Statistical Association, 58(303):728–736, September 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282721.

Grant:1966:BRB

[Gra66] Eugene L. Grant. Book review: [untitled]. Journal of the American Statistical Association, 61(313):264–265, March 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283062.

Gregory:1962:BRB

[Gre62] G. Gregory. Book review: [untitled]. Journal of the American Statistical Association, 57(297):237–238, March 1962. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2282469.

Greenberg:1963:BRB

[Gre63a] Edward Greenberg. Book review: [untitled]. Journal of the American Statistical Association, 58(304):1191–1192, December 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283374.

Greene:1963:BRB

[Gre63b] Mark R. Greene. Book review: [untitled]. Journal of the American Statistical Association, 58(302):584–585, June 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283307.

Greenberg:1965:ADI

[Gre65a] Edward Greenberg. Appropriations data and the investment decision. Journal of the American Statistical Association, 60(310): 503–515, June 1965. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2282684.

Greenberger:1965:BRB

[Gre65b] Martin Greenberger. Book review: [untitled]. Journal of the American Statistical Association, 60(309):363–364, March 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283161.

Greenberg:1969:PSI

[Gre69] B. G. Greenberg. Problems of statistical inference in health with special reference to the cigarette smoking and lung cancer controversy. Journal of the American Statistical Association, 64(327): 739–758, September 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2283458.

Griffin:1961:BRB

[Gri61] John I. Griffin. Book review: [untitled]. Journal of the American Statistical Association, 56(295):749–750, September 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282101.

Grizzle:1962:APT

[Gri62] James E. Grizzle. Asymptotic power of tests of linear hypotheses using the probit and logit transformations. Journal of the American Statistical Association, 57(300):877–894, December 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281816. See corrigenda [Gri64].

Grizzle:1964:CAP

[Gri64] James E. Grizzle. Corrigenda: Asymptotic Power of Tests of Linear Hypothesis Using the Probit and Logit Transformations. Journal of the American Statistical Association, 59(308):1297, December 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2282647. See

[Gri62].

Grigg:1966:BRB

[Gri66] Charles M. Grigg. Book review: [untitled]. Journal of the American Statistical Association, 61(313):281–282, March 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283072.

Gringorten:1968:EFT

[Gri68] Irving I. Gringorten. Estimating finite-time maxima and minima of a stationary Gaussian Ornstein–Uhlenbeck process by Monte Carlo simulation. Journal of the American Statistical Association, 63 (324):1517–1521, December 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2285900.

Grove:1960:BRB

[Gro60] Ernest W. Grove. Book review: [untitled]. Journal of the American Statistical Association, 55(291):587–589, September 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281921.

Gallaway:1961:QEM

[GS61] Lowell E. Gallaway and Paul E. Smith. A quarterly econometric model of the United States. Journal of the American Statistical Association, 56(294):379–383, June 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282264.

Gaylor:1965:DOP

[GS65] D. W. Gaylor and H. C. Sweeny. Design for optimal prediction in simple linear regression. Journal of the American Statistical Association, 60(309):205–216, March 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283146.

Gray:1968:EDF

[GS68] H. L. Gray and W. R. Schucany. On the evaluation of distribution functions. Journal of the American Statistical Association, 63 (322):715–720, June 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284041.

Guenther:1965:SGU

[GT65] William C. Guenther and P. O. Thomas. Some graphs useful for statistical inference. Journal of the American Statistical Association, 60(309):334–343, March 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283157.

Guenther:1964:ADN

[Gue64] William C. Guenther. Another derivation of the non-central chisquare distribution. Journal of the American Statistical Association, 59(307):957–960, September 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283113.

Guenther:1965:BRB

[Gue65] William C. Guenther. Book review: [untitled]. Journal of the American Statistical Association, 60(311):931–932, September 1965. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283264.

Gue:1967:BRB

[Gue67] Ronald L. Gue. Book review: [untitled]. Journal of the American Statistical Association, 62(319):1079–1080, September 1967. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283701.

Guenther:1968:BRB

[Gue68] William C. Guenther. Book review: [untitled]. Journal of the American Statistical Association, 63(321):378–379, March 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283864.

Gumbel:1960:BED

[Gum60] E. J. Gumbel. Bivariate exponential distributions. Journal of the American Statistical Association, 55(292):698–707, December 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281591.

Gumbel:1961:BLD

[Gum61] E. J. Gumbel. Bivariate logistic distributions. Journal of the American Statistical Association, 56(294):335–349, June 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282259.

Gustafson:1960:BRB

[Gus60] Robert L. Gustafson. Book review: [untitled]. Journal of the American Statistical Association, 55(291):599–601, September 1960. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281928.

Gustafson:1961:PCR

[Gus61] Robert L. Gustafson. Partial correlations in regression computations. Journal of the American Statistical Association, 56(294): 363–367, June 1961. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2282262.

Guthrie:1960:CPH

[Gut60] Harold W. Guthrie. Consumer’s propensities to hold liquid assets. Journal of the American Statistical Association, 55(291): 469–490, September 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2281909.

Greenberg:1965:DCB

[GW65] Richard A. Greenberg and Colin White. The detection of a correlation between the sexes of adjacent sibs in human families. Journal of the American Statistical Association, 60(312):1035–1045, December 1965. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2283403.

Guenther:1966:CRT

[GW66] William C. Guenther and Mary G. Whitcomb. Critical regions for tests of interval hypotheses about the variance. Journal of the American Statistical Association, 61(313):204–219, March 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283055.

H:1960:BRB

[H.60] C. H. Book review: [untitled]. Journal of the American Statistical Association, 55(291):611, September 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281938.

H:1961:BRBa

[H.61a] C. H. Book review: [untitled]. Journal of the American Statistical Association, 56(294):469, June 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282317.

H:1961:BRBb

[H.61b] C. H. Book review: [untitled]. Journal of the American Statistical Association, 56(295):777, September 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282131.

H:1961:BRBc

[H.61c] C. H. Book review: [untitled]. Journal of the American Statistical Association, 56(296):1028, December 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282036.

H:1962:BRBa

[H.62a] C. H. Book review: [untitled]. Journal of the American Statistical Association, 57(298):527, June 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281686.

H:1962:BRBb

[H.62b] C. H. Book review: [untitled]. Journal of the American Statistical Association, 57(298):527, June 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281687.

H:1962:BRBc

[H.62c] C. H. Book review: [untitled]. Journal of the American Statistical Association, 57(300):945, December 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281846.

H:1963:BRBa

[H.63a] C. H. Book review: [untitled]. Journal of the American Statistical Association, 58(304):1203, December 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283387.

H:1963:BRBb

[H.63b] C. H. Book review: [untitled]. Journal of the American Statistical Association, 58(304):1204, December 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283391.

H:1964:BRBa

[H.64a] C. H. Book review: [untitled]. Journal of the American Statistical Association, 59(305):315, March 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282899.

H:1964:BRBb

[H.64b] C. H. Book review: [untitled]. Journal of the American Statistical Association, 59(306):628, June 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283035.

H:1964:BRBc

[H.64c] C. H. Book review: [untitled]. Journal of the American Statistical Association, 59(307):979–980, September 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283127.

H:1964:BRBd

[H.64d] C. H. Book review: [untitled]. Journal of the American Statistical Association, 59(307):981, September 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283130.

H:1964:BRBe

[H.64e] C. H. Book review: [untitled]. Journal of the American Statistical Association, 59(308):1324–1325, December 1964. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2282672.

H:1964:BRBf

[H.64f] C. H. Book review: [untitled]. Journal of the American Statistical Association, 59(308):1326, December 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282674.

Hadley:1963:BRB

[Had63] George Hadley. Book review: [untitled]. Journal of the American Statistical Association, 58(303):843–845, September 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282742.

Hader:1966:BRB

[Had66] R. J. Hader. Book review: [untitled]. Journal of the American Statistical Association, 61(316):1230–1231, December 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283212.

Haenszel:1962:BRB

[Hae62] William Haenszel. Book review: [untitled]. Journal of the American Statistical Association, 57(299):698–699, September 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282411.

Hahn:1969:FCT

[Hah69] Gerald J. Hahn. Factors for calculating two-sided prediction intervals for samples from a normal distribution. Journal of the American Statistical Association, 64(327):878–888, September 1969. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283469.

Haitovsky:1966:UMR

[Hai66] Yoel Haitovsky. Unbiased multiple regression coefficients estimated from one-way-classification tables when the cross classifications are unknown. Journal of the American Statistical Association, 61(315): 720–728, September 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2282782.

Halperin:1960:CEW

[Hal60a] Max Halperin. Corrigenda: Extension of the Wilcoxon–Mann– Whitney Test to Samples Censored at the Same Fixed Point. Journal of the American Statistical Association, 55(292):755, December 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2281601. See

[Hal60b].

Halperin:1960:EWM

[Hal60b] Max Halperin. Extension of the Wilcoxon–Mann–Whitney test to samples censored at the same fixed point. Journal of the American Statistical Association, 55(289):125–138, March 1960. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282183. See corrigenda

[Hal60a].

Halperin:1961:ALO

[Hal61a] Max Halperin. Almost linearly-optimum combination of unbiased estimates. Journal of the American Statistical Association, 56(293): 36–43, March 1961. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2282328.

Halperin:1961:FSL

[Hal61b] Max Halperin. Fitting of straight lines and prediction when both variables are subject to error. Journal of the American Statistical Association, 56(295):657–669, September 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282086.

Halperin:1964:IEL

[Hal64a] Max Halperin. Interval estimation in linear regression when both variables are subject to error. Journal of the American Statistical Association, 59(308):1112–1120, December 1964. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2282626.

Halperin:1964:IEN

[Hal64b] Max Halperin. Interval estimation of non-linear parametric functions, II. Journal of the American Statistical Association, 59(305): 168–181, March 1964. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2282868.

Halperin:1965:IEN

[Hal65] Max Halperin. Interval estimation of non-linear parametric functions, III. Journal of the American Statistical Association, 60(312): 1191–1199, December 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2283415.

Halperin:1967:IBS

[Hal67] Max Halperin. An inequality on a bivariate Student’s “t” distribution. Journal of the American Statistical Association, 62(318): 603–606, June 1967. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2283987.

Hamdan:1963:NWC

[Ham63] M. A. Hamdan. The number and width of classes in the chisquare test. Journal of the American Statistical Association, 58 (303):678–689, September 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2282716.

Hamburger:1965:BRB

[Ham65] Michael J. Hamburger. Book review: [untitled]. Journal of the American Statistical Association, 60(310):682, June 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282706.

Hamblen:1968:BRB

[Ham68a] John W. Hamblen. Book review: [untitled]. Journal of the American Statistical Association, 63(321):375, March 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283861.

Hamdan:1968:OCC

[Ham68b] M. A. Hamdan. Optimum choice of classes for contingency tables. Journal of the American Statistical Association, 63(321):291– 297, March 1968. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2283848.

Hansen:1961:CAS

[Han61] Morris H. Hansen. Cooperation among statistical and other societies: Presidential address. Journal of the American Statistical Association, 56(293):1–10, March 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282325.

Hannan:1963:CES

[Han63a] E. J. Hannan. Corrigenda: The Estimation of Seasonal Variation in Economic Time Series. Journal of the American Statistical Association, 58(304):1162, December 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283346. See [Han63b].

Hannan:1963:ESV

[Han63b] E. J. Hannan. The estimation of seasonal variation in economic time series. Journal of the American Statistical Association, 58 (301):31–44, March 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/ stable/2282953. See corrigenda [Han63a].

Hansen:1963:BRB

[Han63c] W. Lee Hansen. Book review: [untitled]. Journal of the American Statistical Association, 58(303):846–847, September 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282744.

Hannan:1964:ECS

[Han64] E. J. Hannan. The estimation of a changing seasonal pattern. Journal of the American Statistical Association, 59(308):1063–1077, December 1964. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/

2282624. See [Han66].

Hannan:1966:CEC

[Han66] E. J. Hannan. Corrigenda: The Estimation of a Changing Seasonal Pattern. Journal of the American Statistical Association, 61(316):1247, December 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283229. See [Han64].

Hanson:1969:MWH

[Han69] Morris H. Hanson. A memorial for William N. Hurwitz. Journal of the American Statistical Association, 64(328):1122–1128, December 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2286053.

Harter:1960:CEP

[Har60a] H. Leon Harter. Circular error probabilities. Journal of the American Statistical Association, 55(292):723–731, December 1960. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281595.

Hartle:1960:BRB

[Har60b] Douglas G. Hartle. Book review: [untitled]. Journal of the American Statistical Association, 55(291):605–607, September 1960. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281933.

Hartley:1960:BRBa

[Har60c] H. O. Hartley. Book review: [untitled]. Journal of the American Statistical Association, 55(292):758–760, December 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281607.

Hartley:1960:BRBb

[Har60d] H. O. Hartley. Book review: [untitled]. Journal of the American Statistical Association, 55(292):779–780, December 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281623.

Harter:1961:USR

[Har61] H. Leon Harter. The use of sample ranges in setting exact confidence bounds for the standard deviation of a rectangular population. Journal of the American Statistical Association, 56(295): 601–609, September 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2282082.

Hardin:1963:BRB

[Har63] Einar Hardin. Book review: [untitled]. Journal of the American Statistical Association, 58(304):1199–1201, December 1963. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283381.

Harter:1964:CBS

[Har64] H. Leon Harter. Criteria for best substitute interval estimators, with an application to the normal distribution. Journal of the American Statistical Association, 59(308):1133–1140, December 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2282628.

Hartley:1966:SSU

[Har66] H. O. Hartley. Systematic sampling with unequal probability and without replacement. Journal of the American Statistical Association, 61(315):739–748, September 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282784.

Hartigan:1967:BRB

[Har67a] J. A. Hartigan. Book review: [untitled]. Journal of the American Statistical Association, 62(320):1506–1507, December 1967. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283801.

Hartigan:1967:RSM

[Har67b] J. A. Hartigan. Representation of similarity matrices by trees. Journal of the American Statistical Association, 62(320):1140– 1158, December 1967. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2283766.

Harville:1967:SDB

[Har67c] David A. Harville. Statistical dependence between random effects and the numbers of observations on the effects for the unbalanced one-way random classification. Journal of the American Statistical Association, 62(320):1375–1386, December 1967. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2283784.

Harris:1968:SIC

[Har68a] Bernard Harris. Statistical inference in the classical occupancy problem unbiased estimation of the number of classes. Journal of the American Statistical Association, 63(323):837–847, September 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2283876.

Harville:1968:SDB

[Har68b] David A. Harville. Statistical dependence between subclass means and the numbers of observations in the subclasses for the two-way completely-random classification. Journal of the American Statistical Association, 63(324):1484–1494, December 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2285897.

Hartigan:1969:BRB

[Har69a] J. A. Hartigan. Book review: [untitled]. Journal of the American Statistical Association, 64(328):1677–1678, December 1969. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2286110.

Hartigan:1969:USV

[Har69b] J. A. Hartigan. Using subsample values as typical values. Journal of the American Statistical Association, 64(328):1303–1317, December 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2286069.

Hastings:1960:BRB

[Has60] Delbet C. Hastings. Book review: [untitled]. Journal of the American Statistical Association, 55(291):618, September 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281945.

Hassanein:1968:AEV

[Has68] Khatab M. Hassanein. Analysis of extreme-value data by sample quantiles for very large samples. Journal of the American Statistical Association, 63(323):877–888, September 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283880.

Hasselblad:1969:EFM

[Has69] Victor Hasselblad. Estimation of finite mixtures of distributions from the exponential family. Journal of the American Statistical Association, 64(328):1459–1471, December 1969. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2286083.

Hauser:1962:BRB

[Hau62] Philip M. Hauser. Book review: [untitled]. Journal of the American Statistical Association, 57(300):926–927, December 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281830.

Hauser:1963:SSP

[Hau63] Philip M. Hauser. Statistics and society: Presidential address. Journal of the American Statistical Association, 58(301): 1–12, March 1963. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2282951.

Havighurst:1967:BRB

[Hav67] Robert J. Havighurst. Book review: [untitled]. Journal of the American Statistical Association, 62(319):1071–1073, September 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2283696.

Hays:1960:NAT

[Hay60] William L. Hays. A note on average tau as a measure of concordance. Journal of the American Statistical Association, 55(290): 331–341, June 1960. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2281746.

Haybittle:1965:TPM

[Hay65] J. L. Haybittle. A two-parameter model for the survival curve of treated cancer patients. Journal of the American Statistical Association, 60(309):16–26, March 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283134.

Han:1968:PMW

[HB68] Chien-Pai Han and T. A. Bancroft. On pooling means when variance is unknown. Journal of the American Statistical Association, 63(324):1333–1342, December 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor. org/stable/2285888. See corrigenda [HB69].

Han:1969:CPM

[HB69] Chien-Pai Han and T. A. Bancroft. Corrigenda: On Pooling Means When Variance Is Unknown. Journal of the American Statistical Association, 64(328):1700–1701, December 1969. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2286137. See [HB68].

Hewett:1969:NDS

[HBA69] John E. Hewett, William G. Bulgren, and D. E. Amos. A note on a double sample test. Journal of the American Statistical Association, 64(328):1542–1548, December 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286087.

Hopkins:1963:SED

[HC63] J. W. Hopkins and P. P. F. Clay. Some empirical distributions of bivariate T^{2} and homoscedasticity criterion M under unequal variance and leptokurtosis. Journal of the American Statistical Association, 58(304):1048–1053, December 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283331.

Holloway:1967:RH

[HD67] Lois Nelson Holloway and Olive Jean Dunn. The robustness of Hotelling’s T^{2}. Journal of the American Statistical Association, 62(317):124–136, March 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2282915.

Heady:1960:BRB

[Hea60] Earl O. Heady. Book review: [untitled]. Journal of the American Statistical Association, 55(292):771–772, December 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281617.

Heer:1960:BRB

[Hee60] David M. Heer. Book review: [untitled]. Journal of the American Statistical Association, 55(292):776–779, December 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281622.

Hege:1967:OPH

[Heg67] Vijaya S. Hege. An optimum property of the Horvitz–Thomson estimate. Journal of the American Statistical Association, 62(319): 1013–1017, September 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2283689.

Heien:1968:NLL

[Hei68] Dale M. Heien. A note on log-linear regression. Journal of the American Statistical Association, 63(323):1034–1038, September 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2283895.

Henderson:1961:BRBa

[Hen61a] John B. Henderson. Book review: [untitled]. Journal of the American Statistical Association, 56(293):212–213, March 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282380.

Henderson:1961:BRBb

[Hen61b] John P. Henderson. Book review: [untitled]. Journal of the American Statistical Association, 56(294):456–457, June 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282303.

Henshaw:1961:BRB

[Hen61c] Richard C. Henshaw, Jr. Book review: [untitled]. Journal of the American Statistical Association, 56(293):194–195, March 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282366.

Henry:1963:BRB

[Hen63] Neil W. Henry. Book review: [untitled]. Journal of the American Statistical Association, 58(303):824, September 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282730.

Hendricks:1964:EPO

[Hen64a] Walter A. Hendricks. Estimation of the probability that an observation will fall into a specified class. Journal of the American Statistical Association, 59(305):225–232, March 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282871.

Henry:1964:CPD

[Hen64b] Anderson Henry. Corrigenda: Product Diversification and the Cost of Living. Journal of the American Statistical Association, 59(308):1296, December 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2282643. See [And63a].

Henshaw:1964:BRB

[Hen64c] Richard C. Henshaw, Jr. Book review: [untitled]. Journal of the American Statistical Association, 59(306):626–627, June 1964. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283032.

Herman:1960:BRB

[Her60] Leon M. Herman. Book review: [untitled]. Journal of the American Statistical Association, 55(291):608–609, September 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281935.

Herzog:1961:BRB

[Her61] Fritz Herzog. Book review: [untitled]. Journal of the American Statistical Association, 56(293):214–215, March 1961. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2282382.

Herrey:1965:CIB

[Her65] Erna M. J. Herrey. Confidence intervals based on the mean absolute deviation of a normal sample. Journal of the American Statistical Association, 60(309):257–269, March 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283150.

Heston:1964:BRB

[Hes64] Alan Heston. Book review: [untitled]. Journal of the American Statistical Association, 59(308):1307–1308, December 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282656.

Hultquist:1965:MSS

[HG65] Robert A. Hultquist and Franklin A. Graybill. Minimal sufficient statistics for the two-way classification mixed model design. Journal of the American Statistical Association, 60(309): 182–192, March 1965. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2283144.

Halperin:1968:CBL

[HG68] Max Halperin and Joan Gurian. Confidence bands in linear regression with constraints on the independent variables. Journal of the American Statistical Association, 63(323):1020–1027, September 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2283893.

Harkness:1968:GHS

[HH68a] W. L. Harkness and M. L. Harkness. Generalized hyperbolic secant distributions. Journal of the American Statistical Association, 63(321):329–337, March 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283852.

Hildreth:1968:SEL

[HH68b] Clifford Hildreth and James P. Houck. Some estimators for a linear model with random coefficients. Journal of the American Statistical Association, 63(322):584–595, June 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284029. See corrigenda [HH69].

Hildreth:1969:CSE

[HH69] Clifford Hildreth and James P. Houck. Corrigenda: Some Estimators for a Linear Model with Random Coefficients. Journal of the American Statistical Association, 64(328):1701, December 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2286138. See [HH68b].

Hickman:1963:PRF

[Hic63] James C. Hickman. Preliminary regional forecasts for the outcome of an estimation problem. Journal of the American Statistical Association, 58(304):1104–1112, December 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283336.

Hickman:1964:NMC

[Hic64] Bert G. Hickman. On a new method of capacity estimation. Journal of the American Statistical Association, 59(306):529–549, June 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2283006.

Hill:1963:IEP

[Hil63a] Bruce M. Hill. Information for estimating the proportions in mixtures of exponential and normal distributions. Journal of the American Statistical Association, 58(304):918–932, December 1963. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283323.

Hill:1963:TPL

[Hil63b] Bruce M. Hill. The three-parameter lognormal distribution and Bayesian analysis of a point-source epidemic. Journal of the American Statistical Association, 58(301):72–84, March 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282955.

Hillier:1964:BRB

[Hil64a] Frederick S. Hillier. Book review: [untitled]. Journal of the American Statistical Association, 59(306):623–624, June 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283030.

Hillier:1964:CSP

[Hil64b] Frederick S. Hillier. Continuous sampling plans under destructive testing. Journal of the American Statistical Association, 59 (306):376–401, June 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2282994.

Hill:1965:BRBa

[Hil65a] Bruce M. Hill. Book review: [untitled]. Journal of the American Statistical Association, 60(309):364–367, March 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283162.

Hill:1965:BRBb

[Hil65b] Bruce M. Hill. Book review: [untitled]. Journal of the American Statistical Association, 60(312):1217–1218, December 1965. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283422.

Hill:1965:IAV

[Hil65c] Bruce M. Hill. Inference about variance components in the oneway model. Journal of the American Statistical Association, 60 (311):806–825, September 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283247.

Hillier:1966:BRB

[Hil66] Frederick S. Hillier. Book review: [untitled]. Journal of the American Statistical Association, 61(313):290, March 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283079.

Hill:1967:CER

[Hil67a] Bruce M. Hill. Correlated errors in the random model. Journal of the American Statistical Association, 62(320):1387–1400, December 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2283785.

Hillier:1967:SSP

[Hil67b] Frederick S. Hillier. Small sample probability limits for the range chart. Journal of the American Statistical Association, 62(320): 1488–1493, December 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/ stable/2283792. See corrigenda [Hil68c].

Hill:1968:PDP

[Hil68a] Bruce M. Hill. Posterior distribution of percentiles: Bayes’ Theorem for sampling from a population. Journal of the American Statistical Association, 63(322):677–691, June 1968. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2284038.

Hillier:1968:BRB

[Hil68b] Frederick S. Hillier. Book review: [untitled]. Journal of the American Statistical Association, 63(321):372–373, March 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283859.

Hillier:1968:CSS

[Hil68c] Frederick S. Hillier. Corrigenda: Small Sample Probability for the Range Charts. Journal of the American Statistical Association, 63 (324):1549–1550, December 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2285917. See [Hil67b].

Hinich:1965:BRB

[Hin65] Melvin Hinich. Book review: [untitled]. Journal of the American Statistical Association, 60(312):1238, December 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283434.

Hodges:1967:MQM

[HL67] J. L. Hodges, Jr. and E. L. Lehmann. On medians and quasi medians. Journal of the American Statistical Association, 62(319): 926–931, September 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2283680.

Halperin:1963:IEN

[HM63] Max Halperin and Nathan Mantel. Interval estimation of nonlinear parametric functions. Journal of the American Statistical Association, 58(303):611–627, September 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282709.

Harter:1966:CLM

[HM66a] H. Leon Harter and Albert H. Moore. Corrigenda: Local-Maximum-Likelihood Estimation of the Parameters of Three-Parameter Lognormal Populations from Complete and Censored Samples. Journal of the American Statistical Association, 61(316): 1247, December 1966. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/

2283230. See [HM66b].

Harter:1966:LML

[HM66b] H. Leon Harter and Albert H. Moore. Local-maximum-likelihood estimation of the parameters of three-parameter lognormal populations from complete and censored samples. Journal of the American Statistical Association, 61(315):842–851, September 1966. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282793. See corrigenda

[HM66a, HM67a, HM68a].

Harter:1967:CLM

[HM67a] H. Leon Harter and Albert H. Moore. Corrigenda: Local-Maximum-Likelihood Estimation of the Parameters of Three-Parameter Lognormal Populations from Complete and Censored Samples. Journal of the American Statistical Association, 62(320): 1519–1520, December 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2283815. See [HM66b].

Harter:1967:MLE

[HM67b] H. Leon Harter and Albert H. Moore. Maximum-likelihood estimation, from censored samples, of the parameters of a logistic distribution. Journal of the American Statistical Association, 62 (318):675–684, June 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2283994.

Harter:1968:CLM

[HM68a] H. Leon Harter and Albert H. Moore. Corrigenda: Local-Maximum-Likelihood Estimation of the Parameters of Three- Parameter Lognormal Population from Complete and Censored Samples. Journal of the American Statistical Association, 63(324): 1549, December 1968. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/

2285916. See [HM66b].

Harter:1968:MLE

[HM68b] H. Leon Harter and Albert H. Moore. Maximum-likelihood estimation, from doubly censored samples, of the parameters of the first asymptotic distribution of extreme values. Journal of the American Statistical Association, 63(323):889–901, September 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283881.

Hoel:1969:CST

[HM69] D. G. Hoel and M. Mazumdar. A class of sequential tests for an exponential parameter. Journal of the American Statistical Association, 64(328):1549–1559, December 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286088.

Ho:1962:BRB

[Ho62] Yu-Chi Ho. Book review: [untitled]. Journal of the American Statistical Association, 57(299):721–722, September 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282429.

Hoadley:1969:CMD

[Hoa69] Bruce Hoadley. The compound multinomial distribution and Bayesian analysis of categorical data from finite populations. Journal of the American Statistical Association, 64(325):216– 229, March 1969. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2283730.

Hochwald:1960:BRB

[Hoc60] Werner Hochwald. Book review: [untitled]. Journal of the American Statistical Association, 55(289):233–234, March 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282198.

Hoch:1962:BRB

[Hoc62] Irving Hoch. Book review: [untitled]. Journal of the American Statistical Association, 57(299):722, September 1962. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2282430.

Hoch:1963:BRB

[Hoc63a] Irving Hoch. Book review: [untitled]. Journal of the American Statistical Association, 58(303):853–857, September 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282748.

Hochwald:1963:BRB

[Hoc63b] Werner Hochwald. Book review: [untitled]. Journal of the American Statistical Association, 58(303):849–851, September 1963. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282746.

Hocking:1965:QRI

[Hoc65] R. R. Hocking. Quadratic regression with inequality restraints on the parameters. Journal of the American Statistical Association, 60(311):914–919, September 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283257.

Hochstim:1967:CCT

[Hoc67] Joseph R. Hochstim. A critical comparison of three strategies of collecting data from households. Journal of the American Statistical Association, 62(319):976–989, September 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283686.

Hodgson:1967:BRB

[Hod67] Vincent Hodgson. Book review: [untitled]. Journal of the American Statistical Association, 62(319):1082–1084, September 1967. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283704.

Hoeber:1962:BRB

[Hoe62] Francis P. Hoeber. Book review: [untitled]. Journal of the American Statistical Association, 57(297):230–231, March 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282466.

Hoeffding:1963:PIS

[Hoe63] Wassily Hoeffding. Probability inequalities for sums of bounded random variables. Journal of the American Statistical Association, 58(301):13–30, March 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2282952.

Hoffnar:1963:BRB

[Hof63] Bernard R. Hoffnar. Book review: [untitled]. Journal of the American Statistical Association, 58(301):257–258, March 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282976.

Hogg:1960:CUS

[Hog60a] Robert V. Hogg. Certain uncorrelated statistics. Journal of the American Statistical Association, 55(290):265–267, June 1960. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281739.

Hogg:1960:CEL

[Hog60b] Robert V. Hogg. On conditional expectations of location statistics. Journal of the American Statistical Association, 55(292): 714–717, December 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2281593.

Hogg:1961:RSH

[Hog61] Robert V. Hogg. On the resolution of statistical hypotheses. Journal of the American Statistical Association, 56(296):978–989, December 1961. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2282008.

Hogg:1962:ITE

[Hog62] Robert V. Hogg. Iterated tests of the equality of several distributions. Journal of the American Statistical Association, 57(299): 579–585, September 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2282395.

Hogg:1965:MHR

[Hog65] Robert V. Hogg. On models and hypotheses with restricted alternatives. Journal of the American Statistical Association, 60 (312):1153–1162, December 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283413.

Hogg:1967:BRB

[Hog67a] Robert V. Hogg. Book review: [untitled]. Journal of the American Statistical Association, 62(319):1091–1092, September 1967. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283709.

Hogg:1967:SOR

[Hog67b] Robert V. Hogg. Some observations on robust estimation. Journal of the American Statistical Association, 62(320):1179–1186, December 1967. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2283768.

Holland:1961:BRB

[Hol61] Daniel M. Holland. Book review: [untitled]. Journal of the American Statistical Association, 56(293):207–209, March 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282377.

Hollander:1962:BRB

[Hol62] Edward Hollander. Book review: [untitled]. Journal of the American Statistical Association, 57(300):935–936, December 1962. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281836.

Hollander:1963:BRB

[Hol63] Stanley C. Hollander. Book review: [untitled]. Journal of the American Statistical Association, 58(302):587–589, June 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283309.

Hollander:1967:AET

[Hol67] Myles Hollander. Asymptotic efficiency of two nonparametric competitors of Wilcoxon’s two sample test. Journal of the American Statistical Association, 62(319):939–949, September 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283682.

Hollander:1968:BRB

[Hol68a] Myles Hollander. Book review: [untitled]. Journal of the American Statistical Association, 63(322):721–722, June 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284042.

Hollander:1968:CUN

[Hol68b] Myles Hollander. Certain uncorrelated nonparametric test statistics. Journal of the American Statistical Association, 63(322): 707–714, June 1968. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2284040.

Hooke:1965:BRB

[Hoo65] Robert Hooke. Book review: [untitled]. Journal of the American Statistical Association, 60(312):1230–1231, December 1965. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283429.

Hooke:1967:BRB

[Hoo67] Robert Hooke. Book review: [untitled]. Journal of the American Statistical Association, 62(318):688–690, June 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283997.

Hopkins:1961:BRB

[Hop61] Carl E. Hopkins. Book review: [untitled]. Journal of the American Statistical Association, 56(293):180–182, March 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282355.

Horowitz:1963:VQI

[Hor63] Ira Horowitz. The varying (?) quality of investment trust management. Journal of the American Statistical Association, 58(304): 1011–1032, December 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2283328.

Horn:1969:BRB

[Hor69] Susan Dadakis Horn. Book review: [untitled]. Journal of the American Statistical Association, 64(327):1094–1095, September 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283495.

Houthakker:1961:BRB

[Hou61] H. S. Houthakker. Book review: [untitled]. Journal of the American Statistical Association, 56(293):205–206, March 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282374.

Houthakker:1962:BRB

[Hou62] Hendrik S. Houthakker. Book review: [untitled]. Journal of the American Statistical Association, 57(299):701–702, September 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2282415.

Howard:1963:BRB

[How63] Ronald A. Howard. Book review: [untitled]. Journal of the American Statistical Association, 58(302):583–584, June 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283306.

Howe:1969:TST

[How69] W. G. Howe. Two-sided tolerance limits for normal populations, some improvements. Journal of the American Statistical Association, 64(326):610–620, June 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283644.

Halperin:1967:SCB

[HRHY67] Max Halperin, Suresh C. Rastogi, Irwin Ho, and Y. Y. Yang. Shorter confidence bands in linear regression. Journal of the American Statistical Association, 62(319):1050–1067, September 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283693.

Hartley:1969:VEO

[HRK69] H. O. Hartley, J. N. K. Rao, and Grace Kiefer. Variance estimation with one unit per stratum. Journal of the American Statistical Association, 64(327):841–851, September 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283465.

Hocking:1968:EPM

[HS68] R. R. Hocking and Wm. B. Smith. Estimation of parameters in the multivariate normal distribution with missing observations. Journal of the American Statistical Association, 63(321):

159–173, March 1968. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2283837.

Hajek:1969:CTR

[HS69] J. Hajek and Z. Sidak. Corrigenda: The Theory of Rank Tests. Journal of the American Statistical Association, 64(328):1699, December 1969. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2286124.

Hess:1966:SPI

[HSB66] Irene Hess, V. K. Sethi, and T. R. Balakrishnan. Stratification: a practical investigation. Journal of the American Statistical Association, 61(313):74–90, March 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283045.

Hsi:1969:MSM

[Hsi69] Bartholomew P. Hsi. The multiple sample up-and-down method in bioassay. Journal of the American Statistical Association, 64 (325):147–162, March 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2283725.

Hogg:1963:IPT

[HT63] Robert V. Hogg and Elliot A. Tanis. An iterated procedure for testing the equality of several exponential distributions. Journal of the American Statistical Association, 58(302):435–443, June 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283277.

Hall:1967:MC

[HT67] Marshall Hall and Nicolaus Tideman. Measures of concentration. Journal of the American Statistical Association, 62(317): 162–168, March 1967. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2282919.

Huang:1963:ISC

[Hua63] David S. Huang. Initial stock and consumer investment in automobiles. Journal of the American Statistical Association, 58 (303):789–798, September 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2282726.

Huang:1964:BRB

[Hua64] David S. Huang. Book review: [untitled]. Journal of the American Statistical Association, 59(308):1323–1324, December 1964. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282670.

Hudson:1966:FSC

[Hud66] Derek J. Hudson. Fitting segmented curves whose join points have to be estimated. Journal of the American Statistical Association, 61 (316):1097–1129, December 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283203.

Hunter:1964:BRB

[Hun64] Holland Hunter. Book review: [untitled]. Journal of the American Statistical Association, 59(306):604–605, June 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283014.

Hurwitz:1962:CCC

[Hur62] Abner Hurwitz. Constants and compromise in the consumer price index. Journal of the American Statistical Association, 57(300): 813–825, December 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2281812.

Hyman:1961:BRB

[Hym61] Ray Hyman. Book review: [untitled]. Journal of the American Statistical Association, 56(294):446–448, June 1961. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2282294.

Ijiri:1966:BRB

[Iji66] Yuji Ijiri. Book review: [untitled]. Journal of the American Statistical Association, 61(313):292–293, March 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283081.

Ireland:1969:SMH

[IKK69] C. T. Ireland, H. H. Ku, and S. Kullback. Symmetry and marginal homogeneity of an r×r contingency table. Journal of the American Statistical Association, 64(328):1323–1341, December 1969. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2286071.

Intriligator:1968:BRB

[Int68] Michael D. Intriligator. Book review: [untitled]. Journal of the American Statistical Association, 63(322):723–724, June 1968. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2284044.

Ireson:1961:BRB

[Ire61] W. Grant Ireson. Book review: [untitled]. Journal of the American Statistical Association, 56(295):771–774, September 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282122.

J:1963:BRBa

[J.63a] D. W. J. Book review: [untitled]. Journal of the American Statistical Association, 58(301):269, March 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282986.

J:1963:BRBb

[J.63b] D. W. J. Book review: [untitled]. Journal of the American Statistical Association, 58(302):592, June 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283312.

J:1964:BRBa

[J.64a] D. W. J. Book review: [untitled]. Journal of the American Statistical Association, 59(306):627, June 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283033.

J:1964:BRBb

[J.64b] D. W. J. Book review: [untitled]. Journal of the American Statistical Association, 59(306):627, June 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283034.

J:1964:BRBc

[J.64c] D. W. J. Book review: [untitled]. Journal of the American Statistical Association, 59(306):628, June 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283036.

J:1964:BRBd

[J.64d] D. W. J. Book review: [untitled]. Journal of the American Statistical Association, 59(306):628, June 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283037.

J:1964:BRBe

[J.64e] D. W. J. Book review: [untitled]. Journal of the American Statistical Association, 59(306):628, June 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283038.

Jackson:1960:BSA

[Jac60] J. Edward Jackson. Bibliography on sequential analysis. Journal of the American Statistical Association, 55(291):561–580, September 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2281915.

Jacobs:1962:BRB

[Jac62] Walter W. Jacobs. Book review: [untitled]. Journal of the American Statistical Association, 57(300):930–931, December 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281833.

Jacobson:1963:WTS

[Jac63] James E. Jacobson. The Wilcoxon two-sample statistic: Tables and bibliography. Journal of the American Statistical Association, 58 (304):1086–1103, December 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283335.

Jackson:1966:BRB

[Jac66] J. E. Jackson. Book review: Concepts of Statistical Inference, by William C. Guenther. Journal of the American Statistical Association, 61(314):529–530, June 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2282845.

Jaech:1964:NET

[Jae64] J. L. Jaech. A note on the equivalence of two methods of fitting a straight line through cumulative data. Journal of the American Statistical Association, 59(307):863–866, September 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283103.

Jaffe:1963:BRB

[Jaf63] A. J. Jaffe. Book review: [untitled]. Journal of the American Statistical Association, 58(303):862–864, September 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282751.

Jaffe:1969:BRB

[Jaf69] A. J. Jaffe. Book review: [untitled]. Journal of the American Statistical Association, 64(327):1087–1089, September 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283489.

Jensen:1969:ICB

[Jen69] D. R. Jensen. An inequality for a class of bivariate chi-square distributions. Journal of the American Statistical Association, 64 (325):333–336, March 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2283743.

Jessen:1964:BRB

[Jes64] Raymond J. Jessen. Book review: [untitled]. Journal of the American Statistical Association, 59(306):616–617, June 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283023.

Jess:1967:BRB

[Jes67] I. Marcum Jess. Book review: [untitled]. Journal of the American Statistical Association, 62(319):1081–1082, September 1967. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283703.

Jessen:1969:SMP

[Jes69] R. J. Jessen. Some methods of probability non-replacement sampling. Journal of the American Statistical Association, 64(325): 175–193, March 1969. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2283727.

Jensen:1969:SCI

[JJ69] D. R. Jensen and M. Q. Jones. Simultaneous confidence intervals for variances. Journal of the American Statistical Association, 64(325):324–332, March 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283742.

Jain:1969:SMB

[JLA69] A. K. Jain, F. F. Leimkuhler, and V. L. Anderson. A statistical model of book use and its application to the book storage problem. Journal of the American Statistical Association, 64(328): 1211–1224, December 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286062.

Johnson:1960:BRB

[Joh60] D. Gale Johnson. Book review: [untitled]. Journal of the American Statistical Association, 55(290):393–394, June 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281766.

Johnson:1961:BRB

[Joh61a] Glenn L. Johnson. Book review: [untitled]. Journal of the American Statistical Association, 56(296):1010–1011, December 1961. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282017.

Johnson:1961:STA

[Joh61b] N. L. Johnson. A simple theoretical approach to cumulative sum control charts. Journal of the American Statistical Association, 56 (296):835–840, December 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281998.

Johnson:1962:BRB

[Joh62] D. Gale Johnson. Book review: [untitled]. Journal of the American Statistical Association, 57(297):262–263, March 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282487.

Johnson:1967:CSO

[Joh67a] Ellis L. Johnson. Computation and structure of optimal reset policies. Journal of the American Statistical Association, 62(320): 1462–1487, December 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/ stable/2283791. See corrigenda [Joh69a].

Johnson:1967:NUR

[Joh67b] N. L. Johnson. Note on a uniqueness relation in certain accident proneness models. Journal of the American Statistical Association, 62(317):288–289, March 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2282931.

Johnson:1968:BRB

[Joh68] Norman L. Johnson. Book review: [untitled]. Journal of the American Statistical Association, 63(321):371–372, March 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283858.

Johnson:1969:CCS

[Joh69a] Ellis L. Johnson. Corrigenda: Computation and Structure of Optimal Reset Policies. Journal of the American Statistical Association, 64(328):1701–1702, December 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286139. See [Joh67a].

Johnson:1969:BRB

[Joh69b] Richard A. Johnson. Book review: [untitled]. Journal of the American Statistical Association, 64(325):392–394, March 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283751.

Joiner:1969:MSL

[Joi69] Brian L. Joiner. The median significance level and other small sample measures of test efficacy. Journal of the American Statistical Association, 64(327):971–985, September 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283479.

Jones:1968:AVD

[Jon68a] Howard L. Jones. The analysis of variance of data from stratified subsamples. Journal of the American Statistical Association, 63(321):64–86, March 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2283828.

Jones:1968:BRB

[Jon68b] Richard H. Jones. Book review: [untitled]. Journal of the American Statistical Association, 63(322):730–732, June 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284050.

Jones:1969:BRB

[Jon69] Richard H. Jones. Book review: [untitled]. Journal of the American Statistical Association, 64(328):1671–1672, December 1969. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2286103.

Jorgenson:1961:BRB

[Jor61a] Dale W. Jorgenson. Book review: [untitled]. Journal of the American Statistical Association, 56(294):426–427, June 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282280.

Jorgenson:1961:MRA

[Jor61b] Dale W. Jorgenson. Multiple regression analysis of a Poisson process. Journal of the American Statistical Association, 56(294): 235–245, June 1961. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2282248.

Jorgenson:1962:BRB

[Jor62] Dale W. Jorgenson. Book review: [untitled]. Journal of the American Statistical Association, 57(297):255–257, March 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282484.

Jorgenson:1964:MVL

[Jor64] Dale W. Jorgenson. Minimum variance, linear, unbiased seasonal adjustment of economic time series. Journal of the American Statistical Association, 59(307):681–724, September 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283093. See [Jor65].

Jorgenson:1965:CMV

[Jor65] Dale W. Jorgenson. Corrigenda: Minimum Variance, Linear, Unbiased Seasonal Adjustment of Economic Time Series. Journal of the American Statistical Association, 60(312):1250, December 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283450. See [Jor64].

Jorgenson:1967:BRB

[Jor67a] Dale Jorgenson. Book review: [untitled]. Journal of the American Statistical Association, 62(320):1509, December 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283804.

Jorgenson:1967:SAD

[Jor67b] Dale W. Jorgenson. Seasonal adjustment of data for econometric analysis. Journal of the American Statistical Association, 62 (317):137–140, March 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2282916.

Joskow:1961:BRB

[Jos61] Jules Joskow. Book review: [untitled]. Journal of the American Statistical Association, 56(293):190–191, March 1961. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2282363.

Joskow:1965:BRB

[Jos65] Jules Joskow. Book review: [untitled]. Journal of the American Statistical Association, 60(310):678–680, June 1965. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2282703.

Joshi:1969:BAM

[Jos69] Prakash C. Joshi. Bounds and approximations for the moments of order statistics. Journal of the American Statistical Association, 64 (328):1617–1624, December 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286093.

Jaeger:1961:ACR

[JP61] Carol M. Jaeger and Jean L. Pennock. An analysis of consistency of response in household surveys. Journal of the American Statistical Association, 56(294):320–327, June 1961. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282257.

Jorgenson:1969:AIB

[JS69] Dale W. Jorgenson and James A. Stephenson. Anticipations and investment behavior in U.S. manufacturing, 1947–1960. Journal of the American Statistical Association, 64(325):67–89, March 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283720.

Judge:1966:IRR

[JT66] G. G. Judge and T. Takayama. Inequality restrictions in regression analysis. Journal of the American Statistical Association, 61 (313):166–181, March 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2283052.

Judge:1962:BRB

[Jud62] George G. Judge. Book review: [untitled]. Journal of the American Statistical Association, 57(297):246–247, March 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282478.

Judge:1964:BRB

[Jud64] George G. Judge. Book review: [untitled]. Journal of the American Statistical Association, 59(305):287–288, March 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282880.

Juster:1966:CBI

[Jus66] F. Thomas Juster. Consumer buying intentions and purchase probability: An experiment in survey design. Journal of the American Statistical Association, 61(315):658–696, September 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282779.

Koerts:1968:PBP

[KA68] J. Koerts and A. P. J. Abrahamse. On the power of the BLUS procedure. Journal of the American Statistical Association, 63 (324):1227–1236, December 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2285879.

Kabe:1963:ECF

[Kab63a] D. G. Kabe. Extension of Cochran’s formulae for addition or omission of a variate in multiple regression analysis. Journal of the American Statistical Association, 58(302):527–534, June 1963. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283288.

Kabe:1963:NED

[Kab63b] D. G. Kabe. A note on the exact distributions of the GCL estimators in two leading over identified cases. Journal of the American Statistical Association, 58(302):535–537, June 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283289.

Kabe:1963:SML

[Kab63c] D. G. Kabe. Stepwise multivariate linear regression. Journal of the American Statistical Association, 58(303):770–773, September 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2282724.

Kabe:1964:EDG

[Kab64] D. G. Kabe. On the exact distributions of the GCL estimators in a leading three-equation case. Journal of the American Statistical Association, 59(307):881–894, September 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283107.

Kabe:1967:MPI

[Kab67] D. G. Kabe. On multivariate prediction intervals for sample mean and covariance based on partial observations. Journal of the American Statistical Association, 62(318):634–637, June 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283990.

Kaiser:1965:BRB

[Kai65] Henry F. Kaiser. Book review: [untitled]. Journal of the American Statistical Association, 60(310):680–681, June 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282704.

Kakwani:1967:UZS

[Kak67] N. C. Kakwani. The unbiasedness of Zellner’s seemingly unrelated regression equations estimators. Journal of the American Statistical Association, 62(317):141–142, March 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282917.

Kallianpur:1962:BRB

[Kal62] G. B. Kallianpur. Book review: [untitled]. Journal of the American Statistical Association, 57(297):243–244, March 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282475.

Kallianpur:1965:BRB

[Kal65] G. Kallianpur. Book review: [untitled]. Journal of the American Statistical Association, 60(309):367–368, March 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283163.

Kamat:1962:SME

[Kam62] A. R. Kamat. Some more estimates of circular probable error. Journal of the American Statistical Association, 57(297):191– 195, March 1962. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2282450.

Kamerschen:1968:MGI

[Kam68] David R. Kamerschen. Market growth and industry concentration. Journal of the American Statistical Association, 63(321): 228–241, March 1968. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2283843.

Kannappan:1963:BRB

[Kan63] Subbiah Kannappan. Book review: [untitled]. Journal of the American Statistical Association, 58(304):1183–1184, December 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283367.

Kao:1960:BRB

[Kao60] Richard C. Kao. Book review: [untitled]. Journal of the American Statistical Association, 55(289):241–242, March 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282203.

Kaplan:1962:BRB

[Kap62] Norman M. Kaplan. Book review: [untitled]. Journal of the American Statistical Association, 57(297):226–227, March 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282464.

Karpinos:1962:CWH

[Kar62] Bernard D. Karpinos. Current weight-height relationships of youths of military age. Journal of the American Statistical Association, 57(300):895–905, December 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281817.

Karcz:1963:BRB

[Kar63a] Jerzy F. Karcz. Book review: [untitled]. Journal of the American Statistical Association, 58(302):572–575, June 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283296.

Karlin:1963:BRB

[Kar63b] Samuel Karlin. Book review: [untitled]. Journal of the American Statistical Association, 58(303):836–837, September 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282736.

Karson:1968:BRB

[Kar68] Marvin Karson. Book review: [untitled]. Journal of the American Statistical Association, 63(323):1047–1049, September 1968. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283899.

Katz:1961:BRB

[Kat61] Leo Katz. Book review: [untitled]. Journal of the American Statistical Association, 56(296):1024, December 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282027.

Katz:1963:BRB

[Kat63] Leo Katz. Book review: [untitled]. Journal of the American Statistical Association, 58(302):581, June 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283304.

Katti:1967:BRB

[Kat67] S. K. Katti. Book review: [untitled]. Journal of the American Statistical Association, 62(320):1508–1509, December 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283803.

Katz:1969:BRB

[Kat69] Leo Katz. Book review: [untitled]. Journal of the American Statistical Association, 64(327):1095–1096, September 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283496.

Kaufman:1965:BRB

[Kau65] Gordon Kaufman. Book review: [untitled]. Journal of the American Statistical Association, 60(312):1234–1236, December 1965. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283432.

Kaufman:1968:OSS

[Kau68] Gordon M. Kaufman. Optimal sample size in two-action problems when the sample observations are lognormal and the precision h is known. Journal of the American Statistical Association, 63(322):653–659, June 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284036.

Kelejian:1969:MOM

[Kel69] H. H. Kelejian. Missing observations in multivariate regression: Efficiency of a first- order method. Journal of the American Statistical Association, 64(328):1609–1616, December 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286092.

Kempthorne:1962:BRB

[Kem62] Oscar Kempthorne. Book review: [untitled]. Journal of the American Statistical Association, 57(299):719–721, September 1962. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282428.

Kempthorne:1964:BRB

[Kem64] Oscar Kempthorne. Book review: [untitled]. Journal of the American Statistical Association, 59(305):285–287, March 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282879.

Kempthorne:1966:SAE

[Kem66] Oscar Kempthorne. Some aspects of experimental inference. Journal of the American Statistical Association, 61(313):11–34, March 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2283043. See cor-

rigenda [Kem67].

Kempthorne:1967:CSA

[Kem67] Oscar Kempthorne. Corrigenda: Some Aspects of Experimental Inference. Journal of the American Statistical Association, 62(320):1520, December 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283816. See [Kem66].

Keyfitz:1966:LTA

[Key66] Nathan Keyfitz. A life table that agrees with the data. Journal of the American Statistical Association, 61(314):305–312, June 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282820.

Keyfitz:1968:LTA

[Key68] Nathan Keyfitz. A life table that agrees with the data: II. Journal of the American Statistical Association, 63(324):1253–1268, December 1968. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2285882.

Keyfitz:1969:BRB

[Key69] Nathan Keyfitz. Book review: [untitled]. Journal of the American Statistical Association, 64(326):670–671, June 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283653.

Kuehn:1963:BRB

[KF63] Alfred A. Kuehn and John U. Farley. Book review: [untitled]. Journal of the American Statistical Association, 58(302):589–591, June 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2283310.

Kmenta:1968:SSP

[KG68] Jan Kmenta and Roy F. Gilbert. Small sample properties of alternative estimators of seemingly unrelated regressions. Journal of the American Statistical Association, 63(324):1180–1200, December 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2285876.

Kish:1959:VRT

[KH59] Leslie Kish and Irene Hess. On variances of ratios and their differences in multi-stage samples. Journal of the American Statistical Association, 54(286):416–446, June 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281782. See corrigenda [KH63].

Kish:1963:CVR

[KH63] Leslie Kish and Irene Hess. Corrigenda: On Variances of Ratios and Their Differences in Multi-Stage Samples. Journal of the American Statistical Association, 58(304):1162, December 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283347. See [KH59].

Khatri:1960:TEP

[Kha60] C. G. Khatri. On testing the equality of parameters in k rectangular populations. Journal of the American Statistical Association, 55(289):144–147, March 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2282185.

Khatri:1967:TLS

[Kha67] C. G. Khatri. A theorem on least squares in multivariate linear regression. Journal of the American Statistical Association, 62 (320):1494–1495, December 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283793.

Khan:1968:NEM

[Kha68] Rasul A. Khan. A note on estimating the mean of a normal distribution with known coefficient of variation. Journal of the American Statistical Association, 63(323):1039–1041, September 1968. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283896.

Kiefer:1960:BRB

[Kie60] Jack Kiefer. Book review: [untitled]. Journal of the American Statistical Association, 55(291):616, September 1960. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2281942.

Kimball:1960:CPP

[Kim60] Bradford F. Kimball. On the choice of plotting positions on probability paper. Journal of the American Statistical Association, 55 (291):546–560, September 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281914.

Kimball:1962:BRB

[Kim62] A. W. Kimball. Book review: [untitled]. Journal of the American Statistical Association, 57(297):203–205, March 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282454.

Kim:1969:EAS

[Kim69] P. J. Kim. On the exact and approximate sampling distribution of the two sample Kolmogorov–Smirnov criterion D_{mn}, m ≤ n. Journal of the American Statistical Association, 64(328):1625–1637, December 1969. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2286094.

Kincaid:1962:CTB

[Kin62a] W. M. Kincaid. The combination of tests based on discrete distributions. Journal of the American Statistical Association, 57 (297):10–19, March 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2282436.

Kindahl:1962:EMT

[Kin62b] James K. Kindahl. Estimation of means and totals from finite populations of unknown size. Journal of the American Statistical Association, 57(297):61–91, March 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282441. See corrigenda [Kin64a].

Kinney:1963:BRB

[Kin63] Paul T. Kinney. Book review: [untitled]. Journal of the American Statistical Association, 58(302):581–583, June 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283305.

Kindahl:1964:CEM

[Kin64a] James K. Kindahl. Corrigenda: Estimation of Means and Totals from Finite Populations of Unknown Size. Journal of the American Statistical Association, 59(308):1297–1298, December 1964. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282648. See [Kin62b].

King:1964:BRB

[Kin64b] Richard A. King. Book review: [untitled]. Journal of the American Statistical Association, 59(306):617–618, June 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283024.

Kingston:1965:APTa

[Kin65a] Charles R. Kingston. Applications of probability theory in criminalistics. Journal of the American Statistical Association, 60(309): 70–80, March 1965. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2283138.

Kingston:1965:APTb

[Kin65b] Charles R. Kingston. Applications of probability theory in criminalistics, II. Journal of the American Statistical Association, 60 (312):1028–1034, December 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283402.

Kindleberger:1966:BRB

[Kin66] C. P. Kindleberger. Book review: [untitled]. Journal of the American Statistical Association, 61(315):875–876, September 1966. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282803.

King:1967:SWC

[Kin67] Benjamin King. Step-wise clustering procedures. Journal of the American Statistical Association, 62(317):86–101, March 1967.

CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282912.

Kish:1961:BRBa

[Kis61a] Leslie Kish. Book review: [untitled]. Journal of the American Statistical Association, 56(295):740–744, September 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282095.

Kish:1961:BRBb

[Kis61b] Leslie Kish. Book review: [untitled]. Journal of the American Statistical Association, 56(296):1011–1013, December 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282018.

Kish:1962:SIV

[Kis62] Leslie Kish. Studies of interviewer variance for attitudinal variables. Journal of the American Statistical Association, 57(297): 92–115, March 1962. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2282442.

Kish:1968:SEI

[Kis68] Leslie Kish. Standard errors for indexes from complex samples. Journal of the American Statistical Association, 63(322):512–529, June 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2284022.

Kish:1969:CSS

[Kis69] Leslie Kish. Corrigenda: Survey Sampling. Journal of the American Statistical Association, 64(328):1699, December 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286127.

Kitagawa:1960:BRB

[Kit60] Evelyn M. Kitagawa. Book review: [untitled]. Journal of the American Statistical Association, 55(290):379–381, June 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281754.

Khatri:1963:EPT

[KJ63] C. G. Khatri and M. C. Jaiswal. Estimation of parameters of a truncated bivariate normal distribution. Journal of the American Statistical Association, 58(302):519–526, June 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283287.

Klauber:1969:BRB

[Kla69] Melville R. Klauber. Book review: [untitled]. Journal of the American Statistical Association, 64(326):675, June 1969. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2283657.

Klein:1960:BRB

[Kle60] L. R. Klein. Book review: [untitled]. Journal of the American Statistical Association, 55(291):583–584, September 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281917.

Klotz:1964:NST

[Klo64] J. H. Klotz. On the normal scores two-sample rank tests. Journal of the American Statistical Association, 59(307):652–664, September 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2283091.

Klotz:1966:WTC

[Klo66] J. H. Klotz. The Wilcoxon, ties, and the computer. Journal of the American Statistical Association, 61(315):772–787, September 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2282786. See cor-

rigenda [Klo67b].

Klotz:1967:AET

[Klo67a] Jerome Klotz. Asymptotic efficiency of the two sample Kolmogorov– Smirnov test. Journal of the American Statistical Association, 62 (319):932–938, September 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283681.

Klotz:1967:CWT

[Klo67b] Jerome Klotz. Corrigenda: The Wilcoxon, Ties, and the Computer. Journal of the American Statistical Association, 62(320): 1520, December 1967. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/

2283817. See [Klo66].

Klotz:1969:MSE

[KMZ69] J. H. Klotz, R. C. Milton, and S. Zacks. Mean square efficiency of estimators of variance components. Journal of the American Statistical Association, 64(328):1383–1402, December 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286076.

Krishnaswami:1968:BMC

[KN68] P. Krishnaswami and Rajeshwar Nath. Bias in multinomial classification. Journal of the American Statistical Association, 63(321): 298–303, March 1968. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2283849.

Knight:1966:CMC

[Kni66] William R. Knight. A computer method for calculating Kendall’s tau with ungrouped data. Journal of the American Statistical Association, 61(314):436–439, June 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282833.

Kish:1962:RBS

[KNP62] Leslie Kish, N. Krishnan Namboodiri, and R. Krishna Pillai. The ratio bias in surveys. Journal of the American Statistical Association, 57(300):863–876, December 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281815.

Keyfitz:1967:IAD

[KNS67] Nathan Keyfitz, Dhruva Nagnur, and Divakar Sharma. On the interpretation of age distributions. Journal of the American Statistical Association, 62(319):862–874, September 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283676.

Koch:1969:ENS

[Koc69a] Gary G. Koch. The effect of non-sampling errors on measures of association in 2 × 2 contingency tables. Journal of the American Statistical Association, 64(327):852–863, September 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283466.

Koch:1969:SAS

[Koc69b] Gary G. Koch. Some aspects of the statistical analysis of “split plot” experiments in completely randomized layouts. Journal of the American Statistical Association, 64(326):485–505, June 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283634.

Koch:1969:ULP

[Koc69c] Gary G. Koch. A useful lemma for proving the equality of two matrices with applications to least squares type quadratic forms. Journal of the American Statistical Association, 64(327):969–970, September 1969. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2283478.

Koenigsberg:1966:BRB

[Koe66] Ernest Koenigsberg. Book review: [untitled]. Journal of the American Statistical Association, 61(316):1236–1237, December 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283217.

Koerts:1967:SFN

[Koe67] John Koerts. Some further notes on disturbance estimates in regression analysis. Journal of the American Statistical Association, 62(317):169–183, March 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2282920.

Konijn:1962:RAS

[Kon62] H. S. Konijn. Regression analysis in sample surveys. Journal of the American Statistical Association, 57(299):590–606, September 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2282397. See cor-

rigenda [Kon63].

Konijn:1963:CRA

[Kon63] H. S. Konijn. Corrigenda: Regression Analysis in Sample Surveys. Journal of the American Statistical Association, 58(304): 1162, December 1963. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/

2283348. See [Kon62].

Koo:1962:BRB

[Koo62a] Anthony Y. C. Koo. Book review: [untitled]. Journal of the American Statistical Association, 57(297):261–262, March 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282486.

Koopmans:1962:BRB

[Koo62b] L. H. Koopmans. Book review: [untitled]. Journal of the American Statistical Association, 57(298):508–510, June 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281660.

Koo:1964:BRB

[Koo64] Anthony Y. C. Koo. Book review: [untitled]. Journal of the American Statistical Association, 59(305):303–306, March 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282891.

Koopmans:1966:NEA

[Koo66] L. H. Koopmans. A note on the estimation of amplitude spectra for stochastic processes with quasi-linear residuals. Journal of the American Statistical Association, 61(314):397–402, June 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282829.

Korbel:1966:CLF

[Kor66a] John Korbel. Corrigenda: Labor Force Entry and Attachment of Young People. Journal of the American Statistical Association, 61(316):1248, December 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283233. See [Kor66b].

Korbel:1966:LFE

[Kor66b] John Korbel. Labor force entry and attachment of young people. Journal of the American Statistical Association, 61(313):117– 127, March 1966. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/

2283048. See corrigenda [Kor66a].

Kossack:1961:BRBa

[Kos61a] Carl F. Kossack. Book review: [untitled]. Journal of the American Statistical Association, 56(294):443–444, June 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282291.

Kossack:1961:BRBb

[Kos61b] Carl F. Kossack. Book review: [untitled]. Journal of the American Statistical Association, 56(294):455–456, June 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282302.

Kotlarski:1966:CCS

[Kot66] Ignacy Kotlarski. On characterizing the chi square distribution by the Student law. Journal of the American Statistical Association, 61(316):976–981, December 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283193.

Kozelka:1966:BRB

[Koz66] R. M. Kozelka. Book review: Mathematics of Choice, by Ivan Niven. Journal of the American Statistical Association, 61(314): 530–531, June 1966. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2282846.

Kpedekpo:1969:WLT

[Kpe69] G. M. K. Kpedekpo. Working life tables for males in Ghana, 1960. Journal of the American Statistical Association, 64(325): 102–110, March 1969. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2283722.

Kraft:1960:BRB

[Kra60a] Charles H. Kraft. Book review: [untitled]. Journal of the American Statistical Association, 55(292):787, December 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281633.

Kravis:1960:BRB

[Kra60b] Irving Kravis. Book review: [untitled]. Journal of the American Statistical Association, 55(290):382–383, June 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281756.

Kramer:1963:TCC

[Kra63] K. H. Kramer. Tables for constructing confidence limits on the multiple correlation coefficient. Journal of the American Statistical Association, 58(304):1082–1085, December 1963. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2283334.

Kramer:1965:BRB

[Kra65] Clyde Y. Kramer. Book review: [untitled]. Journal of the American Statistical Association, 60(309):385–386, March 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283177.

Kreinin:1960:BRB

[Kre60] Mordechai E. Kreinin. Book review: [untitled]. Journal of the American Statistical Association, 55(291):592–594, September 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2281924.

Kreinin:1964:BRB

[Kre64] Mordechai E. Kreinin. Book review: [untitled]. Journal of the American Statistical Association, 59(305):309–311, March 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282893.

Krishnan:1967:CMD

[Kri67a] Marakatha Krishnan. Corrigenda: The Moments of a Doubly Noncentral t-Distribution. Journal of the American Statistical Association, 62(320):1520, December 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283818. See [Kri67b].

Krishnan:1967:MDN

[Kri67b] Marakatha Krishnan. The moments of a doubly noncentral tdistribution. Journal of the American Statistical Association, 62 (317):278–287, March 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/ stable/2282930. See corrigenda [Kri67a].

Krishnan:1968:SRD

[Kri68] Marakatha Krishnan. Series representations of the doubly noncentral t-distribution. Journal of the American Statistical Association, 63(323):1004–1012, September 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283891. See corrigenda [Kri69].

Krishnan:1969:CSR

[Kri69] Marakatha Krishnan. Corrigenda: Series Representations of the Doubly Noncentral t-Distribution. Journal of the American Statistical Association, 64(328):1702, December 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286140. See [Kri68].

Krueger:1960:BRB

[Kru60] Anne Krueger. Book review: [untitled]. Journal of the American Statistical Association, 55(292):772–773, December 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281618.

Krafft:1969:NHI

[KS69] O. Krafft and N. Schmitz. A note on Hoeffding’s inequality. Journal of the American Statistical Association, 64(327):907–912, September 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2283471.

Kronmal:1968:EPD

[KT68] R. Kronmal and M. Tarter. The estimation of probability densities and cumulatives by Fourier series methods. Journal of the American Statistical Association, 63(323):925–952, September 1968. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283885.

Kuebler:1965:BRB

[Kue65] Roy R. Kuebler, Jr. Book review: [untitled]. Journal of the American Statistical Association, 60(312):1207–1210, December 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283418.

Kuh:1960:BRB

[Kuh60] Edwin Kuh. Book review: [untitled]. Journal of the American Statistical Association, 55(291):594–595, September 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281925.

Kullback:1965:BRB

[Kul65] Solomon Kullback. Book review: [untitled]. Journal of the American Statistical Association, 60(310):677–678, June 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282702.

Kullback:1967:TCI

[Kul67] S. Kullback. The two concepts of information. Journal of the American Statistical Association, 62(318):685–686, June 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283995.

Kumar:1967:MCA

[Kum67] Joginder Kumar. Method of construction of attrition life tables for the single population based on two successive censuses. Journal of the American Statistical Association, 62(320):1433–1451, December 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2283788. See

corrigenda [Kum68].

Kumar:1968:CMC

[Kum68] Joginder Kumar. Corrigenda: Method of Construction of Attrition Life Tables for the Single Population Based on Two Successive Censuses. Journal of the American Statistical Association, 63(324):1550, December 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2285918. See [Kum67].

Kurnow:1960:BRB

[Kur60] Ernest Kurnow. Book review: [untitled]. Journal of the American Statistical Association, 55(289):231–233, March 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282197.

Kurz:1962:BRB

[Kur62] Mordecai Kurz. Book review: [untitled]. Journal of the American Statistical Association, 57(300):936–938, December 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281837.

Kapadia:1964:AGD

[KW64] C. H. Kapadia and D. L. Weeks. On the analysis of group divisible designs. Journal of the American Statistical Association, 59 (308):1217–1219, December 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2282637.

Ladd:1961:SMF

[Lad61] George W. Ladd. On some measures of food marketing services. Journal of the American Statistical Association, 56(293): 65–69, March 1961. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2282331.

Ladd:1963:BRB

[Lad63] George W. Ladd. Book review: [untitled]. Journal of the American Statistical Association, 58(304):1198–1199, December 1963. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283380.

Ladd:1964:RAS

[Lad64] George W. Ladd. Regression analysis of seasonal data. Journal of the American Statistical Association, 59(306):402–421, June 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282995.

Lampman:1964:BRB

[Lam64] Robert J. Lampman. Book review: [untitled]. Journal of the American Statistical Association, 59(306):606–607, June 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283016.

Lancaster:1961:STD

[Lan61] H. O. Lancaster. Significance tests in discrete distributions. Journal of the American Statistical Association, 56(294):223–234, June 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2282247. See cor-

rigenda [Lan62].

Lancaster:1962:CST

[Lan62] H. O. Lancaster. Corrigenda: Significance Tests in Discrete Distributions. Journal of the American Statistical Association, 57 (300):919, December 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2281824. See [Lan61].

Lancaster:1966:NEV

[Lan66] Tony Lancaster. A note on an ’errors in variables’ model. Journal of the American Statistical Association, 61(313):128–135, March 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2283049.

Lansing:1967:BRB

[Lan67] John B. Lansing. Book review: [untitled]. Journal of the American Statistical Association, 62(320):1496–1497, December 1967. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283794.

Lancaster:1968:CGE

[Lan68a] Tony Lancaster. Corrigenda: Grouping Estimators on Heteroscedastic Data. Journal of the American Statistical Association, 63(324):1550, December 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2285919. See [Lan68b].

Lancaster:1968:GEH

[Lan68b] Tony Lancaster. Grouping estimators on heteroscedastic data. Journal of the American Statistical Association, 63(321):182– 191, March 1968. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/

2283839. See corrigenda [Lan68a].

Latta:1961:BRB

[Lat61] Gordon E. Latta. Book review: [untitled]. Journal of the American Statistical Association, 56(294):466–467, June 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282311.

Latane:1963:BRB

[Lat63] Henry A. Latane. Book review: [untitled]. Journal of the American Statistical Association, 58(302):567–568, June 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283294.

Laubscher:1961:SBN

[Lau61] Nico F. Laubscher. On stabilizing the binomial and negative binomial variances. Journal of the American Statistical Association, 56(293):143–150, March 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2282340.

Laurent:1963:CLD

[Lau63a] Andre G. Laurent. Corrigenda: The Lognormal Distribution and the Translation Method: Description and Estimation Problems. Journal of the American Statistical Association, 58(304): 1163, December 1963. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/

2283349. See [Lau63b].

Laurent:1963:LDT

[Lau63b] Andre G. Laurent. The lognormal distribution and the translation method: Description and estimation problems. Journal of the American Statistical Association, 58(301):231–235, March 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2282966. See cor-

rigenda [Lau63a].

Lave:1966:BRB

[Lav66] Roy G. Lave, Jr. Book review: [untitled]. Journal of the American Statistical Association, 61(315):879–881, September 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282809.

Lave:1967:BRB

[Lav67] Roy E. Lave, Jr. Book review: Quality Control and Industrial Statistics, by Acheson J. Duncan. Journal of the American Statistical Association, 62(317):303–304, March 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282940.

LaValle:1968:CEIa

[LaV68a] Irving H. LaValle. On cash equivalents and information evaluation in decisions under uncertainty: Part I: Basic theory. Journal of the American Statistical Association, 63(321):252–276, March 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2283845.

LaValle:1968:CEIb

[LaV68b] Irving H. LaValle. On cash equivalents and information evaluation in decisions under uncertainty: Part II: Incremental information decisions. Journal of the American Statistical Association, 63(321):277–284, March 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283846.

LaValle:1968:CEIc

[LaV68c] Irving H. LaValle. On cash equivalents and information evaluation in decisions under uncertainty: Part III: Exchanging partition-J for partition-K information. Journal of the American Statistical Association, 63(321):285–290, March 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283847.

Lazerwitz:1961:CMU

[Laz61] Bernard Lazerwitz. A comparison of major United States religious groups. Journal of the American Statistical Association, 56 (295):568–579, September 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2282080.

Lentner:1963:SIA

[LB63] M. M. Lentner and R. J. Buehler. Some inferences about gamma parameters with an application to a reliability problem. Journal of the American Statistical Association, 58(303):670–677, September 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2282715.

Leadbetter:1969:BRB

[Lea69] M. R. Leadbetter. Book review: [untitled]. Journal of the American Statistical Association, 64(326):678, June 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283661.

Lechner:1964:OOC

[Lec64] James A. Lechner. Optimality and the OC curve for the Wald SPRT. Journal of the American Statistical Association, 59(306): 464–468, June 1964. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2283000.

Leeuw:1962:CC

[Lee62] Frank De Leeuw. The concept of capacity. Journal of the American Statistical Association, 57(300):826–840, December 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281813.

Lee:1964:BRB

[Lee64a] Ivan M. Lee. Book review: [untitled]. Journal of the American Statistical Association, 59(305):289–291, March 1964. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2282881.

Lee:1964:IIC

[Lee64b] Maw Lin Lee. Income, income change, and durable goods demand. Journal of the American Statistical Association, 59(308): 1194–1202, December 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2282634.

Lee:1964:IWD

[Lee64c] Tong Hun Lee. Income, wealth, and the demand for money: Some evidence from cross-section data. Journal of the American Statistical Association, 59(307):746–762, September 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283096.

Lee:1968:BRB

[Lee68] Wayne Lee. Book review: [untitled]. Journal of the American Statistical Association, 63(321):370–371, March 1968. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2283857.

Lehman:1961:EAD

[Leh61] Shirley Young Lehman. Exact and approximate distributions for the Wilcoxon statistic with ties. Journal of the American Statistical Association, 56(294):293–298, June 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282254.

Leipnik:1964:BRB

[Lei64] Roy B. Leipnik. Book review: [untitled]. Journal of the American Statistical Association, 59(308):1300–1301, December 1964. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282652.

Leong:1962:UIM

[Leo62] Y. S. Leong. The use of an iterated moving average in measuring seasonal variations. Journal of the American Statistical Association, 57(297):149–171, March 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2282445.

Lerner:1962:BRBa

[Ler62a] Abba P. Lerner. Book review: [untitled]. Journal of the American Statistical Association, 57(297):211–220, March 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282458.

Lerner:1962:BRBb

[Ler62b] Abba P. Lerner. Book review: [untitled]. Journal of the American Statistical Association, 57(299):704–709, September 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282417.

Leser:1963:EQL

[Les63] C. E. V. Leser. Estimation of quasi-linear trend and seasonal variation. Journal of the American Statistical Association, 58(304): 1033–1043, December 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2283329.

Levine:1962:BRB

[Lev62] Herbert S. Levine. Book review: [untitled]. Journal of the American Statistical Association, 57(300):928–930, December 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281832.

Lazerfeld:1969:CRM

[LH69] Paul F. Lazerfeld and Neil W. Henry. Corrigenda: Readings in Mathematical Social Sciences. Journal of the American Statistical Association, 64(328):1700, December 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286133.

Li:1962:SMS

[Li62] Chou Hsiung Li. A sequential method for screening experimental variables. Journal of the American Statistical Association, 57 (298):455–477, June 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2281652.

Lieberman:1960:BRB

[Lie60] G. J. Lieberman. Book review: [untitled]. Journal of the American Statistical Association, 55(290):383–384, June 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281757.

Lieberman:1961:BRB

[Lie61] Gerald J. Lieberman. Book review: [untitled]. Journal of the American Statistical Association, 56(296):1025, December 1961. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282029.

Lieberman:1963:BRB

[Lie63] Gerald J. Lieberman. Book review: [untitled]. Journal of the American Statistical Association, 58(302):566, June 1963. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2283293.

Lieberman:1964:BRB

[Lie64] G. J. Lieberman. Book review: [untitled]. Journal of the American Statistical Association, 59(308):1301–1302, December 1964. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282653.

Likert:1960:DFS

[Lik60] Rensis Likert. The dual function of statistics: Presidential address. Journal of the American Statistical Association, 55(289): 1–7, March 1960. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2282175.

Likes:1967:DSS

[Lik67] Jiří Likeš. Distributions of some statistics in samples from exponential and power-function populations. Journal of the American Statistical Association, 62(317):259–271, March 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282928.

Lilliefors:1967:KST

[Lil67] Hubert W. Lilliefors. On the Kolmogorov–Smirnov test for normality with mean and variance unknown. Journal of the American Statistical Association, 62(318):399–402, June 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283970.

Lilliefors:1969:CKS

[Lil69a] Hubert W. Lilliefors. Corrigenda: On the Kolmogorov–Smirnov Test for Normality with Mean and Variance Unknown. Journal of the American Statistical Association, 64(328):1702, December 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2286141. See

[Lil69b].

Lilliefors:1969:KST

[Lil69b] Hubert W. Lilliefors. On the Kolmogorov–Smirnov test for the exponential distribution with mean unknown. Journal of the American Statistical Association, 64(325):387–389, March 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283748. See corrigenda [Lil69a].

Lindsey:1961:PSD

[Lin61] G. R. Lindsey. The progress of the score during a baseball game. Journal of the American Statistical Association, 56(295):703–728, September 1961. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2282091.

Lindeman:1962:BRB

[Lin62a] Richard Lindeman. Book review: [untitled]. Journal of the American Statistical Association, 57(298):517–518, June 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281668.

Lindley:1962:BRB

[Lin62b] D. V. Lindley. Book review: [untitled]. Journal of the American Statistical Association, 57(300):922–924, December 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281828.

Lindeman:1963:BRBa

[Lin63a] Richard Lindeman. Book review: [untitled]. Journal of the American Statistical Association, 58(301):251, March 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282971.

Lindeman:1963:BRBb

[Lin63b] Richard H. Lindeman. Book review: [untitled]. Journal of the American Statistical Association, 58(302):576–577, June 1963. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283299.

Lindgren:1963:BRB

[Lin63c] B. W. Lindgren. Book review: [untitled]. Journal of the American Statistical Association, 58(301):256–257, March 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282975.

Littauer:1961:BRB

[Lit61] S. B. Littauer. Book review: [untitled]. Journal of the American Statistical Association, 56(293):203–204, March 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282372.

Littauer:1963:BRB

[Lit63] Sebastian B. Littauer. Book review: [untitled]. Journal of the American Statistical Association, 58(301):262–263, March 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282980.

Littauer:1965:BRB

[Lit65] Sebastian B. Littauer. Book review: [untitled]. Journal of the American Statistical Association, 60(312):1233–1234, December 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2283431.

Lee:1968:MLB

[LJZ68] T. C. Lee, G. G. Judge, and A. Zellner. Maximum likelihood and Bayesian estimation of transition probabilities. Journal of the American Statistical Association, 63(324):1162–1179, December 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2285875.

Lee:1960:IMS

[LL60] Everett S. Lee and Anne S. Lee. Internal migration statistics for the United States. Journal of the American Statistical Association, 55(292):664–697, December 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281590.

Lees:1961:NCP

[LL61] Ruth W. Lees and Frederic M. Lord. A nomograph for computing partial correlation coefficients. Journal of the American Statistical Association, 56(296):995–997, December 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282010. See corrigenda [LL62].

Lees:1962:CNC

[LL62] Ruth W. Lees and Frederic M. Lord. Corrigenda: A Nomograph for Computing Partial Correlation Coefficients. Journal of the American Statistical Association, 57(300):917–918, December 1962. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281822. See [LL61].

LEsperance:1969:GSP

[LNF69] W. L. L’Esperance, G. Nestel, and D. Fromm. Gross state product and an econometric model of a state. Journal of the American Statistical Association, 64(327):787–807, September 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283461.

Lewis:1966:GGM

[LO66] T. O. Lewis and P. L. Odell. A generalization of the Gauss–Markov theorem. Journal of the American Statistical Association, 61(316): 1063–1066, December 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2283200.

Locks:1960:BRB

[Loc60] Mitchell O. Locks. Book review: [untitled]. Journal of the American Statistical Association, 55(291):603–604, September 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281931.

Long:1960:BRB

[Lon60] Clarence D. Long. Book review: [untitled]. Journal of the American Statistical Association, 55(290):390–391, June 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281763.

Longley:1967:ALS

[Lon67] James W. Longley. An appraisal of least squares programs for the electronic computer from the point of view of the user. Journal of the American Statistical Association, 62(319):819–841, September 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2283673.

Lord:1960:LSC

[Lor60] Frederic M. Lord. Large-sample covariance analysis when the control variable is fallible. Journal of the American Statistical Association, 55(290):307–321, June 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281743.

Lovell:1963:SAE

[Lov63] Michael C. Lovell. Seasonal adjustment of economic time series and multiple regression analysis. Journal of the American Statistical Association, 58(304):993–1010, December 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283327.

Lovell:1966:AAS

[Lov66a] Michael C. Lovell. Alternative axiomatizations of seasonal adjustment. Journal of the American Statistical Association, 61(315): 800–802, September 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/ stable/2282788. See corrigenda [Lov67].

Lovell:1966:BRB

[Lov66b] Michael C. Lovell. Book review: [untitled]. Journal of the American Statistical Association, 61(315):865–867, September 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282796.

Lovell:1967:CAA

[Lov67] Michael C. Lovell. Corrigenda: Alternative Axiomatizations of Seasonal Adjustment. Journal of the American Statistical Association, 62(320):1520–1522, December 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283819. See [Lov66a].

Low:1969:EVC

[Low69] Leone Y. Low. Estimators of variance components in the balanced incomplete block design. Journal of the American Statistical Association, 64(327):1014–1030, September 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283482.

Leone:1961:USQ

[LRT61] F. C. Leone, Y. H. Rutenberg, and C. W. Topp. The use of sample quasi-ranges in setting confidence intervals for the population standard deviation. Journal of the American Statistical Association, 56(294):260–272, June 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2282251.

Lubell:1961:BRB

[Lub61] H. Lubell. Book review: [untitled]. Journal of the American Statistical Association, 56(294):448–449, June 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282296.

Lydall:1961:BRB

[Lyd61] Harold Lydall. Book review: [untitled]. Journal of the American Statistical Association, 56(293):169–171, March 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282346.

M:1960:BRBa

[M.60a] W. G. M. Book review: [untitled]. Journal of the American Statistical Association, 55(292):785, December 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281630.

M:1960:BRBb

[M.60b] W. G. M. Book review: [untitled]. Journal of the American Statistical Association, 55(292):786, December 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281631.

M:1961:BRBa

[M.61a] W. G. M. Book review: [untitled]. Journal of the American Statistical Association, 56(293):206–207, March 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282376.

M:1961:BRBb

[M.61b] W. G. M. Book review: [untitled]. Journal of the American Statistical Association, 56(293):215, March 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282383.

M:1961:BRBc

[M.61c] W. G. M. Book review: [untitled]. Journal of the American Statistical Association, 56(293):216, March 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282386.

M:1961:BRBd

[M.61d] W. G. M. Book review: [untitled]. Journal of the American Statistical Association, 56(293):216–217, March 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282387.

M:1961:BRBe

[M.61e] W. G. M. Book review: [untitled]. Journal of the American Statistical Association, 56(293):217–218, March 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282389.

M:1961:BRBf

[M.61f] W. G. M. Book review: [untitled]. Journal of the American Statistical Association, 56(295):776, September 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282126.

M:1961:BRBg

[M.61g] W. G. M. Book review: [untitled]. Journal of the American Statistical Association, 56(296):1028–1029, December 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282037.

M:1962:BRBa

[M.62a] W. G. M. Book review: [untitled]. Journal of the American Statistical Association, 57(297):264, March 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282491.

M:1962:BRBb

[M.62b] W. G. M. Book review: [untitled]. Journal of the American Statistical Association, 57(299):723, September 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282433.

M:1966:BRBa

[M.66a] A. M. Book review: [untitled]. Journal of the American Statistical Association, 61(313):296, March 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283085.

M:1966:BRBb

[M.66b] A. M. Book review: [untitled]. Journal of the American Statistical Association, 61(313):296, March 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283086.

M:1967:BRBc

[M.67a] A. M. Book review: Finite Graphs and Networks: An Introduction with Applications, by Robert G. Busacker; Thomas L. Saaty. Journal of the American Statistical Association, 62(317):309, March 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2282946.

M:1967:BRBd

[M.67b] A. M. Book review: Omnitab, A Computer Program for Statistical and Numerical Analysis, by Joseph Hilsenrath. Journal of the American Statistical Association, 62(317):309–310, March 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282947.

M:1967:BRBb

[M.67c] A. M. Book review: Programming, Games and Transportation Networks, by Claude Berge; A. Ghouila-Houri. Journal of the American Statistical Association, 62(317):309, March 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282945.

M:1967:BRBe

[M.67d] C. M. Book review: [untitled]. Journal of the American Statistical Association, 62(319):1093, September 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283711.

M:1967:BRBa

[M.67e] W. G. M. Book review: Bibliography of Statistical Literature, 1940– 1949, by M. G. Kendall; A. G. Doig. Journal of the American Statistical Association, 62(317):309, March 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282944.

M:1969:BRB

[M.69] C. L. M. Book review: [untitled]. Journal of the American Statistical Association, 64(327):1099, September 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283501.

MacKinnon:1964:TBS

[Mac64] William J. MacKinnon. Table for both the sign test and distribution-free confidence intervals of the median for sample sizes to 1,000. Journal of the American Statistical Association, 59(307): 935–956, September 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2283112.

Madow:1960:BRBa

[Mad60a] William G. Madow. Book review: [untitled]. Journal of the American Statistical Association, 55(290):394, June 1960. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2281767.

Madow:1960:BRBb

[Mad60b] William G. Madow. Book review: [untitled]. Journal of the American Statistical Association, 55(292):780, December 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281624.

Madansky:1962:MLC

[Mad62] Albert Madansky. More on length of confidence intervals. Journal of the American Statistical Association, 57(299):586–589, September 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2282396.

Madansky:1963:THC

[Mad63] Albert Madansky. Tests of homogeneity for correlated samples. Journal of the American Statistical Association, 58(301): 97–119, March 1963. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2282957.

Maddala:1968:BRB

[Mad68] G. S. Maddala. Book review: [untitled]. Journal of the American Statistical Association, 63(321):380–381, March 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283866.

Madansky:1969:BRB

[Mad69a] Albert Madansky. Book review: [untitled]. Journal of the American Statistical Association, 64(328):1675–1676, December 1969. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2286107.

Maddala:1969:BRB

[Mad69b] G. S. Maddala. Book review: [untitled]. Journal of the American Statistical Association, 64(328):1674–1675, December 1969. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2286106.

Maeshiro:1966:SMR

[Mae66] Asatoshi Maeshiro. A simple mathematical relationship among k-class estimators. Journal of the American Statistical Association, 61(314):368–374, June 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2282825.

Mahamunulu:1967:NRM

[Mah67] D. M. Mahamunulu. A note on regression in the multivariate Poisson distribution. Journal of the American Statistical Association, 62(317):251–258, March 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2282927.

Maisel:1960:CRC

[Mai60] Sherman J. Maisel. Changes in the rate and components of household formation. Journal of the American Statistical Association, 55(290):268–283, June 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2281740.

Mainland:1961:BRB

[Mai61a] Donald Mainland. Book review: [untitled]. Journal of the American Statistical Association, 56(295):758–759, September 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282108.

Mair:1961:BRB

[Mai61b] George F. Mair. Book review: [untitled]. Journal of the American Statistical Association, 56(296):1016–1017, December 1961. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282021.

Maisel:1963:BRB

[Mai63] Sherman J. Maisel. Book review: [untitled]. Journal of the American Statistical Association, 58(301):259–260, March 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282978.

Maisel:1966:BC

[Mai66] H. Maisel. Best k of 2k − 1 comparisons. Journal of the American Statistical Association, 61(314):329–344, June 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282822.

Mallios:1967:SRA

[Mal67a] W. S. Mallios. A structural regression approach to covariance analysis when the covariable is uncontrolled. Journal of the American Statistical Association, 62(319):1037–1049, September 1967. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283692.

Mallows:1967:BRB

[Mal67b] C. L. Mallows. Book review: [untitled]. Journal of the American Statistical Association, 62(319):1075–1076, September 1967. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283698.

Mandel:1960:BRB

[Man60] John Mandel. Book review: [untitled]. Journal of the American Statistical Association, 55(289):217, March 1960. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2282189.

Mandel:1961:NAT

[Man61] John Mandel. Non-additivity in two-way analysis of variance. Journal of the American Statistical Association, 56(296):878–888, December 1961. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2282001.

Mansfield:1963:BRB

[Man63a] Edwin Mansfield. Book review: [untitled]. Journal of the American Statistical Association, 58(304):1166–1167, December 1963. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283354.

Mantel:1963:CST

[Man63b] Nathan Mantel. Chi-square tests with one degree of freedom; extensions of the Mantel–Haenszel procedure. Journal of the American Statistical Association, 58(303):690–700, September 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282717.

Manderscheid:1964:BRB

[Man64a] Lester V. Manderscheid. Book review: [untitled]. Journal of the American Statistical Association, 59(308):1313–1314, December 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2282661.

Mansfield:1964:BRB

[Man64b] Edwin Mansfield. Book review: [untitled]. Journal of the American Statistical Association, 59(306):605–606, June 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283015.

Mandrekar:1969:BRB

[Man69a] V. Mandrekar. Book review: [untitled]. Journal of the American Statistical Association, 64(326):672–673, June 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283655.

Mann:1969:ETO

[Man69b] Nancy R. Mann. Exact three-order-statistic confidence bounds on reliable life for a Weibull model with progressive censoring. Journal of the American Statistical Association, 64(325):306–315, March 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2283740.

Marcus:1960:BRB

[Mar60a] Edward Marcus. Book review: [untitled]. Journal of the American Statistical Association, 55(289):228–229, March 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282194.

Margolis:1960:BRB

[Mar60b] Julius Margolis. Book review: [untitled]. Journal of the American Statistical Association, 55(291):601–602, September 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281929.

Markham:1960:BRB

[Mar60c] Jesse W. Markham. Book review: [untitled]. Journal of the American Statistical Association, 55(292):763–764, December 1960. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281610.

Martin:1960:BRB

[Mar60d] Margaret Martin. Book review: [untitled]. Journal of the American Statistical Association, 55(291):615–616, September 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281941.

Martinez:1962:MDR

[Mar62a] Charles Martinez. A method to determine the reliability of telemetry systems reports. Journal of the American Statistical Association, 57(299):686–689, September 1962. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282406.

Marx:1962:BRB

[Mar62b] Daniel Marx, Jr. Book review: [untitled]. Journal of the American Statistical Association, 57(300):940–941, December 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281839.

Marks:1963:BRB

[Mar63a] Eli S. Marks. Book review: [untitled]. Journal of the American Statistical Association, 58(304):1193–1195, December 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283375.

Marschak:1963:BRB

[Mar63b] Thomas A. Marschak. Book review: [untitled]. Journal of the American Statistical Association, 58(303):851–853, September 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2282747.

Marsaglia:1964:CMC

[Mar64] George Marsaglia. Conditional means and covariances of normal variables with singular covariance matrix. Journal of the American Statistical Association, 59(308):1203–1204, December 1964. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282635.

Marc:1965:CCM

[Mar65a] Nerlove Marc. Corrigenda: A Comparison of a Modified “Hannan” and the BLS Seasonal Adjustment Filters. Journal of the American Statistical Association, 60(312):1250–1251, December 1965. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283452. See [Ner65b].

Markowitz:1965:BRB

[Mar65b] Harry Markowitz. Book review: [untitled]. Journal of the American Statistical Association, 60(309):381, March 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283173.

Marsaglia:1965:RNV

[Mar65c] George Marsaglia. Ratios of normal variables and ratios of sums of uniform variables. Journal of the American Statistical Association, 60(309):193–204, March 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283145.

Massy:1962:BRB

[Mas62] William F. Massy. Book review: [untitled]. Journal of the American Statistical Association, 57(300):943–944, December 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281842.

Masucci:1963:BRB

[Mas63] Robert H. Masucci. Book review: [untitled]. Journal of the American Statistical Association, 58(301):254–255, March 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282974.

Massell:1964:BRB

[Mas64] B. F. Massell. Book review: [untitled]. Journal of the American Statistical Association, 59(308):1308–1310, December 1964. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282657.

Massy:1965:PCR

[Mas65] William F. Massy. Principal components regression in exploratory statistical research. Journal of the American Statistical Association, 60(309):234–256, March 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283149.

Masani:1966:BRB

[Mas66] P. Masani. Book review: [untitled]. Journal of the American Statistical Association, 61(313):268–273, March 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283065.

Matuszewski:1962:CSP

[Mat62a] T. I. Matuszewski. Corrigenda: Some Properties of Pascal Distribution for Finite Population. Journal of the American Statistical Association, 57(300):919, December 1962. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281825. See [Mat62b].

Matuszewski:1962:SPP

[Mat62b] T. I. Matuszewski. Some properties of Pascal distribution for finite population. Journal of the American Statistical Association, 57(297):172–174, March 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor. org/stable/2282446. See corrigenda [Mat62a].

Maxwell:1963:BRB

[Max63] A. E. Maxwell. Book review: [untitled]. Journal of the American Statistical Association, 58(302):577–578, June 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283300.

Mayer:1960:BRB

[May60] Thomas Mayer. Book review: [untitled]. Journal of the American Statistical Association, 55(292):781–782, December 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281626.

Maynes:1961:BRB

[May61] E. Scott Maynes. Book review: [untitled]. Journal of the American Statistical Association, 56(296):1018–1021, December 1961. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282022.

Maynes:1962:BRB

[May62] E. Scott Maynes. Book review: [untitled]. Journal of the American Statistical Association, 57(299):710–711, September 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282419.

Maynes:1963:BRB

[May63] E. Scott Maynes. Book review: [untitled]. Journal of the American Statistical Association, 58(302):569–572, June 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283295.

Maynes:1968:MRE

[May68] E. Scott Maynes. Minimizing response errors in financial data: The possibilities. Journal of the American Statistical Association, 63(321):214–227, March 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283842.

Mazur:1963:DME

[Maz63] D. Peter Mazur. A demographic model for estimating age-order specific fertility rates. Journal of the American Statistical Association, 58(303):774–788, September 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282725.

Mawaziny:1967:CLR

[MB67a] A. H. El Mawaziny and R. J. Buehler. Confidence limits for the reliability of series systems. Journal of the American Statistical Association, 62(320):1452–1459, December 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283789.

Morse:1967:CES

[MB67b] Pamela M. Morse and A. Bickle. The combination of estimates from similar experiments, allowing for inter-experiment variation. Journal of the American Statistical Association, 62(317): 241–250, March 1967. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2282926.

Meng:1966:PCS

[MC66] Rosa C. Meng and Douglas G. Chapman. The power of chi square tests for contingency tables. Journal of the American Statistical Association, 61(316):965–975, December 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283192.

McCarthy:1965:SSD

[McC65a] Philip J. McCarthy. Stratified sampling and distribution-free confidence intervals for a median. Journal of the American Statistical Association, 60(311):772–783, September 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283244.

McCornack:1965:ETW

[McC65b] Robert L. McCornack. Extended tables of the Wilcoxon matched pair signed rank statistic. Journal of the American Statistical Association, 60(311):864–871, September 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283253.

McCarthy:1968:BRB

[McC68] Philip J. McCarthy. Book review: [untitled]. Journal of the American Statistical Association, 63(323):1058–1059, September 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283904.

McCall:1969:BRB

[McC69a] J. J. McCall. Book review: [untitled]. Journal of the American Statistical Association, 64(328):1668, December 1969. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2286099.

McClain:1969:BRBa

[McC69b] John McClain. Book review: [untitled]. Journal of the American Statistical Association, 64(327):1097, September 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283498.

McClain:1969:BRBb

[McC69c] John McClain. Book review: [untitled]. Journal of the American Statistical Association, 64(327):1098, September 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283499.

McDonough:1969:BRB

[McD69] R. N. McDonough. Book review: [untitled]. Journal of the American Statistical Association, 64(326):680–681, June 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283663.

McElroy:1967:NSC

[McE67] F. W. McElroy. A necessary and sufficient condition that ordinary least-squares estimators be best linear unbiased. Journal of the American Statistical Association, 62(320):1302–1304, December 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2283779.

McFadden:1967:BRB

[McF67] Daniel McFadden. Book review: Manufacturing Production Functions in the U. S., 1957: An Interindustry and Interstate Comparison of Productivity, by George H. Hildebrand; Ta-Chung Liu. Journal of the American Statistical Association, 62(317):295– 300, March 1967. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2282936.

McGilchrist:1969:DDE

[McG69] C. A. McGilchrist. Discrete distribution estimators from the recurrence equation for probabilities. Journal of the American Statistical Association, 64(326):602–609, June 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283643.

McKenzie:1960:BRB

[McK60] Lionel McKenzie. Book review: [untitled]. Journal of the American Statistical Association, 55(290):394–396, June 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281769.

McKinnon:1961:BRB

[McK61] Ronald I. McKinnon. Book review: [untitled]. Journal of the American Statistical Association, 56(293):201–202, March 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282370.

McKinnon:1962:BRBa

[McK62a] Ronald I. McKinnon. Book review: [untitled]. Journal of the American Statistical Association, 57(297):249–250, March 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282480.

McKinnon:1962:BRBb

[McK62b] Ronald I. McKinnon. Book review: [untitled]. Journal of the American Statistical Association, 57(299):711–712, September 1962. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282420.

McKinnon:1962:BRBc

[McK62c] Ronald I. McKinnon. Book review: [untitled]. Journal of the American Statistical Association, 57(300):938–940, December 1962. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281838.

McKenzie:1964:BRB

[McK64] Lionel McKenzie. Book review: [untitled]. Journal of the American Statistical Association, 59(307):968–969, September 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283117.

McNeil:1967:ELD

[McN67] D. R. McNeil. Efficiency loss due to grouping in distribution-free tests. Journal of the American Statistical Association, 62(319): 954–965, September 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2283684.

Meeter:1967:BRB

[Mee67] Duane Meeter. Book review: [untitled]. Journal of the American Statistical Association, 62(320):1504–1506, December 1967. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283800.

Meissner:1960:BRB

[Mei60] Frank Meissner. Book review: [untitled]. Journal of the American Statistical Association, 55(292):776, December 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281621.

Messick:1967:BRB

[Mes67] David M. Messick. Book review: [untitled]. Journal of the American Statistical Association, 62(318):704–707, June 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284004.

Meyer:1961:BRB

[Mey61] Paul Meyer. Book review: [untitled]. Journal of the American Statistical Association, 56(295):771, September 1961. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2282121.

Meyer:1962:BRB

[Mey62] John R. Meyer. Book review: [untitled]. Journal of the American Statistical Association, 57(297):227–230, March 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282465.

Meyer:1967:MLE

[Mey67] Paul L. Meyer. The maximum likelihood estimate of the noncentrality parameter of a non-central χ^{2} variate. Journal of the American Statistical Association, 62(320):1258–1264, December 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2283775.

Myers:1963:DNC

[MF63] James H. Myers and Edward W. Forgy. The development of numerical credit evaluation systems. Journal of the American Statistical Association, 58(303):799–806, September 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282727.

McGuire:1968:EIL

[MFLR68] Timothy W. McGuire, John U. Farley, Robert E. Lucas, Jr., and L. Winston Ring. Estimation and inference for linear models in which subsets of the dependent variable are constrained. Journal of the American Statistical Association, 63(324):1201–1213, December 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2285877.

Mehta:1969:CUE

[MG69] J. S. Mehta and John Gurland. Combinations of unbiased estimators of the mean which consider inequality of unknown variances. Journal of the American Statistical Association, 64(327): 1042–1055, September 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2283484.

Mosteller:1963:BRBa

[MH63] Frederick Mosteller and E. A. Hammel. Book review: [untitled]. Journal of the American Statistical Association, 58(303):835–836, September 1963. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2282735.

Mihram:1967:BWT

[MH67] George A. Mihram and Robert A. Hultquist. A bivariate warningtime/failure-time distribution. Journal of the American Statistical Association, 62(318):589–599, June 1967. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283985.

Mickey:1959:SFP

[Mic59] M. R. Mickey. Some finite population unbiased ratio and regression estimators. Journal of the American Statistical Association, 54 (287):594–612, September 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor. org/stable/2282540. See corrigenda [Mic60].

Mickey:1960:CSF

[Mic60] M. R. Mickey. Corrigenda: Some Finite Population Unbiased Ratio and Regression Estimators. Journal of the American Statistical Association, 55(292):755, December 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281602. See [Mic59].

Miller:1960:EFL

[Mil60] Rupert G. Miller, Jr. Early failures in life testing. Journal of the American Statistical Association, 55(291):491–502, September 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2281910.

Miller:1961:BRB

[Mil61a] Roger Miller. Book review: [untitled]. Journal of the American Statistical Association, 56(293):192–193, March 1961. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2282364.

Milliman:1961:BRB

[Mil61b] J. W. Milliman. Book review: [untitled]. Journal of the American Statistical Association, 56(295):769, September 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282119.

Mills:1962:BRBa

[Mil62a] Edwin S. Mills. Book review: [untitled]. Journal of the American Statistical Association, 57(297):222–224, March 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282461.

Mills:1962:BRBb

[Mil62b] Edwin S. Mills. Book review: [untitled]. Journal of the American Statistical Association, 57(299):709–710, September 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282418.

Miller:1963:BRBa

[Mil63a] Irwin Miller. Book review: [untitled]. Journal of the American Statistical Association, 58(301):263–265, March 1963. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2282981.

Miller:1963:BRBb

[Mil63b] Robert L. Miller. Book review: [untitled]. Journal of the American Statistical Association, 58(304):1178–1179, December 1963. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283362.

Milder:1964:RSN

[Mil64a] D. Michael Milder. Regression with systematic noise. Journal of the American Statistical Association, 59(306):422–428, June 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282996.

Milton:1964:ETC

[Mil64b] Roy C. Milton. An extended table of critical values for the Mann– Whitney (Wilcoxon) two-sample statistic. Journal of the American Statistical Association, 59(307):925–934, September 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283111.

Miller:1965:BRB

[Mil65] Rupert G. Miller, Jr. Book review: [untitled]. Journal of the American Statistical Association, 60(312):1218–1220, December 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283423.

Miller:1967:MEP

[Mil67] Ann R. Miller. The migration of employed persons to and from metropolitan areas of the United States. Journal of the American Statistical Association, 62(320):1418–1432, December 1967. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283787.

Milch:1969:BRB

[Mil69] Paul R. Milch. Book review: [untitled]. Journal of the American Statistical Association, 64(326):677, June 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283659.

Minton:1968:BRB

[Min68] Paul D. Minton. Book review: [untitled]. Journal of the American Statistical Association, 63(323):1064, September 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283910.

Minton:1969:ICE

[Min69a] George Minton. Inspection and correction error in data processing. Journal of the American Statistical Association, 64(328):1256– 1275, December 1969. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2286066.

Minton:1969:BRB

[Min69b] Paul D. Minton. Book review: [untitled]. Journal of the American Statistical Association, 64(328):1681–1682, December 1969. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2286114.

McHugh:1968:NVE

[MM68] Richard B. McHugh and Paul W. Mielke, Jr. Negative variance estimates and statistical dependence in nested sampling. Journal of the American Statistical Association, 63(323):1000–1003, September 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2283890.

Marshall:1967:MED

[MO67] Albert W. Marshall and Ingram Olkin. A multivariate exponential distribution. Journal of the American Statistical Association, 62(317):30–44, March 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2282907.

Mode:1963:PC

[Mod63] Elmer B. Mode. Probability and criminalistics. Journal of the American Statistical Association, 58(303):628–640, September 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2282710.

Mohring:1961:BRB

[Moh61] Herbert D. Mohring. Book review: [untitled]. Journal of the American Statistical Association, 56(294):458–460, June 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282304.

Montalbano:1966:BRB

[Mon66] Michael S. Montalbano. Book review: [untitled]. Journal of the American Statistical Association, 61(315):877, September 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282805.

Monroe:1967:BRB

[Mon67] Robert J. Monroe. Book review: [untitled]. Journal of the American Statistical Association, 62(319):1090–1091, September 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283708.

Moore:1964:BRB

[Moo64] P. G. Moore. Book review: [untitled]. Journal of the American Statistical Association, 59(308):1317–1318, December 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282664.

Moore:1969:FST

[Moo69] Geoffrey H. Moore. Forecasting short-term economic change. Journal of the American Statistical Association, 64(325):1–22, March 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2283717.

Moranda:1959:CEC

[Mor59] P. B. Moranda. Comparison of estimates of circular probable error. Journal of the American Statistical Association, 54(288):794– 800, December 1959. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/

2282503. See corrigenda [Mor60a].

Moranda:1960:CCE

[Mor60a] P. B. Moranda. Corrigenda: Comparison of Estimates of Circular Probable Error. Journal of the American Statistical Association, 55(292):755–756, December 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281603. See [Mor59].

Moranda:1960:EBE

[Mor60b] P. B. Moranda. Effect of bias on estimates of the circular probable error. Journal of the American Statistical Association, 55 (292):732–735, December 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281596.

Morag:1961:BRB

[Mor61a] Amotz Morag. Book review: [untitled]. Journal of the American Statistical Association, 56(295):751–752, September 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282103.

Morrison:1961:BRB

[Mor61b] Donald F. Morrison. Book review: [untitled]. Journal of the American Statistical Association, 56(294):444–446, June 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282293.

Morton:1961:BRB

[Mor61c] Newton E. Morton. Book review: [untitled]. Journal of the American Statistical Association, 56(295):760–761, September 1961. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282111.

Moroney:1966:BRB

[Mor66a] M. J. Moroney. Book review: [untitled]. Journal of the American Statistical Association, 61(313):285–286, March 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283076.

Morrison:1966:BRB

[Mor66b] Donald G. Morrison. Book review: [untitled]. Journal of the American Statistical Association, 61(313):286–288, March 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283077.

Morrison:1968:BRB

[Mor68] Donald F. Morrison. Book review: [untitled]. Journal of the American Statistical Association, 63(324):1528–1529, December 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2285903.

Moses:1961:BRB

[Mos61] Lincoln E. Moses. Book review: [untitled]. Journal of the American Statistical Association, 56(293):179–180, March 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282354.

Mosbaek:1962:BRB

[Mos62] Ernest J. Mosbaek. Book review: [untitled]. Journal of the American Statistical Association, 57(297):225–226, March 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282463.

Moses:1963:BRB

[Mos63a] Lincoln E. Moses. Book review: [untitled]. Journal of the American Statistical Association, 58(303):840–842, September 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282740.

Mosteller:1963:BRBb

[Mos63b] Frederick Mosteller. Book review: [untitled]. Journal of the American Statistical Association, 58(304):1180–1181, December 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283364.

Moses:1964:OSL

[Mos64] Lincoln E. Moses. One sample limits of some two-sample rank tests. Journal of the American Statistical Association, 59(307): 645–651, September 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2283090.

Mosteller:1968:AEC

[Mos68] Frederick Mosteller. Association and estimation in contingency tables. Journal of the American Statistical Association, 63(321): 1–28, March 1968. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2283825.

Mantel:1966:CLB

[MP66a] Nathan Mantel and Bernard S. Pasternack. Corrigenda: Light Bulb Statistics. Journal of the American Statistical Association, 61(316):1248, December 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283234. See [MP66b].

Mantel:1966:LBS

[MP66b] Nathan Mantel and Bernard S. Pasternack. Light bulb statistics. Journal of the American Statistical Association, 61(315): 633–639, September 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/ stable/2282776. See corrigenda [MP66a].

Murthy:1961:RRM

[MS61] M. N. Murthy and V. K. Sethi. Randomized rounded-off multipliers in sampling theory. Journal of the American Statistical Association, 56(294):328–334, June 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2282258.

Morgan:1963:PAS

[MS63] James N. Morgan and John A. Sonquist. Problems in the analysis of survey data, and a proposal. Journal of the American Statistical Association, 58(302):415–434, June 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283276.

Mielke:1965:CTI

[MS65] Paul W. Mielke, Jr. and M. M. Siddiqui. A combinatorial test for independence of dichotomous responses. Journal of the American Statistical Association, 60(310):437–441, June 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282681.

Miller:1962:LCA

[MSD62] C. R. Miller, G. Sabagh, and H. F. Dingman. Latent class analysis and differential mortality. Journal of the American Statistical Association, 57(298):430–438, June 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281649.

Mellinger:1965:MMA

[MSGM65] Glen D. Mellinger, David L. Sylwester, William R. Gaffey, and Dean I. Manheimer. A mathematical model with applications to a study of accident repeatedness among children. Journal of the American Statistical Association, 60(312):1046–1059, December 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2283404.

Mueller:1960:BRB

[Mue60] Eva Mueller. Book review: [untitled]. Journal of the American Statistical Association, 55(291):597–599, September 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281927.

Mueller:1963:TYC

[Mue63] Eva Mueller. Ten years of consumer attitude surveys: Their forecasting record. Journal of the American Statistical Association, 58(304):899–917, December 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283322.

Mueller:1964:BRB

[Mue64] M. G. Mueller. Book review: [untitled]. Journal of the American Statistical Association, 59(306):610–612, June 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283019.

Murthy:1963:SRA

[Mur63] M. N. Murthy. Some recent advances in sampling theory. Journal of the American Statistical Association, 58(303):737–755, September 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2282722.

Murphy:1966:BRB

[Mur66a] James L. Murphy. Book review: [untitled]. Journal of the American Statistical Association, 61(313):288–289, March 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283078.

Murray:1966:BRB

[Mur66b] George R. Murray, Jr. Book review: [untitled]. Journal of the American Statistical Association, 61(315):878, September 1966.

CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282807.

Mustafi:1967:ITH

[Mus67] Chandan K. Mustafi. On iterated tests of hypotheses. Journal of the American Statistical Association, 62(318):520–524, June 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283979.

Musgrave:1969:MPC

[Mus69a] John C. Musgrave. The measurement of price changes in construction. Journal of the American Statistical Association, 64 (327):771–786, September 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283460.

Mustafi:1969:RRD

[Mus69b] Chandan Kumar Mustafi. A recurrence relation for distribution functions of order statistics from bivariate distributions. Journal of the American Statistical Association, 64(326):600–601, June 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283642.

Muth:1960:BRBb

[Mut60a] John F. Muth. Book review: [untitled]. Journal of the American Statistical Association, 55(292):782–783, December 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281627.

Muth:1960:OPE

[Mut60b] John F. Muth. Optimal properties of exponentially weighted forecasts. Journal of the American Statistical Association, 55(290): 299–306, June 1960. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/

2281742. See corrigenda [Mut62a].

Muth:1960:BRBa

[Mut60c] Richard F. Muth. Book review: [untitled]. Journal of the American Statistical Association, 55(292):768–769, December 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281615.

Muth:1962:COP

[Mut62a] John F. Muth. Corrigenda: Optimal Properties of Exponentially Weighted Forecasts. Journal of the American Statistical Association, 57(300):919–920, December 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281826. See [Mut60b].

Muth:1962:BRB

[Mut62b] Richard F. Muth. Book review: [untitled]. Journal of the American Statistical Association, 57(298):512–513, June 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281663.

Mosteller:1963:IAP

[MW63] Frederick Mosteller and David L. Wallace. Inference in an authorship problem. Journal of the American Statistical Association, 58(302):275–309, June 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283270.

Margolin:1967:EMO

[MW67a] Barry H. Margolin and Herbert S. Winokur, Jr. Exact moments of the order statistics of the geometric distribution and their relation to inverse sampling and reliability of redundant systems. Journal of the American Statistical Association, 62(319):915–925, September 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2283679.

McCulloch:1967:LTR

[MW67b] A. J. McCulloch and John E. Walsh. Life-testing results based on a few heterogeneous lognormal observations. Journal of the American Statistical Association, 62(317):45–47, March 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282908.

Myers:1966:BRB

[Mye66] Robert J. Myers. Book review: [untitled]. Journal of the American Statistical Association, 61(313):293–295, March 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283082.

N:1961:BRBa

[N.61a] M. N. Book review: [untitled]. Journal of the American Statistical Association, 56(294):468–469, June 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282316.

N:1961:BRBb

[N.61b] M. N. Book review: [untitled]. Journal of the American Statistical Association, 56(294):469, June 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282319.

N:1961:BRBc

[N.61c] M. N. Book review: [untitled]. Journal of the American Statistical Association, 56(294):469–470, June 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282320.

N:1961:BRBd

[N.61d] M. N. Book review: [untitled]. Journal of the American Statistical Association, 56(295):776, September 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282128.

N:1961:BRBe

[N.61e] M. N. Book review: [untitled]. Journal of the American Statistical Association, 56(295):776–777, September 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282129.

N:1961:BRBf

[N.61f] M. N. Book review: [untitled]. Journal of the American Statistical Association, 56(295):777, September 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282130.

N:1961:BRBg

[N.61g] M. N. Book review: [untitled]. Journal of the American Statistical Association, 56(296):1029, December 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282038.

N:1962:BRB

[N.62] M. N. Book review: [untitled]. Journal of the American Statistical Association, 57(297):263–264, March 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282489.

Nagnur:1969:LHN

[Nag69] B. N. Nagnur. Lamst and the hypotheses of no three factor interaction in contingency tables. Journal of the American Statistical Association, 64(325):207–215, March 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283729.

Nair:1964:NCD

[Nai64] C. Ramankutty Nair. A new class of designs. Journal of the American Statistical Association, 59(307):817–833, September 1964. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283100. See corrigenda

[Ram65].

Nair:1967:SBP

[Nai67] C. Ramankutty Nair. Sequences balanced for pairs of residual effects. Journal of the American Statistical Association, 62(317): 205–225, March 1967. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2282924.

Naik:1969:EPT

[Nai69] Umesh D. Naik. The equal probability test and its applications to some simultaneous inference problems. Journal of the American Statistical Association, 64(327):986–998, September 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283480.

Nathan:1967:OPR

[Nat67] Gad Nathan. Outcome probabilities for a record matching process with complete invariant information. Journal of the American Statistical Association, 62(318):454–469, June 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283974.

Nath:1968:PMF

[Nat68] S. N. Nath. On product moments from a finite universe. Journal of the American Statistical Association, 63(322):535–541, June 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2284024.

Nath:1969:MRP

[Nat69] S. N. Nath. More results on product moments from a finite universe. Journal of the American Statistical Association, 64(327): 864–869, September 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2283467.

Naus:1965:DSM

[Nau65] Joseph I. Naus. The distribution of the size of the maximum cluster of points on a line. Journal of the American Statistical Association, 60(310):532–538, June 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2282688.

Naus:1966:SPE

[Nau66] J. I. Naus. Some probabilities, expectations and variances for the size of largest clusters and smallest intervals. Journal of the American Statistical Association, 61(316):1191–1199, December 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283208.

Naus:1969:DLS

[Nau69] J. I. Naus. The distribution of the logarithm of the sum of two log-normal variates. Journal of the American Statistical Association, 64(326):655–659, June 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283649.

Naylor:1967:CSE

[NBS67] Thomas H. Naylor, Donald S. Burdick, and W. Earl Sasser. Computer simulation experiments with economic systems: The problem of experimental design. Journal of the American Statistical Association, 62(320):1315–1337, December 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283781.

Neave:1966:DTQ

[Nea66] Henry R. Neave. A development of Tukey’s quick test of location. Journal of the American Statistical Association, 61(316): 949–964, December 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/ stable/2283191. See corrigenda [Nea67].

Neave:1967:CDT

[Nea67] Henry R. Neave. Corrigenda: A Development of Tukey’s Quick Test of Location. Journal of the American Statistical Association, 62(320):1522, December 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283820. See [Nea66].

Neisser:1960:BRB

[Nei60] Hans Neisser. Book review: [untitled]. Journal of the American Statistical Association, 55(289):220–223, March 1960. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2282192.

Neisser:1964:BRB

[Nei64] H. Neisser. Book review: [untitled]. Journal of the American Statistical Association, 59(305):297–301, March 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282889.

Nelson:1960:MGC

[Nel60] Ralph L. Nelson. Market growth, company diversification and product concentration 1947–1954. Journal of the American Statistical Association, 55(292):640–649, December 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281587.

Nelson:1961:BRB

[Nel61] Richard R. Nelson. Book review: [untitled]. Journal of the American Statistical Association, 56(294):423–424, June 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282278.

Nerlove:1964:BRB

[Ner64] Marc Nerlove. Book review: [untitled]. Journal of the American Statistical Association, 59(308):1311–1312, December 1964. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282659.

Nerlove:1965:BRB

[Ner65a] Marc Nerlove. Book review: [untitled]. Journal of the American Statistical Association, 60(312):1239–1240, December 1965. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283435.

Nerlove:1965:CMH

[Ner65b] Marc Nerlove. A comparison of a modified “Hannan” and the BLS seasonal adjustment filters. Journal of the American Statistical Association, 60(310):442–491, June 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www. jstor.org/stable/2282682. See corrections [Mar65a, Ner66].

Nerlove:1966:CCM

[Ner66] Marc Nerlove. Corrigenda: A Comparison of a Modified “Hannan” and the BLS Seasonal Adjustment Filters. Journal of the American Statistical Association, 61(316):1248, December 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283235. See [Ner65b].

Neter:1962:BRB

[Net62] John Neter. Book review: [untitled]. Journal of the American Statistical Association, 57(297):240–241, March 1962. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2282472.

Nettheim:1965:FME

[Net65] Nigel F. Nettheim. Fourier methods for evolving seasonal patterns. Journal of the American Statistical Association, 60(310):492–502, June 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2282683.

Neter:1966:BRB

[Net66] John Neter. Book review: [untitled]. Journal of the American Statistical Association, 61(313):273–275, March 1966. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2283067.

Neter:1968:BRB

[Net68] John Neter. Book review: [untitled]. Journal of the American Statistical Association, 63(323):1045–1047, September 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283898.

Neuts:1964:BRB

[Neu64] Marcel F. Neuts. Book review: [untitled]. Journal of the American Statistical Association, 59(307):972–973, September 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283121.

Neuberger:1965:BRB

[Neu65] Egon Neuberger. Book review: [untitled]. Journal of the American Statistical Association, 60(309):379–381, March 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283172.

Neudecker:1969:NBE

[Neu69a] H. Neudecker. A note on BLUS estimation. Journal of the American Statistical Association, 64(327):949–952, September 1969. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283475.

Neudecker:1969:STM

[Neu69b] H. Neudecker. Some theorems on matrix differentiation with special reference to Kronecker matrix products. Journal of the American Statistical Association, 64(327):953–963, September 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283476.

Newcomb:1961:BRB

[New61] Robinson Newcomb. Book review: [untitled]. Journal of the American Statistical Association, 56(294):449–451, June 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282297.

Newman:1968:BRB

[New68] Thomas G. Newman. Book review: [untitled]. Journal of the American Statistical Association, 63(321):376–378, March 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283863.

Neyman:1960:ISN

[Ney60] Jerzy Neyman. Indeterminism in science and new demands on statisticians. Journal of the American Statistical Association, 55 (292):625–639, December 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281586.

Neyman:1969:SPS

[Ney69] Jerzy Neyman. Statistical problems in science. the symmetric test of a composite hypothesis. Journal of the American Statistical Association, 64(328):1154–1171, December 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286059.

Novick:1965:BAA

[NG65] Melvin R. Novick and James E. Grizzle. A Bayesian approach to the analysis of data from clinical trials. Journal of the American Statistical Association, 60(309):81–96, March 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283139.

Novick:1965:BIP

[NH65] Melvin R. Novick and W. J. Hall. A Bayesian indifference procedure. Journal of the American Statistical Association, 60(312): 1104–1117, December 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2283409.

Nicholson:1961:BRB

[Nic61] George E. Nicholson, Jr. Book review: [untitled]. Journal of the American Statistical Association, 56(294):412–413, June 1961. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282269.

Nicholson:1962:BRB

[Nic62] George E. Nicholson, Jr. Book review: [untitled]. Journal of the American Statistical Association, 57(299):719, September 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282427.

Nichols:1966:BRB

[Nic66] Eugene D. Nichols. Book review: A Programmed Introduction to Probability, by John R. Dixon. Journal of the American Statistical Association, 61(314):531, June 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282847.

Nielsen:1967:BRB

[Nie67] Norman R. Nielsen. Book review: [untitled]. Journal of the American Statistical Association, 62(318):710–711, June 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284008.

Nisselson:1961:BRBa

[Nis61a] Harold Nisselson. Book review: [untitled]. Journal of the American Statistical Association, 56(293):193–194, March 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282365.

Nisselson:1961:BRBb

[Nis61b] Harold Nisselson. Book review: [untitled]. Journal of the American Statistical Association, 56(293):195–197, March 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282367.

Naya:1969:AIT

[NM69] Seiji Naya and Theodore Morgan. The accuracy of international trade data: The case of Southeast Asian countries. Journal of the American Statistical Association, 64(326):452–467, June 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283632.

Neter:1965:EMM

[NMR65] John Neter, E. Scott Maynes, and R. Ramanathan. The effect of mismatching on the measurement of response error. Journal of the American Statistical Association, 60(312):1005–1027, December 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2283401.

Niederhoffer:1966:MMR

[NO66] Victor Niederhoffer and M. F. M. Osborne. Market making and reversal on the stock exchange. Journal of the American Statistical Association, 61(316):897–916, December 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283188.

Noether:1963:BRB

[Noe63a] Gottfried E. Noether. Book review: [untitled]. Journal of the American Statistical Association, 58(301):265–266, March 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282982.

Noether:1963:EWT

[Noe63b] Gottfried E. Noether. Efficiency of the Wilcoxon two-sample statistic for randomized blocks. Journal of the American Statistical Association, 58(304):894–898, December 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283321.

Noether:1964:BRB

[Noe64] Gottfried E. Noether. Book review: [untitled]. Journal of the American Statistical Association, 59(308):1314–1315, December 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282662.

Noether:1965:BRB

[Noe65] Gottfried E. Noether. Book review: [untitled]. Journal of the American Statistical Association, 60(309):368, March 1965. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283164.

Noether:1967:WCI

[Noe67] Gottfried E. Noether. Wilcoxon confidence intervals for location parameters in the discrete case. Journal of the American Statistical Association, 62(317):184–188, March 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282921.

Norman:1965:BRB

[Nor65] Warren T. Norman. Book review: [untitled]. Journal of the American Statistical Association, 60(312):1227–1228, December 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283427.

Nourse:1965:BRB

[Nou65] Hugh O. Nourse. Book review: [untitled]. Journal of the American Statistical Association, 60(309):374–375, March 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283169.

Neyman:1961:FCF

[NS61] Jerzy Neyman and Elizabeth L. Scott. Further comments on the “Final Report of the Advisory Committee on Weather Control”. Journal of the American Statistical Association, 56(295): 580–600, September 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2282081. See [Bro60a].

Neyman:1965:AOT

[NS65] Jerzy Neyman and Elizabeth L. Scott. Asymptotically optimal tests of composite hypotheses for randomized experiments with noncontrolled predictor variables. Journal of the American Statistical Association, 60(311):699–721, September 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283240.

Neter:1964:SRE

[NW64] John Neter and Joseph Waksberg. A study of response errors in expenditures data from household interviews. Journal of the American Statistical Association, 59(305):18–55, March 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282857.

Naylor:1967:CSM

[NWS67] Thomas H. Naylor, William H. Wallace, and W. Earl Sasser. A computer simulation model of the textile industry. Journal of the American Statistical Association, 62(320):1338–1364, December 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2283782.

O:1961:BRBa

[O.61a] I. O. Book review: [untitled]. Journal of the American Statistical Association, 56(294):468, June 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282314.

O:1961:BRBb

[O.61b] I. O. Book review: [untitled]. Journal of the American Statistical Association, 56(295):776, September 1961. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282127.

O:1961:BRBc

[O.61c] I. O. Book review: [untitled]. Journal of the American Statistical Association, 56(296):1029–1030, December 1961. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2282040.

O:1962:BRBa

[O.62a] I. O. Book review: [untitled]. Journal of the American Statistical Association, 57(298):524, June 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281675.

O:1962:BRBb

[O.62b] I. O. Book review: [untitled]. Journal of the American Statistical Association, 57(298):524–525, June 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281676.

O:1962:BRBc

[O.62c] I. O. Book review: [untitled]. Journal of the American Statistical Association, 57(298):527, June 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281685.

O:1962:BRBd

[O.62d] I. O. Book review: [untitled]. Journal of the American Statistical Association, 57(300):946, December 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281850.

O:1963:BRBa

[O.63a] I. O. Book review: [untitled]. Journal of the American Statistical Association, 58(304):1202–1203, December 1963. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2283386.

O:1963:BRBb

[O.63b] I. O. Book review: [untitled]. Journal of the American Statistical Association, 58(304):1204, December 1963. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283390.

O:1963:BRBc

[O.63c] I. O. Book review: [untitled]. Journal of the American Statistical Association, 58(304):1204, December 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283392.

O:1964:BRBa

[O.64a] I. O. Book review: Error Propagation for Difference Methods, by Peter Henrici. Journal of the American Statistical Association, 59(305):314, March 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2282896;   https://www.jstor.org/stable/2282897.

O:1964:BRBb

[O.64b] I. O. Book review: [untitled]. Journal of the American Statistical Association, 59(305):314, March 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282897.

O:1964:BRBc

[O.64c] I. O. Book review: [untitled]. Journal of the American Statistical Association, 59(305):315–316, March 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282900.

O:1964:BRBd

[O.64d] I. O. Book review: [untitled]. Journal of the American Statistical Association, 59(305):316, March 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282901.

O:1964:BRBe

[O.64e] I. O. Book review: [untitled]. Journal of the American Statistical Association, 59(307):980, September 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283128.

O:1964:BRBf

[O.64f] I. O. Book review: [untitled]. Journal of the American Statistical Association, 59(307):980, September 1964. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283129.

O:1964:BRBg

[O.64g] I. O. Book review: [untitled]. Journal of the American Statistical Association, 59(307):981, September 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283131.

O:1964:BRBh

[O.64h] I. O. Book review: [untitled]. Journal of the American Statistical Association, 59(308):1325–1326, December 1964. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2282673.

O:1964:BRBi

[O.64i] I. O. Book review: [untitled]. Journal of the American Statistical Association, 59(308):1326, December 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282675.

O:1966:BRBa

[O.66a] I. O. Book review: Group Theory, by Eugene Schenkman. Journal of the American Statistical Association, 61(314):535, June 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282851.

O:1966:BRBb

[O.66b] I. O. Book review: [untitled]. Journal of the American Statistical Association, 61(315):886, September 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282814.

O:1966:BRBc

[O.66c] I. O. Book review: [untitled]. Journal of the American Statistical Association, 61(315):886, September 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282815.

O:1967:BRBa

[O.67a] I. O. Book review: [untitled]. Journal of the American Statistical Association, 62(319):1093, September 1967. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283712.

O:1967:BRBb

[O.67b] I. O. Book review: [untitled]. Journal of the American Statistical Association, 62(320):1509, December 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283805.

O:1967:BRBc

[O.67c] I. O. Book review: [untitled]. Journal of the American Statistical Association, 62(320):1510, December 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283806.

O:1968:BRB

[O.68] D. B. O. Book review: [untitled]. Journal of the American Statistical Association, 63(323):1064–1065, September 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283911.

Owen:1964:NUC

[OCH64] D. B. Owen, K. J. Craswell, and D. L. Hanson. Nonparametric upper confidence bounds for Pr{Y < X} and confidence limits for Pr{Y < X} when X and Y are normal. Journal of the American Statistical Association, 59(307):906–924, September 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283110.

Odeh:1969:BRB

[Ode69] Robert E. Odeh. Book review: [untitled]. Journal of the American Statistical Association, 64(325):401, March 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283757.

Oehmke:1960:BRB

[Oeh60] Robert H. Oehmke. Book review: [untitled]. Journal of the American Statistical Association, 55(292):773–774, December 1960. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281619.

Odell:1966:NPG

[OF66a] P. L. Odell and A. H. Feiveson. A numerical procedure to generate a sample covariance matrix. Journal of the American Statistical Association, 61(313):199–203, March 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283054. See corrigenda [OF66b].

Odell:1966:CNP

[OF66b] P. L. Odell and A. L. Feiveson. Corrigenda: A Numerical Procedure to Generate a Sample Covariance Matrix. Journal of the American Statistical Association, 61(316):1248–1249, December 1966. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283236. See [OF66a].

Ogus:1969:NNB

[Ogu69] Jack L. Ogus. A note on the “necessary best estimator”. Journal of the American Statistical Association, 64(328):1350–1352, December 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2286073.

Oi:1962:BRB

[Oi62] Walter Y. Oi. Book review: [untitled]. Journal of the American Statistical Association, 57(297):224–225, March 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282462.

Oliver:1966:AML

[Oli66] F. R. Oliver. Aspects of maximum likelihood estimation of the logistic growth function. Journal of the American Statistical Association, 61(315):697–705, September 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282780.

Olshen:1968:BRB

[Ols68] Richard A. Olshen. Book review: [untitled]. Journal of the American Statistical Association, 63(324):1532–1533, December 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2285906.

OReagan:1969:RCC

[O’R69] Robert T. O’Reagan. Relative costs of computerized error inspection plans. Journal of the American Statistical Association, 64 (328):1245–1255, December 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286065.

Orey:1961:BRBa

[Ore61a] Steven Orey. Book review: [untitled]. Journal of the American Statistical Association, 56(296):1023, December 1961. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2282025.

Orey:1961:BRBb

[Ore61b] Steven Orey. Book review: [untitled]. Journal of the American Statistical Association, 56(296):1023, December 1961. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2282026.

Orey:1962:BRB

[Ore62] Steven Orey. Book review: [untitled]. Journal of the American Statistical Association, 57(297):263, March 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282488.

Orr:1963:BRBb

[Orr63a] Daniel Orr. Book review: [untitled]. Journal of the American Statistical Association, 58(304):1196–1197, December 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283378.

Orr:1963:BRBa

[Orr63b] Lloyd D. Orr. Book review: [untitled]. Journal of the American Statistical Association, 58(304):1187–1188, December 1963. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283371.

Osborne:1967:SQT

[Osb67] M. F. M. Osborne. Some quantitative tests for stock price generating models and trading folklore. Journal of the American Statistical Association, 62(318):321–340, June 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283966.

Ostle:1969:BRB

[Ost69] Bernard Ostle. Book review: [untitled]. Journal of the American Statistical Association, 64(327):1092–1094, September 1969. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283494.

Owen:1965:CPS

[Owe65a] D. B. Owen. Corrigenda: The Power of Student’s t-Test. Journal of the American Statistical Association, 60(312):1251, December 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2283453. See

[Owe65b].

Owen:1965:PST

[Owe65b] D. B. Owen. The power of Student’s t-test. Journal of the American Statistical Association, 60(309):320–333, March 1965. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283156. See corrigenda

[Owe65a].

Owen:1969:BRB

[Owe69] D. B. Owen. Book review: [untitled]. Journal of the American Statistical Association, 64(328):1684–1686, December 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286117.

P:1962:BRBa

[P.62a] J. W. P. Book review: [untitled]. Journal of the American Statistical Association, 57(298):525, June 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281677.

P:1962:BRBb

[P.62b] J. W. P. Book review: [untitled]. Journal of the American Statistical Association, 57(298):525, June 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281678.

P:1962:BRBc

[P.62c] J. W. P. Book review: [untitled]. Journal of the American Statistical Association, 57(298):525–526, June 1962. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281679.

P:1962:BRBd

[P.62d] J. W. P. Book review: [untitled]. Journal of the American Statistical Association, 57(298):526, June 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281680.

P:1962:BRBe

[P.62e] J. W. P. Book review: [untitled]. Journal of the American Statistical Association, 57(298):526, June 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281681.

P:1962:BRBf

[P.62f] J. W. P. Book review: [untitled]. Journal of the American Statistical Association, 57(298):526, June 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281682.

P:1962:BRBg

[P.62g] J. W. P. Book review: [untitled]. Journal of the American Statistical Association, 57(298):526–527, June 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281683.

P:1963:BRBa

[P.63a] J. W. P. Book review: [untitled]. Journal of the American Statistical Association, 58(301):268, March 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282984.

P:1963:BRBb

[P.63b] J. W. P. Book review: [untitled]. Journal of the American Statistical Association, 58(301):269, March 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282987.

P:1963:BRBc

[P.63c] J. W. P. Book review: [untitled]. Journal of the American Statistical Association, 58(301):269–270, March 1963. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282989.

P:1963:BRBd

[P.63d] J. W. P. Book review: [untitled]. Journal of the American Statistical Association, 58(304):1203, December 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283388.

P:1963:BRBe

[P.63e] J. W. P. Book review: [untitled]. Journal of the American Statistical Association, 58(304):1203–1204, December 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283389.

P:1964:BRB

[P.64] J. W. P. Book review: [untitled]. Journal of the American Statistical Association, 59(305):314–315, March 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282898.

P:1965:BRBa

[P.65a] J. W. P. Book review: [untitled]. Journal of the American Statistical Association, 60(312):1242, December 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283439.

P:1965:BRBb

[P.65b] J. W. P. Book review: [untitled]. Journal of the American Statistical Association, 60(312):1243, December 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283442.

P:1966:BRB

[P.66] J. W. P. Book review: [untitled]. Journal of the American Statistical Association, 61(315):886, September 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282816.

P:1967:BRB

[P.67] J. W. P. Book review: [untitled]. Journal of the American Statistical Association, 62(318):715, June 1967. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284011.

Pachares:1960:TCL

[Pac60] James Pachares. Tables of confidence limits for the binomial distribution. Journal of the American Statistical Association, 55(291): 521–533, September 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2281912.

Page:1963:OHM

[Pag63] Ellis Batten Page. Ordered hypotheses for multiple treatments: A significance test for linear ranks. Journal of the American Statistical Association, 58(301):216–230, March 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282965.

Page:1964:BRB

[Pag64] Ellis Batten Page. Book review: [untitled]. Journal of the American Statistical Association, 59(306):612–613, June 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283020.

Palda:1967:BRB

[Pal67] Kristian S. Palda. Book review: [untitled]. Journal of the American Statistical Association, 62(319):1078–1079, September 1967. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283700.

Palmour:1969:BRB

[Pal69] Vernon E. Palmour. Book review: [untitled]. Journal of the American Statistical Association, 64(328):1673, December 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286104.

Papi:1960:BRB

[Pap60] G. Ugo Papi. Book review: [untitled]. Journal of the American Statistical Association, 55(292):761–763, December 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281609.

Parzen:1961:BRBa

[Par61a] Emanuel Parzen. Book review: [untitled]. Journal of the American Statistical Association, 56(294):465–466, June 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282309.

Parzen:1961:BRBb

[Par61b] Emanuel Parzen. Book review: [untitled]. Journal of the American Statistical Association, 56(295):744–745, September 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282096.

Parzen:1962:BRB

[Par62] Emanuel Parzen. Book review: [untitled]. Journal of the American Statistical Association, 57(298):505–506, June 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281657.

Parzen:1964:BRBa

[Par64a] Emanuel Parzen. Book review: [untitled]. Journal of the American Statistical Association, 59(307):973–974, September 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283122.

Parzen:1964:BRBb

[Par64b] Emanuel Parzen. Book review: [untitled]. Journal of the American Statistical Association, 59(308):1305–1307, December 1964. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282655.

Parzen:1966:BRB

[Par66] Emanuel Parzen. Book review: [untitled]. Journal of the American Statistical Association, 61(315):868–869, September 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282798.

Parks:1967:EES

[Par67] Richard W. Parks. Efficient estimation of a system of regression equations when disturbances are both serially and contemporaneously correlated. Journal of the American Statistical Association, 62(318):500–509, June 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283977.

Parzen:1969:BRB

[Par69] Emanuel Parzen. Book review: [untitled]. Journal of the American Statistical Association, 64(327):1089, September 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283490.

Patchen:1961:BRB

[Pat61] Martin Patchen. Book review: [untitled]. Journal of the American Statistical Association, 56(294):462–463, June 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282307.

Patil:1969:BRB

[Pat69] G. P. Patil. Book review: [untitled]. Journal of the American Statistical Association, 64(328):1681, December 1969. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2286113.

Pavlov:1963:BRB

[Pav63] Nicolai Pavlov. Book review: [untitled]. Journal of the American Statistical Association, 58(303):864–865, September 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282752.

Preston:1961:SAI

[PB61] Lee E. Preston and Earl J. Bell. The statistical analysis of industry structure: an application to food industries. Journal of the American Statistical Association, 56(296):925–932, December 1961. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282004.

Patil:1967:MLS

[PB67] Ganapati P. Patil and Sheela Bildikar. Multivariate logarithmic series distribution as a probability model in population and community ecology and some of its statistical properties. Journal of the American Statistical Association, 62(318):655–674, June 1967. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283993. See corrigenda

[PB69].

Patil:1969:CML

[PB69] Ganapati P. Patil and Sheela Bildikar. Corrigenda: Multivariate Logarithmic Series Distribution as a Probability Model in Population and Community Ecology and Some of Its Statistical Properties. Journal of the American Statistical Association, 64(328): 1702, December 1969. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/

2286143. See [PB67].

Peach:1961:BPR

[Pea61] Paul Peach. Bias in pseudo-random numbers. Journal of the American Statistical Association, 56(295):610–618, September 1961. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282083.

Pearson:1969:BRB

[Pea69] Judson B. Pearson. Book review: [untitled]. Journal of the American Statistical Association, 64(325):399, March 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283754.

Perlman:1961:BRB

[Per61a] Mark Perlman. Book review: [untitled]. Journal of the American Statistical Association, 56(295):770–771, September 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282120.

Perry:1961:BRB

[Per61b] C. L. Perry. Book review: [untitled]. Journal of the American Statistical Association, 56(293):179, March 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282353.

Pesek:1964:BRB

[Pes64] Boris P. Pesek. Book review: [untitled]. Journal of the American Statistical Association, 59(305):291–292, March 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282882.

Petersen:1962:BRB

[Pet62] William Petersen. Book review: [untitled]. Journal of the American Statistical Association, 57(299):714–716, September 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282422.

Peterson:1967:BRB

[Pet67] D. E. Peterson. Book review: [untitled]. Journal of the American Statistical Association, 62(318):700–702, June 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284002.

Phelps:1962:BRB

[Phe62] Edmund S. Phelps. Book review: [untitled]. Journal of the American Statistical Association, 57(297):221–222, March 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282460.

Phillips:1962:BRB

[Phi62] Joseph D. Phillips. Book review: [untitled]. Journal of the American Statistical Association, 57(297):251–253, March 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282482.

Pierson:1968:BRB

[Pie68] James H. S. Pierson. Book review: [untitled]. Journal of the American Statistical Association, 63(323):1057–1058, September 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283903.

Pike:1969:NLC

[Pik69] Eugene W. Pike. A note on “learning curves”. Journal of the American Statistical Association, 64(328):1276–1277, December 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2286067.

Pinkham:1969:BRB

[Pin69] Roger S. Pinkham. Book review: [untitled]. Journal of the American Statistical Association, 64(327):1089–1090, September 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283491.

Plackett:1965:CBD

[Pla65] R. L. Plackett. A class of bivariate distributions. Journal of the American Statistical Association, 60(310):516–522, June 1965. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282685.

Pasternack:1961:PRA

[PO61] Bernard S. Pasternack and Junjiro Ogawa. The probability of reversal associated with a test procedure when data are incomplete. Journal of the American Statistical Association, 56(293): 125–134, March 1961. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2282338.

Polemis:1960:BRB

[Pol60] Bernice W. Polemis. Book review: [untitled]. Journal of the American Statistical Association, 55(289):239–241, March 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282202.

Polemis:1961:BRB

[Pol61] Zane M. Polemis. Book review: [untitled]. Journal of the American Statistical Association, 56(293):188–189, March 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282361.

Potthoff:1965:SST

[Pot65] Richard F. Potthoff. Some Scheffé-type tests for some Behrens– Fisher-type regression problems. Journal of the American Statistical Association, 60(312):1163–1190, December 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283414.

Powell:1964:BRB

[Pow64] James H. Powell. Book review: [untitled]. Journal of the American Statistical Association, 59(308):1322–1323, December 1964. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282669.

Powers:1965:BRB

[Pow65] Mary G. Powers. Book review: [untitled]. Journal of the American Statistical Association, 60(309):381–383, March 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283174.

Powell:1969:AET

[Pow69] Alan Powell. Aitken estimators as a tool in allocating predetermined aggregates. Journal of the American Statistical Association, 64(327):913–922, September 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283472.

Peizer:1968:NAB

[PP68] David B. Peizer and John W. Pratt. A normal approximation for binomial, F, beta, and other common, related tail probabilities, I. Journal of the American Statistical Association, 63(324): 1416–1456, December 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285895.

Pratt:1961:BRB

[Pra61a] John W. Pratt. Book review: [untitled]. Journal of the American Statistical Association, 56(293):163–167, March 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282344.

Pratt:1961:LCI

[Pra61b] John W. Pratt. Length of confidence intervals. Journal of the American Statistical Association, 56(295):549–567, September 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2282079.

Pratt:1964:RSP

[Pra64] John W. Pratt. Robustness of some procedures for the two-sample location problem. Journal of the American Statistical Association, 59(307):665–680, September 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283092.

Pratt:1968:NAB

[Pra68] John W. Pratt. A normal approximation for binomial, F, beta, and other common, related tail probabilities, II. Journal of the American Statistical Association, 63(324):1457–1483, December 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2285896.

Posner:1969:AEH

[PRAL69] Edward C. Posner, Eugene R. Rodemich, John C. Ashlock, and Sandra Lurie. Application of an estimator of high efficiency in bivariate extreme value theory. Journal of the American Statistical Association, 64(328):1403–1414, December 1969. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2286077.

Press:1965:BRB

[Pre65] S. J. Press. Book review: [untitled]. Journal of the American Statistical Association, 60(311):927–928, September 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283261.

Press:1966:CIC

[Pre66] S. James Press. A confidence interval comparison of two test procedures proposed for the Behrens–Fisher problems. Journal of the American Statistical Association, 61(314):454–466, June 1966. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282835.

Press:1968:EMD

[Pre68a] S. James Press. Estimating from misclassified data. Journal of the American Statistical Association, 63(321):123–133, March 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283833.

Press:1968:MCP

[Pre68b] S. James Press. A modified compound Poisson process with normal compounding. Journal of the American Statistical Association, 63(322):607–613, June 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284031.

Press:1969:RD

[Pre69] S. James Press. The t-ratio distribution. Journal of the American Statistical Association, 64(325):242–252, March 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283732.

Proschan:1964:BRB

[Pro64] Frank Proschan. Book review: [untitled]. Journal of the American Statistical Association, 59(306):619, June 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283025.

Projector:1968:BRB

[Pro68] Dorothy S. Projector. Book review: [untitled]. Journal of the American Statistical Association, 63(322):733–736, June 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284052.

Pratt:1964:FDU

[PRS64] John W. Pratt, Howard Raiffa, and Robert Schlaifer. The foundations of decision under uncertainty: An elementary exposition. Journal of the American Statistical Association, 59(306):353–375, June 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2282993.

Puri:1967:SON

[PS67] Madan Lal Puri and Pranab Kumar Sen. On some optimum nonparametric procedures in two-way layouts. Journal of the American Statistical Association, 62(320):1214–1229, December 1967. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283771.

Puri:1968:NCR

[PS68] Madan Lal Puri and Pranab Kumar Sen. Nonparametric confidence regions for some multivariate location problems. Journal of the American Statistical Association, 63(324):1373–1378, December 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2285890.

Putter:1967:OBE

[Put67] Joseph Putter. Orthonormal bases of error spaces and their use for investigating the normality and variances of residuals. Journal of the American Statistical Association, 62(319):1022–1036, September 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2283691.

Prairie:1964:FEF

[PZ64] R. R. Prairie and W. J. Zimmer. 2^{p} factorial experiments with the factors applied sequentially. Journal of the American Statistical Association, 59(308):1205–1216, December 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282636.

Prairie:1968:FRF

[PZ68] R. R. Prairie and W. J. Zimmer. Fractional replications of 2^{p} factorial experiments with the factors applied sequentially. Journal of the American Statistical Association, 63(322):644–652, June 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2284035.

Quandt:1960:THL

[Qua60] Richard E. Quandt. Tests of the hypothesis that a linear regression system obeys two separate regimes. Journal of the American Statistical Association, 55(290):324–330, June 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281745.

Quandt:1964:BRB

[Qua64] Richard E. Quandt. Book review: [untitled]. Journal of the American Statistical Association, 59(305):294–296, March 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282886.

Quade:1967:RAC

[Qua67] Dana Quade. Rank analysis of covariance. Journal of the American Statistical Association, 62(320):1187–1200, December 1967. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283769.

Quirk:1965:BRB

[Qui65] James Quirk. Book review: [untitled]. Journal of the American Statistical Association, 60(309):387–388, March 1965. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2283179.

Rabinowitz:1969:NCP

[Rab69] Philip Rabinowitz. New Chebyshev polynomial approximations to Mills’ ratio. Journal of the American Statistical Association, 64 (326):647–654, June 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2283648.

Rahman:1964:SGD

[Rah64] N. A. Rahman. Some generalisations of the distributions of product statistics arising from rectangular populations. Journal of the American Statistical Association, 59(306):557–563, June 1964. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283008.

Raj:1961:MLS

[Raj61] Des Raj. On matching lists by samples. Journal of the American Statistical Association, 56(293):151–155, March 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282341.

Raj:1964:NVR

[Raj64a] Des Raj. A note on the variance of the ratio estimate. Journal of the American Statistical Association, 59(307):895–898, September 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2283108.

Raj:1964:FSE

[Raj64b] Des Raj. On forming strata of equal aggregate size. Journal of the American Statistical Association, 59(306):481–486, June 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283002.

Raj:1964:USS

[Raj64c] Des Raj. The use of systematic sampling with probability proportionate to size in a large scale survey. Journal of the American Statistical Association, 59(305):251–255, March 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282873.

Raj:1965:MUM

[Raj65a] Des Raj. On a method of using multi-auxiliary information in sample surveys. Journal of the American Statistical Association, 60(309):270–277, March 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283151.

Raj:1965:VER

[Raj65b] Des Raj. Variance estimation in randomized systematic sampling with probability proportionate to size. Journal of the American Statistical Association, 60(309):278–284, March 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283152.

Raj:1966:SRS

[Raj66] Des Raj. Some remarks on a simple procedure of sampling without replacement. Journal of the American Statistical Association, 61 (314):391–396, June 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2282828.

Ramankutty:1965:CNC

[Ram65] C. Nair Ramankutty. Corrigenda: A New Class of Designs. Journal of the American Statistical Association, 60(312):1250, December 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2283451. See

[Nai64].

Ramanathan:1969:EEI

[Ram69] R. Ramanathan. An econometric exploration of Indian saving behavior. Journal of the American Statistical Association, 64(325): 90–101, March 1969. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2283721.

Rao:1963:BRB

[Rao63a] C. R. Rao. Book review: [untitled]. Journal of the American Statistical Association, 58(304):1185–1186, December 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283369.

Rao:1963:TPU

[Rao63b] J. N. K. Rao. On three procedures of unequal probability sampling without replacement. Journal of the American Statistical Association, 58(301):202–215, March 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2282964.

Rao:1965:AGS

[Rao65] M. Bhaskar Rao. Application of Greenberg and Sarhan’s method of inversion of partitioned matrices in the analysis of non-orthogonal data. Journal of the American Statistical Association, 60(312): 1200–1202, December 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2283416.

Rao:1968:BRB

[Rao68a] B. L. S. Prakasa Rao. Book review: [untitled]. Journal of the American Statistical Association, 63(323):1061–1062, September 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283907.

Rao:1968:SNS

[Rao68b] J. N. K. Rao. Some nonresponse sampling theory when the frame contains an unknown amount of duplication. Journal of the American Statistical Association, 63(321):87–90, March 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283829.

Rao:1969:CFR

[Rao69] Poduri S. R. S. Rao. Comparison of four ratio-type estimates under a model. Journal of the American Statistical Association, 64 (326):574–580, June 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2283639.

Rapoport:1961:BRB

[Rap61] Anatol Rapoport. Book review: [untitled]. Journal of the American Statistical Association, 56(294):433–438, June 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282284.

Rapoport:1963:BRB

[Rap63] Anatol Rapoport. Book review: [untitled]. Journal of the American Statistical Association, 58(303):829–832, September 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282732.

Rapoport:1967:BRB

[Rap67] Amnon Rapoport. Book review: [untitled]. Journal of the American Statistical Association, 62(318):702–704, June 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284003.

Rathmell:1962:BRB

[Rat62] John M. Rathmell. Book review: [untitled]. Journal of the American Statistical Association, 57(298):522–523, June 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281673.

Ray:1968:BRB

[Ray68] D. Michael Ray. Book review: [untitled]. Journal of the American Statistical Association, 63(322):726–728, June 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284047.

Rutemiller:1968:EHR

[RB68] Herbert C. Rutemiller and David A. Bowers. Estimation in a heteroscedastic regression model. Journal of the American Statistical Association, 63(322):552–557, June 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284026.

Rao:1969:ESS

[RB69] J. N. K. Rao and D. L. Bayless. An empirical study of the stabilities of estimators and variance estimators in unequal probability sampling of two units per stratum. Journal of the American Statistical Association, 64(326):540–559, June 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283637.

Reder:1960:BRB

[Red60] M. W. Reder. Book review: [untitled]. Journal of the American Statistical Association, 55(290):388–389, June 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281761.

Regier:1968:TSM

[Reg68] Mary H. Regier. A two-state Markov model for behavioral change. Journal of the American Statistical Association, 63(323):993–999, September 1968. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2283889.

Reid:1963:BRB

[Rei63] Margaret G. Reid. Book review: [untitled]. Journal of the American Statistical Association, 58(303):824–829, September 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282731.

Reiter:1966:BRB

[Rei66] Stanley Reiter. Book review: [untitled]. Journal of the American Statistical Association, 61(313):283–285, March 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283074.

Reich:1967:BRB

[Rei67] Edgar Reich. Book review: Probability and its Engineering Uses, by Thornton C. Fry. Journal of the American Statistical Association, 62(317):304–305, March 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2282941.

Remington:1965:BRB

[Rem65] Richard D. Remington. Book review: [untitled]. Journal of the American Statistical Association, 60(309):371–372, March 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283167.

Resnikoff:1961:BRB

[Res61] George J. Resnikoff. Book review: [untitled]. Journal of the American Statistical Association, 56(294):448, June 1961. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2282295.

Resnikoff:1968:BRB

[Res68] George J. Resnikoff. Book review: [untitled]. Journal of the American Statistical Association, 63(322):737–738, June 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284055.

Rothenberg:1964:NEC

[RFT64] Thomas J. Rothenberg, Franklin M. Fisher, and C. B. Tilanus. A note on estimation from a Cauchy sample. Journal of the American Statistical Association, 59(306):460–463, June 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282999.

Rao:1964:RDS

[RG64] J. N. K. Rao and Jack E. Graham. Rotation designs for sampling on repeated occasions. Journal of the American Statistical Association, 59(306):492–509, June 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283004.

Rao:1969:SSP

[RG69] Potluri Rao and Zvi Griliches. Small-sample properties of several two-stage regression methods in the context of auto-correlated errors. Journal of the American Statistical Association, 64(325): 253–272, March 1969. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2283733.

Rohde:1965:ULS

[RH65] C. A. Rohde and J. R. Harvey. Unified least squares analysis. Journal of the American Statistical Association, 60(310):523–527, June 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2282686.

Rhomberg:1962:BRB

[Rho62] Rudolf R. Rhomberg. Book review: [untitled]. Journal of the American Statistical Association, 57(297):253–255, March 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282483.

Rhomberg:1963:BRB

[Rho63] Rudolf R. Rhomberg. Book review: [untitled]. Journal of the American Statistical Association, 58(304):1189–1190, December 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283372.

Richmond:1964:BRB

[Ric64] Samuel B. Richmond. Book review: [untitled]. Journal of the American Statistical Association, 59(306):622–623, June 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283029.

Richter:1967:BRB

[Ric67] Donald Richter. Book review: [untitled]. Journal of the American Statistical Association, 62(320):1503–1504, December 1967. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283799.

Richardson:1968:EDS

[Ric68] David H. Richardson. The exact distribution of a structural coefficient estimator. Journal of the American Statistical Association, 63 (324):1214–1226, December 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2285878.

Rider:1960:VMSb

[Rid60a] Paul R. Rider. Variance of the median of samples from a Cauchy distribution. Journal of the American Statistical Association, 55 (290):322–323, June 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2281744.

Rider:1960:VMSa

[Rid60b] Paul R. Rider. Variance of the median of small samples from several special populations. Journal of the American Statistical Association, 55(289):148–150, March 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2282186.

Rider:1962:EVS

[Rid62] Paul R. Rider. Expected values and standard deviations of the reciprocal of a variable from a decapitated negative binomial distribution. Journal of the American Statistical Association, 57(298): 439–445, June 1962. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2281650.

Rider:1963:STP

[Rid63] Paul R. Rider. Sampling from a triangular population. Journal of the American Statistical Association, 58(302):509–512, June 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283285.

Rider:1964:DPQ

[Rid64] Paul R. Rider. Distribution of product and of quotient of maximum values in samples from a power-function population. Journal of the American Statistical Association, 59(307):877–880, September 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2283106.

Rieber:1967:BRB

[Rie67a] Michael Rieber. Book review: [untitled]. Journal of the American Statistical Association, 62(318):697–700, June 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284001.

Riedwyl:1967:GF

[Rie67b] Hans Riedwyl. Goodness of fit. Journal of the American Statistical Association, 62(318):390–398, June 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283969.

Riley:1961:SAS

[Ril61] H. E. Riley. Some aspects of seasonality in the consumer price index. Journal of the American Statistical Association, 56(293): 27–35, March 1961. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2282327.

Riordan:1967:BRB

[Rio67] John Riordan. Book review: Symmetric Function and Allied Tables, by F. N. David; M. G. Kendall; D. E. Barton. Journal of the American Statistical Association, 62(317):301–303, March 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282938.

Rao:1967:TPC

[RK67] P. V. Rao and L. L. Kupper. Ties in paired-comparison experiments: a generalization of the Bradley–Terry model. Journal of the American Statistical Association, 62(317):194–204, March 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2282923. See cor-

rigenda [RK68].

Rao:1968:CTP

[RK68] P. V. Rao and L. L. Kupper. Corrigenda: Ties in Paired-Comparison Experiments: a Generalization of the Bradley–Terry Model. Journal of the American Statistical Association, 63(324): 1550–1551, December 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285921. See [RK67].

Rao:1967:GME

[RM67] Poduri S. R. S. Rao and Govind S. Mudholkar. Generalized multivariate estimator for the mean of finite populations. Journal of the American Statistical Association, 62(319):1009–1012, September 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2283688.

Roberts:1962:BRBb

[Rob62a] Donald M. Roberts. Book review: [untitled]. Journal of the American Statistical Association, 57(298):516–517, June 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281666.

Roberts:1962:BRBa

[Rob62b] Harry V. Roberts. Book review: [untitled]. Journal of the American Statistical Association, 57(297):199–202, March 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282452.

Roberts:1963:BRBa

[Rob63a] David R. Roberts. Book review: [untitled]. Journal of the American Statistical Association, 58(304):1174–1175, December 1963. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283358.

Roberts:1963:BRBc

[Rob63b] Donald M. Roberts. Book review: [untitled]. Journal of the American Statistical Association, 58(304):1196, December 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283377.

Roberts:1963:BRBb

[Rob63c] Harry V. Roberts. Book review: [untitled]. Journal of the American Statistical Association, 58(304):1190–1191, December 1963. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283373.

Roberts:1964:BRB

[Rob64a] Harry V. Roberts. Book review: [untitled]. Journal of the American Statistical Association, 59(306):607–609, June 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283017.

Robison:1964:EPI

[Rob64b] D. E. Robison. Estimates for the points of intersection of two polynomial regressions. Journal of the American Statistical Association, 59(305):214–224, March 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2282870.

Roberts:1965:PP

[Rob65a] Harry V. Roberts. Probabilistic prediction. Journal of the American Statistical Association, 60(309):50–62, March 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283136.

Robichek:1965:BRB

[Rob65b] Alexander A. Robichek. Book review: [untitled]. Journal of the American Statistical Association, 60(310):675–676, June 1965. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282700.

Roberts:1966:CMU

[Rob66] Charles Roberts. A correlation model useful in the study of twins. Journal of the American Statistical Association, 61(316): 1184–1190, December 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2283207.

Roberts:1967:ISR

[Rob67] Harry V. Roberts. Informative stopping rules and inferences about population size. Journal of the American Statistical Association, 62(319):763–775, September 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283670.

Rogot:1961:NME

[Rog61] Eugene Rogot. A note on measurement errors and detecting real differences. Journal of the American Statistical Association, 56 (294):314–319, June 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2282256.

Rogers:1965:SAS

[Rog65] Andrei Rogers. A stochastic analysis of the spatial clustering of retail establishments. Journal of the American Statistical Association, 60(312):1094–1103, December 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283408.

Rolph:1963:BRB

[Rol63] Earl R. Rolph. Book review: [untitled]. Journal of the American Statistical Association, 58(303):845–846, September 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282743.

Roose:1961:BRB

[Roo61] Kenneth D. Roose. Book review: [untitled]. Journal of the American Statistical Association, 56(295):753–755, September 1961. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282105.

Rosner:1961:BRB

[Ros61a] Benjamin Rosner. Book review: [untitled]. Journal of the American Statistical Association, 56(293):185–187, March 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282359.

Ross:1961:VEO

[Ros61b] Alan Ross. Variance estimates in “optimum” sample designs. Journal of the American Statistical Association, 56(293):135– 142, March 1961. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2282339.

Rosenbaum:1965:STS

[Ros65] S. Rosenbaum. On some two-sample non-parametric tests. Journal of the American Statistical Association, 60(312):1118–1126, December 1965. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/

2283410. See corrigenda [Ros66a].

Rosenbaum:1966:CST

[Ros66a] S. Rosenbaum. Corrigenda: On Some Two-Sample Non-Parametric Tests. Journal of the American Statistical Association, 61(316): 1249, December 1966. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/

2283237. See [Ros65].

Rosenthal:1966:DSV

[Ros66b] Irene Rosenthal. Distribution of the sample version of the measure of association, gamma. Journal of the American Statistical Association, 61(314):440–453, June 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2282834.

Rosenwaike:1966:SVD

[Ros66c] Ira Rosenwaike. Seasonal variation of deaths in the United States, 1951–1960. Journal of the American Statistical Association, 61 (315):706–719, September 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2282781.

Rosenblatt:1968:SEB

[Ros68a] Harry M. Rosenblatt. Spectral evaluation of BLS and census revised seasonal adjustment procedures. Journal of the American Statistical Association, 63(322):472–501, June 1968. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2284020.

Rosenwaike:1968:MEA

[Ros68b] Ira Rosenwaike. On measuring the extreme aged in the population. Journal of the American Statistical Association, 63(321): 29–40, March 1968. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2283826.

Royall:1968:OAF

[Roy68] Richard Royall. An old approach to finite population sampling theory. Journal of the American Statistical Association, 63(324): 1269–1279, December 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285883.

Ruggles:1960:BRB

[Rug60] Nancy D. Ruggles. Book review: [untitled]. Journal of the American Statistical Association, 55(291):590–591, September 1960. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281923.

Rutemiller:1966:PER

[Rut66] Herbert C. Rutemiller. Point estimation of reliability of a system comprised of k elements from the same exponential distribution. Journal of the American Statistical Association, 61(316): 1029–1032, December 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2283197.

Rutemiller:1967:EPZ

[Rut67] Herbert C. Rutemiller. Estimation of the probability of zero failures in m binomial trials. Journal of the American Statistical Association, 62(317):272–277, March 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2282929.

Robson:1961:UCR

[RV61] D. S. Robson and C. Vithayasai. Unbiased componentwise ratio estimation. Journal of the American Statistical Association, 56 (294):350–358, June 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/ stable/2282260. See corrigenda [RV63].

Robson:1963:CUC

[RV63] D. S. Robson and C. Vithayasai. Corrigenda: Unbiased Componentwise Ratio Estimation. Journal of the American Statistical Association, 58(304):1163, December 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283350. See [RV61].

S:1961:BRB

[S.61] R. S. Book review: [untitled]. Journal of the American Statistical Association, 56(296):1029, December 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282039.

S:1965:BRBa

[S.65a] I. R. S. Book review: [untitled]. Journal of the American Statistical Association, 60(312):1242, December 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283438.

S:1965:BRBb

[S.65b] I. R. S. Book review: [untitled]. Journal of the American Statistical Association, 60(312):1242–1243, December 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283440.

S:1965:BRBc

[S.65c] I. R. S. Book review: [untitled]. Journal of the American Statistical Association, 60(312):1243, December 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283441.

S:1966:BRB

[S.66] I. R. S. Book review: Nonparametric Trend Analysis, by George Andrew Ferguson. Journal of the American Statistical Association, 61(314):535, June 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2282850.

Samuel:1966:EPB

[Sam66] Ester Samuel. Estimators with prescribed bound on the variance for the parameters in the binomial and Poisson distributions, based on two-stage sampling. Journal of the American Statistical Association, 61(313):220–227, March 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283056.

Samuelson:1968:BRB

[Sam68a] Paul Samuelson. Book review: [untitled]. Journal of the American Statistical Association, 63(323):1049–1051, September 1968. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283900.

Samuelson:1968:CUR

[Sam68b] Paul A. Samuelson. Constructing an unbiased random sequence. Journal of the American Statistical Association, 63(324):1526– 1527, December 1968. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2285902.

Samuelson:1968:HDC

[Sam68c] Paul A. Samuelson. How deviant can you be? Journal of the American Statistical Association, 63(324):1522–1525, December 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2285901.

Sandiford:1960:NBA

[San60] Peter J. Sandiford. A new binomial approximation for use in sampling from finite populations. Journal of the American Statistical Association, 55(292):718–722, December 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281594.

Sandee:1967:BRB

[San67] J. Sandee. Book review: [untitled]. Journal of the American Statistical Association, 62(319):1084–1085, September 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283705.

Sarndal:1964:UDS

[Sär64] Carl-Erik Särndal. A unified derivation of some nonparametric distributions. Journal of the American Statistical Association, 59 (308):1042–1053, December 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2282622.

Sarndal:1968:USV

[Sär68] Carl-Erik Särndal. The use of a stratification variable in estimation by proportional stratified sampling. Journal of the American Statistical Association, 63(324):1310–1320, December 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2285886.

Sasaki:1964:BRB

[Sas64] Kyohei Sasaki. Book review: [untitled]. Journal of the American Statistical Association, 59(305):313–314, March 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282895.

Savage:1961:BRB

[Sav61] Leonard J. Savage. Book review: [untitled]. Journal of the American Statistical Association, 56(295):746–747, September 1961. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282098.

Savage:1962:BRB

[Sav62] L. J. Savage. Book review: [untitled]. Journal of the American Statistical Association, 57(300):921–922, December 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281827.

Savage:1964:BRB

[Sav64] I. Richard Savage. Book review: [untitled]. Journal of the American Statistical Association, 59(307):975–976, September 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283123.

Sawa:1969:ESD

[Saw69] Takamitsu Sawa. The exact sampling distribution of ordinary least squares and two-stage least squares estimators. Journal of the American Statistical Association, 64(327):923–937, September 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2283473.

Stekler:1968:SED

[SB68] H. O. Stekler and Susan W. Burch. Selected economic data, accuracy vs. reporting speed. Journal of the American Statistical Association, 63(322):436–444, June 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284018.

Siddiqui:1969:AJD

[SB69] M. M. Siddiqui and Calvin Butler. Asymptotic joint distribution of linear systematic statistics from multivariate distributions. Journal of the American Statistical Association, 64(325):300–305, March 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2283739.

Savage:1962:FSI

[SBC^{+}62] L. J. Savage, George Barnard, Jerome Cornfield, Irwin Bross, George E. P. Box, I. J. Good, D. V. Lindley, C. W. Clunies-Ross, John W. Pratt, Howard Levene, Thomas Goldman, A. P. Dempster, Oscar Kempthorne, and Allan Birnbaum. On the foundations of statistical inference: Discussion. Journal of the American Statistical Association, 57(298):307–326, June 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281641. See [Bir62].

Stilson:1962:NCT

[SC62] Donald W. Stilson and Vincent N. Campbell. A note on calculating tau and average tau and on the sampling distribution of average tau with a criterion ranking. Journal of the American Statistical Association, 57(299):567–571, September 1962. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282393.

Speckman:1965:EOP

[SC65] Janace A. Speckman and Richard G. Cornell. Estimation for a one-parameter exponential model. Journal of the American Statistical Association, 60(310):560–572, June 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282691.

Scarf:1961:BRB

[Sca61] Herbert Scarf. Book review: [untitled]. Journal of the American Statistical Association, 56(294):411, June 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282267.

Scarf:1963:BRB

[Sca63] Herbert Scarf. Book review: [untitled]. Journal of the American Statistical Association, 58(304):1173, December 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283357.

Schnore:1960:TSD

[Sch60] Leo F. Schnore. Three sources of data on commuting: Problems and possibilities. Journal of the American Statistical Association, 55(289):8–22, March 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2282176.

Schaller:1961:BRB

[Sch61] Howard G. Schaller. Book review: [untitled]. Journal of the American Statistical Association, 56(293):204–205, March 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282373.

Scheuer:1962:MRE

[Sch62] Ernest M. Scheuer. Moments of the radial error. Journal of the American Statistical Association, 57(297):187–190, March 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282449. See corrigenda

[Sch65].

Schuh:1963:BRB

[Sch63] G. Edward Schuh. Book review: Surplus Labor in Greek Agriculture, 1953–60, by Adam A. Pepelasis; Pan A. Yotopoulos. Journal of the American Statistical Association, 58(303):857–859, September 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2282749.

Scheuer:1965:CMR

[Sch65] Ernest M. Scheuer. Corrigenda: Moments of the Radial Error. Journal of the American Statistical Association, 60(312): 1251, December 1965. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/

2283454. See [Sch62].

Schaller:1966:BRB

[Sch66a] Howard G. Schaller. Book review: Elements of Regional Accounts. Papers Presented at the Conference on Regional Accounts, 1962, by Werner Z. Hirsch. Journal of the American Statistical Association, 61(314):533–535, June 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2282849.

Schatzoff:1966:SCA

[Sch66b] Martin Schatzoff. Sensitivity comparisons among tests of the general linear hypothesis. Journal of the American Statistical Association, 61(314):415–435, June 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2282832.

Schleifer:1966:BRB

[Sch66c] Arthur Schleifer, Jr. Book review: [untitled]. Journal of the American Statistical Association, 61(316):1231–1232, December 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283213.

Schatzoff:1968:ATS

[Sch68] Martin Schatzoff. Applications of time-shared computers in a statistics curriculum. Journal of the American Statistical Association, 63(321):192–208, March 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283840.

Schleifer:1969:TSN

[Sch69] Arthur Schleifer, Jr. Two-stage normal sampling in two-action problems with linear economics. Journal of the American Statistical Association, 64(328):1504–1541, December 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286086.

Sclove:1968:IEC

[Scl68] Stanley L. Sclove. Improved estimators for coefficients in linear regression. Journal of the American Statistical Association, 63 (322):596–606, June 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284030.

Schaeffer:1963:CMS

[SD63] Esther Schaeffer and Paul S. Dwyer. Computation with multiple k-statistics. Journal of the American Statistical Association, 58(301):120–151, March 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2282958.

Searls:1963:PWD

[Sea63] Donald T. Searls. On the probability of winning with different tournament procedures. Journal of the American Statistical Association, 58(304):1064–1081, December 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283333.

Searls:1964:UKC

[Sea64] Donald T. Searls. The utilization of a known coefficient of variation in the estimation procedure. Journal of the American Statistical Association, 59(308):1225–1226, December 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282639.

Searls:1966:EPM

[Sea66] Donald T. Searls. An estimator for a population mean which reduces the effect of large true observations. Journal of the American Statistical Association, 61(316):1200–1204, December 1966. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283209.

Seal:1969:BRB

[Sea69] Hilary L. Seal. Book review: [untitled]. Journal of the American Statistical Association, 64(328):1669–1671, December 1969. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2286102.

Sedransk:1965:DSS

[Sed65] J. Sedransk. A double sampling scheme for analytical surveys. Journal of the American Statistical Association, 60(312):985–1004, December 1965. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2283400.

Sedransk:1967:DSM

[Sed67] J. Sedransk. Designing some multi-factor analytical studies. Journal of the American Statistical Association, 62(320):1121–1139, December 1967. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2283765.

Sedransk:1969:BRB

[Sed69] J. Sedransk. Book review: [untitled]. Journal of the American Statistical Association, 64(326):678–680, June 1969. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2283662.

Sen:1961:BRB

[Sen61] Amartya Kumar Sen. Book review: [untitled]. Journal of the American Statistical Association, 56(296):1008–1009, December 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282016.

Sen:1965:SNP

[Sen65] Pranab Kumar Sen. Some non-parametric tests for m-dependent time series. Journal of the American Statistical Association, 60 (309):134–147, March 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2283141.

Sen:1967:SMP

[Sen67] Pranab Kumar Sen. On some multisample permutation tests based on a class of U-statistics. Journal of the American Statistical Association, 62(320):1201–1213, December 1967. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283770.

Sen:1968:ERC

[Sen68] Pranab Kumar Sen. Estimates of the regression coefficient based on Kendall’s tau. Journal of the American Statistical Association, 63 (324):1379–1389, December 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2285891.

Sen:1969:GMM

[Sen69] Pranab Kumar Sen. A generalization of the T -method of multiple comparisons for interactions. Journal of the American Statistical Association, 64(325):290–295, March 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283737.

Serfling:1968:AOS

[Ser68] R. J. Serfling. Approximately optimal stratification. Journal of the American Statistical Association, 63(324):1298–1309, December 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2285885.

Seshadri:1963:CUB

[Ses63] Vanamamalai Seshadri. Constructing uniformly better estimators. Journal of the American Statistical Association, 58(301): 172–175, March 1963. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2282961.

Severo:1963:BRB

[Sev63] Norman C. Severo. Book review: [untitled]. Journal of the American Statistical Association, 58(304):1176–1177, December 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283360.

Severo:1968:BRB

[Sev68] Norman C. Severo. Book review: [untitled]. Journal of the American Statistical Association, 63(324):1541, December 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2285911.

Shaffer:1960:BRB

[Sha60] James Duncan Shaffer. Book review: [untitled]. Journal of the American Statistical Association, 55(292):766, December 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281612.

Shah:1961:AVM

[Sha61a] S. M. Shah. The asymptotic variances of method of moments estimates of the parameters of the truncated binomial and negative binomial distributions. Journal of the American Statistical Association, 56(296):990–994, December 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282009.

Shah:1961:NGP

[Sha61b] S. M. Shah. A note on Griffin’s paper “graphic computation of tau as a coefficient of disarray”. Journal of the American Statistical Association, 56(295):736, September 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282093.

Shakun:1965:MAS

[Sha65] Melvin F. Shakun. Multivariate acceptance sampling procedures for general specification ellipsoids. Journal of the American Statistical Association, 60(311):905–913, September 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283256.

Shah:1966:EPD

[Sha66] S. M. Shah. On estimating the parameter of a doubly truncated binomial distribution. Journal of the American Statistical Association, 61(313):259–263, March 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283061.

Shah:1967:UBC

[Sha67] K. R. Shah. Uniformly better combined estimators in factorial arrangements with confounding. Journal of the American Statistical Association, 62(318):638–642, June 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283991.

Shell:1967:BRB

[She67] Karl Shell. Book review: [untitled]. Journal of the American Statistical Association, 62(318):708–709, June 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284006.

Sheth:1969:UFA

[She69] Jagdish N. Sheth. Using factor analysis to estimate parameters. Journal of the American Statistical Association, 64(327):808–822, September 1969. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2283462.

Shishido:1964:CPI

[Shi64a] Shuntaro Shishido. Corrigenda: Problems in the International Standardization of Interindustry Tables. Journal of the American Statistical Association, 59(308):1299, December 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282650. See [Shi64b].

Shishido:1964:PIS

[Shi64b] Shuntaro Shishido. Problems in the international standardization of interindustry tables. Journal of the American Statistical Association, 59(305):256–272, March 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2282874. See [Shi64a].

Shorack:1968:RGD

[Sho68] Roger A. Shorack. Recursive generation of the distribution of several non-parametric test statistics under censoring. Journal of the American Statistical Association, 63(321):353–366, March 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283855.

Shorack:1969:TER

[Sho69] Galen R. Shorack. Testing and estimating ratios of scale parameters. Journal of the American Statistical Association, 64(327): 999–1013, September 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2283481.

Shryock:1960:BRB

[Shr60] Henry S. Shryock, Jr. Book review: [untitled]. Journal of the American Statistical Association, 55(290):389–390, June 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281762.

Shryock:1964:BRB

[Shr64] Henry S. Shryock. Book review: [untitled]. Journal of the American Statistical Association, 59(308):1312, December 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282660.

Shubik:1960:BSG

[Shu60] Martin Shubik. Bibliography on simulation, gaming, artificial intelligence and allied topics. Journal of the American Statistical Association, 55(292):736–751, December 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281597.

Shupp:1961:BRB

[Shu61] Franklin R. Shupp. Book review: [untitled]. Journal of the American Statistical Association, 56(293):210–212, March 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282379.

Shubik:1964:BRB

[Shu64] Martin Shubik. Book review: [untitled]. Journal of the American Statistical Association, 59(305):294, March 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282885.

Shubik:1965:BRB

[Shu65] Martin Shubik. Book review: [untitled]. Journal of the American Statistical Association, 60(309):378–379, March 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283171.

Shuster:1968:IGD

[Shu68] Jonathan Shuster. On the inverse Gaussian distribution function. Journal of the American Statistical Association, 63(324): 1514–1516, December 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285899.

Sidak:1967:RCR

[Sid67] Zbynek Sidak. Rectangular confidence regions for the means of multivariate normal distributions. Journal of the American Statistical Association, 62(318):626–633, June 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283989.

Siddiqui:1968:BRB

[Sid68] M. M. Siddiqui. Book review: [untitled]. Journal of the American Statistical Association, 63(323):1060–1061, September 1968. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283906.

Siddiqui:1969:BRB

[Sid69] M. M. Siddiqui. Book review: [untitled]. Journal of the American Statistical Association, 64(328):1669, December 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286101.

Silver:1966:BRB

[Sil66] Edward A. Silver. Book review: [untitled]. Journal of the American Statistical Association, 61(315):884–885, September 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282812.

Simmons:1962:BRB

[Sim62a] Walt R. Simmons. Book review: [untitled]. Journal of the American Statistical Association, 57(299):699–700, September 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282412.

Simpson:1962:BRB

[Sim62b] Paul B. Simpson. Book review: [untitled]. Journal of the American Statistical Association, 57(298):515–516, June 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281665.

Simon:1966:BRB

[Sim66] Herbert A. Simon. Book review: [untitled]. Journal of the American Statistical Association, 61(316):1232–1233, December 1966. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283214.

Singh:1963:PMV

[Sin63] S. N. Singh. Probability models for the variation in the number of births per couple. Journal of the American Statistical Association, 58(303):721–727, September 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2282720.

Singh:1966:PCR

[Sin66] H. R. Singh. Producer and consumer risks for asymmetrical test and specification limits. Journal of the American Statistical Association, 61(314):505–513, June 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2282840.

Singh:1968:ESS

[Sin68a] D. Singh. Estimates in successive sampling using a multi-stage design. Journal of the American Statistical Association, 63(321): 99–112, March 1968. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2283831.

Singh:1968:CMV

[Sin68b] S. N. Singh. A chance mechanism of the variation in the number of births per couple. Journal of the American Statistical Association, 63(321):209–213, March 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283841.

Singer:1969:BRB

[Sin69] Burton Singer. Book review: [untitled]. Journal of the American Statistical Association, 64(325):394–397, March 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283752.

Siotani:1964:IEL

[Sio64] Minoru Siotani. Interval estimation for linear combinations of means. Journal of the American Statistical Association, 59(308): 1141–1164, December 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2282629.

Sitgreaves:1960:BRBa

[Sit60a] Rosedith Sitgreaves. Book review: [untitled]. Journal of the American Statistical Association, 55(289):218, March 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282190.

Sitgreaves:1960:BRBb

[Sit60b] Rosedith Sitgreaves. Book review: [untitled]. Journal of the American Statistical Association, 55(290):384–385, June 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281758.

Sivamurthy:1969:EEN

[Siv69] M. Sivamurthy. Errors in the estimation of net migration rate in the studies of internal migration. Journal of the American Statistical Association, 64(328):1434–1438, December 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286080.

Sprott:1969:ELC

[SK69] D. A. Sprott and John D. Kalbfleisch. Examples of likelihoods and comparison with point estimates and large sample approximations. Journal of the American Statistical Association, 64(326):468–484, June 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2283633.

Slater:1962:BRB

[Sla62] David Slater. Book review: [untitled]. Journal of the American Statistical Association, 57(299):718, September 1962. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2282425.

Slakter:1965:CPC

[Sla65] Malcolm J. Slakter. A comparison of the Pearson chi-square and Kolmogorov goodness-of-fit tests with respect to validity. Journal of the American Statistical Association, 60(311):854–858, September 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2283251. See

corrigenda [Sla66].

Slakter:1966:CCP

[Sla66] Malcolm J. Slakter. Corrigenda: A Comparison of the Pearson Chi-Square and Kolmogorov Goodness-of-Fit Tests with Respect to Validity. Journal of the American Statistical Association, 61 (316):1249–1252, December 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283238. See [Sla65].

Slakter:1968:AAP

[Sla68] Malcolm J. Slakter. Accuracy of an approximation to the power of the chi-square goodness of fit test with small but equal expected frequencies. Journal of the American Statistical Association, 63 (323):912–918, September 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283883.

Swamy:1969:TMR

[SM69] P. A. V. B. Swamy and J. S. Mehta. On Theil’s mixed regression estimator. Journal of the American Statistical Association, 64(325):273–276, March 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283734.

Smith:1961:BRB

[Smi61] Nicholas M. Smith. Book review: [untitled]. Journal of the American Statistical Association, 56(296):1013–1015, December 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282019.

Smith:1962:BRBa

[Smi62a] John H. Smith. Book review: [untitled]. Journal of the American Statistical Association, 57(299):718–719, September 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282426.

Smith:1962:BRBb

[Smi62b] Victor E. Smith. Book review: [untitled]. Journal of the American Statistical Association, 57(300):931–932, December 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281834.

Smith:1964:BRBc

[Smi64a] Barnard E. Smith. Book review: [untitled]. Journal of the American Statistical Association, 59(308):1320–1321, December 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282667.

Smith:1964:BRBb

[Smi64b] Paul E. Smith. Book review: [untitled]. Journal of the American Statistical Association, 59(308):1310–1311, December 1964. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282658.

Smith:1964:BRBa

[Smi64c] Vernon L. Smith. Book review: [untitled]. Journal of the American Statistical Association, 59(306):619–620, June 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283026.

Smith:1966:BRB

[Smi66] Harry Smith, Jr. Book review: [untitled]. Journal of the American Statistical Association, 61(315):878–879, September 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282808.

Smith:1967:BRB

[Smi67] H. Fairfield Smith. Book review: [untitled]. Journal of the American Statistical Association, 62(319):1076–1078, September 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283699.

Smith:1968:BRB

[Smi68] James C. Smith. Book review: [untitled]. Journal of the American Statistical Association, 63(322):725–726, June 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284046.

Smuts:1960:FLF

[Smu60] Robert W. Smuts. The female labor force: a case study in the interpretation of historical statistics. Journal of the American Statistical Association, 55(289):71–79, March 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282179.

Stockwell:1963:ITS

[SN63] Edward G. Stockwell and Charles B. Nam. Illustrative tables of school life. Journal of the American Statistical Association, 58 (304):1113–1124, December 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283337. See [SN64].

Stockwell:1964:CIT

[SN64] Edward G. Stockwell and Charles B. Nam. Corrigenda: Illustrative Tables of School Life. Journal of the American Statistical Association, 59(308):1299, December 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2282651. See [SN63].

Snell:1961:BRBa

[Sne61a] J. Laurie Snell. Book review: [untitled]. Journal of the American Statistical Association, 56(293):167–169, March 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282345.

Snell:1961:BRBb

[Sne61b] J. Laurie Snell. Book review: [untitled]. Journal of the American Statistical Association, 56(294):413–414, June 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282270.

Snell:1962:BRB

[Sne62] J. Laurie Snell. Book review: [untitled]. Journal of the American Statistical Association, 57(297):202–203, March 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282453.

Somers:1962:SBG

[Som62] Robert H. Somers. A similarity between Goodman and Kruskal’s tau and Kendall’s tau, with a partial interpretation of the latter. Journal of the American Statistical Association, 57(300):804– 812, December 1962. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2281811.

Somers:1965:BRB

[Som65] Harold M. Somers. Book review: [untitled]. Journal of the American Statistical Association, 60(309):375–378, March 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283170.

Somers:1969:BRB

[Som69] Robert H. Somers. Book review: [untitled]. Journal of the American Statistical Association, 64(326):673–675, June 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283656.

Sparks:1961:BRB

[Spa61] Willard Sparks. Book review: [untitled]. Journal of the American Statistical Association, 56(294):453–454, June 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282299.

Spiegelglas:1960:SII

[Spi60] Stephen Spiegelglas. A statistical investigation of the industrialization controversy. Journal of the American Statistical Association, 55(290):284–298, June 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281741.

Spitzer:1964:BRB

[Spi64] F. Spitzer. Book review: [untitled]. Journal of the American Statistical Association, 59(305):292–293, March 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282883.

Spivey:1966:BRB

[Spi66] W. Allen Spivey. Book review: [untitled]. Journal of the American Statistical Association, 61(315):872–874, September 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282800.

Sprowls:1964:BRB

[Spr64] R. Clay Sprowls. Book review: [untitled]. Journal of the American Statistical Association, 59(308):1319–1320, December 1964. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282666.

Spurr:1964:BRB

[Spu64] William A. Spurr. Book review: [untitled]. Journal of the American Statistical Association, 59(307):979, September 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283126.

Siddiqui:1967:ARE

[SR67] M. M. Siddiqui and K. Raghunandanan. Asymptotically robust estimators of location. Journal of the American Statistical Association, 62(319):950–953, September 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283683.

Scheuer:1963:CSP

[SS63a] Ernest M. Scheuer and Robert A. Spurgeon. Corrigenda: Some Percentage Points of the Non-Central t-Distribution. Journal of the American Statistical Association, 58(304):1163, December 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283351. See [SS63b].

Scheuer:1963:SPP

[SS63b] Ernest M. Scheuer and Robert A. Spurgeon. Some percentage points of the non-central t-distribution. Journal of the American Statistical Association, 58(301):176–182, March 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282962. See corrigenda [SS63a].

Singh:1965:DSS

[SS65] D. Singh and B. D. Singh. Double sampling for stratification on successive occasions. Journal of the American Statistical Association, 60(311):784–792, September 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283245.

Scott:1969:EMS

[SS69] Alastair Scott and T. M. F. Smith. Estimation in multi-stage surveys. Journal of the American Statistical Association, 64(327): 830–840, September 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2283464.

Siegel:1960:NSR

[ST60] Sidney Siegel and John W. Tukey. A nonparametric sum of ranks procedure for relative spread in unpaired samples. Journal of the American Statistical Association, 55(291):429–445, September 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2281906. See cor-

rigenda [ST61].

Siegel:1961:CNS

[ST61] Sidney Siegel and John W. Tukey. Corrigenda: A Nonparametric Sum of Ranks Procedure for Relative Spread in Unpaired Samples. Journal of the American Statistical Association, 56(296): 1005, December 1961. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/

2282014. See [ST60].

Saxena:1969:IEL

[ST69] K. M. Lal Saxena and Yung Liang Tong. Interval estimation of the largest mean of k normal populations with known variances. Journal of the American Statistical Association, 64(325): 296–299, March 1969. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2283738.

Staley:1961:BRB

[Sta61] Eugene Staley. Book review: [untitled]. Journal of the American Statistical Association, 56(295):764–765, September 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282115.

Stapleton:1962:BRB

[Sta62] J. H. Stapleton. Book review: [untitled]. Journal of the American Statistical Association, 57(297):239–240, March 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282471.

Starr:1964:BRB

[Sta64] Martin K. Starr. Book review: [untitled]. Journal of the American Statistical Association, 59(308):1321–1322, December 1964. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282668.

Staff:1967:DPD

[Sta67] P. J. Staff. The displaced Poisson distribution. region B. Journal of the American Statistical Association, 62(318):643–654, June 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283992.

Stekler:1961:FIP

[Ste61a] H. O. Stekler. Forecasting industrial production. Journal of the American Statistical Association, 56(296):869–877, December 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282000.

Stemberger:1961:BRB

[Ste61b] Anthony P. Stemberger. Book review: [untitled]. Journal of the American Statistical Association, 56(293):197–198, March 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282368.

Steinmueller:1962:BRB

[Ste62] Milton H. Steinmueller. Book review: [untitled]. Journal of the American Statistical Association, 57(299):712–714, September 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2282421.

Stephenson:1963:BRB

[Ste63a] John L. Stephenson. Book review: [untitled]. Journal of the American Statistical Association, 58(302):576, June 1963. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2283298.

Stewart:1963:BRB

[Ste63b] John B. Stewart. Book review: [untitled]. Journal of the American Statistical Association, 58(303):848–849, September 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282745.

Stegler:1964:BRB

[Ste64a] George J. Stegler. Book review: [untitled]. Journal of the American Statistical Association, 59(307):969–970, September 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283118.

Steiner:1964:BRB

[Ste64b] Peter O. Steiner. Book review: [untitled]. Journal of the American Statistical Association, 59(306):624–626, June 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283031.

Steinkamp:1964:IEI

[Ste64c] Stanley W. Steinkamp. The identification of effective interviewers. Journal of the American Statistical Association, 59(308):1165– 1174, December 1964. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2282630.

Stekler:1964:BRB

[Ste64d] H. O. Stekler. Book review: [untitled]. Journal of the American Statistical Association, 59(307):965–967, September 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283115.

Stekler:1964:VPS

[Ste64e] H. O. Stekler. The variability of profitability with size of firm, 1947–1958. Journal of the American Statistical Association, 59 (308):1183–1193, December 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2282633.

Stephens:1964:TUV

[Ste64f] M. A. Stephens. The testing of unit vectors for randomness. Journal of the American Statistical Association, 59(305):160– 167, March 1964. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2282867.

Sterrett:1964:BRB

[Ste64g] Andrew Sterrett. Book review: [untitled]. Journal of the American Statistical Association, 59(308):1324, December 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282671.

Stekler:1967:DRE

[Ste67a] H. O. Stekler. Data revisions and economic forecasting. Journal of the American Statistical Association, 62(318):470–483, June 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283975.

Stephan:1967:QSI

[Ste67b] Frederick F. Stephan. The quality of statistical information and statistical inference in a rapidly changing world. Journal of the American Statistical Association, 62(317):1–9, March 1967. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282905.

Steinberg:1968:BRB

[Ste68] Joseph Steinberg. Book review: [untitled]. Journal of the American Statistical Association, 63(324):1531–1532, December 1968. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2285905.

Steffens:1969:CCV

[Ste69a] F. E. Steffens. Corrigenda: Critical Values for Bivariate Student t-Tests. Journal of the American Statistical Association, 64(328): 1702–1703, December 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286144. See [Ste69b].

Steffens:1969:CVB

[Ste69b] F. E. Steffens. Critical values for bivariate Student t-tests. Journal of the American Statistical Association, 64(326):637–646, June 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2283647. See cor-

rigenda [Ste69a].

Stephens:1969:TRD

[Ste69c] M. A. Stephens. Tests for randomness of directions against two circular alternatives. Journal of the American Statistical Association, 64(325):280–289, March 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283736.

Sterrett:1969:BRB

[Ste69d] Andrew Sterrett. Book review: [untitled]. Journal of the American Statistical Association, 64(328):1676–1677, December 1969. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2286109.

Stieber:1963:BRB

[Sti63] Jack Stieber. Book review: [untitled]. Journal of the American Statistical Association, 58(303):859–862, September 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282750.

Stephan:1965:SW

[STM^{+}65] Frederick F. Stephan, John W. Tukey, Frederick Mosteller, Alex M. Mood, Morris H. Hansen, Leslie E. Simon, and W. J. Dixon. Samuel S. Wilks. Journal of the American Statistical Association, 60(312):939–966, December 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283397.

Stone:1960:BRB

[Sto60] Richard Stone. Book review: [untitled]. Journal of the American Statistical Association, 55(290):375–376, June 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281750.

Stone:1963:RNI

[Sto63] Mervyn Stone. Robustness of non-ideal decision procedures. Journal of the American Statistical Association, 58(302):480–486, June 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2283282.

Stoller:1966:BRBa

[Sto66a] David S. Stoller. Book review: Statistica Economica, by G. Lasorsa; Metodologia Statistica, by G. Lasorsa. Journal of the American Statistical Association, 61(314):526, June 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282842.

Stoller:1966:BRBb

[Sto66b] David S. Stoller. Book review: [untitled]. Journal of the American Statistical Association, 61(315):877, September 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282806.

Stone:1969:BRB

[Sto69] Charles J. Stone. Book review: [untitled]. Journal of the American Statistical Association, 64(328):1676, December 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286108.

Strassmann:1960:BRB

[Str60] W. Paul Strassmann. Book review: [untitled]. Journal of the American Statistical Association, 55(289):237–239, March 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282201.

Striner:1961:BRB

[Str61a] Herbert E. Striner. Book review: [untitled]. Journal of the American Statistical Association, 56(294):461–462, June 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282306.

Strotz:1961:BRB

[Str61b] Robert H. Strotz. Book review: [untitled]. Journal of the American Statistical Association, 56(294):422–423, June 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282277.

Stuart:1966:CES

[Stu66] Walter J. Stuart. Computer editing of survey data. five years of experience in BLS manpower surveys. Journal of the American Statistical Association, 61(314):375–383, June 1966. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2282826.

Subrahmaniam:1967:BRB

[Sub67] K. Subrahmaniam. Book review: [untitled]. Journal of the American Statistical Association, 62(319):1085–1090, September 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283707.

Sudman:1966:PSQ

[Sud66] Seymour Sudman. Probability sampling with quotas. Journal of the American Statistical Association, 61(315):749–771, September 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2282785.

Sukhatme:1962:SRT

[Suk62] Balkrishna V. Sukhatme. Some ratio-type estimators in two-phase sampling. Journal of the American Statistical Association, 57(299): 628–632, September 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2282400.

Suppes:1961:BRB

[Sup61] Patrick Suppes. Book review: [untitled]. Journal of the American Statistical Association, 56(293):172–174, March 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282347.

Suzuki:1968:KST

[Suz68] Gitiro Suzuki. Kolmogorov–Smirnov tests of fit based on some general bounds. Journal of the American Statistical Association, 63(323):919–924, September 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283884.

Sveshnikov:1969:CPP

[Sve69] A. A. Sveshnikov. Corrigenda: Problems in Probability Theory, Mathematical Statistics and Theory of Random Functions. Journal of the American Statistical Association, 64(328):1699–1700, December 1969. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2286131.

Swamy:1962:JEE

[Swa62] P. S. Swamy. On the joint efficiency of the estimates of the parameters of normal populations based on singly and doubly truncated samples. Journal of the American Statistical Association, 57(297):46–53, March 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2282439.

Shapiro:1968:CSV

[SWC68] S. S. Shapiro, M. B. Wilk, and H. J. Chen. A comparative study of various tests for normality. Journal of the American Statistical Association, 63(324):1343–1372, December 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285889.

Sykes:1969:SSV

[Syk69] Z. M. Sykes. Some stochastic versions of the matrix model for population dynamics. Journal of the American Statistical Association, 64(325):111–130, March 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283723.

Sylwester:1969:BRB

[Syl69] David Sylwester. Book review: [untitled]. Journal of the American Statistical Association, 64(326):677–678, June 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283660.

Taeuber:1967:BRB

[Tae67] Conrad Taeuber. Book review: [untitled]. Journal of the American Statistical Association, 62(320):1499, December 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283796.

Takacs:1962:GBP

[Tak62] Lajos Takacs. A generalization of the ballot problem and its application in the theory of queues. Journal of the American Statistical Association, 57(298):327–337, June 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281642.

Takacs:1967:MIE

[Tak67] Lajos Takács. On the method of inclusion and exclusion. Journal of the American Statistical Association, 62(317):102–113, March 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2282913.

Takacs:1969:CRP

[Tak69] Lajos Takacs. On the classical ruin problems. Journal of the American Statistical Association, 64(327):889–906, September 1969. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283470.

Tan:1969:NMG

[Tan69] W. Y. Tan. Note on the multivariate and the generalized multivariate beta distributions. Journal of the American Statistical Association, 64(325):230–241, March 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283731.

Tarver:1962:ECS

[Tar62] James D. Tarver. Evaluation of census survival rates in estimating intercensal state net migration. Journal of the American Statistical Association, 57(300):841–862, December 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281814.

Tarter:1966:EMP

[Tar66] Michael E. Tarter. Exact moments and product moments of the order statistics from the truncated logistic distribution. Journal of the American Statistical Association, 61(314):514–525, June 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282841.

Tate:1961:UPO

[Tat61] R. F. Tate. On the use of partially ordered observations in measuring the support for a complete order. Journal of the American Statistical Association, 56(294):299–313, June 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282255.

Tate:1966:CNR

[Tat66a] R. F. Tate. Conditional-normal regression models. Journal of the American Statistical Association, 61(314):477–489, June 1966. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282837.

Tatsuoka:1966:BRB

[Tat66b] Maurice M. Tatsuoka. Book review: [untitled]. Journal of the American Statistical Association, 61(315):869–872, September 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2282799.

Taylor:1962:BRB

[Tay62] William F. Taylor. Book review: [untitled]. Journal of the American Statistical Association, 57(297):207–209, March 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282456.

Taylor:1964:CAR

[Tay64] Wilson L. Taylor. Correcting the average rank correlation coefficient for ties in rankings. Journal of the American Statistical Association, 59(307):872–876, September 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283105.

Taylor:1965:BRB

[Tay65] Howard Taylor. Book review: [untitled]. Journal of the American Statistical Association, 60(312):1231–1233, December 1965. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283430.

Tiao:1965:IDD

[TC65] George G. Tiao and Irwin Cuttman. The inverted Dirichlet distribution with applications. Journal of the American Statistical Association, 60(311):793–805, September 1965. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283246. See corrigenda [TG65].

Taylor:1962:MSP

[TD62] R. J. Taylor and H. A. David. A multi-stage procedure for the selection of the best of several populations. Journal of the American Statistical Association, 57(300):785–796, December 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281809.

Tracy:1969:MMM

[TD69] Derrick S. Tracy and Paul S. Dwyer. Multivariate maxima and minima with matrix derivatives. Journal of the American Statistical Association, 64(328):1576–1594, December 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286090.

Teichroew:1961:BRB

[Tei61] Daniel Teichroew. Book review: [untitled]. Journal of the American Statistical Association, 56(293):206, March 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282375.

Teichroew:1965:CHD

[Tei65a] Daniel Teichroew. Corrigenda: A History of Distribution Sampling Prior to the Era of the Computer and its Relevance to Simulation. Journal of the American Statistical Association, 60(312): 1251, December 1965. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/

2283455. See [Tei65b].

Teichroew:1965:HDS

[Tei65b] Daniel Teichroew. A history of distribution sampling prior to the era of the computer and its relevance to simulation. Journal of the American Statistical Association, 60(309):27–49, March 1965. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283135. See corrigenda

[Tei65a].

Telser:1964:IES

[Tel64] Lester G. Telser. Iterative estimation of a set of linear regression equations. Journal of the American Statistical Association, 59 (307):845–862, September 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283102.

Telser:1967:DSM

[Tel67a] L. G. Telser. Discrete samples and moving sums in stationary stochastic processes. Journal of the American Statistical Association, 62(318):484–499, June 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283976.

Telser:1967:BRB

[Tel67b] Lester G. Telser. Book review: [untitled]. Journal of the American Statistical Association, 62(318):687–688, June 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283996.

Temin:1966:BRB

[Tem66] Peter Temin. Book review: [untitled]. Journal of the American Statistical Association, 61(315):885, September 1966. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2282813.

Tepping:1968:BRB

[Tep68a] Benjamin J. Tepping. Book review: [untitled]. Journal of the American Statistical Association, 63(322):728–730, June 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284049.

Tepping:1968:MOL

[Tep68b] Benjamin J. Tepping. A model for optimum linkage of records. Journal of the American Statistical Association, 63(324):1321– 1332, December 1968. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2285887.

Terry:1963:BRB

[Ter63] M. E. Terry. Book review: [untitled]. Journal of the American Statistical Association, 58(304):1202, December 1963. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2283385.

Taylor:1963:SCA

[TF63] Wilson L. Taylor and Ching Fong. Some contributions to average rank correlation methods and to the distribution of the average rank correlation coefficient. Journal of the American Statistical Association, 58(303):756–769, September 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282723.

Tiao:1965:CID

[TG65] C. George Tiao and Irwin Guttman. Corrigenda: The Inverted Dirichlet Distribution with Applications. Journal of the American Statistical Association, 60(312):1251–1252, December 1965. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283456. See [TC65].

Thompson:1967:DPA

[TGD67] Rory Thompson, Z. Govindarajulu, and K. A. Doksum. Distribution and power of the absolute normal scores test. Journal of the American Statistical Association, 62(319):966–975, September 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2283685.

Theil:1963:UIP

[The63] H. Theil. On the use of incomplete prior information in regression analysis. Journal of the American Statistical Association, 58 (302):401–414, June 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2283275.

Theil:1965:ADR

[The65a] H. Theil. The analysis of disturbances in regression analysis. Journal of the American Statistical Association, 60(312):1067–1079, December 1965. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2283406.

Theophilos:1965:CRB

[The65b] Cacoullos Theophilos. Corrigenda: A Relation Between t and F - Distributions. Journal of the American Statistical Association, 60 (312):1249, December 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2283446. See [Cac65].

Theil:1968:SBP

[The68] Henri Theil. A simplification of the BLUS procedure for analyzing regression disturbances. Journal of the American Statistical Association, 63(321):242–251, March 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283844.

Thimm:1963:BRB

[Thi63] Alfred L. Thimm. Book review: [untitled]. Journal of the American Statistical Association, 58(304):1201–1202, December 1963. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283383.

Thomlinson:1961:MMA

[Tho61a] Ralph Thomlinson. A model for migration analysis. Journal of the American Statistical Association, 56(295):675–686, September 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2282088.

Thompson:1961:BRB

[Tho61b] Gerald L. Thompson. Book review: [untitled]. Journal of the American Statistical Association, 56(293):176–178, March 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282351.

Thompson:1962:BRB

[Tho62] Gerald L. Thompson. Book review: [untitled]. Journal of the American Statistical Association, 57(300):927–928, December 1962. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281831.

Thompson:1963:PSM

[Tho63a] W. A. Thompson, Jr. Precision of simultaneous measurement procedures. Journal of the American Statistical Association, 58(302): 474–479, June 1963. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2283281.

Thorndike:1963:BRB

[Tho63b] Robert L. Thorndike. Book review: [untitled]. Journal of the American Statistical Association, 58(304):1182–1183, December 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283366.

Thornton:1964:BRB

[Tho64] Judith Thornton. Book review: [untitled]. Journal of the American Statistical Association, 59(305):306–308, March 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282892.

Thompson:1966:BOS

[Tho66] Rory Thompson. Bias of the one-sample Cramer–von Mises test. Journal of the American Statistical Association, 61(313):246– 247, March 1966. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2283059.

Thompson:1967:BRB

[Tho67a] G. L. Thompson. Book review: [untitled]. Journal of the American Statistical Association, 62(318):709–710, June 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284007.

Thornber:1967:FSM

[Tho67b] Hodson Thornber. Finite sample Monte Carlo studies: An autoregressive illustration. Journal of the American Statistical Association, 62(319):801–818, September 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283672.

Thompson:1968:ABE

[Tho68a] James R. Thompson. Accuracy borrowing in the estimation of the mean by shrinkage to an interval. Journal of the American Statistical Association, 63(323):953–963, September 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283886.

Thompson:1968:SST

[Tho68b] James R. Thompson. Some shrinkage techniques for estimating the mean. Journal of the American Statistical Association, 63 (321):113–122, March 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2283832.

Thoni:1969:TEM

[Tho69] Hanspeter Thoni. A table for estimating the mean of a lognormal distribution. Journal of the American Statistical Association, 64 (326):632–636, June 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2283646.

Taeuber:1961:RHE

[THS61] Karl E. Taeuber, William Haenszel, and Monroe G. Sirken. Residence histories and exposure residences for the United States population. Journal of the American Statistical Association, 56(296): 824–834, December 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2281997.

Tiao:1967:BRB

[Tia67] George C. Tiao. Book review: Inference and Disputed Authorship: The Federalist, by Frederick Mosteller; David L. Wallace. Journal of the American Statistical Association, 62(317):306–309, March 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2282943.

Tiebout:1962:BRB

[Tie62] Charles M. Tiebout. Book review: [untitled]. Journal of the American Statistical Association, 57(297):250–251, March 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282481.

Tiku:1964:NNM

[Tik64] M. L. Tiku. A note on the negative moments of a truncated Poisson variate. Journal of the American Statistical Association, 59 (308):1220–1224, December 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2282638.

Tiku:1967:TPT

[Tik67] M. L. Tiku. Tables of the power of the F -test. Journal of the American Statistical Association, 62(318):525–539, June 1967. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283980. See corrigenda

[Tik68a].

Tiku:1968:CTP

[Tik68a] M. L. Tiku. Corrigenda: Tables of the Power of the F Test. Journal of the American Statistical Association, 63(324):1551, December 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2285922. See

[Tik67].

Tiku:1968:EPL

[Tik68b] M. L. Tiku. Estimating the parameters of log-normal distribution from censored samples. Journal of the American Statistical Association, 63(321):134–140, March 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283834.

Tilanus:1965:BRB

[Til65] C. B. Tilanus. Book review: [untitled]. Journal of the American Statistical Association, 60(312):1237, December 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283433.

Tintner:1961:SWO

[Tin61] Gerhard Tintner. The statistical work of oskar Anderson. Journal of the American Statistical Association, 56(294):273–280, June 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2282252.

Tin:1965:CSR

[Tin65] Myint Tin. Comparison of some ratio estimators. Journal of the American Statistical Association, 60(309):294–307, March 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283154.

Tinsley:1967:AVW

[Tin67] P. A. Tinsley. An application of variable weight distributed lags. Journal of the American Statistical Association, 62(320):1277– 1289, December 1967. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2283777.

Thompson:1961:NFS

[TK61] D. J. Thompson and D. Kodlin. A note on follow-up for survival in the presence of movement. Journal of the American Statistical Association, 56(293):119–124, March 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282337.

Theil:1961:TIR

[TN61] H. Theil and A. L. Nagar. Testing the independence of regression disturbances. Journal of the American Statistical Association, 56 (296):793–806, December 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281995.

Tobler:1969:BRB

[Tob69] W. R. Tobler. Book review: [untitled]. Journal of the American Statistical Association, 64(325):391–392, March 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283750.

Tomasson:1961:BEU

[Tom61] Richard F. Tomasson. Bias in estimates of the U.S. nonwhite population as indicated by trends in death rates. Journal of the American Statistical Association, 56(293):44–51, March 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282329.

Trawinski:1967:AOZ

[Tra67] Irene M. Trawinski. An algorithm for obtaining the zero of a function of the dispersion matrix in multivariate analysis. Journal of the American Statistical Association, 62(317):114–123, March 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2282914.

Trinkl:1967:BRB

[Tri67] Frank H. Trinkl. Book review: [untitled]. Journal of the American Statistical Association, 62(318):707–708, June 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284005.

Troy:1963:BRB

[Tro63] Leo Troy. Book review: [untitled]. Journal of the American Statistical Association, 58(302):591–592, June 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283311.

Tsutakawa:1967:RWD

[Tsu67] R. K. Tsutakawa. Random walk design in bio-assay. Journal of the American Statistical Association, 62(319):842–856, September 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2283674.

Tukey:1960:WDW

[Tuk60] John W. Tukey. Where do we go from here? Journal of the American Statistical Association, 55(289):80–93, March 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282180.

Turner:1965:BRB

[Tur65] Malcolm E. Turner. Book review: [untitled]. Journal of the American Statistical Association, 60(312):1223–1225, December 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283425.

Tweedie:1968:ICO

[TV68] M. C. K. Tweedie and A. Veevers. The inversion of cumulant operators for power-series distributions, and the approximate stabilization of variance by transformations. Journal of the American Statistical Association, 63(321):321–328, March 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283851. See corrigenda [TV69].

Tweedie:1969:CIC

[TV69] M. C. K. Tweedie and A. Veevers. Corrigenda: The Inversion of Cumulant Operators for Power-Series Distributions, and the Approximate Stabilization of Variance by Transformations. Journal of the American Statistical Association, 64(328):1703, December 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2286145. See

[TV68].

Toro-Vizcarrondo:1968:TMS

[TVW68] Carlos Toro-Vizcarrondo and T. D. Wallace. A test of the mean square error criterion for restrictions in linear regression. Journal of the American Statistical Association, 63(322):558–572, June 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2284027.

Thorp:1966:FSB

[TW66] Edward O. Thorp and William E. Walden. A favorable side bet in Nevada baccarat. Journal of the American Statistical Association, 61(314):313–328, June 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2282821.

Ury:1966:NTC

[Ury66] Hans K. Ury. A note on taking a covariable into account. Journal of the American Statistical Association, 61(314):490–495, June 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282838.

Vajda:1964:BRB

[Vaj64] S. Vajda. Book review: [untitled]. Journal of the American Statistical Association, 59(306):620–621, June 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283027.

Vajda:1966:BRB

[Vaj66] S. Vajda. Book review: Management and Mathematics, by Allan Fletcher; Geoffrey Clarke. Journal of the American Statistical Association, 61(314):526–527, June 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282843.

Vance:1963:BRB

[Van63] Lawrence L. Vance. Book review: [untitled]. Journal of the American Statistical Association, 58(302):578–579, June 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283301.

VanRyzin:1969:BRU

[Van69] J. Van Ryzin. Book review: [untitled]. Journal of the American Statistical Association, 64(326):675–676, June 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283658.

Varde:1969:LTR

[Var69] S. D. Varde. Life testing and reliability estimation for the two parameter exponential distribution. Journal of the American Statistical Association, 64(326):621–631, June 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283645.

vanEeden:1964:NCS

[vE64] Constance van Eeden. Note on the consistency of some distributionfree tests for dispersion. Journal of the American Statistical Association, 59(305):105–119, March 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282862.

Veinott:1964:BRBb

[Vei64a] Arthur F. Veinott. Book review: [untitled]. Journal of the American Statistical Association, 59(307):976–977, September 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283124.

Veinott:1964:BRBa

[Vei64b] Arthur F. Veinott, Jr. Book review: [untitled]. Journal of the American Statistical Association, 59(305):283–285, March 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282878.

Veinott:1968:BRB

[Vei68] Arthur F. Veinott, Jr. Book review: [untitled]. Journal of the American Statistical Association, 63(323):1063, September 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283909.

Venneman:1961:BRB

[Ven61] Harry Venneman. Book review: [untitled]. Journal of the American Statistical Association, 56(295):749, September 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282100.

Vinod:1969:IPT

[Vin69] Hrishikesh D. Vinod. Integer programming and the theory of grouping. Journal of the American Statistical Association, 64(326): 506–519, June 1969. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2283635.

Whitman:1968:BRB

[vNW68] Marina v. N. Whitman. Book review: [untitled]. Journal of the American Statistical Association, 63(323):1051–1054, September 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2283901.

W:1960:BRB

[W.60] D. L. W. Book review: [untitled]. Journal of the American Statistical Association, 55(290):394, June 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281768.

Waugh:1967:FPM

[WA67] Frederick V. Waugh and Martin E. Abel. On fractional powers of a matrix. Journal of the American Statistical Association, 62 (319):1018–1021, September 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283690.

Wagner:1960:BRB

[Wag60] Harvey M. Wagner. Book review: [untitled]. Journal of the American Statistical Association, 55(291):610–611, September 1960. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281937.

Wagner:1961:BRB

[Wag61] Harvey M. Wagner. Book review: [untitled]. Journal of the American Statistical Association, 56(293):175–176, March 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282350.

Wagner:1962:BRB

[Wag62a] Harvey M. Wagner. Book review: [untitled]. Journal of the American Statistical Association, 57(297):257–261, March 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282485.

Wagner:1962:NLR

[Wag62b] Harvey M. Wagner. Non-linear regression with minimal assumptions. Journal of the American Statistical Association, 57(299): 572–578, September 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2282394.

Wahi:1969:BRB

[Wah69a] Pran N. Wahi. Book review: [untitled]. Journal of the American Statistical Association, 64(327):1098–1099, September 1969. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283500.

Wahl:1969:BRB

[Wah69b] Pran N. Wahl. Book review: [untitled]. Journal of the American Statistical Association, 64(328):1684, December 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286116.

Walker:1961:BRB

[Wal61] Francis Walker. Book review: [untitled]. Journal of the American Statistical Association, 56(296):1021–1022, December 1961. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282023.

Walsh:1962:STS

[Wal62] John E. Walsh. Some two-sided distribution-free tolerance intervals of a general nature. Journal of the American Statistical Association, 57(300):775–784, December 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281808.

Walsh:1963:BRB

[Wal63] John E. Walsh. Book review: [untitled]. Journal of the American Statistical Association, 58(301):251–253, March 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282972.

Wallace:1964:ESR

[Wal64a] T. D. Wallace. Efficiencies for stepwise regressions. Journal of the American Statistical Association, 59(308):1179–1182, December 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2282632.

Walsh:1964:ADE

[Wal64b] John E. Walsh. Approximate distribution of extremes for nonsample cases. Journal of the American Statistical Association, 59 (306):429–436, June 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2282997.

Waldorp:1965:DMS

[Wal65] William H. Waldorp. Demand for manufacturers’ services for bakery products and fruits and vegetables. Journal of the American Statistical Association, 60(311):740–749, September 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283242.

Wales:1966:EAD

[Wal66a] Terence J. Wales. Estimation of an accelerated depreciation learning function. Journal of the American Statistical Association, 61 (316):995–1009, December 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283195.

Wallis:1966:ESE

[Wal66b] W. Allen Wallis. Economic statistics and economic policy. Journal of the American Statistical Association, 61(313):1–10, March 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283042.

Wallenius:1967:SC

[Wal67] K. T. Wallenius. Sampling for confidence. Journal of the American Statistical Association, 62(318):540–547, June 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283981.

Ward:1963:HGO

[War63a] Joe H. Ward, Jr. Hierarchical grouping to optimize an objective function. Journal of the American Statistical Association, 58(301):236–244, March 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2282967.

Warner:1963:MRD

[War63b] Stanley L. Warner. Multivariate regression of dummy variates under normality assumptions. Journal of the American Statistical Association, 58(304):1054–1063, December 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283332.

Warner:1965:RRS

[War65] Stanley L. Warner. Randomized response: a survey technique for eliminating evasive answer bias. Journal of the American Statistical Association, 60(309):63–69, March 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283137.

Warner:1967:AVD

[War67] Stanley L. Warner. Asymptotic variances for dummy variate regression under normality assumptions. Journal of the American Statistical Association, 62(320):1305–1314, December 1967. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283780.

Watson:1960:BRB

[Wat60a] Walter B. Watson. Book review: [untitled]. Journal of the American Statistical Association, 55(291):613–614, September 1960. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281940.

Watts:1960:BRB

[Wat60b] Harold W. Watts. Book review: [untitled]. Journal of the American Statistical Association, 55(292):766–767, December 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281613.

Watts:1963:BRB

[Wat63] Harold W. Watts. Book review: [untitled]. Journal of the American Statistical Association, 58(303):842–843, September 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282741.

Watts:1965:BRB

[Wat65] Harold W. Watts. Book review: [untitled]. Journal of the American Statistical Association, 60(311):930–931, September 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283263.

Waud:1966:SSB

[Wau66] Roger N. Waud. Small sample bias due to misspecification in the “partial adjustment” and “adaptive expectations” models. Journal of the American Statistical Association, 61(316):1130–1152, December 1966. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2283204.

Waller:1969:BRS

[WD69] Ray A. Waller and David B. Duncan. A Bayes rule for the symmetric multiple comparisons problems. Journal of the American Statistical Association, 64(328):1484–1503, December 1969. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2286085.

Webster:1967:BRB

[Web67] J. T. Webster. Book review: [untitled]. Journal of the American Statistical Association, 62(319):1092–1093, September 1967. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283710.

Weiss:1960:BRB

[Wei60] Lionel Weiss. Book review: [untitled]. Journal of the American Statistical Association, 55(291):584–585, September 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281918.

Weiler:1961:BRB

[Wei61a] E. T. Weiler. Book review: [untitled]. Journal of the American Statistical Association, 56(293):175, March 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282349.

Weingarten:1961:BRB

[Wei61b] Harry Weingarten. Book review: [untitled]. Journal of the American Statistical Association, 56(293):187–188, March 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282360.

Weiss:1962:STW

[Wei62a] Lionel Weiss. On sequential tests which minimize the maximum expected sample size. Journal of the American Statistical Association, 57(299):551–566, September 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2282392.

Weiss:1962:STE

[Wei62b] Lionel Weiss. A sequential test of the equality of probabilities in a multinomial distribution. Journal of the American Statistical Association, 57(300):769–774, December 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281807.

Weiss:1963:BRBa

[Wei63a] Lionel Weiss. Book review: [untitled]. Journal of the American Statistical Association, 58(301):250, March 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282970.

Weiss:1963:BRBb

[Wei63b] Lionel Weiss. Book review: [untitled]. Journal of the American Statistical Association, 58(302):579–580, June 1963. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2283302.

Weiss:1963:ESL

[Wei63c] Lionel Weiss. On estimating scale and location parameters. Journal of the American Statistical Association, 58(303):658–659, September 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2282713.

Weiss:1966:BRB

[Wei66] Lionel Weiss. Book review: [untitled]. Journal of the American Statistical Association, 61(313):273, March 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283066.

Weiss:1967:BRB

[Wei67] Lionel Weiss. Book review: [untitled]. Journal of the American Statistical Association, 62(318):694–695, June 1967. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2283999.

Wells:1963:BRB

[Wel63] John D. Wells. Book review: [untitled]. Journal of the American Statistical Association, 58(304):1186–1187, December 1963. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283370.

Wendel:1965:BRB

[Wen65] J. G. Wendel. Book review: [untitled]. Journal of the American Statistical Association, 60(312):1210–1213, December 1965. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283419.

Weston:1960:BRB

[Wes60] J. Fred Weston. Book review: [untitled]. Journal of the American Statistical Association, 55(291):618–619, September 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281946.

Westfield:1967:BRB

[Wes67] Fred M. Westfield. Book review: [untitled]. Journal of the American Statistical Association, 62(318):713–715, June 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284010.

Wilk:1963:EEV

[WGF63] M. B. Wilk, R. Gnandesikan, and Anne E. Freeny. Estimation of error variance from smallest ordered contrasts. Journal of the American Statistical Association, 58(301):152–160, March 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282959.

Wharton:1960:PUD

[Wha60] Clifton R. Wharton, Jr. Processing underdeveloped data from an underdeveloped area. Journal of the American Statistical Association, 55(289):23–37, March 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2282177.

Whinston:1966:BRB

[Whi66a] Andrew Whinston. Book review: [untitled]. Journal of the American Statistical Association, 61(315):874–875, September 1966. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282801.

White:1966:EHC

[Whi66b] Leon S. White. The evaluation of H 106 continuous sampling plans under the assumption of worst conditions. Journal of the American Statistical Association, 61(315):833–841, September 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282792.

Whitney:1966:BRB

[Whi66c] William G. Whitney. Book review: [untitled]. Journal of the American Statistical Association, 61(313):275–276, March 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283068.

Wilkins:1961:PCW

[Wil61a] Coleridge A. Wilkins. A problem concerned with weighting of distributions. Journal of the American Statistical Association, 56 (294):281–292, June 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2282253.

Williams:1961:BRB

[Wil61b] Frank J. Williams. Book review: [untitled]. Journal of the American Statistical Association, 56(294):460–461, June 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282305.

Williams:1962:TMU

[Wil62a] W. H. Williams. On two methods of unbiased estimation with auxiliary variates. Journal of the American Statistical Association, 57(297):184–186, March 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2282448.

Williams:1962:VEP

[Wil62b] W. H. Williams. The variance of an estimator with post-stratified weighting. Journal of the American Statistical Association, 57 (299):622–627, September 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2282399.

Wilber:1963:MEU

[Wil63] George L. Wilber. Migration expectancy in the United States. Journal of the American Statistical Association, 58(302):444–453, June 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2283278.

Williams:1964:SSC

[Wil64] W. H. Williams. Sample selection and the choice of estimator in two-way stratified populations. Journal of the American Statistical Association, 59(308):1054–1062, December 1964. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2282623.

Willke:1965:BRB

[Wil65a] T. A. Willke. Book review: [untitled]. Journal of the American Statistical Association, 60(310):676–677, June 1965. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2282701.

Wilson:1965:BRB

[Wil65b] Robert Wilson. Book review: [untitled]. Journal of the American Statistical Association, 60(311):932–933, September 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283265.

Wilkerson:1967:SEC

[Wil67a] Marvin Wilkerson. Sampling error in the consumer price index. Journal of the American Statistical Association, 62(319):899–914, September 1967. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2283678.

Williams:1967:VWR

[Wil67b] J. S. Williams. The variance of weighted regression estimators. Journal of the American Statistical Association, 62(320):1290– 1301, December 1967. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2283778.

Wilson:1967:BRB

[Wil67c] Robert Wilson. Book review: [untitled]. Journal of the American Statistical Association, 62(318):695–697, June 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284000.

Winter:1961:BRB

[Win61] Sidney G. Winter, Jr. Book review: [untitled]. Journal of the American Statistical Association, 56(296):1022–1023, December 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282024.

Winer:1963:BRB

[Win63] B. J. Winer. Book review: [untitled]. Journal of the American Statistical Association, 58(302):580–581, June 1963. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2283303.

Winer:1964:BRB

[Win64] B. J. Winer. Book review: [untitled]. Journal of the American Statistical Association, 59(308):1302–1305, December 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282654.

Winkler:1967:APD

[Win67a] Robert L. Winkler. The assessment of prior distributions in Bayesian analysis. Journal of the American Statistical Association, 62(319):776–800, September 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283671.

Winkler:1967:QJS

[Win67b] Robert L. Winkler. The quantification of judgment: Some methodological suggestions. Journal of the American Statistical Association, 62(320):1105–1120, December 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283764.

Winkler:1969:SRE

[Win69] Robert L. Winkler. Scoring rules and the evaluation of probability assessors. Journal of the American Statistical Association, 64 (327):1073–1078, September 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283486.

Witt:1962:BRBa

[Wit62a] Lawrence W. Witt. Book review: [untitled]. Journal of the American Statistical Association, 57(298):518, June 1962. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2281669.

Witt:1962:BRBb

[Wit62b] Lawrence W. Witt. Book review: [untitled]. Journal of the American Statistical Association, 57(300):944, December 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281843.

Wasserman:1966:PAL

[WN66] William Wasserman and John Neter. Potentials in applying linear programming to the consumer price index. Journal of the American Statistical Association, 61(316):982–994, December 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283194.

Wold:1960:BRB

[Wol60] Herman Wold. Book review: [untitled]. Journal of the American Statistical Association, 55(290):376–377, June 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281751.

Wolfbein:1961:BRB

[Wol61] Seymour L. Wolfbein. Book review: [untitled]. Journal of the American Statistical Association, 56(293):174–175, March 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282348.

Wold:1965:BRB

[Wol65] Herman O. A. Wold. Book review: [untitled]. Journal of the American Statistical Association, 60(312):1213–1215, December 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283420.

Wolberg:1969:CPA

[Wol69] John R. Wolberg. Corrigenda: Prediction Analysis. Journal of the American Statistical Association, 64(328):1700, December 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2286132.

Woodruff:1963:URS

[Woo63] Ralph S. Woodruff. The use of rotating samples in the census Bureau’s monthly surveys. Journal of the American Statistical Association, 58(302):454–467, June 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283279.

Woodruff:1966:URT

[Woo66] Ralph S. Woodruff. Use of a regression technique to produce area breakdowns of the monthly national estimates of retail trade. Journal of the American Statistical Association, 61(314):496–504, June 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2282839.

Workman:1963:BRB

[Wor63] P. L. Workman. Book review: [untitled]. Journal of the American Statistical Association, 58(302):575, June 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283297.

Waksberg:1969:CCM

[WP69] Joseph Waksberg and Leon Pritzker. Changes in census methods. Journal of the American Statistical Association, 64(328): 1141–1149, December 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286056.

Wright:1960:BRB

[Wri60] C. Ashley Wright. Book review: [untitled]. Journal of the American Statistical Association, 55(289):230–231, March 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282196.

Wallace:1969:TMS

[WTV69] T. D. Wallace and C. E. Toro-Vizcarrondo. Tables for the mean square error test for exact linear restrictions in regression. Journal of the American Statistical Association, 64(328):1649–1663, December 1969. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2286097.

Yancey:1960:BRB

[Yan60] T. A. Yancey. Book review: [untitled]. Journal of the American Statistical Association, 55(291):602–603, September 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281930.

Yancey:1962:BRB

[Yan62] T. A. Yancey. Book review: [untitled]. Journal of the American Statistical Association, 57(298):518–519, June 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281670.

Yaspan:1969:CEP

[Yas69] Arthur Yaspan. Corrigenda: Essentials of Probability. Journal of the American Statistical Association, 64(328):1699, December 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2286130.

Yen:1968:BRB

[Yen68] Elizabeth H. Yen. Book review: [untitled]. Journal of the American Statistical Association, 63(322):728, June 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284048.

Yerushalmy:1961:BRB

[Yer61] J. Yerushalmy. Book review: [untitled]. Journal of the American Statistical Association, 56(294):440–441, June 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282287.

Youden:1962:MSR

[You62] W. J. Youden. Memorial to Sir Ronald Aylmer Fisher: 1890– 1962. Journal of the American Statistical Association, 57(300): 727–728, December 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2281804.

Young:1968:LAC

[You68] Allan H. Young. Linear approximations to the census and BLS seasonal adjustment methods. Journal of the American Statistical Association, 63(322):445–471, June 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284019.

Zabel:1966:BRB

[Zab66] Edward Zabel. Book review: [untitled]. Journal of the American Statistical Association, 61(315):867–868, September 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282797.

Zachariah:1962:NCS

[Zac62] K. C. Zachariah. A note on the census survival ratio method of estimating net migration. Journal of the American Statistical Association, 57(297):175–183, March 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2282447.

Zacks:1966:UEC

[Zac66] S. Zacks. Unbiased estimation of the common mean of two normal distributions based on small samples of equal size. Journal of the American Statistical Association, 61(314):467–476, June 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282836.

Zacks:1969:BSD

[Zac69] Shelley Zacks. Bayes sequential designs of fixed size samples from finite populations. Journal of the American Statistical Association, 64(328):1342–1349, December 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286072.

Zaring:1960:BRB

[Zar60] Wilson M. Zaring. Book review: [untitled]. Journal of the American Statistical Association, 55(292):783–784, December 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281628.

Zarnowitz:1967:BRB

[Zar67] Victor Zarnowitz. Book review: [untitled]. Journal of the American Statistical Association, 62(320):1500–1502, December 1967. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283797.

Zarembka:1968:FFD

[Zar68] Paul Zarembka. Functional form in the demand for money. Journal of the American Statistical Association, 63(322):502–511, June 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2284021.

Zellner:1965:PDP

[ZC65] Arnold Zellner and V. Karuppan Chetty. Prediction and decision problems in regression models from the Bayesian point of view. Journal of the American Statistical Association, 60(310):608–616, June 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2282695. See

corrigenda [ZZK68, ZC69].

Zellner:1969:CPD

[ZC69] Arnold Zellner and V. Karuppan Chetty. Corrigenda: Prediction and Decision Problems in Regression Models from the Bayesian Point of View. Journal of the American Statistical Association, 64(328):1703, December 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286146. See [ZC65].

Zacks:1966:ESS

[ZE66a] S. Zacks and M. Even. The efficiencies in small samples of the maximum likelihood and best unbiased estimators of reliability functions. Journal of the American Statistical Association, 61(316): 1033–1051, December 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2283198.

Zacks:1966:MVU

[ZE66b] S. Zacks and M. Even. Minimum variance unbiased and maximum likelihood estimators of reliability functions for systems in series and in parallel. Journal of the American Statistical Association, 61 (316):1052–1062, December 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283199.

Zehna:1961:BRB

[Zeh61] Peter Zehna. Book review: [untitled]. Journal of the American Statistical Association, 56(296):1025–1026, December 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282030.

Zeigler:1965:UTC

[Zei65] R. K. Zeigler. A uniqueness theorem concerning moment distributions. Journal of the American Statistical Association, 60(312): 1203–1206, December 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2283417.

Zellner:1962:EME

[Zel62] Arnold Zellner. An efficient method of estimating seemingly unrelated regressions and tests for aggregation bias. Journal of the American Statistical Association, 57(298):348–368, June 1962. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281644.

Zellner:1963:ESU

[Zel63] Arnold Zellner. Estimators for seemingly unrelated regression equations: Some exact finite sample results. Journal of the American Statistical Association, 58(304):977–992, December 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283326.

Zelnik:1964:ECE

[Zel64] Melvin Zelnik. Errors in the 1960 census enumeration of native whites. Journal of the American Statistical Association, 59(306): 437–459, June 1964. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2282998.

Zelen:1965:BRB

[Zel65] Marvin Zelen. Book review: [untitled]. Journal of the American Statistical Association, 60(310):681–682, June 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282705.

Zelen:1969:PWR

[Zel69a] M. Zelen. Play the winner rule and the controlled clinical trial. Journal of the American Statistical Association, 64(325): 131–146, March 1969. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2283724.

Zelnik:1969:APM

[Zel69b] Melvin Zelnik. Age patterns of mortality of American Negroes: 1900–02 to 1959–61. Journal of the American Statistical Association, 64(326):433–451, June 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283631.

Ziegler:1962:BRB

[Zie62] Raymond J. Ziegler. Book review: [untitled]. Journal of the American Statistical Association, 57(297):242–243, March 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282474.

Zinnes:1961:BRB

[Zin61] Joseph L. Zinnes. Book review: [untitled]. Journal of the American Statistical Association, 56(294):430–433, June 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282283.

Zorn:1960:BRB

[Zor60] Eugene C. Zorn, Jr. Book review: [untitled]. Journal of the American Statistical Association, 55(291):607–608, September 1960. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281934.

Zellner:1964:BAR

[ZT64] Arnold Zellner and George C. Tiao. Bayesian analysis of the regression model with autocorrelated errors. Journal of the American Statistical Association, 59(307):763–778, September 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283097.

Zyskind:1963:NRA

[Zys63] George Zyskind. A note on residual analysis. Journal of the American Statistical Association, 58(304):1125–1132, December 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283338.

Zyskind:1969:PAE

[Zys69] George Zyskind. Parametric augmentations and error structures under which certain simple least squares and analysis of variance procedures are also best. Journal of the American Statistical Association, 64(328):1353–1368, December 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286074.

Zellner:1968:CPD

[ZZK68] Arnold Zellner, Chetty Zellner, and V. Karuppan. Corrigenda: Prediction and Decision Problems in Regression Models from the Bayesian Point of View. Journal of the American Statistical Association, 63(324):1551, December 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285923. See [ZC65].